### v1

# Stability of Non-Hyperbolic Sets

Abstract
In families of unfoldings we find codimension-one laminations consisting of maps exibiting an invariant, non-hyperbolic Cantor set containing a Collet–Eckmann point whose orbit is dense in the set. The topology of this Cantor set remains unchanged along the leaves of the lamination, ensuring codimension-one stability. Furthermore, within this lamination we identify a sublamination whose leaves contain maps with infinitely many sinks accumulating on the non-hyperbolic Cantor set containing the Collet–Eckmann point. These results provide a new perspective on the stability of chaotic non-hyperbolic dynamics in two-dimensional systems and reveal surprising coexistence phenomena that highlight the richness and complexity of such dynamics.
1 Introduction
This paper addresses two central questions in dynamical systems: the existence and stability of dynamically relevant sets. Given a dynamical system $(\mathcal{M},f)$ where $\mathcal{M}$ is a manifold and $f$ an evolution acting on it, one aims to identify the regions where the dynamics effectively take place—that is, the sets on which most orbits spend most of their time. Establishing the existence of such sets can already be a challenging task. Once their existence is established, one naturally seeks to understand whether these sets, together with their dynamical properties, persist under perturbations of the system. Are these sets isolated? Are their properties robust and observable across nearby systems? In other words, are they stable? Let us now make the notion of stability more precise.
Given two dynamical systems $(M,f)$ and $(M,g)$ with relevant sets $A(f)$ and $A(g)$, we say that they are conjugate if there exists a homeomorphism $h:A(f)\to A(g)$ such that $h\circ f=g\circ h.$ One can think of this homeomorphism as a change of coordinates. For instance, if $A(f)$ and $A(g)$ are Cantor sets, the conjugacy map “rearranges” the pieces of one Cantor set to produce the other. Roughly speaking, this means that the two sets share the same topological structure, they are topologically equivalent.
Within the space of dynamical systems, one can thus consider the class of maps whose relevant sets are topologically equivalent in this sense. The geometry of this class provides information about the stability of the corresponding sets. We say that the set $A(f)$ is codimension-$k$ stable if the class of systems conjugate to it forms a codimension-$k$ submanifold of the space of dynamical systems. This manifold is referred to as the manifold of stability.
The sets exhibiting the strongest form of stability are sinks or closed attracting periodic orbits. Indeed, such sets are codimension-0 stable: their topology does not change under small perturbations of the system. More interesting examples arise in one-dimensional dynamics, such as $\mathcal{C}^{3}$ circle maps. In this class of systems, the topology of the relevant set (in this case, the entire circle) persists on codimension-one manifolds. The set is therefore codimension-$1$ stable. This phenomenon follows from one of the most profound results in one-dimensional dynamics, namely Herman’s Theorem, see [28]; for more modern expositions, see [20] and [30].
The stability of sets in one-dimensional dynamics is by now largely understood, and no major surprises are expected. Obtaining this complete picture required several decades of work and the efforts of many leading mathematicians. Although the main questions have been settled, a number of subtle and intriguing problems remain open.
In contrast, the situation in two dimensions is far richer and less complete. We now turn to the family of systems at the center of our study, namely the Hénon family (see Section 3 for a precise definition). In fact, in this paper we consider a broader class of families, of which the Hénon family is only one example. Specifically, we study unfoldings of maps with a homoclinic tangency and a transversal intersection. The precise definition is given in Subsection 2.1.
Which relevant sets are known to exist in such systems and what do we know about their stability?
As expected, finitely many sinks can exist, and these are codimension-0 stable. Indeed, sinks are hyperbolic sets. The existence of infinitely many sinks has also been established (see [33]), together with some preliminary results concerning their stability (see [23]). More intricate dynamics have likewise been shown to occur. In particular, the existence of a period-doubling Cantor attractor—an invariant, non-hyperbolic, minimal set carrying a measure with zero Lyapunov exponent—has been proved (see [14, 31]). This set is known to be codimension- 1 stable. Furthermore, the existence of two period-doubling Cantor attractors has been demonstrated, and these are codimension-2 stable (see [36]).
Even more complex invariant sets, known as strange attractors, have been shown to exist in this family. Strange attractors are invariant, non-hyperbolic, and chaotic sets (they carry a measure having positive Lyapunov exponent). The seminal work establishing their existence is [3], where a remarkably intricate and ingenious proof was developed. Similar constructions adapted to more general systems and further studies were later given in [7, 4, 6, 27, 32, 43, 44]. Coexistence of these sets is established in [5]. What makes these sets chaotic is the existence of a point satisfying the Collet–Eckmann condition, which, roughly speaking, asserts that the derivative of the map grows at least exponentially fast along the orbit of this point (see Definition 14 for a precise formulation).
A natural question then arises: what can be said about the stability of strange attractors? Their chaotic nature makes this question extremely challenging. It is known that, in parameter space, there exists a positive Lebesgue measure set of maps exhibiting such attractors; however, their topology changes as the parameter varies. So far, we lack the tools to clearly address this problem. The present paper aims to provide a partial answer to this question. A new construction and new tools are given. Furthermore, while trying to answer this question, we found new fascinating phenomena, namely the existence of maps with infinitely many sinks accumulating on a set containing a Collet-Eckmann point.
We now summarize our main results, which together describe the existence and stability of non-hyperbolic Cantor sets in families of unfoldings.”
Theorem A.
Let $\mathcal{M}$ be a two-dimensional manifold and let $F:\left([-t_{0},t_{0}]\times[-a_{0},a_{0}]\right)\times\mathcal{M}\to\mathcal{M}$ be a family of unfolding. There exists a codimension-$1$ lamination $\mathfrak{L}$ in parameter space such that, each leave $\ell\subset\mathfrak{L}$ and each map $f\in\ell$ have a set $A(f)$ satisfying the following.
  • (i)
    The set $A(f)$ in an invariant, non-trivial, non-hyperbolic Cantor set.
  • (ii)
    The set $A(f)$ contains a point with dense orbit.
  • (iii)
    The set $A(f)$ contains a point whose limit set coincides with the orbit of a transversal homoclinic intersection point.
  • (iv)
    The set $A(f)$ has a unique invariant probability measure. This invariant measure is the point mass on the fixed point in $A(f)$ .
  • (v)
    Every leaves of $\mathfrak{L}$ is the graph of a smooth function $\ell:[-t_{0},t_{0}]\mapsto[-a_{0},a_{0}]$. Moreover if $t\in[-t_{0},t_{0}]$, the closure of the set $L_{t}=\left\{(t,a)\in\mathfrak{L}\right\}$ is a Cantor set.
  • (vi)
    If $f$ and $\tilde{f}$ are in the same leave $\ell$ of the lamination $\mathfrak{L}$ then they are conjugate on their respective Cantor sets.
  • (vii)
    The set $A(f)$ is codimension-$1$ stable and $\ell$ is the smooth manifold of stability.
Theorem $A$ is the summary of several results proved along the paper. In order: Lemma 22, Corollary 1, Lemma 24, Lemma 25, Theorem 6, Proposition 9, Theorem 4, Theorem 5.
Theorem A establishes the existence of an invariant, non-hyperbolic Cantor set that is codimension-1 stable. In particular, any two maps lying on the same leaf of the lamination are conjugate on their corresponding Cantor sets. In other words, along each leaf of the lamination, the corresponding Cantor sets are topologically conjugate, and their structure remains invariant.
These results are obtained through the introduction of dynamical partitions and the associated directed graphs. The inverse limit of these graphs is homeomorphic to the Cantor set. This combinatorial structure provides effective control over both the topological dynamics along the leaves and the measure-theoretical properties of the Cantor set. These inverse limits and directed graphs were introduced in [21]. In the present work, we adapt these ideas to the setting of families of maps exhibiting homoclinic tangencies.
To formally establish the codimension-one stability of the Cantor set, we introduce the notion of infinitely renormalizable maps (see Subsection 10.3). For simplicity, we refer to this process as renormalization, although it is, strictly speaking, a pre-renormalization. In particular, we study the first-return maps to the sets generating the dynamical partition, without performing any rescaling. Since our main focus lies on topological properties, this pre-renormalization suffices for our purposes. To the best of our knowledge, this represents the first attempt to formulate a generalized renormalization framework with two branches in the setting of two-dimensional systems. A formulation of renormalization for two-dimensional systems using only one branch has been explored in [8, 14, 16, 17, 36], but our construction differs in that it employs two branches and more closely resembles generalized renormalization as used for one-dimensional systems.
Let us go back to analyze the statements in Theorem A. In particular, one should compare these results with those of [14], where the period-doubling Cantor attractor was shown to be codimension-1 stable. In that case, the topology of the Cantor set also persists along a codimension-1 leave; however, the constructed leave is very short. Numerical experiments suggest that the period-doubling locus should, in fact, form a long global curve, but proving this remains a challenging open problem. In contrast, our approach allows for the construction of global curves that are genuine manifolds of stability.
We next strengthen these results by showing the existence of a sub-lamination such that the corresponding non-hyperbolic invariant Cantor sets contain a Collet–Eckmann point with a dense orbit in the Cantor set. These sets, which exhibit chaotic dynamics, are likewise codimension-1 stable, i.e. their topology is preserved along the leaves of the lamination. This result represents an initial step toward the study of stability of strange attractors. To the best of our knowledge, no prior results are available in this direction. The proof of the theorem relies on the dynamical structure described above, together with a detailed analytical control along the tangency loci.
Theorem B.
There exists a sub-lamination $\mathfrak{L}_{CE}\subset\mathfrak{L}$ satisfying the following property. For each function $f\in\ell\subset\mathfrak{L}_{CE}$, the invariant non-hyperbolic Cantor set $A(f)$ contains a Collet-Eckmann point $c$ such that $\omega(c)=A(f)$. Moreover the closure $\overline{\mathfrak{L}_{CE}}=\overline{\mathfrak{L}}$.
Theorem B is proved in Theorem 7.
In pursuing the stability of chaotic non-hyperbolic sets, we encounter a striking coexistence phenomenon that further illustrates the richness of the dynamics in these systems. We prove the existence of a sub-lamination in which each leaf contains a map with infinitely many sinks of increasing period, accumulating on a non-hyperbolic Cantor set that contains a Collet–Eckmann point. This coexistence of infinitely many sinks with chaotic dynamics is both unusual and unexpected. The formal statement of the theorem is given below.
Theorem C.
There exists a sub-lamination $\mathfrak{L}_{N}\subset\mathfrak{L}_{CE}\subset\mathfrak{L}$ satisfying the following property. Each leave $\ell\subset\mathfrak{L}_{N}$ contains a map $f_{\ell}$ with infinitely many attracting periodic points, $p_{k}$ of arbitrarily high period, such that,
$\bigcap_{n}\overline{\bigcup_{k\geq n}\text{Orb}(p_{k})}=A(f_{\ell}).$
Moreover if $NC=\left\{f_{\ell}\right\}_{f_{\in}\ell\subset\mathfrak{L}_{N}}$, then the closure $\overline{{NC}}=\overline{\mathfrak{L}_{CE}}=\overline{\mathfrak{L}}$.
Theorem C is proved in Theorem 8. The proof relies on an intermediate result, Theorem 3, which establishes the existence of maps with infinitely many sinks in families of unfoldings. Maps exhibiting the Newhouse phenomenon have been constructed in various contexts. In particular, numerous examples of Baire [^1] A Baire set in a locally compact Hausdorff space X X is a countable intersection of open and dense subsets. sets of maps with infinitely many sinks in the space of dynamical systems are known; see, for instance, [1, 9, 13, 11, 15, 18, 19, 24, 25, 29, 33, 34, 41, 42].
In the classical Newhouse construction, the coexistence of infinitely many sinks is produced by the persistence of homoclinic tangencies. The underlying mechanism relies on the intersection of stable and unstable Cantor sets whose thicknesses exceed one, ensuring that tangencies survive under perturbation.
Our approach is fundamentally different. To prove Theorem C, we avoid using the persistence of tangencies altogether. Instead, we introduce a quantitative condition (2.2) that involves only the eigenvalues of a saddle point. This spectral condition replaces the geometric thickness criterion of Newhouse and provides an explicit, verifiable rule for when the construction can be iterated. It is worth noting that, in the Hénon family, the set of maps satisfying the thickness condition and the set satisfying condition (2.2) intersect, but neither is contained in the other.
The proof proceeds inductively: starting from a map with a sink and a homoclinic tangency, we show that the sink persists under small parameter changes, while the tangency gives rise to a new tangency and an additional sink. Iterating this mechanism yields infinitely many sinks accumulating on a non-hyperbolic Cantor set containing a Collet–Eckmann point.
The most significant difference between our approach and Newhouse’s original method [33] lies in the creation of the secondary tangency. In Newhouse’s construction, the persistence of tangencies guarantees that every nearby map possesses a tangency somewhere, but this dependence on the map is highly discontinuous, making the analysis delicate. Our method—inspired by [3, 5]—replaces this mechanism with the use of critical points and binding points, thereby avoiding any dependence on the persistence of tangencies.
Thus, our eigenvalue-based and constructive method offers an alternative route to the Newhouse phenomenon and clarifies how coexistence of chaos and stability can occur.
Acknowledgements.
The authors express their gratitude to Michael Benedicks for his continued support throughout the preparation of this paper and for the many invaluable discussions at both the inspirational and technical levels. The second author was partially supported by the starting grant SFB 67575 and by the VR grant 67578, Renormalization in dynamics.
2 Preliminaries
To define the families of maps that will concern us, we first need a suitable linearization theorem. The following well-known result, due to Sternberg (see [39]), provides exactly this.
Theorem 1.
Given $\left(\mu,\lambda\right)\in\mathbb{R}^{2}$, there exists $k\in{\mathbb{N}}$ such that the following holds. Let $\mathcal{M}$ be a two dimensional $\mathcal{C}$$^{\infty}$} manifold and let $f:\mathcal{M}\to\mathcal{M}$ be a diffeomorphism with saddle point $p\in\mathcal{M}$ having unstable eigenvalue $|\mu|>1$ and stable eigenvalue $|\lambda|$. If
$\lambda\neq\mu^{k_{1}}\text{ and }\mu\neq\lambda^{k_{2}}$ (2.1)
for $k=(k_{1},k_{2})\in{\mathbb{N}}^{2}$ and $2\leq|k|=k_{1}+k_{2}\leq k$, then $f$ is ${\mathcal{C}}^{4}$} linearizable.
Definition 1.
Let $\mathcal{M}$ be an two dimensional $\mathcal{C}$$^{\infty}$} manifold and let $f:\mathcal{M}\to\mathcal{M}$ be a diffeomorphism with saddle point $p\in\mathcal{M}$. We say that $p$ satisfies the ${\mathcal{C}}^{4}$} non-resonance condition if (2.1) holds.
Theorem 2.
Let $\mathcal{M}$ be a two dimensional $\mathcal{C}$$^{\infty}$} manifold and let $f:\mathcal{M}\to\mathcal{M}$ be a diffeomorphism with a saddle point $p\in\mathcal{M}$ which satisfies the ${\mathcal{C}}^{4}$} non-resonance condition. Let $0\in\mathcal{P}\subset\mathbb{R}^{n}$ and let $F:\mathcal{P}\times\mathcal{M}\to\mathcal{M}$ be a $\mathcal{C}$$^{\infty}$} family with $F_{0}=f$. Then, there exists a neighborhood $U$ of $p$ and a neighborhood $V$ of 0 such that, for every $t\in V$, $F_{t}$ has a saddle point $p_{t}\in U$ satisfying the ${\mathcal{C}}^{4}$} non-resonance condition. Moreover $p_{t}$ is ${\mathcal{C}}^{4}$} linearizable in the neighborhood $U$ and the linearization depends ${\mathcal{C}}^{4}$} on the parameters.
The proofs of Theorem 1 and Theorem 2 can be found in [12, 26].
2.1 Families of unfoldings
In the sequel we define a map with a strong homoclinic tangency by listing several conditions. To summarize, a map with a strong homoclinic tangency has a non degenerate homoclinic tangency $q_{1}$ and a transversal homoclinic intersection $q_{2}$. Conditions $(f6)$, $(f7)$ and $(f8)$ ensure that $q_{1}$ and the unstable local manifolds of $q_{1}$ and $q_{2}$ accumulate on the leg of the unstable manifold of the saddle point containing the transversal homoclinic intersection $q_{2}$. These conditions are redundant if $\mu<-1$. Moreover to implement the construction, a condition on the eigenvalues is also required, see $(f2)$. The basic properties of a map with a strong homoclinic tangency are summarized in Figure 1.
Definition 2.
Let $\mathcal{M}$ be a two dimensional $\mathcal{C}$$^{\infty}$} manifold and let $f:\mathcal{M}\to\mathcal{M}$ be a local diffeomorphism satisfying the following conditions:
  • $(f1)$
    $f$ has a saddle point $p\in\mathcal{M}$ with unstable eigenvalue $|\mu|>1$ stable eigenvalue $\lambda$,
  • $(f2)$
    $|\lambda||\mu|^{3}<1$,
  • $(f3)$
    $p$ satisfies the ${\mathcal{C}}^{4}$} non-resonance condition,
  • $(f4)$
    $f$ has a non degenerate homoclinic tangency, $q_{1}\in W^{u}(p)\cap W^{s}(p)$,
  • $(f5)$
    $f$ has a transversal homoclinic intersection, $q_{2}\in W^{u}(p)\pitchfork W^{s}(p)$,
  • $(f6)$
    let $p,q_{2}]^{u}\subset W^{u}(p)$ be the arc connecting $p$ to $q_{2}$, then there exist arcs $W^{u}_{\text{\rm loc},n}(q_{2})=[q_{2},u_{n}]^{u}\subset W^{u}(q_{2})$ such that $p,q_{2}]^{u}\cap[q_{2},u_{n}]^{u}=\left\{q_{2}\right\}$ and
    $\lim_{n\to\infty}f^{n}\left(W^{u}_{\text{\rm loc},n}(q_{2})\right)=[p,q_{2}]^{u},$
  • $(f7)$
    there exist neighborhoods $W^{u}_{\text{\rm loc},n}(q_{1})\subset W^{u}(q_{1})$ such that
    $\lim_{n\to\infty}f^{n}\left(W^{u}_{\text{\rm loc},n}(q_{1})\right)=[p,q_{2}]^{u},$
  • $(f8)$
    there exists $N,M\in{\mathbb{N}}$ such that
    $f^{-N}(q_{1})\in[p,q_{2}]^{u}$
    and
    $f^{M}(q_{2})\in W^{s}_{\text{\rm loc}}(p).$
A map $f$ with these properties is called a map with a strong homoclinic tangency, see Figure 1.
Figure 1: A map with a strong homoclinic tangency
Remark 1.
The conditions defining a map with a strong homoclinic tangency are natural, except $(f2)$. This condition requires that the contraction at the saddle is strong enough. It plays a crucial role in many fundamental places towards the paper.
Remark 2.
Observe that all conditions are open in the space of maps with an homoclinic and transversal tangency. Also, except for $(f2)$, all conditions are dense.
Remark 3.
If the unstable eigenvalue is negative, $\mu<-1$, then $(f6)$, $(f7)$, and $(f8)$ are redundant.
Following [35], we construct now a family of unfolding of a map $f$ with a homoclinic tangency and a transversal intersection. Let $\mathcal{P}=[-r,r]\times[-r,r]$ with $r>0$. Given a map $f$ with a strong homoclinic tangency, we consider a $\mathcal{C}$$^{\infty}$} family $F:\mathcal{P}\times\mathcal{M}\to\mathcal{M}$ through $f$ with the following properties:
  • $(F1)$
    $F_{0,0}=f$,
  • $(F2)$
    $F_{t,a}$ has a saddle point $p(t,a)$ with unstable eigenvalue $|\mu(t,a)|>1$, with stable eigenvalue $\lambda(t,a)$, and
    $\frac{\partial\mu}{\partial t}\neq 0,$
  • $(F3)$
    let $\mu_{\text{max}}=\max_{(t,a)}|\mu(t,a)|$, $\lambda_{\text{max}}=\max_{(t,a)}|\lambda(t,a)|$ and assume
    $\lambda_{\text{max}}\mu_{\text{max}}^{3}<1,$ (2.2)
  • $(F4)$
    there exists a ${\mathcal{C}}^{2}$} function $r,r]\ni t\mapsto q_{1}(t)\in W^{u}(p(t,0))\cap W^{s}(p(t,0))$ such that $q_{1}(t)$ is a non degenerate homoclinic tangency of $F_{t,0}$,
  • $(F5)$
    there exists a ${\mathcal{C}}^{2}$} function $r,r]^{2}\ni(t,a)\mapsto q_{2}(t,a)\in W^{u}(p(t,a))\cap W^{s}(p(t,a))$ such that $q_{2}(t,a)$ is a transversal homoclinic intersection of $F_{t,a}$.
Remark 4.
Without loss of generality, we may assume that
$\frac{\partial\mu}{\partial t}>0,\mu(t,a)>0\text{ and }\lambda(t,a)>0.$
Observe that, by $(F3)$ and by continuity, for $t_{0}$ small enough,
$0<\frac{\log\mu_{\text{max}}}{\log{\frac{1}{\lambda_{\text{max}}}}}<\frac{3}{2}\frac{\log\mu_{\text{min}}}{\log{\frac{1}{\lambda_{\text{min}}}}}<\frac{1}{2}.$ (2.3)
2.1.1 Rescaled Families
According to Theorem 2 we may make a change of coordinates to ensure that the family $F$ is ${\mathcal{C}}^{4}$} and for all $(t,a)\in[-r_{0},r_{0}]\times[-r_{0},r_{0}]$ with $0<r_{0}<r$ and by rescaling we can assume that $F_{t,a}$ is linear on the ball $2,2]^{2}$, namely
$F_{t,a}=\left(\begin{matrix}\lambda(t,a)&0\\ 0&\mu(t,a)\\ \end{matrix}\right).$ (2.4)
Moreover the saddle point $p(t,a)=(0,0)$ and the local stable and unstable manifolds satisfy:
    $W^{s}_{\text{loc}}(0)=[-2,2]\times\left\{0\right\}$,
    $W^{u}_{\text{loc}}(0)=\left\{0\right\}\times[-2,2]$ ,
    $q_{1}(t)\in W^{s}_{\text{loc}}(0)$,
    $q_{2}(t,a)\in\left\{0\right\}\times\left(\frac{1}{\mu},2\right)\subset W^{u}_{\text{loc}}(0)$,
    there exists $N$ such that $f^{N}(q_{3}(t))=q_{1}(t)$ where $q_{3}(t)=(0,1)$.
2.1.2 Critical Points and Critical Values
In this subsection we aim to introduce the concept of critical points and critical values in our setting. This will lead to the formal definition of unfoldings.
In the next lemma we prove that $q_{3}$ is contained in a curve of points whose vertical expanding tangent vectors are mapped by $DF^{N}$ to horizontal contracting ones. Let $(x,y)$ be in a neighborhood of $q_{3}$ and consider the point
$(X_{t,a}(x,y),Y_{t,a}(x,y))=F^{N}_{t,a}(x,y).$
The following lemma says that $F^{N}_{t,a}(x,y)$ produces an unfolding in the sense of [35]. The formal definition of unfolding is given in Definition 4 below.
Lemma 1.
There exist $x_{0},a_{0}>0$, a ${\mathcal{C}}^{2}$} function $c:[-x_{0},x_{0}]\times[-t_{0},t_{0}]\times[-a_{0},a_{0}]\to\mathbb{R}$ and a positive constant $Q$ such that
$\frac{\partial Y_{t,a}}{\partial y}\left(x,c(x,t,a)\right)=0,$ (2.5)
and
$\frac{\partial^{2}Y_{t,a}}{\partial y^{2}}\left(x,c(x,t,a)\right)\geq Q.$
Moreover
$|c(x,t,a)-c(0,t,a)|=O\left(|x|\right).$ (2.6)
Proof.
Let $\Phi:[-\frac{1}{2},\frac{3}{2}]\times\left([-1,1]\times[-t_{0},t_{0}]\times[-a_{0},a_{0}]\right)\to\mathbb{R}$ be defined by
$\Phi(y,x,t,a)=\frac{\partial Y_{t,a}}{\partial y}\left(x,y\right).$
Observe that $\Phi$ is ${\mathcal{C}}^{2}$}. Let $q_{3}(t)=F^{-N}_{t,0}(q_{1}(t))$. Because $q_{1}(t)$ is an homoclinic tangency, see $(F4)$, we have
$\Phi(q_{3}(t),0,t,0)=0,$
and because $q_{1}(t)$ is a non degenerate tangency, we get
$\frac{\partial\Phi}{\partial y}(q_{3}(t),0,t,0)>0.$
For every $t\in[-t_{0},t_{0}]$ there exist, by the implicit function theorem, $\epsilon>0$ and a unique ${\mathcal{C}}^{2}$} function $c:[-\epsilon,\epsilon]\times[-t_{0}-\epsilon,t_{0}+\epsilon]\times[-\epsilon,\epsilon]\to\mathbb{R}$ locally satisfying the requirements of the lemma. These local functions extend to global ones because of the compactness of the interval $t_{0},t_{0}]$ and the local uniqueness. ∎
Remark 5.
Without lose of generality, by a smooth change of coordinates in the $y$ direction, we may assume that $c(0,t,a)=1$.
Definition 3.
Let $\left(t,a\right)\in[-t_{0},t_{0}]\times[-a_{0},a_{0}]$. We call the point
$c_{t,a}=\left(0,c(0,t,a)\right),$
the primary critical point and
$z_{t,a}=F_{t,a}^{N}\left(c_{t,a}\right)=(z_{x}(t,a),z_{y}(t,a)),$
the primary critical value of $F_{t,a}$.
Observe that, near the saddle point, vertical vectors are expanding and horizontal ones are contracting. The critical points are defined to have the property that the expanding vertical vector is sent to the contracting horizontal one under $DF^{N}$. Let us briefly discuss the concept of critical points for dissipative maps in order to compare our definition with the ones previously used.
Let us start by recalling that critical points are fundamental in the study of one-dimensional dynamics and they are easy to detect: those are the points where the map is not locally a diffeomorphism. This definition has no meaning for diffeomorphism of higher dimensional manifolds. The formal definition in this setting is given in [37], where the authors define the critical points as homoclinic tangencies, i.e. a tangency between the stable and unstable manifold of a saddle point. One can also, which is the basis of the corresponding construction in [3], define a critical point as a tangency between a local stable manifold and an unstable manifold of a periodic point.
Homoclinic tangencies play a crucial topological role in general. However, they are difficult to detect. That is why in varies studies, starting with [3], there are notions of approximate critical points which share with homoclinic tangencies the property that expanding vectors are mapped into contracting ones. In particular, in [3], critical points are rather tangencies between the unstable manifold of the saddle point and an approximate local stable manifold, not necessarily of the saddle point. In our situation, similarly the tangent vector of the unstable manifold at the critical point $c_{t,a}$ is mapped into the contracting horizontal vector at the critical value $z_{t,a}$.
The crucial fact, which comes from the fundamental property of the approximate critical points (and homoclinic tangencies), is that in a neighborhood, the return map is an Hénon-like map. The local Hénon behavior is what is important and allows the analysis. From a technical point of view, the notion of critical point itself is less important; the local Hénon behavior is all what is needed.
Another instance where critical points occur but play only a secondary role is in the context of strongly dissipative Hénon maps at the boundary of chaos. These maps have a period doubling Cantor attractor, see [14]. The Cantor attractors are studied using renormalization zooming in to the so-called tip of the Cantor set as in [14]. Indeed, return maps to neighborhoods of the tip are Hénon-like maps. A posterior one shows that the stable manifold at the tip is tangent to the direction of the neutral Lyapunov exponent. The tip plays the role of a critical point. However, this fact did not play any role during the renormalization analysis. This is another instance where, from a technical point of view, the notion of critical point is not that important.
Definition 4.
A family $F_{t,a}$ is called an unfolding of $f$ if it can be reparametrized such that
  • $(P1)$
    $z_{y}(t,0)=0$,
  • $(P2)$
    $\frac{\partial z_{y}(t,0)}{\partial a}\neq 0.$
Figure 2: Unfolding
Remark 6.
Without lose of generality by a suitable coordinate change we may assume that if $F$ is an unfolding then $z_{y}(t,a)=a$, the primary critical value is at height $a$ and the primary critical point $c(t,a)=(0,1)$, see Figure 2.
Remark 7.
A generic two-dimensional family trough f can locally be reparametrized to become an unfolding.
Remark 8.
The results and the techniques presented in this paper can be generalized to higher dimensional families, both in parameter and phase space.
2.2 Properties of unfoldings
In this section, we outline several fundamental properties of unfoldings.
Estimates of the stable manifold at $q_{2}$.
Consider the transversal homoclinic intersection $q_{2}$ defined in $(f5)$. Let $W=W^{s}_{\text{loc}}(q_{2})$ and for all $n\in{\mathbb{N}}$ let $W_{n}=F^{-n}_{t,a}(W)$. Because $W\pitchfork W^{u}(p)$ we can apply the $\lambda$-Lemma which implies that $W_{n}$ converges to $W^{s}_{\text{loc}}(0)$. In particular $W_{n}$ is the graph of a function which will also be denoted by $W_{n}$. Moreover, because
$W_{n}(x)=\mu^{-n}W\left(\lambda^{n}x\right),$ (2.7)
then
$\frac{1}{2\mu}\mu^{-n}\leq|W_{n}|\leq 2\mu^{-n},$ (2.8)
$\left|\frac{\partial W_{n}}{\partial x}\right|=O\left(\frac{\lambda}{\mu}\right)^{n},$ (2.9)
and
$\left|\frac{\partial^{2}W_{n}}{\partial x^{2}}\right|=O\left(\frac{\lambda^{2}}{\mu}\right)^{n}.$ (2.10)
We estimate now the speed of the stable manifold $W_{n}$ as the graph of a function in the phase space.
Lemma 2.
$\frac{\partial W_{n}}{\partial t}=-\frac{n}{\mu}\frac{\partial\mu}{\partial t}W_{n}+\left(\frac{\lambda}{\mu}\right)^{n}\frac{n}{\lambda}\frac{\partial\lambda}{\partial t}\frac{\partial W}{\partial x}x+\frac{1}{\mu^{n}}\frac{\partial W}{\partial t}=O\left(\frac{n}{\mu^{n}}\right),$
$\frac{\partial W_{n}}{\partial a}=-\frac{n}{\mu}\frac{\partial\mu}{\partial a}W_{n}+\left(\frac{\lambda}{\mu}\right)^{n}\frac{n}{\lambda}\frac{\partial\lambda}{\partial a}\frac{\partial W}{\partial x}x+\frac{1}{\mu^{n}}\frac{\partial W}{\partial a}=O\left(\frac{n}{\mu^{n}}\right).$
Proof.
Fix a point $x\in[-2,2]$ and a parameter $(t,a)\in[-r_{0},r_{0}]^{2}$. We denote by $W_{n}$ the manifold corresponding to $(t,a)$ and by $W_{n}+\Delta W_{n}$ the manifold to $(t,a+\Delta a)$. Then, since the maps $F^{n}_{t,a}$ are linear, by differentiating (2.7) we obtain
$\left(\mu+\frac{\partial\mu}{\partial a}\Delta a\right)^{n}\left(W_{n}+\Delta W_{n}\right)(x)=\mu^{n}W_{n}(x)+\frac{\partial W}{\partial x}\left(\left(\left(\lambda+\frac{\partial\lambda}{\partial a}\Delta a\right)^{n}-\lambda^{n}\right)x\right)+\frac{\partial W}{\partial a}\Delta a.$
Similarly, one gets the same bound for ${\partial W_{n}}/{\partial t}$. ∎
The $\Gamma$ curve and the $a_{n}$ curve.
Fix an unfolding $F$ and for each $\left(t,a\right)\in[-t_{0},t_{0}]\times[-a_{0},a_{0}]$ let
$\Gamma_{t,a}=\left\{(x,c(x,t,a))|x\in[-x_{0},x_{0}]\right\}.$
In the next lemma we build a curve $a_{n}$ of points, in the parameter space, whose corresponding critical values are mapped after $n$ steps into $\Gamma_{t,a}$.
Lemma 3.
For $n$ large enough, there exists a ${\mathcal{C}}^{2}$} function $a_{n}:[-t_{0},t_{0}]\to(0,a_{0}]$ such that
$F^{n}_{t,a_{n}\left(t\right)}\left(z_{\left(t,a_{n}\left(t\right)\right)}\right)\in\Gamma_{\left(t,a_{n}\left(t\right)\right)}.$
Moreover
$\frac{da_{n}}{dt}=-n\frac{\partial\mu}{\partial t}\frac{1}{\mu^{n+1}}\left[1+O\left(\lambda^{n}\right)\right].$ (2.11)
Proof.
Let $\Gamma=\text{graph}(c)$, namely,
$\Gamma=\left\{\left(x,c(x,t,a),t,a\right)|\left(x,t,a\right)\in[-x_{0},x_{0}]\times[-t_{0},t_{0}]\times[-a_{0},a_{0}]\right\}.$
Then $\Gamma$ is a ${\mathcal{C}}^{2}$} codimension $1$ manifold transversal to $W^{u}(0)$. For $n\geq 0$ let
$\Gamma_{n}=\left\{\left(x,y,t,a\right)|F^{n}_{t,a}\left(x,y\right)\in\Gamma\right\},$ (2.12)
and the limit of the $\Gamma_{n}$:s as $n\to\infty$ is,
$\Gamma_{\infty}=\left\{\left(x,0,t,a\right)|\left(x,t,a\right)\in[-x_{0},x_{0}]\times[-t_{0},t_{0}]\times[-a_{0},a_{0}]\right\}.$
This follows since, for each $(t,a)\in[-t_{0},t_{0}]\times[-a_{0},a_{0}]$, $F_{t,a}$ is linear, $\Gamma_{n}$ converges to $\Gamma_{\infty}$ in the ${\mathcal{C}}^{2}$} topology. Namely for large $n$, $\Gamma_{n}$ is a graph of a function also denoted by $\Gamma_{n}$ and
$\left\|\Gamma_{n}-\Gamma_{\infty}\right\|_{{{\mathcal{C}}^{2}}}\to 0.$
Let $z:[-t_{0},t_{0}]\times[-a_{0},a_{0}]\to\mathbb{R}^{2}\times[-t_{0},t_{0}]\times[-a_{0},a_{0}]$ be the ${\mathcal{C}}^{2}$} critical value function defined as
$z(t,a)=\left(z_{x}(t,a),z_{y}(t,a),t,a\right)=\left(z_{x}(t,a),a,t,a\right),$
where $z_{t,a}=\left(z_{x}(t,a),z_{y}(t,a)\right)$. Observe that
$\frac{\partial z_{y}}{\partial a}=1\text{ and }z_{y}(t,0)=0.$ (2.13)
Let $Z=\text{Image}(z)$. Because of (2.13), $Z$ is a manifold transversal to $\Gamma_{\infty}$. Hence, there exists $n_{1}>0$ such that, for all $n\geq n_{1}$, $Z$ is transversal to $\Gamma_{n}$. As consequence, for all $n\geq n_{1}$,
$A_{n}=z^{-1}\left(\Gamma_{n}\right)$
is a ${\mathcal{C}}^{2}$} codimension $1$ manifold. We define $a_{n}:t\mapsto a_{n}(t)$ as a function whose graph is $A_{n}$. Observe that, by Lemma 1 and Remark 5, the ${\mathcal{C}}^{2}$} function $c$ satisfies, $c(0,t,a)=1$. Hence $\partial c/\partial t=O(x)$ and $\partial c/\partial a=O(x)$. Recall that
$\mu^{n}a_{n}=c\left(\lambda^{n}z_{x},t,a_{n}\right).$ (2.14)
By differentiating (2.14) and using $\partial c/\partial a,\partial c/\partial t=O(x)$ we have
$\mu^{n}\frac{da_{n}}{dt}+na_{n}\mu^{n-1}\frac{\partial\mu}{\partial t}=O\left(\lambda^{n}\right).$
The lemma follows. ∎
A proof similar to that of Lemma 2.7 gives the following.
Lemma 4.
Let $\Gamma_{n}$ be as in (2.12), then
$\frac{\partial\Gamma_{n}}{\partial t}=-\frac{n}{\mu}\frac{\partial\mu}{\partial t}\Gamma_{n}+\left(\frac{\lambda}{\mu}\right)^{n}\frac{n}{\lambda}\frac{\partial\lambda}{\partial t}\frac{\partial\Gamma}{\partial x}x+\frac{1}{\mu^{n}}\frac{\partial\Gamma}{\partial t}=O\left(\frac{n}{\mu^{n}}\right),$
$\frac{\partial\Gamma_{n}}{\partial a}=-\frac{n}{\mu}\frac{\partial\mu}{\partial a}\Gamma_{n}+\left(\frac{\lambda}{\mu}\right)^{n}\frac{n}{\lambda}\frac{\partial\lambda}{\partial a}\frac{\partial\Gamma}{\partial x}x+\frac{1}{\mu^{n}}\frac{\partial\Gamma}{\partial a}=O\left(\frac{n}{\mu^{n}}\right).$
Exact form of the map $F^{N}$.
For families of unfoldings the following holds.
Lemma 5.
There exist $x^{\prime}_{0}<x_{0}$, $a^{\prime}_{0}<a_{0}$, $b>0$ and $Q>0$, such that, for all $(t,a)\in[-t_{0},t_{0}]\times[-a^{\prime}_{0},a^{\prime}_{0}]$ and for every $(x,y)\in\Gamma_{t,a}$ with $|x|<|x^{\prime}_{0}|$ the following holds. There exist matrices $A_{x,y}$, $B_{x,y}\neq 0$ and $C_{x,y}$ such that $F^{N}_{t,a}$ in coordinates centered in $(x,y)$ and $F^{N}_{t,a}(x,y)$ has the form
$F^{N}_{t,a}\left(\begin{matrix}\Delta x\\ \Delta y\end{matrix}\right)$ $=$ $\left(\begin{matrix}A_{x,y}&B_{x,y}\\ C_{x,y}&0\end{matrix}\right)\left(\begin{matrix}\Delta x\\ \Delta y\end{matrix}\right)$ (2.15)
$+$ $\left(\begin{matrix}O_{1,1}\Delta x^{2}+O_{1,2}\Delta x\Delta y+O_{1,3}\Delta y^{2}\\ Q_{x,y}\Delta y^{2}+O_{2,1}\Delta x^{2}+O_{2,2}\Delta x\Delta y+O_{2,3}\Delta y^{3}\end{matrix}\right),$
where the matrices $A_{x,y}$, $C_{x,y}$ and the vector $B_{x,y}$ are ${\mathcal{C}}^{2}$} dependent on $x$ and $y$, $Q_{x,y}>Q$, $|B_{x,y}|>b>0$ and the vector valued functions $O_{i,j}$ are ${\mathcal{C}}^{2}$} dependent on $x$ and $y$ and they are uniformly bounded.
Proof.
The lemma gives the Taylor expansions of $F^{N}_{t,a}$ when $(x,y)\in\Gamma_{t,a}$. It follows immediately from Lemma 1 where we state that a vertical curve trough $(x,y)$ in $\Gamma_{t,a}$ is mapped to a curve with a non degenerate horizontal tangency. In particular $\partial Y_{t,a}/\partial y=0$ by (2.5) and the horizontal tangency,
$DF^{N}_{t,a}(x,y)\left(\begin{matrix}0\\ 1\end{matrix}\right)=\left(\begin{matrix}B_{x,y}\\ 0\end{matrix}\right)$
is not degenerate for all $(x,y)\in\Gamma_{t,a}$. This is a consequence of the following argument. Let $t\in[-t_{0},t_{0}]$. Because $F^{N}_{t,0}(q_{3}(t))=q_{1}(t)$ is a non degenerate homoclinic tangency, we know that the vector $B_{q_{3}(t)}\neq 0$ and $Q_{q_{3}(t)}>0$. By taking $|a|<|a^{\prime}_{0}|$, $|x|<|x^{\prime}_{0}|$ small enough, the lower bounds on $Q_{x,y}$ and $B_{x,y}$ follow. ∎
2.3 Choice of $\theta$
In this subsection, we introduce a small number $\theta$ that is going to play a fundamental role in the construction of our combinatorics.
The condition $\lambda_{\text{max}}\mu_{\text{max}}^{3}<1$, see $(F3)$, allows us to choose $\theta\in(0,\frac{1}{2})$ such that
$1<\lambda_{\text{min}}^{2\theta}\mu_{\text{min}}^{3}\text{ and }\lambda_{\text{max}}^{3\theta}\mu^{4}_{\text{max}}<1.$ (2.16)
We choose any $\theta$ satisfying
$0<\theta_{0}=\frac{4}{3}\frac{\log\mu_{\text{max}}}{\log{\frac{1}{\lambda_{\text{max}}}}}<\theta<\frac{3}{2}\frac{\log\mu_{\text{min}}}{\log{\frac{1}{\lambda_{\text{min}}}}}=\theta_{1}<\frac{1}{2},$ (2.17)
where we used $(F3)$, the initial condition $\lambda_{\text{max}}\mu_{\text{max}}^{3}<1$ and (2.3).
Moreover, by restricting the parameter domain, we choose $\theta$ such that
$\left(\frac{\theta}{\lambda_{\text{max}}}\frac{\partial\lambda}{\partial t}+\frac{1}{\mu_{\text{max}}}\frac{\partial\mu}{\partial t}\right)>0.$ (2.18)
3 The Hénon family
Consider the real Hénon family $F:\mathbb{R}^{2}\times\mathbb{R}^{2}\to\mathbb{R}^{2}$,
$F_{a,b}\left(\begin{matrix}x\\ y\end{matrix}\right)=\left(\begin{matrix}a-x^{2}-by\\ x\end{matrix}\right),$
a two parameter family. In this section we are going to prove that the Hénon family is an unfolding of a map with a strong homoclinic tangency and a transversal intersection.
Proposition 1.
The real Hénon family is an unfolding.
Proof.
Consider the map $f(x)=2-x^{2}$. Then x=2x=-2 is an expanding fixed point and $f^{2}(0)=-2$ where 0 is the critical point. For given $b>0$ and $a$ large enough, the Hénon map $F_{a,b}$ has an horse-shoe. By decreasing $a$ to $a(b)$, one arrives at the first homoclinic tangency. Hence, there exists an analytic curve $b\mapsto a(b)$ with $a(0)=2$ such that the parameter $(a(b),b)$ corresponds to a Hénon map with an homoclinic tangency of the saddle point $p(b)$ which is a continuation of p(0)=(2,2)p(0)=(-2,-2). For all $b$ positive and small enough $F_{a(b),b}$ has a strong homoclinic tangency and the Hénon family $F_{a,b}$ restricted to a small neighborhood of $(a(b),b)$ is an unfolding. ∎
4 Existence of secondary tangencies
In this section we prove the existence of secondary tangencies in a certain strip of parameters that we introduce below.
Remember the choice of $\theta$ in (2.16) and (2.17). Denote by $C$ the entry of the matrix in Lemma 5 corresponding to the critical point $c$, see Definition 3. Let $\alpha\in(0,1)$ and let $n_{0}=[n\alpha]+O(1)$. Choose $\epsilon>0$, $T$ a positive large constant and define, for large $n$,
$\mathcal{B}_{n}=\left\{(t,a)\in[-t_{0},t_{0}]\times[-a_{0},a_{0}]\left|\right.\left(\frac{\epsilon}{\mu^{\frac{n_{0}}{2}}}-C\right)\lambda^{\theta n}\leq(a-a_{n}(t))\leq\frac{T}{\mu^{n+\frac{n_{0}}{2}}}\right\}.$ (4.1)
The strip $\mathcal{B}_{n}$ is built to contain the curves of tangencies. The exact choice of $n_{0}$ is made in (4.16).
Denote by $E=C-\frac{\epsilon}{\mu^{{n_{0}}/{2}}}$. Observe that in this definition of the strip, $\mu$ and $\lambda$ are taken in the parameter $(t,a_{n}(t))$. The following remark points out the distortion of the eigenvalues when changing the parameters in $\mathcal{B}_{n}$.
Remark 9.
[…]
$\left(\frac{\mu(t,a_{n}(t))}{\mu(\tilde{t},\tilde{a})}\right)^{n}=1+O\left(n\lambda(t,a_{n}(t))^{\theta n}\right),\left(\frac{\lambda(t,a_{n}(t))}{\lambda(\tilde{t},\tilde{a})}\right)^{n}=1+O\left(n\lambda(t,a_{n}(t))^{\theta n}\right)$
where we used the fact that $\partial\mu/\partial t$ and $\partial\mu/\partial a$ are bounded. Moreover if $(t,a)\in\mathcal{B}_{n}$, then
$\frac{\mu(t,a_{n}(t))^{n}(a-a_{n}(t))}{\lambda(t,a)^{\theta n}\mu(t,a)^{n}}\geq{E},$
and
${\mu(t,a)^{n+\frac{n_{0}}{2}}(a-a_{n}(t))}\leq{T}.$
4.1 The combinatorics
Fix $(t,a)\in\mathcal{B}_{n}$ and in the sequel when there is no possibility of confusion, we will suppress this choice in the notation, for example, $z=z_{t,a}$, $c=c_{t,a}$ (see Definition 3), $\lambda=\lambda(t,a)$ and $\mu=\mu(t,a)$.
The first loop: $N+\theta n+N$ iterates.
By the initial conditions on the family, for $n$ large enough and $W^{u}_{\text{loc}}(z)$ small enough, $F^{\theta n}_{t,a}\left(W^{u}_{\text{loc}}(z)\right)$ intersects $\Gamma$ in exactly two points. Choose one of these points
$c^{\prime}\in F^{\theta n}_{t,a}\left(W^{u}_{\text{loc}}(z)\right)\cap\Gamma,$
and a local unstable manifold $W^{u}_{\text{loc}}(c^{\prime})$ of diameter $0<L^{\frac{1}{2}}\lambda^{\frac{\theta n}{2}}$ where is a large constant that will be chosen later.
Let $z^{(1)}=(z_{x}^{(1)},z_{y}^{(1)})$ be the lowest point of $F^{N}\left(W^{u}_{\text{loc}}(c^{\prime})\right)$, see Figure 3.
Figure 3: $N+\theta n$ iterates
Lemma 6.
For $n$ large enough and for all $(t,a)\in\mathcal{B}_{n}$,
$C\lambda^{\theta n}+O\left(\lambda^{2\theta n}\right)\leq|z_{y}^{(1)}(t,a)-z_{y}(t,a)|\leq\left(C+{QL}\right)\lambda^{\theta n}+O\left(\lambda^{2\theta n}\right),$
and the tangent space $T_{z^{(1)}}W^{u}_{\text{loc}}(z^{(1)})\subset\mathbb{R}^{m-1}\times\left\{0\right\}$ is horizontal. Moreover, the limit $\lim_{n\to\infty}{|z_{y}^{(1)}(t,a)-z_{y}(t,a)|}/{\lambda^{\theta n}}$ exists.
In particular,
$|z_{y}^{(1)}(t,a_{n}(t))-z_{y}(t,a_{n}(t))|=C\lambda^{\theta n}+O\left(\lambda^{2\theta n}\right).$ (4.2)
Proof.
Use coordinates centered at the critical point $c$ of the parameter $(t,a)$ and let $(\Delta x,\Delta y)\in W^{u}_{\text{loc}}(c^{\prime})$ with $(\Delta\tilde{x},\Delta\tilde{y})\in F_{t,a}^{N}\left(W^{u}_{\text{loc}}(c^{\prime})\right)$ centered in $z$. By Lemma 5 we have
$\Delta\tilde{y}$ $=$ $C\Delta x+Q(\Delta y)^{2}+O\left(|\Delta x|^{2}+|\Delta x||\Delta y|+|\Delta y|^{3}\right).$
Because the first coordinate of $q_{1}$ is equal to $1$, see $(F4)$ and the following normalization, we have that $\Delta x=\lambda^{\theta n}+O\left(\lambda^{2\theta n}\right)$ and since $0\leq|\Delta y|\leq L^{\frac{1}{2}}\lambda^{\frac{\theta n}{2}}$,
$C\lambda^{\theta n}+O\left(\lambda^{2\theta n}\right)\leq\Delta\tilde{y} \displaystyle\leq$ $\left(C+{QL}\right)\lambda^{\theta n}+O\left(\lambda^{2\theta n}\right).$
The first claim of the lemma is then proved. The fact that, in the parameter $(t,a_{n}(t))$, $\Delta y$ coincides with $c^{\prime}$, ensures that 4.2 is also true.
For the second claim of the lemma, use the coordinates centered in $c^{\prime}$. Let $(\Delta x,\Delta y)\in W^{u}_{\text{loc}}(c^{\prime})$ with $(\Delta\tilde{x},\Delta\tilde{y})=F_{t,a}^{N}\left((\Delta x,\Delta y)\right)$ written in coordinates centered in $F^{N}(c^{\prime})$. Consider $W^{u}_{\text{loc}}(c^{\prime})$ as a graph on the $y$-axis and take a tangent vector $(\Delta u,\Delta v)\in T_{c^{\prime}}W^{u}_{\text{loc}}(c^{\prime})$. Define the slope-vector at $c^{\prime}$ as $s=\Delta u/\Delta v$. Then
$s=O\left(\left(\frac{\lambda^{2}}{\mu}\right)^{\frac{\theta n}{2}}\right).$ (4.3)
This estimate is obtained as follows. Let $(z_{x}+\Delta z_{x},z_{y}+\Delta z_{y})\in W^{u}_{\text{loc}}(z)$ such that $F^{\theta n}\left(z_{x}+\Delta z_{x},z_{y}+\Delta z_{y}\right)=c^{\prime}$ and $(\Delta u_{0},\Delta v_{0})$ be a tangent vector to $W^{u}_{\text{loc}}(z)$ at the point $\left(z_{x}+\Delta z_{x},z_{y}+\Delta z_{y}\right)$. Then there exist uniform constants $K_{1},K_{2}$ such that
${K_{1}^{-1}}{\mu^{-\theta n}}\leq\Delta z_{y}\leq K_{1}{\mu^{-\theta n}},$
and because of the quadratic behavior of $W^{u}_{\text{loc}}(z)$,
$K_{2}^{-1}{\mu^{-\frac{\theta n}{2}}}\leq|\Delta z_{x}|\leq K_{2}{\mu^{-\frac{\theta n}{2}}}.$
In particular
$K_{2}^{-1}{\mu^{-\frac{\theta n}{2}}}|\Delta u_{0}|\leq|\Delta v_{0}|\leq K_{2}{\mu^{-\frac{\theta n}{2}}}|\Delta u_{0}|.$
Observe that
$s=O\left(\frac{\lambda^{\theta n}}{\mu^{\theta n}}\frac{\Delta u_{0}}{\Delta v_{0}}\right).$
The estimate for the slope follows. Use now coordinates $\left(\Delta x^{\prime},\Delta y^{\prime}\right)$ centered around $z$ with $F^{\theta n}\left(\Delta x^{\prime},\Delta y^{\prime}\right)=\left(\Delta x,\Delta y\right)$. Then there is a non zero smooth function $R$ such that
$|\Delta x^{\prime}|=R\left(\Delta y^{\prime}\right)\left(\Delta y^{\prime}\right)^{\frac{1}{2}},$
and
$\Delta y^{\prime}=O\left(\frac{1}{\mu^{\theta n}}\right).$
In particular,
$\frac{d^{2}|\Delta x^{\prime}|}{d\left(\Delta y^{\prime}\right)^{2}}=O\left(\left(\Delta y^{\prime}\right)^{-\frac{3}{2}}\right)=O\left(\mu^{\frac{3}{2}\theta n}\right).$
This implies that
$\frac{d^{2}|\Delta x|}{d\left(\Delta y\right)^{2}}=O\left(\mu^{\frac{3}{2}\theta n}\frac{\lambda^{\theta n}}{\mu^{2\theta n}}\right)=O\left(\left(\frac{\lambda^{2}}{\mu}\right)^{\frac{\theta n}{2}}\right).$
Hence,
$\Delta x=s\Delta y+O\left(\left(\frac{\lambda^{2}}{\mu}\right)^{\frac{\theta n}{2}}\right)(\Delta y)^{2}.$ (4.4)
From (4.4) and (4.3) we have that $|\Delta x|=O\left(s\Delta y\right)=o\left(\Delta y\right)$. From Lemma 5,
$\Delta\tilde{x}$ $=$ $B\Delta y+O_{x}\left(\Delta y\right)(\Delta y)^{2}$
$\Delta\tilde{y}$ $=$ $s\Delta y+Q(\Delta y)^{2}+O_{y}\left(\Delta y\right)(\Delta y)^{3},$
where $B,Q\neq 0$, $O_{x}$ and $O_{y}$ are bounded differentiable functions in $\Delta y$. After differentiating one obtains
$\frac{d\tilde{y}}{d\tilde{x}}=\frac{2Q}{B}\Delta y+O\left(\left(\frac{\lambda^{2}}{\mu}\right)^{\frac{\theta n}{2}}\right),$ (4.5)
where we used (4.3). As consequence $z^{(1)}$ which is the point where ${d\tilde{y}}/{d\tilde{x}}=0$ corresponds to
$\Delta y=O\left(\left(\frac{\lambda^{2}}{\mu}\right)^{\frac{\theta n}{2}}\right)\leq L^{\frac{1}{2}}\lambda^{\frac{\theta n}{2}}.$
Indeed $W^{u}_{\text{loc}}(z^{(1)})$ has a horizontal tangency. ∎
Figure 4: $2N+n$ iterates
The second loop: $N+\theta n+N+n+N$ iterates.
Let $z^{(2)}=(z_{x}^{(2)},z_{y}^{(2)})=F_{t,a}^{n}(z^{(1)})$ and take a local unstable manifold $W^{u}_{\text{loc}}(z^{(2)})$ of vertical size $L\left(\lambda^{\theta}\mu\right)^{n}$ with $L$ a large constant which is going to be chosen later. Observe that $z^{(2)}$ is located close to $c$, see Figure 4.
The following lemma gives information on the size and the position of $W^{u}_{\text{\rm loc}}(z^{(2)})$. In particular Lemma 7 states how much $W^{u}_{\text{loc}}(z^{(2)})$ is shifted up with respect to the critical point $c$, see Figure 4. Its proof follows from Lemma 6 and from the fact that we are in the domain of linearization.
Lemma 7.
For $n$ large enough and $(x,y)\in W^{u}_{\text{\rm loc}}$, if $(t,a)\in\mathcal{B}_{n}$,
$(C-E)\left(\lambda^{\theta}\mu\right)^{n}+O\left(\mu^{n}\lambda^{2\theta n}\right)$ $\leq$
$y-\mu_{0}^{n}z_{n,y}+\mu^{n}(a-a_{n}(t))$ $\leq$
$\frac{T}{\mu^{n_{0}/2}}+(C+QL)\left(\lambda^{\theta}\mu\right)^{n}+O\left(\mu^{n}\lambda^{2\theta n}\right),$
where $z_{n}=z(t,a_{n}(t))$, $\mu_{0}=\mu(t,a_{n}(t))$ and $\mu=\mu(t,a)$. Moreover
$x=O\left(\lambda^{n}\right).$
In particular, when $a=a_{n}(t)$,
$y-\mu_{0}^{n}z_{n,y}=C\left(\lambda^{\theta}\mu\right)^{n}+O\left(\mu^{n}\lambda^{2\theta n}\right).$
The curves $W^{u}_{\text{\rm loc}}(z^{(1)})$, $W^{u}_{\text{\rm loc}}(z^{(2)})$ and $W^{u}_{\text{\rm loc}}(z^{(3)})$ (which will be introduced later) are graphs over smooth curves in the $x$-plane. We can describe these curves in coordinates centered at $z^{(i)}$. In particular, for each component of $W^{u}_{\text{\rm loc}}(z^{(i)})\setminus z^{(i)}$, there exists a smooth function $w$ such that if $(\Delta x,\Delta y)\in W^{u}_{\text{\rm loc}}(z^{(i)})$ then $\Delta y=w\left(|\Delta x|\right)$.
In the following lemma we study the curvature of $W^{u}_{\text{\rm loc}}(z^{(2)})$.
Lemma 8.
For $n\geq 1$ large enough, in a coordinate system centered in $z^{(2)}$ the ${\mathcal{C}}^{4}$} curve $W^{u}_{\text{\rm loc}}(z^{(2)})$ is given by
$\Delta y=Q_{2}\left(\frac{\mu}{\lambda^{2}}\right)^{n}|\Delta x|^{2}+O\left(|\Delta x|^{3}\left(\frac{\mu}{\lambda^{3}}\right)^{n}\right),$
where $(\Delta x,\Delta y)\in W^{u}_{\text{\rm loc}}(z^{(2)})$, $0\leq\Delta y<L\left(\lambda^{\theta}\mu\right)^{n}$, $|\Delta x|=O\left(\lambda^{2+\theta}\right)^{\frac{n}{2}}$ and the constant $Q_{2}>0$ is a uniformly bounded and bounded away from zero.
Proof.
Consider the $W^{u}_{\text{loc}}(z^{(1)})$ of vertical size $L\lambda^{\theta n}$ and let $B\in T_{z^{(1)}}W^{u}_{\text{loc}}(z^{(1)})$ be a unit vector. Then in coordinate centered in $z^{(1)}$ the ${\mathcal{C}}^{4}$} curve $W^{u}_{\text{loc}}(z^{(1)})$ is given by
$\Delta y$ $=$ $Q_{1}|\Delta x|^{2}+O(|\Delta x|^{3}).$
where $Q_{1}$ is a uniform constant, see (4.5). Given the fact that $B$ has a non zero first coordinate, see $(F5)$, then the linear map $F^{n}_{t,a}$ turns the curve $W^{u}_{\text{loc}}(z^{(1)})$ into a curve of the form as stated in the lemma. Namely, recall that the curves $W^{u}_{\text{loc}}(z^{(1)})$ and $W^{u}_{\text{loc}}(z^{(2)})$ are graphs of functions over smooth curves in the $x$-plane. The curve $W^{u}_{\text{loc}}(z^{(1)})$ is parametrized by $|\Delta x|$ and by iteration by the linear map we obtain that the curve $W^{u}_{\text{loc}}(z^{(2)})$ is parametrized by $\lambda^{n}|\Delta x|$.
Observe also that for $n$ large enough and for all $(t,a)\in\mathcal{B}_{n}$, the highest vertical point of $W^{u}_{\text{loc}}(z^{(2)})$ is still in the domain of linearization. The factor $\left({\mu}/{\lambda^{2}}\right)^{n}$ comes from the change of curvature by $n$ linear iterates. ∎
Let $z^{(3)}=(z_{x}^{(3)},z_{y}^{(3)})$ be the lowest point of $F^{N}_{t,a}\left(W^{u}_{\text{loc}}(z^{(2)})\right)$. In the next lemma we prove that the new critical value $z^{(3)}$ is far from the critical value $z$ and from now on they will have independent behavior. In particular, while $z$ is contained in the basin of the sink, the new critical value $z^{(3)}$ it is outside of the basin and it will be involved in the creation of the new tangency.
Lemma 9.
For all $(t,a)\in\mathcal{B}_{n}$,
$0\leq Q(C-E)^{2}\left(\lambda^{\theta}\mu\right)^{2n}+O\left(\left(\lambda^{\theta}\mu\right)^{3n}\right)$
$\leq z_{y}^{(3)}(t,a)-z_{y}(t,a)\leq$
$Q(C+QL)^{2}\left(\lambda^{\theta}\mu\right)^{2n}+\frac{QT^{2}}{\mu^{n_{0}}}+O\left(\frac{\left(\lambda^{\theta}\mu\right)^{n}}{\mu^{n_{0}/2}}\right)+O\left(\left(\lambda^{\theta}\mu\right)^{3n}\right),$
Moreover $T_{z^{(3)}}W^{u}_{\text{loc}}(z^{(3)})\subset\mathbb{R}^{m-1}\times\left\{0\right\}$ is horizontal. In particular,
$z_{y}^{(3)}(t,a_{n}(t))-z_{y}(t,a_{n}(t))=QC^{2}\left(\lambda^{\theta}\mu\right)^{2n}+O\left(\left(\lambda^{\theta}\mu\right)^{3n}\right).$
Proof.
Use coordinates centered at $c$ and let $(\Delta x,\Delta y)\in W^{u}_{\text{loc}}(z^{(2)})=F_{t,a}^{n}\left(W^{u}_{\text{loc}}(z^{(1)})\right)$. Observe that
$\Delta y$ $=$ $y-\mu_{0}^{n}z_{n,y}+\mu_{0}^{n}z_{n,y}-1$ (4.6)
$=$ $y-\mu_{0}^{n}z_{n,y}+O\left(\lambda(t,a_{n}(t))|^{n}\right),$
where we used that $F^{n}_{t,a_{n}(t)}(z_{n})\in\Gamma$ which implies that
$|\mu_{0}^{n}z_{n,y}-1|=O\left(\left|F^{n}_{t,a_{n}(t)}(z_{n})_{x}\right|\right)=O\left(\lambda(t,a_{n}(t))^{n}\right).$
From Lemma 7 and for all $(t,a)\in\mathcal{B}_{n}$ we get
$(C-E)\left(\lambda^{\theta}\mu\right)^{n}+O\left(\lambda(t,a_{n}(t))^{n}\right)$ $\leq\Delta y\leq$ (4.7)
$\frac{T}{\mu^{n_{0}/2}}+(C+QL)\left(\lambda^{\theta}\mu\right)^{n}+O\left(\lambda(t,a_{n}(t))^{n}\right),$ (4.8)
$|\Delta x|=O\left(\lambda^{n}\right).$ (4.9)
[…]
$\Delta\tilde{y}=O\left(\lambda^{n}\right)+Q\left(\Delta y\right)^{2}+O\left(\left(\lambda^{\theta}\mu\right)^{3n}\right)=Q\left(\Delta y\right)^{2}+O\left(\left(\lambda^{\theta}\mu\right)^{3n}\right),$
[…]
In the following lemma we study the curvature of $W^{u}_{\text{\rm loc}}(z^{(3)})$. One can observe that the curvature of $W^{u}_{\text{\rm loc}}(z^{(3)})$ has grown considerably compared to that of $W^{u}_{\text{\rm loc}}(z)$.
Lemma 10.
Let $(t,a)\in\mathcal{B}_{n}$ then $W^{u}_{\text{loc}}(z^{(3)}(t,a))$ is the graph of a function and its curvature satisfies
$\text{\rm curv}\left(W^{u}_{\text{loc}}(z^{(3)}(t,a))\right)\geq T\left(\frac{\mu^{4}}{\lambda^{2-3\theta}}\right)^{n},$
with $T>0$ a uniform constant and ${\mu^{4}}/{\lambda^{2-3\theta}}>1$.
Proof.
Use coordinates centered at $z^{(2)}$ and let $m_{5}=(\Delta x,\Delta y)\in W^{u}_{\text{\rm loc}}(z^{(2)})$ be the preimage of $z^{(3)}$ under $N$ iterates, i.e. $F^{N}_{t,a}(\Delta x,\Delta y)=z^{(3)}$. Then $0\leq\Delta y\leq L\left(\lambda^{\theta}\mu\right)^{n}$. Let $v=\left(\begin{matrix}v_{1}\\ v_{2}\end{matrix}\right)$ be the tangent vector at $m_{5}$ to $W^{u}_{\text{loc}}(z^{(2)})$. We use Lemma 8 and we obtain that
$|v_{2}|=\left[2Q_{2}\left(\frac{\mu}{\lambda^{2}}\right)^{n}|\Delta x|+O\left(|\Delta x|^{2}\left(\frac{\mu}{\lambda^{3}}\right)^{n}\right)\right]|v_{1}|.$ (4.10)
By Lemma 1 we have that ${\partial^{2}Y}/{\partial y^{2}}$ is bounded away from zero in a neighbourhood of $c$. As a consequence, by Lemma 5 we have
$DF^{N}_{t,a}(m_{5})=\left(\begin{matrix}A&B\\ C&D\end{matrix}\right),$
where $A$, $B$, $C$ are matrices and $D$ is a number. Observe that, from Lemma 1, there exists a uniform constant $d^{\prime}>0$ such that
$D=d^{\prime}\left(m_{5,y}-1\right)+O(m_{5,x}),$
and by Lemma 7, $m_{5,x}=O\left(\lambda^{n}\right)$. By using again the Lemma 7 and the bound on $m_{5,x}$, by taking $\epsilon_{1}$ small enough, we get that $D\geq d\left(\lambda^{\theta}\mu\right)^{n}$ where $d>0$ is a uniform constant. Because $F^{N}_{t,a}(m_{5})$ is the minimum of the curve $W^{u}_{\text{loc}}(z^{(3)})$, then
$Cv_{1}+Dv_{2}=0.$ (4.11)
Use coordinates $\left(\Delta x^{\prime},\Delta y^{\prime}\right)$ along $W^{u}_{\text{loc}}\left(z^{(1)}\right)$ centered around $z^{(1)}$ such that $F^{n}\left(\Delta x^{\prime},\Delta y^{\prime}\right)=\left(\Delta x,\Delta y\right)$. Then, from (4.7), $\Delta y^{\prime}=O\left(\lambda^{\theta n}\right)$ and by the quadratic behavior, $|\Delta x^{\prime}|=O\left(\lambda^{\frac{\theta}{2}n}\right)$. This implies that
$|\Delta x|=O\left(\lambda^{\left(1+\frac{\theta}{2}\right)n}\right).$ (4.12)
By (4.10), (4.11), Lemma 8, the fact that $D\geq d\left(\lambda^{\theta}\mu\right)^{n}$ and that $C$ is a uniform constant, we have
$2dQ_{2}|\Delta x|+O\left(|\Delta x|^{2}\left(\frac{1}{\lambda}\right)^{n}\right)=O\left(\left(\frac{\lambda^{2-\theta}}{\mu^{2}}\right)^{n}\right).$
Moreover by (4.12), $|\Delta x|\left(\frac{1}{\lambda}\right)^{n}=O\left(\lambda^{\frac{\theta}{2}n}\right)$ and because $\lambda^{\frac{\theta}{2}n}<1$ we obtain
$|\Delta x|=O\left(\left(\frac{\lambda^{2-\theta}}{\mu^{2}}\right)^{n}\right).$ (4.13)
Consider the projection of the curve $W^{u}_{\text{loc}}(z^{(2)})$ on the the $(x,y)$-plane. The curvature of this projected curve is proportional to the curvature of the curve $W^{u}_{\text{loc}}(z^{(2)})$. Because $F^{n}$ is linear, we can estimate the curvature of the projection of the curve $W^{u}_{\text{loc}}(z^{(2)})$ to the $(x,y)$-plane by calculating how the linear map, which preserves the $(x,y)$-plane, changes the curvature of $W^{u}_{\text{loc}}(z^{(1)})$ on the $(x,y)$-plane. We denote the $x$-coordinate by $x$. We calculate now the second derivatives using again Lemma 8 and the fact that the curve $W^{u}_{\text{loc}}(z^{(2)})$ is in $n$ ${\mathcal{C}}^{2}$} exponentially close to the projected curve to the $xy$-plane. We get
$\frac{d^{2}y}{dx^{2}}\geq\frac{3}{2}Q_{2}\left(\frac{\mu}{\lambda^{2}}\right)^{n}+O\left(|\Delta x|\left(\frac{\mu}{\lambda^{3}}\right)^{n}\right)\geq Q_{2}\left(\frac{\mu}{\lambda^{2}}\right)^{n},$ (4.14)
where we used (4.13). Observe that $W^{u}_{\text{loc}}(m_{5})$ is a graph over an $x$-direction with second derivative at $m_{5}$ given by (4.14). By (4.10) and (4.13), the slope satisfies
$\frac{dy}{dx}=O\left(\left(\frac{1}{\lambda^{\theta}\mu}\right)^{n}\right).$ (4.15)
Consider now the same curve as a graph over the $y$-axis. Then, by (4.14), (4.15) and using $d^{2}x/dy^{2}=-1/\left(dy/dx\right)^{3}d^{2}y/dx^{2}$, we have
$\frac{d^{2}x}{dy^{2}}\geq C\left(\frac{\mu^{4}}{\lambda^{2-3\theta}}\right)^{n}.$
The map $F^{N}_{t,a}$ will preserve this order of curvature. As final remark observe that, by (2.16), ${\mu^{4}}/{\lambda^{2-3\theta}}=\left({\mu}/{\lambda^{2-\theta}}\right){\mu^{3}}{\lambda^{2\theta}}>1$. ∎
The full cycle: $N+\theta n+N+n+N+n_{0}$ iterates.
We are now ready to chose $n_{0}$. Let
$S_{n}=\left\{(x,y)\in[-2,2]^{m}|y\geq W_{n}(x)\right\}.$
Observe that $z\in S_{n}\setminus S_{n-1}$. Fix $t$ and let $n_{0}$ be the maximal $n\in{\mathbb{N}}$ such that
$W^{u}_{\text{loc}}(z^{(3)}(t,a_{n}(t)))\subset S_{n-n_{0}}.$ (4.16)
Observe that such a choice of $n_{0}$ gives information on the position of $W^{u}_{\text{loc}}(z^{(3)}(t,a_{n}(t)))$ with respect to the pull-backs of the stable manifold, see Figure 5. A more precise estimate on the size of $n_{0}$ comes from the following lemma. Let $\alpha={\log\left(\lambda^{2\theta}\mu^{3}\right)}/{\log\mu}$. By (2.17), $\alpha\in\left(0,{1}\right)$.
Lemma 11.
The integer $n_{0}$ satisfies the following:
$n_{0}=[n\alpha]+O(1).$
Proof.
Fiz $t$ as in the definition of $n_{0}$. From Lemma 9, (2.8) and the fact that $z\in S_{n}\setminus S_{n-1}$, there exist uniform constants $K_{1},K_{2}>0$ such that
$z_{y}^{(3)}(t,a_{n}(t))=K_{1}\left(\frac{1}{\mu^{n}}+\left(\lambda^{\theta}\mu\right)^{2n}\right).$
Moreover the definition of $n_{0}$ implies that
$\frac{1}{K_{2}}\frac{1}{\mu^{n-n_{0}}}\leq z_{y}^{(3)}\leq{K_{2}}\frac{1}{\mu^{n-n_{0}}}.$
The two previous inequalities imply the existence of a positive constant $K$ such that
$\frac{1}{K}\leq\frac{1}{\mu^{n_{0}}}\left(1+\left(\lambda^{2\theta}\mu^{3}\right)^{n}\right)\leq K.$ (4.17)
The lemma follows from (2.16). ∎
Remark 10.
Observe that by combining (4.17) and (2.16), one gets
$\lambda^{2\theta}\mu^{2}<\frac{1}{\mu^{\alpha}}.$ (4.18)
This simplifies the formulas in Lemmas 7 and 9.
4.2 The secondary tangencies
In this subsection we construct secondary tangencies along the curve $a_{n}$. We start by giving the formal definition of secondary tangency.
Definition 5.
A tangency between $W^{u}_{\text{loc}}(z^{(3)})$ and $W_{n-n_{0}}$ is called a secondary tangency of type $n_{0}$.
In order to prove the existence of secondary tangencies, we need to estimate the variation of $z_{y}^{(3)}$ with respect to the parameters $a$ and $t$. This gives the speed of $W^{u}_{\text{\rm loc}}(z^{(3)})$ in the phase space. Comparing this speed with the one of the pullbacks of the stable manifold at $q_{2}$ allows finding the new tangencies.
We take a vector at a point $m_{1}\in W^{u}_{\text{\rm loc}}\left(c\right)$ and calculate its variation along its orbit at crucial moments, step by step, until we get a vector at the point $F^{3N+\theta n+n}(m_{1})\in W^{u}_{\text{\rm loc}}(z^{(3)})$. The steps will be defined precisely in the proof of the proposition. The reader can refer to Figure 4.
A crucial application of Proposition 2 is the speed of $W^{u}_{\text{\rm loc}}(z^{(3)})$ when varying the parameters along the curve $a_{n}$. The dominating terms in the estimates for the partial derivatives will cancel in the calculation of this speed; see (4.48). This is the reason why we cannot suppress more terms in the Taylor development. This cancellation is also crucial in the proof of Proposition 4.
Proposition 2.
For $n\geq 1$ large enough and for $(t,a)\in\mathcal{B}_{n}$,
$D(C-E)\left(\lambda^{\theta}\mu\right)^{n}\left\{\frac{n}{\mu}\frac{\partial\mu}{\partial t}+n\left(\lambda^{\theta}\mu\right)^{n}\left[C\frac{\theta}{\lambda}\frac{\partial\lambda}{\partial t}+\frac{(C-E)}{\mu}\frac{\partial\mu}{\partial t}\right]\right\}+O\left(\left(\lambda^{\theta}\mu\right)^{2n}\right)$
$\leq\frac{\partial z_{y}^{(3)}}{\partial t}\leq$
$\frac{DT}{\mu^{n_{0}/2}}\left\{\frac{n}{\mu}\frac{\partial\mu}{\partial t}+\frac{n}{\mu}\frac{\partial\mu}{\partial t}\frac{T}{\mu^{n_{0}/2}}\right\}+O\left(\frac{1}{\mu^{n_{0}}}\right),$
and
$D(C-E)\left(\lambda^{\theta}\mu^{2}\right)^{n}+1+D(C-E)\left(\lambda^{\theta}\mu\right)^{n}\left\{\frac{n}{\mu}\frac{\partial\mu}{\partial a}+n\left(\lambda^{\theta}\mu\right)^{n}\left[C\frac{\theta}{\lambda}\frac{\partial\lambda}{\partial a}+\frac{(C-E)}{\mu}\frac{\partial\mu}{\partial a}\right]\right\}$
$+O\left(\left(\lambda^{\theta}\mu\right)^{2n}\right)\leq\frac{\partial z_{y}^{(3)}}{\partial a}\leq$
$DT\mu^{n-n_{0}/2}+1+\frac{DT}{\mu^{n_{0}/2}}\left\{\frac{n}{\mu}\frac{\partial\mu}{\partial a}+\frac{n}{\mu}\frac{\partial\mu}{\partial a}\frac{T}{\mu^{n_{0}/2}}\right\}+O\left(\frac{1}{\mu^{n_{0}}}\right).$
Moreover $\lambda^{\theta}\mu^{2}>1$, CE,D,T>0C-E,D,T>0 and they converge to a non zero limit as $n\to\infty$.
In particular, when $a=a_{n}$,
$\frac{\partial z_{y}^{(3)}}{\partial t}=DC\left(\lambda^{\theta}\mu\right)^{n}\left\{\frac{n}{\mu}\frac{\partial\mu}{\partial t}+Cn\left(\lambda^{\theta}\mu\right)^{n}\left[\frac{\theta}{\lambda}\frac{\partial\lambda}{\partial t}+\frac{1}{\mu}\frac{\partial\mu}{\partial t}\right]\right\}+O\left(\left(\lambda^{\theta}\mu\right)^{2n}\right),$
and
$\frac{\partial z_{y}^{(3)}}{\partial a}$ $=$ $DC\left(\lambda^{\theta}\mu^{2}\right)^{n}+1+DC\left(\lambda^{\theta}\mu\right)^{n}\left\{\frac{n}{\mu}\frac{\partial\mu}{\partial a}+Cn\left(\lambda^{\theta}\mu\right)^{n}\left[\frac{\theta}{\lambda}\frac{\partial\lambda}{\partial a}+\frac{1}{\mu}\frac{\partial\mu}{\partial a}\right]\right\}$
$+$ $O\left(\left(\lambda^{\theta}\mu\right)^{2n}\right).$
Proof.
Choose $t_{1}\in[-t_{0},t_{0}]$. Observe that, by Lemma 5, the definition of unfolding, by using coordinates $(x,y,t,a)$ centered around $(0,1,t_{1},0)$ in the domain and coordinates centered around $F_{t_{1},0}^{N}(0,1)$ we get the following expression for $F_{t,a}^{N}(x,y)$,
$F_{t,a}^{N}(x,y)$ $=$ $\left(\begin{matrix}Ax+By+tE+aF\\ Cx+Qy^{2}+a\left[1+x\varphi_{x}+y^{2}\varphi_{y}\right]\end{matrix}\right),$ (4.19)
where the uniformly bounded coefficients satisfy
    $A(x,y,t,a)$ is a $(n-1)\times(n-1)$ matrix valued ${\mathcal{C}}^{3}$} function,
    $B(x,y,t,a),E(x,y,t,a),F(x,y,t,a)$ are n1n-1 vector valued ${\mathcal{C}}^{3}$} functions,
    $C(x,y,t,a)$ is a $1\times(n-1)$ matrix valued ${\mathcal{C}}^{3}$} function,
    $Q(x,y,t,a),\varphi_{x}(x,y,t,a),\varphi_{y}(x,y,t,a)$ are ${\mathcal{C}}^{3}$} functions.
Moreover $B,Q\neq 0$. The $x$ component of $F_{t,a}^{N}(x,y)$ is a general expression and the fact that $B\neq 0$ follows from $(f5)$. For the $y$ component of $F_{t,a}^{N}(x,y)$ observe that the linear terms in $t$ and $y$ are absent because in $a=0$, the point $F^{N}_{t,0}(0,1)$ is a non-degenerate homoclinic tangency, see Remark 6. The $a$-dependence of the second component of $F^{N}_{t,a}$ follows from Remark 6.
Recall that $c^{\prime}\in F^{\theta n}_{t,a}\left(W^{u}_{\text{loc}}(z)\right)\cap\Gamma$, $z^{(1)}$ is defined as the lowest point of $F^{N}_{t,a}\left(W^{u}_{\text{loc}}(c^{\prime})\right)$, $z^{(2)}$ as the lowest point of $F^{n}_{t,a}\left(W^{u}_{\text{loc}}(z^{(1)})\right)$ and $z^{(3)}$ as the lowest point of $F^{N}_{t,a}\left(W^{u}_{\text{loc}}(z^{(2)})\right)$. We fix the points $m_{6},m_{5},m_{4},m_{3},m_{2},m_{1}$ so that they satisfy the following, see Figure 4.
    $m_{6}\in W^{u}_{\text{loc}}(z^{(3)})$ sufficiently close to $z^{(3)}$.
    $m_{5}=(m_{5,x},m_{5,y})\in W^{u}_{\text{loc}}(z^{(2)})$, $F_{t,a}^{-N}(m_{6})=m_{5}$. By Lemma 7, Remark 9 and Remark 10,
    $(C-E)\left(\lambda^{\theta}\mu\right)^{n}+O\left(\mu^{n}\lambda^{2\theta n}\right)$ $\leq$
    $|m_{5,y}-c_{n,y}|$ $\leq$ (4.20)
    $\frac{T}{\mu^{n_{0}/2}}+O\left(\mu^{n}\lambda^{\theta n}\right),$
    and by (4.13)
    $|m_{5,x}-z^{(2)}_{x}|=O\left(\left(\frac{\lambda^{2-\theta}}{\mu^{2}}\right)^{n}\right).$ (4.21)
    In particular, when $a=a_{n}(t)$,
    $|m_{5,y}-c_{n,y}|=C\left(\lambda^{\theta}\mu\right)^{n}+O\left(\mu^{n}\lambda^{2\theta n}\right).$ (4.22)
    $m_{4}=(m_{4,x},m_{4,y})\in W^{u}_{\text{loc}}(z^{(1)})$, $F_{t,a}^{-n}(m_{5})=m_{4}$. By Lemma 6 and the fact that $z_{n,y}$ can be estimated by $\mu^{-n}$, see Remark 9,
    $|m_{4,y}|=O\left(\frac{1}{\mu^{n}}\right),$ (4.23)
    and by (4.21), using the linear map backward and $(F5)$ (which ensures that the vector $m_{4,x}-z^{(1)}_{x}$ has a dominating first component) we get,
    $|m_{4,x}-z^{(1)}_{x}|=O\left(\left(\frac{\lambda^{1-\theta}}{\mu^{2}}\right)^{n}\right).$ (4.24)
    $m_{3}=(m_{3,x},m_{3,y})\in W^{u}_{\text{loc}}(c^{\prime})$, $F_{t,a}^{-N}(m_{4})=m_{3}$. By Lemma 6 (see Figure 3) and (4.24),
    $|m_{3}-c|$ $=$ $O\left({\lambda^{\theta{n}}}\right),$ (4.25)
    $|m_{3}-c^{\prime}|$ $=$ $O\left(\left(\frac{\lambda^{1-\theta}}{\mu^{2}}\right)^{n}\right).$ (4.26)
    $m_{2}=(m_{2,x},m_{2,y})\in W^{u}_{\text{loc}}(z)$, $F_{t,a}^{-\theta n}(m_{3})=m_{2}$. Therefore, there exists a uniform constant $K_{2}>0$ such that
    $\frac{1}{K_{2}}\frac{1}{\mu^{\theta{n}}}\leq|m_{2,y}|\leq K_{2}\frac{1}{\mu^{\theta{n}}}.$ (4.27)
    $m_{1}=(m_{1,x},m_{1,y})\in W^{u}_{\text{loc}}(c)$, $F_{t,a}^{-N}(m_{2})=m_{1}$. By (4.27) and because of the quadratic behavior, there exists a uniform constant $K_{1}>0$ such that
    $\frac{1}{K_{1}}\frac{1}{\mu^{\theta\frac{n}{2}}}\leq|m_{1,y}-c_{y}|\leq K_{1}\frac{1}{\mu^{\theta\frac{n}{2}}}.$ (4.28)
Let us recall that, $m_{2}=F^{N}(m_{1})$, $m_{3}=F^{\theta n}(m_{2})$, $m_{4}=F^{N}(m_{3})$, $m_{5}=F^{n}(m_{4})$, and $m_{6}=F^{N}(m_{5})$. Moreover, $t_{1}\in[-t_{0},t_{0}]$ was chosen arbitrary. Take $(t_{1}+t,a)\in\mathcal{B}_{n}$ and consider the ${\mathcal{C}}^{4}$} map
$(\Delta t,\Delta a,\Delta y)\longmapsto F^{3N+(\theta+1)n}_{t+\Delta t,a+\Delta a}{(m_{1}+(0,\Delta y))}=m_{6}+(\Delta m_{6,x},\Delta m_{6,y}).$
We are interested in the partial derivatives of $m_{6}$. Observe that for $i=1,3,5$,
$dm_{i+1}$ $=$ $DF^{N}_{t,a}\left(m_{i}\right)dm_{i}+\frac{\partial F^{N}_{t,a}\left(m_{i}\right)}{\partial t}dt+\frac{\partial F^{N}_{t,a}\left(m_{i}\right)}{\partial a}da,$ (4.29)
for $i=4$,
$dm_{i+1}$ $=$ $DF^{n}_{t,a}\left(m_{i}\right)dm_{i}+\frac{\partial F^{n}_{t,a}\left(m_{i}\right)}{\partial t}dt+\frac{\partial F^{n}_{t,a}\left(m_{i}\right)}{\partial a}da,$ (4.30)
and for $i=2$,
$dm_{i+1}$ $=$ $DF^{\theta n}_{t,a}\left(m_{i}\right)dm_{i}+\frac{\partial F^{\theta n}_{t,a}\left(m_{i}\right)}{\partial t}dt+\frac{\partial F^{\theta n}_{t,a}\left(m_{i}\right)}{\partial a}da.$ (4.31)
All partial derivatives in (4.29) are uniformly bounded. However we will need more careful estimate for $\Delta m_{i+1,y}$. Namely, for $i=1,3,5$,
$DF^{N}_{t,a}\left(m_{i}\right)$ $=$ $\left(\begin{matrix}A_{i}&B_{i}\\ C_{i}&O\left(|m_{i,x}|\right)+D_{i}(m_{i,y}-1)\end{matrix}\right),$
which follows from (4.19). Similarly from (4.19) one obtains
$\frac{\partial m_{i+1,y}}{\partial t}$ $=$ $\frac{\partial C}{\partial t}m_{i,x}+\frac{\partial Q}{\partial t}\left(m_{i,x}-1\right)^{2}+am_{i,x}\frac{\partial\varphi_{x}}{\partial t}+a\left(m_{i,y}-1\right)^{2}\frac{\partial\varphi_{y}}{\partial t}$
$=$ $d_{i}\left|m_{i,y}-1\right|^{2}+O\left(|m_{i,x}|\right),$
where $d_{i}\neq 0$ is uniformly bounded away from zero. By differentiation of (4.19) with respect of $\Delta a$ we obtain
$\frac{\partial m_{i+1,y}}{\partial a}$ $=$ $1+O\left(\left|m_{i,y}-1\right|^{2}\right)+O\left(|m_{i,x}|\right).$
For $i=3$ we refine the estimate
$DF^{N}_{t,a}\left(m_{3}\right)$ $=$ $\left(\begin{matrix}A_{3}&B_{3}\\ C_{3}&O\left(|m_{3}-c^{\prime}|\right)\end{matrix}\right),$
where we estimate $DF^{N}(m_{3})$ with $DF^{N}(c^{\prime})$ evaluated at a point at distance $|m_{3}-c^{\prime}|$. For $i=1,5$ we get in an analogous way
$dm_{i+1}$ $=$ $\left(\begin{matrix}A_{i}&B_{i}\\ C_{i}&O\left(|m_{i,x}|\right)+D_{i}(m_{i,y}-1)\end{matrix}\right)dm_{i}$
$+$ $\left(\begin{matrix}O\left(dt\right)+O\left(da\right)\\ \left[d_{i}|m_{i,y}-1|^{2}+O\left(|m_{i,x}|\right)\right]dt+\left[1+O\left(|m_{i,y}-1|^{2}+|m_{i,x}|\right)\right]da\end{matrix}\right),$
and
$dm_{4}$ $=$ $\left(\begin{matrix}A_{3}&B_{3}\\ C_{3}&O\left(|m_{3}-c^{\prime}|\right)\end{matrix}\right)dm_{3}$
$+$ $\left(\begin{matrix}O\left(dt\right)+O\left(da\right)\\ \left[d_{3}|m_{3,y}-1|^{2}+O\left(|m_{3,x}|\right)\right]dt+\left[1+O\left(|m_{3,y}-1|^{2}+|m_{3,x}|\right)\right]da\end{matrix}\right),$
where $D_{i},B_{i},C_{i}\neq 0$ (because $q_{1}$ is a non degenerate tangency in general direction and $DF^{N}(0,1)$ is non singular). Moreover, by construction,
$C_{3}=C(1+\lambda^{\theta_{n}}).$ (4.34)
From $dm_{4}$ to $dm_{5}$ we use the linear map $F^{n}$. From (4.30) and using that $m_{5,y}=\mu^{n}m_{4,y}$ we get
$dm_{5}$ $=$ $\left(\begin{matrix}\lambda^{n}&0\\ 0&\mu^{n}\end{matrix}\right)dm_{4}+\left(\begin{matrix}X_{4}\frac{n\lambda^{n}}{\lambda}\left[\frac{\partial\lambda}{\partial t}dt+\frac{\partial\lambda}{\partial a}da\right]\\ m_{5,y}\frac{n}{\mu}\left[\frac{\partial\mu}{\partial t}dt+\frac{\partial\mu}{\partial a}da\right]\end{matrix}\right),$ (4.35)
with $0\neq X_{4}=m_{4,x}\approx 1$. A similar formula holds going from $dm_{2}$ to $dm_{3}$ where we use the linear map $F^{\theta n}$, (4.31) and the fact that $m_{3,y}=\mu^{\theta n}m_{2,y}$. Namely,
$dm_{3}$ $=$ $\left(\begin{matrix}\lambda^{\theta n}&0\\ 0&\mu^{\theta n}\end{matrix}\right)dm_{2}+\left(\begin{matrix}X_{2}\frac{\theta n\lambda^{\theta n}}{\lambda}\left[\frac{\partial\lambda}{\partial t}dt+\frac{\partial\lambda}{\partial a}da\right]\\ m_{3,y}\frac{\theta n}{\mu}\left[\frac{\partial\mu}{\partial t}dt+\frac{\partial\mu}{\partial a}da\right]\end{matrix}\right)$ (4.36)
where
$0\neq X_{2}=m_{2,x}\approx\left(\begin{matrix}1&0&\ldots&0\end{matrix}\right)$ (4.37)
Observe that $m_{1,x}=0$ and recall that $\left(dm_{1,x},dm_{1,y}\right)=\left(0,dy\right)$. By (4.2) and by (4.28),
$\left(\begin{matrix}dm_{2,x}\\ dm_{2,y}\end{matrix}\right)=\left(\begin{matrix}A_{1}&B_{1}\\ C_{1}&D_{1}\tilde{K}_{1}{\mu^{\frac{-\theta{n}}{2}}}\end{matrix}\right)\left(\begin{matrix}0\\ dy\end{matrix}\right)+\left(\begin{matrix}O\left(dt\right)+O\left(da\right)\\ \left[d_{1}\tilde{K}_{1}^{2}{\mu^{{-\theta{n}}}}\right]dt+\left[1+O\left({\mu^{{-\theta{n}}}}\right)\right]da\end{matrix}\right),$
where $\tilde{K}_{1}$ is the constant which gives an equality in (4.28). As a consequence
$dm_{2,x}$ $=$ $O\left(dt\right)+O\left(da\right)+O\left(dy\right),$
$dm_{2,y}$ $=$ $\frac{\tilde{d}_{1}}{\mu^{\theta{n}}}dt+\left[1+O\left(\frac{1}{\mu^{\theta{n}}}\right)\right]da+\frac{\tilde{D}_{1}}{\mu^{\theta\frac{n}{2}}}dy,$
where $\tilde{d}_{1}=d_{1}\tilde{K}_{1}^{2}$ and $\tilde{D}_{1}=D_{1}\tilde{K}_{1}$. By using the fact that $F^{\theta n}_{t,a}$ is linear and using (4.36),
$\left(\begin{matrix}dm_{3,x}\\ dm_{3,y}\end{matrix}\right)$ $=$ $\left(\begin{matrix}\lambda^{\theta n}&0\\ 0&\mu^{\theta n}\end{matrix}\right)\left(\begin{matrix}dm_{2,x}\\ dm_{2,y}\end{matrix}\right)+\left(\begin{matrix}X_{2}\frac{\theta n\lambda^{\theta n}}{\lambda}\left[\frac{\partial\lambda}{\partial t}dt+\frac{\partial\lambda}{\partial a}da\right]\\ \left(\left(m_{3,y}-c_{n,y}\right)+c_{n,y}\right)\frac{\theta n}{\mu}\left[\frac{\partial\mu}{\partial t}dt+\frac{\partial\mu}{\partial a}da\right]\end{matrix}\right),$
[…]
$dm_{3,x}$ $=$ $X_{2}\frac{\theta n\lambda^{\theta n}}{\lambda}\left[\frac{\partial\lambda}{\partial t}dt+\frac{\partial\lambda}{\partial a}da\right]+O\left(\lambda^{\theta n}dy\right)$
$+$ $O\left(\lambda^{\theta n}dt\right)+O\left(\lambda^{\theta n}da\right),$
$dm_{3,y}$ $=$ $\left[\frac{\theta n}{\mu}\frac{\partial\mu}{\partial t}+\tilde{d}_{1}+O\left(n\lambda^{\theta n}\right)\right]dt$
$+$ $\left[\mu^{\theta n}+\frac{\theta n}{\mu}\frac{\partial\mu}{\partial a}+O\left(1\right)\right]da+\tilde{D}_{1}{\mu^{\theta\frac{n}{2}}}dy,$
where $X_{2}\neq 0$ is pointing in the direction of $e_{1}$, see Figure 3. By (4.2) and (4.25) (observe that $\left\{|m_{3,y}-c_{n,y}|,|m_{3,x}|\right\}=O\left(\lambda^{\theta n}\right)$), we get
$\left(\begin{matrix}dm_{4,x}\\ dm_{4,y}\end{matrix}\right)=\left(\begin{matrix}A_{3}&B_{3}\\ C_{3}&O\left(\left(\frac{\lambda^{1-\theta}}{\mu^{2}}\right)^{n}\right)\end{matrix}\right)\left(\begin{matrix}dm_{3,x}\\ dm_{3,y}\end{matrix}\right)+\left(\begin{matrix}O\left(dt\right)+O\left(da\right)\\ O\left(\lambda^{\theta n}\right)dt+\left[1+O\left(\lambda^{\theta n}\right)\right]da\end{matrix}\right).$
As consequence, using also (4.37) and (4.34),
$dm_{4,x}$ $=$ $\left[B_{3}\frac{\theta n}{\mu}\frac{\partial\mu}{\partial t}+O\left(1\right)\right]dt$
$+$ $\left[B_{3}\mu^{\theta n}+B_{3}\frac{\theta n}{\mu}\frac{\partial\mu}{\partial a}+O\left(1\right)\right]da$
$+$ $\left[B_{3}\tilde{D}_{1}{\mu^{\theta\frac{n}{2}}}+O\left({\lambda^{\theta{n}}}\right)\right]dy,$
$dm_{4,y}$ $=$ $\left[C\frac{\theta n\lambda^{\theta n}}{\lambda}\frac{\partial\lambda}{\partial t}+O\left(\lambda^{\theta n}\right)\right]dt$
$+$ $\left[1+C\frac{\theta n\lambda^{\theta n}}{\lambda}\frac{\partial\lambda}{\partial a}+O\left(\lambda^{\theta n}\right)\right]da$
$+$ $O\left(\lambda|^{\theta{n}}dy\right).$
By (4.35) we get
$dm_{5,x}$ $=$ $\left[B_{3}\frac{\theta n\lambda^{n}}{\mu}\frac{\partial\mu}{\partial t}+X_{4}\frac{n\lambda^{n}}{\lambda}\frac{\partial\lambda}{\partial t}+O\left(\lambda^{n}\right)\right]dt$
$+$ $\left[B_{3}\left(\lambda\mu^{\theta}\right)^{n}+B_{3}\frac{\theta n\lambda^{n}}{\mu}\frac{\partial\mu}{\partial a}+X_{4}\frac{n\lambda^{n}}{\lambda}\frac{\partial\lambda}{\partial a}+O\left(\lambda^{n}\right)\right]da$
$+$ $\left[B_{3}\tilde{D}_{1}\left(\lambda\mu^{\frac{\theta}{2}}\right)^{n}+O\left(\left({\lambda^{\theta+1}}\right)^{n}\right)\right]dy,$
$dm_{5,y}$ $\geq$ $\left[\frac{n}{\mu}\frac{\partial\mu}{\partial t}+n\left(\lambda^{\theta}\mu\right)^{n}\left[C\frac{\theta}{\lambda}\frac{\partial\lambda}{\partial t}+\frac{(C-E)}{\mu}\frac{\partial\mu}{\partial t}\right]+O\left(\left(\lambda^{\theta}\mu\right)^{n}\right)\right]dt$
$+$ $\left[\mu^{n}+\frac{n}{\mu}\frac{\partial\mu}{\partial a}+n\left(\lambda^{\theta}\mu\right)^{n}\left[C\frac{\theta}{\lambda}\frac{\partial\lambda}{\partial a}+\frac{(C-E)}{\mu}\frac{\partial\mu}{\partial a}\right]+O\left(\left(\lambda^{\theta}\mu\right)^{n}\right)\right]da$
$+$ $O\left(\left(\lambda|^{\theta}\mu\right)^{{n}}dy\right),$
where we used that $m_{5,y}=(m_{5,y}-c_{n,y})+c_{n,y}$, the fact that the point $c$ is at height 11 and (-4.2). Similarly,
$dm_{5,y}$ $\leq$ $\left[\frac{n}{\mu}\frac{\partial\mu}{\partial t}+\frac{n}{\mu}\frac{\partial\mu}{\partial t}\frac{T}{\mu^{n_{0}/2}}+n\left(\lambda^{\theta}\mu\right)^{n}C\frac{\theta}{\lambda}\frac{\partial\lambda}{\partial t}+O\left(\left(\lambda^{\theta}\mu\right)^{n}\right)\right]dt$
$+$ $\left[\mu^{n}+\frac{n}{\mu}\frac{\partial\mu}{\partial a}+\frac{n}{\mu}\frac{\partial\mu}{\partial a}\frac{T}{\mu^{n_{0}/2}}+n\left(\lambda^{\theta}\mu\right)^{n}C\frac{\theta}{\lambda}\frac{\partial\lambda}{\partial a}+O\left(\left(\lambda^{\theta}\mu\right)^{n}\right)\right]da$
$+$ $O\left(\left(\lambda|^{\theta}\mu\right)^{{n}}dy\right).$
In particular, when $a=a_{n}$, using the same equalities as before and 4.22,
$dm_{5,y}$ $=$ $\left[\frac{n}{\mu}\frac{\partial\mu}{\partial t}+n\left(\lambda^{\theta}\mu\right)^{n}C\left[\frac{\theta}{\lambda}\frac{\partial\lambda}{\partial t}+\frac{1}{\mu}\frac{\partial\mu}{\partial t}\right]+O\left(\left(\lambda^{\theta}\mu\right)^{n}\right)\right]dt$
$+$ $\left[\mu^{n}+\frac{n}{\mu}\frac{\partial\mu}{\partial a}+n\left(\lambda^{\theta}\mu\right)^{n}C\left[\frac{\theta}{\lambda}\frac{\partial\lambda}{\partial a}+\frac{1}{\mu}\frac{\partial\mu}{\partial a}\right]+O\left(\left(\lambda^{\theta}\mu\right)^{n}\right)\right]da$
$+$ $O\left(\left(\lambda|^{\theta}\mu\right)^{{n}}dy\right),$
By (4.2), the lower bound in (-4.2) and (4.21) we get
$\left(\begin{matrix}dm_{6,x}\\ dm_{6,y}\end{matrix}\right)$ $\geq$ $\left(\begin{matrix}A_{5}&B_{5}\\ C_{5}&O\left(\lambda^{n}\right)+D_{5}(C-E)\left(\lambda|^{\theta}\mu\right)^{{n}}\end{matrix}\right)\left(\begin{matrix}dm_{5,x}\\ dm_{5,y}\end{matrix}\right)$
$+$ $\left(\begin{matrix}O\left(dt\right)+O\left(da\right)\\ \left[d_{5}(C-E)^{2}\left(\lambda|^{\theta}\mu\right)^{{2n}}+O\left(\lambda^{n}\right)\right]dt+\left[1+O\left(\left(\lambda|^{\theta}\mu\right)^{{2n}}\right)\right]da\end{matrix}\right).$
As consequence,
$dm_{6,x}$ $\geq$ $\left[B_{5}\frac{n}{\mu}\frac{\partial\mu}{\partial t}+O\left(1\right)\right]dt+\left[B_{5}\mu^{n}+B_{5}\frac{n}{\mu}\frac{\partial\mu}{\partial a}+O\left(1\right)\right]da$
$+$ $O\left(\left(\lambda|^{\theta}\mu\right)^{{n}}dy\right),$
$dm_{6,y}$ $\geq$ $\left[D_{5}(C-E)\left(\lambda^{\theta}\mu\right)^{n}\frac{n}{\mu}\frac{\partial\mu}{\partial t}+D_{5}(C-E)n\left(\lambda^{\theta}\mu\right)^{2n}\left[C\frac{\theta}{\lambda}\frac{\partial\lambda}{\partial t}+\frac{(C-E)}{\mu}\frac{\partial\mu}{\partial t}\right]\right$.
$+$ $\left.O\left(\left(\lambda^{\theta}\mu\right)^{2n}\right)\right]dt$
$+$ $\left[D_{5}(C-E)\left(\lambda^{\theta}\mu^{2}\right)^{n}+1+D_{5}(C-E)\left(\lambda^{\theta}\mu\right)^{n}\frac{n}{\mu}\frac{\partial\mu}{\partial a}\right$.
$+$ $\left.D_{5}(C-E)n\left(\lambda^{\theta}\mu\right)^{2n}\left[C\frac{\theta}{\lambda}\frac{\partial\lambda}{\partial a}+\frac{(C-E)}{\mu}\frac{\partial\mu}{\partial a}\right]\right$.
$+$ $\left.O\left(\left(\lambda^{\theta}\mu\right)^{2n}\right)\right]da+O\left(\left(\lambda|^{\theta}\mu\right)^{{2n}}\right)dy,$
where $D_{5},B_{5},(C-E)\neq 0$ and $\lambda^{\theta}\mu^{2}>1$ (see (2.16)). Similarly, By (4.2), the upper bound in (-4.2) and (4.21) we get
$\left(\begin{matrix}dm_{6,x}\\ dm_{6,y}\end{matrix}\right)$ $\leq$ $\left(\begin{matrix}A_{5}&B_{5}\\ C_{5}&{D_{5}T}/{\mu^{n_{0}/2}}+O\left(\lambda^{\theta}\mu\right)^{n}\end{matrix}\right)\left(\begin{matrix}dm_{5,x}\\ dm_{5,y}\end{matrix}\right)$
$+$ $\left(\begin{matrix}O\left(dt\right)+O\left(da\right)\\ \left[{d_{5}T^{2}}/{\mu^{n_{0}}}+O\left(\lambda^{\theta}\mu\right)^{2n}\right]dt+\left[1+O\left({1}/{\mu^{n_{0}}}\right)\right]da\end{matrix}\right)$
As consequence, if $(t,a)\in\mathcal{B}_{n}$
$dm_{6,x}$ $\leq$ $\left[B_{5}\frac{n}{\mu}\frac{\partial\mu}{\partial t}+O\left(1\right)\right]dt+\left[B_{5}\mu^{n}+B_{5}\frac{n}{\mu}\frac{\partial\mu}{\partial a}+O\left(1\right)\right]da$
$+$ $O\left(\left(\lambda|^{\theta}\mu\right)^{{n}}dy\right),$
$dm_{6,y}$ $\leq$ $\left[\frac{D_{5}T}{\mu^{n_{0}/2}}\frac{n}{\mu}\frac{\partial\mu}{\partial t}+\frac{n}{\mu}\frac{\partial\mu}{\partial t}\frac{D_{5}T^{2}}{\mu^{n_{0}}}+O\left(\frac{1}{\mu^{n_{0}}}\right)\right]dt$ (4.43)
$+$ $\left[D_{5}T\mu^{n-n_{0}/2}+1+\frac{D_{5}T}{\mu^{n_{0}/2}}\frac{n}{\mu}\frac{\partial\mu}{\partial a}+\frac{n}{\mu}\frac{\partial\mu}{\partial a}\frac{D_{5}T^{2}}{\mu^{n_{0}}}+O\left(\frac{1}{\mu^{n_{0}}}\right)\right]da$
$+$ $O\left(\left(\lambda|^{\theta}\mu\right)^{{2n}}\right)dy,$
where $D_{5},B_{5},T\neq 0$ and $\mu^{n-n_{0}/2}>1$ (see (2.16)).
In particular, when $a=a_{n}$, using the same equalities as before, (4.22) and (4.2) we get,
$dm_{6,x}$ $=$ $\left[B_{5}\frac{n}{\mu}\frac{\partial\mu}{\partial t}+O\left(1\right)\right]dt+\left[B_{5}\mu^{n}+B_{5}\frac{n}{\mu}\frac{\partial\mu}{\partial a}+O\left(1\right)\right]da$
$+$ $O\left(\left(\lambda|^{\theta}\mu\right)^{{n}}dy\right),$
and
$dm_{6,y}$ $=$ $\left[D_{5}C\left(\lambda^{\theta}\mu\right)^{n}\frac{n}{\mu}\frac{\partial\mu}{\partial t}+D_{5}Cn\left(\lambda^{\theta}\mu\right)^{2n}C\left[\frac{\theta}{\lambda}\frac{\partial\lambda}{\partial t}+\frac{1}{\mu}\frac{\partial\mu}{\partial t}\right]\right$.
$+$ $\left.O\left(\left(\lambda^{\theta}\mu\right)^{2n}\right)\right]dt$
$+$ $\left[D_{5}C\left(\lambda^{\theta}\mu^{2}\right)^{n}+1+D_{5}C\left(\lambda^{\theta}\mu\right)^{n}\frac{n}{\mu}\frac{\partial\mu}{\partial a}\right$.
$+$ $\left.D_{5}Cn\left(\lambda^{\theta}\mu\right)^{2n}C\left[\frac{\theta}{\lambda}\frac{\partial\lambda}{\partial a}+\frac{1}{\mu}\frac{\partial\mu}{\partial a}\right]\right$.
$+$ $\left.O\left(\left(\lambda^{\theta}\mu\right)^{2n}\right)\right]da+O\left(\left(\lambda|^{\theta}\mu\right)^{{2n}}\right)dy,$
where, by (2.18)
$\left[\frac{\theta}{\lambda}\frac{\partial\lambda}{\partial t}+\frac{1}{\mu}\frac{\partial\mu}{\partial t}\right]>0.$
Refer to Figure 4. The formula for $m_{6,y}$ holds in general for all $\Delta t,\Delta a$ and $\Delta y$. However at the point $(t_{1},a)$ when $\Delta t=\Delta a=0$ we have ${\partial m_{6,y}}/{\partial y}=0$. Hence, the Taylor polynomial of second order for $\Delta m_{6,y}$ does not contain a linear term in $\Delta y$. As consequence ${\partial z^{(3)}_{y}}/{\partial t}={\partial m_{6,y}}/{\partial t}$ and ${\partial z^{(3)}_{y}}/{\partial a}={\partial m_{6,y}}/{\partial a}$. It is left to prove that $C,D_{5}$ converge. This is achieved since they are part of the derivative $DF^{N}$ which converges to $DF^{N}(0,1)$. The proposition follows. ∎
In the next proposition we are going to prove that secondary tangencies exist for certain parameters in $a_{n}$ and they are at distance of order $1/n$ to each other. The result is achieved by comparing the rate of speed of $W^{u}_{\text{loc}}(z^{(3)})$ and $W_{n-n_{0}}$ when changing parameters in the phase space and it relies on the cancellation of the main term.
Proposition 3.
For all $t\in\left(-t_{0},t_{0}\right)$ and for $n$ large enough there exists $t_{n}\in\left(-t_{0},t_{0}\right)$ such that the following holds:
    $|t-t_{n}|=O\left(\frac{1}{n}\right)$,
    $F_{t_{n},a_{n}(t_{n})}$ has a secondary tangency $q_{1,n,n_{0}}$.
Proof.
Choose $t^{*}\in\left(-t_{0},t_{0}\right)$. Use the notation of Proposition 2 and notice that, by the choice of $\theta$, see (2.18),
$DC^{2}\left[\frac{\theta}{\lambda}\frac{\partial\lambda}{\partial t}+\frac{1}{\mu}\frac{\partial\mu}{\partial t}\right]=v>0$
Take $n$ large enough and a point $(t,a_{n}(t))\in A_{n}$ near $(t^{*},a_{n}(t^{*}))\in A_{n}$. By Proposition 2 and Lemma 3
$z^{(3)}_{y}\left(t,a_{n}(t)\right)$ $=$ $z^{(3)}_{y}\left(t^{*},a_{n}(t^{*})\right)+\int_{0}^{(t-t^{*})}\left[\frac{\partial z^{(3)}}{\partial t}+\frac{\partial z^{(3)}}{\partial a}\frac{da_{n}}{dt}\right]dt=z^{(3)}_{y}\left(t^{*},a_{n}(t^{*})\right)$ (4.48)
$+$ $\left\{DC\left(\lambda^{\theta}\mu\right)^{n}\frac{n}{\mu}\frac{\partial\mu}{\partial t}+DC^{2}n\left(\lambda^{\theta}\mu\right)^{2n}\left[\frac{\theta}{\lambda}\frac{\partial\lambda}{\partial t}+\frac{1}{\mu}\frac{\partial\mu}{\partial t}\right]\right$.
- $\left.DC\left(\lambda^{\theta}\mu\right)^{n}\frac{n}{\mu}\frac{\partial\mu}{\partial t}+O\left(\left(\lambda^{\theta}\mu\right)^{2n}\right)\right\}(t-t^{*})$
$=$ $z^{(3)}_{y}\left(t^{*},a_{n}(t^{*})\right)+n\left(\lambda^{\theta}\mu\right)^{2n}\frac{v}{2}(t-t^{*}),$
where we used that $\lambda^{2\theta}\mu^{3}>1$ and $\lambda^{\theta}\mu^{2}>1$ see (2.16). Observe that we have a cancellation of the dominant terms in the partial derivatives obtained by combining Proposition 2 and Lemma 3. Let $n^{*}_{0}$ be such that $W^{u}_{\text{loc}}\left(z^{(3)}\left(t^{*},a_{n}(t^{*})\right)\right)\subset S_{n-n_{0}^{*}}$, then, by (2.8),
$z^{(3)}_{y}\left(t^{*},a_{n}(t^{*})\right)\geq\frac{2}{\mu}\frac{1}{\mu^{n-n^{*}_{0}}}.$ (4.49)
By (4.48) and (4.49)
$z^{(3)}_{y}\left(t,a_{n}(t)\right)\geq\frac{2}{\mu}\frac{1}{\mu^{n-n^{*}_{0}}}+n\left(\lambda^{\theta}\mu\right)^{2n}\frac{v}{2}(t-t^{*}).$ (4.50)
Choose $\kappa\geq 2$. Then by (2.8), in a neighborhood of $(z^{(3)}\left(t^{*},a_{n}(t^{*})\right),t^{*},a_{n}(t^{*}))$
$\max W_{n-n_{0}^{*}-\kappa}$ $\leq$ $\frac{2}{\mu^{n-n^{*}_{0}-\kappa}}.$
From the previous inequality, (4.50), we get that $W^{u}_{\text{loc}}(z^{(3)}(t,a_{n}(t)))$ is above $W_{n-n_{0}^{*}-\kappa}$ if
$n\left(\lambda^{\theta}\mu\right)^{2n}\frac{v}{2}(t-t^{*})\geq\frac{2}{\mu^{n-n^{*}_{0}}}\left[\mu^{\kappa}-\frac{2}{\mu}\right]\geq 2C^{\prime}\left(\lambda^{\theta}\mu\right)^{2n}\left[\mu^{\kappa}-\frac{2}{\mu}\right],$
where we also used (4.17) and $C^{\prime}$ is a uniform constant. As consequence, if
$t-t^{*}\geq\frac{4C^{\prime}}{n}\frac{\mu^{\kappa}}{v},$
then, $W^{u}_{\text{loc}}(z^{(3)}(t,a_{n}(t)))$ is above $W_{n-n_{0}^{*}-\kappa}$. Because $W^{u}_{\text{loc}}(z^{(3)}(t^{*},a_{n}(t^{*})))$ contains a point below $W_{n-n_{0}^{*}-1}$, there exists a parameter between $(t,a_{n}(t))$ and $(t^{*},a_{n}(t^{*}))$ for which a secondary homoclinic tangency of type $n_{0}=n_{0}^{*}+\kappa$ occurs and $|t^{*}-t|=O\left(\frac{1}{n}\right)$. Moreover by Lemma 2, (2.8), Lemma 11 and the definition of $\alpha$ we get
$\frac{\partial W_{n-n_{0}}}{\partial t}(q_{1,n,n_{0}})=-2n\left(\lambda^{\theta}\mu\right)^{2n}\mu^{\kappa-1}\left(1-\alpha-\frac{\kappa}{n}\right)\frac{\partial\mu}{\partial t}+O\left(\left(\lambda^{\theta}\mu\right)^{2n}\right).$ (4.51)
Observe that, from $(F6)$, the value of $\alpha$ and the lower bound for $\theta$, see (2.17), we have
$\left(1-\alpha-\frac{\kappa}{n}\right)\frac{\partial\mu}{\partial t}>0.$ (4.52)
Figure 5: Position of $z^{(3)}$
The next lemma estimates the curvature of any secondary tangency point of type $n_{0}$ in $\mathcal{B}_{n}$ satisfies the following estimate.
Lemma 12.
Let $(t,a)\in\mathcal{B}_{n}$ be such that $W^{u}_{\text{\rm loc}}(z^{(3)}(t,a))$ has a secondary tangency $q_{1,n,n_{0}}$ of type $n_{0}$, then $q_{1,n,n_{0}}$ is non degenerate. Namely, $W^{u}_{\text{\rm loc}}(q_{1,n,n_{0}})$ is the graph of a function and its curvature satisfies
$\text{\rm curv}\left(W^{u}_{\text{\rm loc}}(q_{1,n,n_{0}})\right)\geq K\left(\frac{\mu^{4}}{\lambda^{2-3\theta}}\right)^{n},$
with $K>0$ a uniform positive constant.
Proof.
Use coordinates centered at $z^{(2)}$ and let $(\Delta x,\Delta y)\in W^{u}_{\text{loc}}(z^{(2)})$ such that $F^{N}_{t,a}(\Delta x,\Delta y)=q_{1,n,n_{0}}$. Then $0\leq\Delta y\leq L\left(\lambda^{\theta}\mu\right)^{n}$. Let $v=\left(\begin{matrix}v_{1}\\ v_{2}\end{matrix}\right)$ be the tangent vector at $(\Delta x,\Delta y)$ to $W^{u}_{\text{loc}}(z^{(2)})$. We use Lemma 8 and we obtain that
$|v_{2}|=\left[2Q_{2}\left(\frac{\mu}{\lambda^{2}}\right)^{n}|\Delta x|+O\left(|\Delta x|^{2}\left(\frac{\mu}{\lambda^{3}}\right)^{n}\right)\right]|v_{1}|.$ (4.53)
By Lemma 1 we have that ${\partial^{2}Y}/{\partial y^{2}}$ is bounded away from zero in a neighborhood of $c$. As consequence, by Lemma 7, Lemma 7 and Lemma 5 we have $DF^{N}_{t,a}(\Delta x,\Delta y)=\left(\begin{matrix}A&B\\ C&D\end{matrix}\right)$ where $D\geq d\left(\lambda^{\theta}\mu\right)^{n}$ and $d>0$ is a uniform constant. Because $F^{N}_{t,a}(\Delta x,\Delta y)$ is a tangency at $q_{1,n,n_{0}}$, then
$Cv_{1}+Dv_{2}=O\left(\left(\frac{\lambda}{\mu}\right)^{n-n_{0}}\left(|v_{1}|+|v_{2}|\right)\right),$ (4.54)
where we used (2.9). The proof of the lemma is completed by following exactly that of Lemma 10. ∎
5 Curves of secondary tangencies
In the previous section, we proved the existence of tangency points on the curve $A_{n}$. In the sequel, we extend these points to create global curves. All points along these curves have a secondary tangency; refer to Figure 6. In particular, we prove that the curves of secondary tangencies are actually long uniform curves that never leave the strip $\mathcal{B}_{n}$.
Figure 6: Curves of Secondary Tangencies
5.1 Continuation of secondary tangencies: the implicit function theorem
The next lemma is the first step in proving the existence of local tangency curves. In particular, we prove that each parameter in $\mathcal{B}_{n}$ corresponding to a map having a secondary tangency of type $n_{0}$ can be locally enlarged to create a curve. Each point on this curve corresponds to a map with a $n_{0}$ secondary tangency. This result is achieved by implementing and using the implicit function theorem.
Lemma 13.
Let $(t^{*},a^{*})\in{\mathcal{B}_{n}}$ be a parameter corresponding to a map with a secondary tangency of type $n_{0}$. Then there exist $\epsilon>0$ and a smooth curve
$b_{n,n_{0}}:[t^{*}-\epsilon,t^{*}+\epsilon]\mapsto\mathbb{R}$
such that, for all $t\in[t^{*}-\epsilon,t^{*}+\epsilon]$ the parameter $(t,b_{n,n_{0}}(t))$ corresponds to a map with a secondary tangency of type $n_{0}$. Moreover
$\frac{(C-E)}{T}\frac{n}{\mu^{n+1}}\frac{\partial\mu}{\partial t}+O\left(n\lambda^{\theta n}\right)\leq\frac{db}{dt}\leq-\frac{T}{(C-E)}\frac{n}{\mu^{n+1}}\frac{\partial\mu}{\partial t}+O\left(\frac{n}{\mu^{n_{0}/2}}\right).$
In particular, there exists a positive constant $v$ such that,
$\frac{db}{dt}=\frac{da_{n}}{dt}+v{n\lambda^{\theta n}}+O\left(\lambda^{\theta n}\right)$
for all $t\in[\tilde{t}^{*}-\epsilon,\tilde{t}^{*}+\epsilon]$, with $(\tilde{t}^{*},b_{n,n_{0}}(\tilde{t}^{*}))\in a_{n}$.
Proof.
Let $(t^{*},a^{*})\in\mathcal{B}_{n}\cap A_{n}$ be a parameter corresponding to a map with a secondary tangency of type $n_{0}$, $q_{1,n,n_{0}}$. Its existence is proved in Proposition 3. We construct a local function whose graph is contained in a neighborhood of $t^{*}$. Let $m_{1}\in W^{u}_{\text{loc}}(c)$ such that $F_{t,a}^{3N+(1+\theta)n}(m_{1})=q_{1,n,n_{0}}$. To describe the perturbation of $W^{u}_{\text{loc}}(q_{1,n,n_{0}})$ we define the following function by choosing coordinates centered in the image at $q_{1,n,n_{0}}$. Take some $\epsilon>0$ and consider the ${\mathcal{C}}^{4}$} function $(\tilde{x},\tilde{y}):(-\epsilon,\epsilon)^{3}\mapsto\mathbb{R}^{m-1}\times\mathbb{R}$ defined by
$\left(\tilde{x}(\Delta y,\Delta t,\Delta a),\tilde{y}(\Delta y,\Delta t,\Delta a)\right)=F_{t+\Delta t,a+\Delta a}^{3N+(1+\theta)n}\left(m_{1}+(0,\Delta y)\right),$
which describes $W^{u}_{\text{loc}}\left(F_{t+\Delta t,a+\Delta a}^{3N+(1+\theta)n}(m_{1})\right)$. Observe that the manifolds $W_{n-n_{0}}$ of $F_{t+\Delta t,a+\Delta a}$ are described locally, near $q_{1,n,n_{0}}$, as the graph of a ${\mathcal{C}}^{4}$} function
$w_{n-n_{0}}:(\Delta x,\Delta t,\Delta a)\longmapsto w_{n-n_{0}}(\Delta x,\Delta t,\Delta a)\in\mathbb{R}.$
The curves of secondary tangencies will be constructed using the implicit function theorem. For this aim define the ${\mathcal{C}}^{2}$} function $\Psi:[-\epsilon,\epsilon]^{3}\to\mathbb{R}^{2}$ as
$\Psi\left(\Delta y,\Delta t,\Delta a\right)=\left(\begin{matrix}\tilde{y}(\Delta y,\Delta t,\Delta a)-w_{n-n_{0}}\left(\tilde{x}(\Delta y,\Delta t,\Delta a),\Delta t,\Delta a\right)\\ \frac{d\tilde{y}}{d\tilde{x}}\left(\Delta y,\Delta t,\Delta a\right)-\frac{dw_{n-n_{0}}}{d\tilde{x}}\left(\tilde{x}(\Delta y,\Delta t,\Delta a),\Delta t,\Delta a\right)\end{matrix}\right)=\left(\begin{matrix}\psi_{1}\\ \psi_{2}\end{matrix}\right).$
Observe that $\Psi^{-1}(0)$ describes locally the perturbation of the secondary tangency $q_{1,n,n_{0}}$ and that
$D\Psi({0,0,0})=\left(\begin{matrix}0&\left(\lambda^{\theta}\mu\right)^{n}\Psi_{1,2}&\left(\lambda^{\theta}\mu^{2}\right)^{n}\Psi_{1,3}\\ \left(\frac{\mu^{3}}{\lambda^{2-2\theta}}\right)^{n}\Psi_{2,1}&\Psi_{2,2}&\Psi_{2,3}\end{matrix}\right).$ (5.1)
In particular, by (4.2), (2.9), (4.2), (4.17) (in the estimation of the order term)
$\Psi_{1,2}$ $=$ $\frac{1}{\left(\lambda^{\theta}\mu\right)^{n}}\left[\frac{\partial m_{6,y}}{\partial t}-\frac{\partial W_{n-n_{0}}}{\partial x}\frac{\partial m_{6,x}}{\partial t}-\frac{\partial W_{n-n_{0}}}{\partial t}\right]$ (5.2)
$=$ $DC\frac{n}{\mu}\frac{\partial\mu}{\partial t}+DC^{2}n\left(\lambda^{\theta}\mu\right)^{n}\left[\frac{\theta}{\lambda}\frac{\partial\lambda}{\partial t}+\frac{1}{\mu}\frac{\partial\mu}{\partial t}\right]$
$+$ $O\left(\left(\lambda^{\theta}\mu\right)^{n}\right)+O\left(n\left(\frac{\lambda}{\mu}\right)^{n-n_{0}}\frac{1}{\left(\lambda^{\theta}\mu\right)^{n}}\right)-\frac{1}{\left(\lambda^{\theta}\mu\right)^{n}}\frac{\partial W_{n-n_{0}}}{\partial t}$
$=$ $DC\frac{n}{\mu}\frac{\partial\mu}{\partial t}+n\left(\lambda^{\theta}\mu\right)^{n}\left\{{DC^{2}}\left[\frac{\theta}{\lambda}\frac{\partial\lambda}{\partial t}+\frac{1}{\mu}\frac{\partial\mu}{\partial t}\right]-\frac{1}{n\left(\lambda^{\theta}\mu\right)^{2n}}\frac{\partial W_{n-n_{0}}}{\partial t}\right\}$
$+$ $O\left(\left(\lambda^{\theta}\mu\right)^{n}\right)=DC\frac{n}{\mu}\frac{\partial\mu}{\partial t}+n\left(\lambda^{\theta}\mu\right)^{n}v+O\left(\left(\lambda^{\theta}\mu\right)^{n}\right)$
$=$ $DC\frac{n}{\mu}\frac{\partial\mu}{\partial t}+O\left(n\left(\lambda^{\theta}\mu\right)^{n}\right),$
and by (4.2), (2.9), (4.2), Lemma 2 and (4.17) (in the estimation of the order term)
$\Psi_{1,3}$ $=$ $\frac{1}{\left(\lambda^{\theta}\mu^{2}\right)^{n}}\left[\frac{\partial m_{6,y}}{\partial a}-\frac{\partial W_{n-n_{0}}}{\partial x}\frac{\partial m_{6,x}}{\partial a}-\frac{\partial W_{n-n_{0}}}{\partial a}\right]$ (5.3)
$=$ $\frac{1}{\left(\lambda^{\theta}\mu^{2}\right)^{n}}\left\{DC\left(\lambda^{\theta}\mu^{2}\right)^{n}+1+DC\left(\lambda^{\theta}\mu\right)^{n}\frac{n}{\mu}\frac{\partial\mu}{\partial a}\right$.
$+$ $\left.DC^{2}n\left(\lambda^{\theta}\mu\right)^{2n}\left[\frac{\theta}{\lambda}\frac{\partial\lambda}{\partial a}+\frac{1}{\mu}\frac{\partial\mu}{\partial a}\right]+O\left(\left(\lambda^{\theta}\mu\right)^{2n}\right)\right$.
$+$ $\left.O\left(\mu^{n}\left(\frac{\lambda}{\mu}\right)^{n-n_{0}}\right)+O\left(\frac{n-n_{0}}{\mu^{n-n_{0}}}\right)\right\}$
$=$ $DC\left[1+O\left({\left({\lambda}^{1+\theta-\alpha}{\mu}\right)^{n}}\right)\right].$
Moreover $\Psi_{1,2},\Psi_{1,3}>0$ and $\Psi_{2,1}\neq 0$, see Lemma 12, (4.2) and (2.10). Because $D\Psi({0,0,0})$ is onto we get that the set of secondary tangencies $\Psi^{-1}(0)$ is locally the graph of a ${\mathcal{C}}^{2}$} function $b$. Moreover the tangent space of $\Psi$ in zero satisfies $T_{(0,0,0)}\Psi^{-1}(0)=\text{Ker}D\Psi$. Hence
$\left(\lambda^{\theta}\mu\right)^{n}\Psi_{1,2}\Delta t+\left(\lambda^{\theta}\mu^{2}\right)^{n}\Psi_{1,3}\Delta b=0,$
and
$\frac{db}{dt}=-\frac{\Psi_{1,2}}{\Psi_{1,3}}\frac{1}{\mu^{n}}=-\frac{n}{\mu^{n+1}}\frac{\partial\mu}{\partial t}-{n\lambda^{\theta n}}v+O\left(\lambda^{\theta n}\right).$ (5.4)
In particular $db/dt$ is a negative number and $b$ is the graph of a smooth function. Furthermore, by the previous estimate on the slope of $b$, by Proposition 3 and by (2.11), the curve $b$ have a angle satisfying the following,
$\frac{db}{dt}=\frac{da_{n}}{dt}-{n\lambda^{\theta n}}v+O\left(\lambda^{\theta n}\right),$ (5.5)
with $v$ a positive constant.
Let $(t_{0},a_{0})$ be any parameter in $\mathcal{B}_{n}$ corresponding to a map with a secondary tangency of type $n_{0}$, $q_{1,n,n_{0}}$. The proof is exactly the same as the previous case. In particular, by (4.2), (2.9), (4.2), (4.17) (in the estimation of the order term)
$\Psi_{1,2}$ $\geq$ $D(C-E))\frac{n}{\mu}\frac{\partial\mu}{\partial t}+O\left(n\left(\lambda^{\theta}\mu\right)^{n}\right),$
and by (4.2), (2.9), (4.2), Lemma 2 and (4.17) (in the estimation of the order term)
$\Psi_{1,3}$ $\geq$ $D(C-E)\left[1+O\left({\left({\lambda}^{1+\theta-\alpha}{\mu}\right)^{n}}\right)\right].$
Moreover, by (4.43), (2.9), (4.2),
$\Psi_{1,2}$ $\leq$ $DT\frac{n}{\mu}\frac{\partial\mu}{\partial t}+O\left(\frac{n}{\mu^{n_{0}/2}}\right),$
and by (4.2), (2.9), (4.2), Lemma 2
$\Psi_{1,3}$ $\leq$ $DT\left[1+O\left(\frac{1}{\mu^{n-n_{0}/2}}\right)\right].$
Since, by Proposition 2 and (F6), $D,C-E,T,{\partial\mu}/{\partial t}>0$, then $\Psi_{1,2}>0$, $\Psi_{1,3}>0$ and $\Psi_{2,1}>0$, see Lemma 12. As before, because $D\Psi({0,0,0})$ is onto we get that the set of secondary tangencies $\Psi^{-1}(0)$ is locally the graph of a ${\mathcal{C}}^{2}$} function $b$. Moreover
$\frac{(C-E)}{T}\frac{n}{\mu^{n+1}}\frac{\partial\mu}{\partial t}+O\left(n\lambda^{\theta n}\right)\leq\frac{db}{dt}\leq-\frac{T}{(C-E)}\frac{n}{\mu^{n+1}}\frac{\partial\mu}{\partial t}+O\left(\frac{n}{\mu^{n_{0}/2}}\right).$ (5.6)
The proof is completed.
5.2 Limitations to the existence of secondary tangencies
We prove now that for any $n_{0}$ by moving in the direction of the $a$ parameter one can find only one tangency of type $n_{0}$ as well as when one move on the $a_{n}$ curve.
Lemma 14.
For $n$ large enough,
$\frac{\partial z_{3,y}}{\partial a}-\frac{\partial W_{n-n_{0}}}{\partial a}>0$
and
$\frac{\partial z_{3,y}}{\partial t}+\frac{\partial z_{3,y}}{\partial a}\frac{da_{n}}{dt}>0.$
Proof.
Using the estimates in Proposition 2, one can notice that the first coefficient of $\partial z_{3,y}/{\partial a}$ is strictly larger than one, while the first coefficient of $\partial W_{n-n_{0}}/{\partial a}$ is strictly smaller than one, see Lemma 2. The first inequality follows.
For the second one, if $(t,a)\in\mathcal{B}_{n}$, by Proposition 2 and (2.11),
$\frac{\partial z_{3,y}}{\partial t}+\frac{\partial z_{3,y}}{\partial a}\frac{da_{n}}{dt}=DC^{2}n\left(\lambda^{\theta}\mu\right)^{2n}\left[\frac{\theta}{\lambda}\frac{\partial\lambda}{\partial t}+\frac{1}{\mu}\frac{\partial\mu}{\partial t}\right]+O\left(\left(\lambda^{\theta}\mu\right)^{2n}\right).$
The statement is proved using Proposition 2 and (2.18). ∎
The following lemma shows that one cannot find a secondary tangency of type $n_{0}$ on the line defining the lower bound of the $\mathcal{B}_{n}$ strip.
Lemma 15.
There are no curve secondary tangencies of type $n_{0}$ along the line $a_{n}-E{\lambda^{\theta n}}$ where $E=C-\frac{\epsilon}{\mu^{n_{0}/2}}$, with $\epsilon>0$ small.
Proof.
By contradiction, there exists a $n_{0}$ tangency on the line $a_{n}-E{\lambda^{\theta n}}$ with $E=C-\frac{\epsilon}{\mu^{n_{0}/2}}$, $\epsilon>0$ small. This means that $z_{3,y}\mu^{n_{0}}=2$. By (4.6) and Lemma 7,
$z_{2,y}=C\left(\lambda^{\theta}\mu\right)^{n}+\mu^{n}(a-a_{n}(t))+O\left(\mu^{n}\lambda^{2\theta n}\right)$
and by (4.19)
$z_{3,y}=Q\left(z_{2,y}\right)^{2}+O\left(\frac{1}{\mu^{n}}\right),$
Hence,
$z_{3,y}\mu^{n_{0}}$ $=$ $Q(C-E)^{2}\left(\lambda^{\theta}\mu\right)^{2n}\mu^{n_{0}}+O\left(\frac{1}{\mu^{n-n_{0}}}\right)$ (5.7)
$=$ $Q\frac{\epsilon^{2}}{\mu_{0}^{n_{0}}}\left(\lambda^{\theta}\mu\right)^{2n}\mu_{0}^{n_{0}}+O\left(\frac{1}{\mu^{n-n_{0}}}\right)$
$=$ $Q{\epsilon^{2}}\left(\lambda^{\theta}\mu\right)^{2n}+O\left(\frac{1}{\mu^{n-n_{0}}}\right).$
Since $\lambda^{\theta}\mu<1$, for $\epsilon$ small enough, this last quantity cannot be equal to $2$. The proof is completed. ∎
The following lemma shows that one cannot find a $n_{0}$ tangency on the line $a_{n}+\frac{T}{\mu^{n+n_{0}/2}}$ with $T$ a large constant.
Lemma 16.
For $T$ large enough, there are no secondary tangencies of type $n_{0}$ along the line $a_{n}+\frac{T}{\mu^{n+n_{0}/2}}$, with $T$ large.
Proof.
As in the proof of the previous lemma, by (4.6) and Lemma 7,
$z_{2,y}=C\left(\lambda^{\theta}\mu\right)^{n}+\mu^{n}(a-a_{n}(t))+O\left(\mu^{n}\lambda^{2\theta n}\right)$
and by (4.19)
$z_{3,y}=Q\left(z_{2,y}\right)^{2}+O\left(\frac{1}{\mu^{n}}\right)$ $=$ $Q\left(C\left(\lambda^{\theta}\mu\right)^{n}+\frac{T}{\mu^{n_{0}/2}}\right)^{2}+O\left(\frac{1}{\mu^{n}}\right)$
$>$ $Q\frac{T^{2}}{\mu^{n_{0}}}+O\left(\frac{1}{\mu^{n}}\right).$
Hence,
$z_{3,y}\mu^{n_{0}}$ $>$ $QT^{2}+O\left(\frac{1}{\mu^{n-n_{0}}}\right).$ (5.8)
For $T$ large enough, this last quantity cannot be equal to $2$, so in particular one cannot have a $n_{0}$ tangency. The proof is completed. ∎
5.3 Global curves of secondary tangencies
In the next proposition we prove the existence of global curves of secondary tangencies contained in $\left([-t_{0},t_{0}]\times\mathbb{R}\right)\cap\mathcal{B}_{n}$, see Figure 6.
Proposition 4.
There exists a smooth function $b_{n,n_{0}}:[-t_{0},t_{0}]\to\mathbb{R}$ such that for all $t\in[-t_{0},t_{0}]$ the parameter point $(t,b_{n,n_{0}}(t))$ corresponds to a map with a secondary tangency of type $n_{0}$. Moreover, the graph of $b_{n,n_{0}}$ is contained in $\mathcal{B}_{n}$.
Proof.
Let $(t^{*},a^{*})\in\mathcal{B}_{n}$ be a parameter corresponding to a map with a secondary tangency of type $n_{0}$. By Lemma 13 there exist $\epsilon>0$ and a smooth curve
$b_{n,n_{0}}:[t^{*}-\epsilon,t^{*}+\epsilon]\mapsto\mathbb{R}$
such that, for all $t\in[t^{*}-\epsilon,t^{*}+\epsilon]$ the parameter $(t,b_{n,n_{0}}(t))$ corresponds to a map with a secondary tangency of type $n_{0}$. We want now to prove that this curve does not stop. Suppose it does.
Consider a sequence $t_{n}\in[t^{*}-\epsilon,t^{*}+\epsilon]$ of secondary tangencies of type $n_{0}$ that converges to a boundary point of $t^{*}-\epsilon,t^{*}+\epsilon]$. Since the set of tangencies is a close set, the limit point is again a tangency. Because of Lemma 14 one cannot have more than one limit point, indeed one can have only one tangency point while moving in the $a$ direction. By lemmas 15 and 16 this unique limit point cannot be on the boundary lines that define $\mathcal{B}_{n}$. Such a limit point is then contained in $\mathcal{B}_{n}$. One can then apply again Proposition 13 to this point and build a smooth continuous curve of maps with secondary tangencies of type $n_{0}$. This new courve $b^{\prime}_{n,n_{0}}$ coincides with the initial one $b_{n,n_{0}}$, because of the continuity of the implicit function theorem used to extend the curve in Proposition 13. In this way one creates a curve which is a graph of a smooth function. This is obtained by using the angle estimate in Proposition 4 and Lemma 14. Because of lemmas 15 and 16 the curve built cannot exit the interior of $\mathcal{B}_{n}$. ∎
In the next lemma we prove that there exists a neighborhood of the curve of secondary tangencies built above such that the initial family restricted to this neighborhood is an unfolding of the secondary tangency.
Lemma 17.
Let $b_{n,n_{0}}:[-t_{0},t_{0}]\to\mathbb{R}$ be a curve of secondary tangencies. There exists a neighborhood $U$ of $b_{n,n_{0}}$ such that $F_{|U}$ can be reparametrized to become an unfolding.
Proof.
Let $U$ be a neighborhood of $b_{n,n_{0}}$ which is strictly contained in $\mathcal{B}_{n}$. Observe that the new family is a restriction of our original family and the secondary tangency curve $b_{n,n_{0}}$ describes homoclinic tangencies associated to the original saddle point. The transversal intersection $q_{2}$ it is still present in the new family. As a consequence, conditions (f1),(f2),(f3),(f4),(f5),(f6),(F1),(F2),(F3),(F5)(f1),(f2),(f3),(f4),(f5),(f6),(F1),(F2),(F3),(F5) are automatically satisfied.
Observe that $f_{n,n_{0}}^{n-n_{0}}\left(q_{1,n,n_{0}}\right)$ converges to $q_{2}$. Property $(f6)$ implies $(f7)$ for the secondary tangency $q_{1,n,n_{0}}$.
Use the notation $m_{1},m_{2},m_{3},m_{4},m_{5}$ from Proposition 2 and observe that $m_{1}\in W^{u}_{\text{loc}}(0)$ close to $q_{3}$. This proves $(f8)$ for the secondary tangency $q_{1,n,n_{0}}$.
For proving $(F4)$ use the ${\mathcal{C}}^{2}$} function $b_{n,n_{0}}$ and observe that the maps on the graph of this curve, $b_{n,n_{0}}$, have a non degenerate homoclinic tangency.
Observe that $(P1)$ follows by the fact that $b_{n,n_{0}}$ is the curve of tangencies and $(P2)$ from the fact that $\Psi_{1,3}\neq 0$, see Lemma 13, Equation (5.1). ∎
6 Newhouse phenomenon
In this section, we identify parameters corresponding to maps that exhibit infinitely many sinks, a phenomenon referred to as the Newhouse phenomenon. Although this result itself is not new, our approach is constructive in nature and departs fundamentally from earlier methods, which in turn enables us to establish a novel and unexpected property of unfoldings.
We start by proving the existence of sinks in two-dimensional unfoldings.
6.1 Cascades of sinks
In this section we are going to prove that, for a two dimensional set of parameters, an unfolding contains maps which have a sink of high period. Choose $\epsilon_{0}>0$ and define, for $n$ large enough,
$\mathcal{A}_{n}=\left\{(t,a)\in[-t_{0},t_{0}]\times[-a_{0},a_{0}]\left|\right.|a-a_{n}(t)|\leq\frac{\epsilon_{0}}{\mu(t,a_{n}(t))^{2n}}\right\}.$
The strip $\mathcal{A}_{n}$ is built to contain the parameters having a sink. The size has to be chosen carefully. The following remark describes the error obtained in the change of the eigenvalues while changing the parameters within $\mathcal{A}_{n}$.
Remark 11.
Since $\partial\mu/\partial a$, $\partial\lambda/\partial t$ and $\partial\lambda/\partial a$ are all bounded we have that, for all $(t,\tilde{a})\in\mathcal{A}_{n}$,
$\left(\frac{\mu(t,a_{n}(t))}{\mu(t,\tilde{a})}\right)^{n}=1+O\left(\frac{n}{\mu(t,a_{n}(t))^{2n}}\right),\left(\frac{\lambda(t,a_{n}(t))}{\lambda(\tilde{t},\tilde{a})|}\right)^{n}=1+O\left(\frac{n}{\mu(t,a_{n}(t))^{2n}}\right).$
In the following we prove, for a properly chosen $\epsilon_{0}$, that for all parameters $(t,a)\in\mathcal{A}_{n}$, $F_{t,a}$ has a sink of period $n+N$ which we call primary sink. The method used appears already in [5, 38]. Namely, we find an invariant box in the phase space and we prove that the $F_{t,a}^{n+N}$ contracts this box. As a consequence, we get a sink.
For $n$ large enough, take $(t,a)\in\mathcal{A}_{n}$ and denote by $c_{n}(t)=c_{t,a_{n}(t)}$ and by $z_{n}(t)=z_{t,a_{n}(t)}=(z_{n,x}(t),z_{n,y}(t))$. When the choice of $t$ is clear, we just use the notation $c_{n}$ and $z_{n}$.
For $\delta>0$ we define the box that is going to contain the sink,
$B^{n}_{\delta}(t,a)=\left\{(x,y)||x-z_{n,x}(t)|\leq\frac{1}{3},|y-z_{n,y}(t)|\leq\frac{\delta}{\mu(t,a_{n}(t))^{2n}}\right\}.$
When the choice of $(t,a)$ is clear we just use the notation $B^{n}_{\delta}$.
In the next lemma we prove that $B^{n}_{\delta}$ returns into itself, see Figure 7. Let $\tilde{Q}=\max Q_{x,y}$ where $Q_{x,y}$ as in Lemma 5.
Figure 7: Invariant Region
Lemma 18.
Choose $\delta=\frac{1}{4\tilde{Q}},\epsilon_{0}=\frac{1}{32\tilde{Q}}$. Then, for $n$ large enough and $(t,a)\in\mathcal{A}_{n}$,
$F^{n+N}_{t,a}\left(B^{n}_{\delta}\right)\subset B^{n}_{\delta}.$
Proof.
[…]
$|\Delta x|\leq\frac{1}{3},|\Delta y|\leq\frac{\delta}{\mu(t,a_{n}(t))^{2n}}\text{ and }|\Delta a|\leq\frac{\epsilon_{0}}{\mu(t,a_{n}(t))^{2n}}.$ (6.1)
Denote by $\left(\Delta x^{\prime},\Delta y^{\prime}\right)=F^{n}_{t,a}\left(\Delta x,\Delta y\right)-c_{t,a}$. Using that $F^{n}_{t,a}$ is linear on $B^{n}_{\delta}$, see (2.4), we get
$|\Delta x^{\prime}|\leq 2\lambda(t,a)^{n},$
and
$|\Delta y^{\prime}|$ $\leq$ $\mu(t,a)^{n}\left|\Delta y\right|+\left|\mu(t,a)^{n}z_{n,y}-c(0,t,a)\right|$
$\leq$ $\mu(t,a)^{n}\left|\Delta y\right|+\left|\left(\left[1+O\left(\frac{n}{\mu(t,a_{n}(t))^{2n}}\right)\right]\mu(t,a_{n}(t))^{n}z_{n,y}\right)-c_{n,y}\right|$
$+$ $\left|c_{n,y}-c(0,t,a)\right|,$
where we used Remark 11. Since $F^{n}_{t,a_{n}(t)}\left(z_{n}\right)\in\Gamma_{t,a_{n}(t)}$, then
$\mu(t,a_{n}(t))^{n}z_{n,y}=c_{n,y}+O\left(\lambda(t,a_{n}(t))^{n}\right).$
Moreover $\left|c_{n,y}-c(0,t,a)\right|=O\left(\Delta a\right)$ (see Lemma 1). Hence, by (6.1) and by Remark 11 we get
$|\Delta y^{\prime}|$ $\leq$ $\mu(t,a)^{n}\left|\Delta y\right|+O\left(\lambda(t,a_{n}(t))^{n}\right)+O\left(\frac{n}{\mu(t,a_{n}(t))^{2n}}\right)+O\left(\frac{1}{\mu(t,a_{n}(t))^{2n}}\right)$
$\leq$ $\frac{\delta}{\mu(t,a_{n}(t))^{n}}\left[1+O\left(\frac{n}{\mu(t,a_{n}(t))^{2n}}\right)\right]+O\left(\frac{n}{\mu(t,a_{n}(t))^{2n}}\right).$
By Lemma 5 (center $F^{N}_{t,a}$ in coordinates around $c_{t,a}$) extended to include also the Taylor expansion in $\Delta a$ we get, for $n$ large enough
[…]
and
[…]
$\leq$ $\frac{\tilde{Q}\delta^{2}}{\mu(t,a_{n}(t))^{2n}}+O\left(\frac{1}{\mu(t,a_{n}(t))^{3n}}\right)+\frac{\epsilon_{0}}{\mu(t,a_{n}(t))^{2n}},$
where we also used $(F3)$ and Remark 11. By our choice of $\epsilon_{0}$ and $\delta$, for $n$ large enough, the lemma follows. ∎
We fix $\epsilon_{0},$ and $\delta$ such that Lemma 18 holds. We are now ready to prove the existence of a sink. This is achieved in the next proposition by proving that $F^{n+N}$ contracts the box $B_{\delta}^{n}$.
Proposition 5.
For $n$ large enough and for all $(t,a)\in\mathcal{A}_{n}$, $B^{n}_{\delta}(t)$ has a unique periodic point which is a sink.
Proof.
Because $F_{t,a}^{n}$ is linear on $B_{\delta}^{n}$, the image $F_{t,a}^{n}\left(B_{\delta}^{n}\right)$ is contained in a neighborhood of $c_{t,a}$ of diameter smaller than $\delta\mu^{-n}\left(1+O\left({n}\mu^{-n}\right)\right)$. From this and by differentiating (2.15) with respect to $\Delta x$ and $\Delta y$ we get
$DF_{t,a}^{N}$ $=$ $\left(\begin{matrix}O(1)&O(1)\\ O(1)&2{\delta\mu^{-n}Q_{x,y}}\left(1+O\left({n}{\mu^{-n}}\right)\right)\\ \end{matrix}\right).$
Note that $N$ is fixed. Using again that $F_{t,a}^{n}$ is linear on $B_{\delta}^{n}$ we obtain that
$DF_{t,a}^{n+N}$ $=$ $\left(\begin{matrix}O(1)&O(1)\\ O(1)&{2\delta}{\mu^{-n}}Q_{x,y}\left(1+O\left({n}{\mu^{-n}}\right)\right)\end{matrix}\right)\left(\begin{matrix}O\left(\lambda^{n}\right)&0\\ 0&\mu^{n}\end{matrix}\right)$
$=$ $\left(\begin{matrix}O\left(\lambda^{n}\right)&O\left(\mu^{n}\right)\\ O\left(\lambda^{n}\right)&2\delta Q_{x,y}\left(1+O\left({n}{\mu^{-n}}\right)\right)\end{matrix}\right).$
Let $D=\left(\begin{matrix}O\left(\lambda^{n}\right)&O\left(\mu^{n}\right)\\ O\left(\lambda^{n}\right)&\frac{3}{4}\end{matrix}\right)$ be a positive matrix and $\left(DF_{t,a}^{n+N}\right)^{k}(\Delta x,\Delta y)=(\Delta x_{k},\Delta y_{k})$, then by the choice of $\delta$
$\left(\begin{matrix}|\Delta x_{k+1}|\\ |\Delta y_{k+1}|\end{matrix}\right)\leq D\left(\begin{matrix}|\Delta x_{k}|\\ |\Delta y_{k}|\end{matrix}\right).$
Observe that $\text{tr}(D)=\frac{3}{4}+O\left(\lambda^{n}\right)$ and $\text{det}(D)=O(\left(\lambda\mu\right)^{n})$. As consequence, for $n$ large enough, $|\Delta x_{k}|,|\Delta y_{k}|\to 0$ exponentially fast. Hence, the periodic point in $B^{n}_{\delta}$ is a sink whose basin of attraction contains $B^{n}_{\delta}$. ∎
6.2 Newhouse phenomenon: infinitely many sinks
We are now ready to set up an induction procedure to prove the existence of maps having infinitely many sinks. This procedure uses the so-called “Newhouse boxes”. In the first generation, the Newhouse boxes are essentially rectangles in $\mathcal{A}_{n}$ whose boundary are defined by the curves of secondary tangencies, see Figure 6. The family restricted to the Newhouse boxes of first generation have one sink and it is an unfolding of a new homoclinic tangency. The propositions and the lemmas proved in the previous sections apply then to these families creating Newhouse boxes of second generation. As a consequence, the family restricted to the Newhouse boxes of second generation have two sinks and it is an unfolding of a new homoclinic tangency. We proceed by induction.
Let $(t_{n,n_{0}},a_{n,n_{0}})$ be the parameters at the intersection point $a_{n}\cap b_{n,n_{0}}$, see Proposition 4, and let
$f_{n,n_{0}}=F_{t_{n,n_{0}},a_{n,n_{0}}}.$
Recall that $b_{n,n_{0}}$ is the graph of a function $b=b_{n,n_{0}}:[-t_{0},t_{0}]\mapsto\mathbb{R}$. Consider its restriction to the interval $t^{-}_{n,n_{0}},t^{+}_{n,n_{0}}$] where $t^{-}_{n,n_{0}}$ and $t^{+}_{n,n_{0}}$ are the two intersection points of the curve $b$ with the boundary of the strip $\mathcal{A}_{n}$, see Figure 6.
The domains
$\mathcal{P}_{n,n_{0}}=\left\{(t,a)\left|\right.t\in[t^{-}_{n,n_{0}},t^{+}_{n,n_{0}}],|a-a_{n}(t)|\leq\frac{\epsilon_{0}}{\mu(t,a_{n}(t))^{2n}}\right\}$
are called the Newhouse boxes of first generation.
The construction in the previous sections started with a map $f:\mathcal{M}\to\mathcal{M}$ with a strong homoclinic tangency and an unfolding $F:\mathcal{P}\times\mathcal{M}\to\mathcal{M}$. The following inductive construction will repeat the discussion of the previous sections starting with the map $f_{n,n_{0}}:\mathcal{M}\to\mathcal{M}$ and an unfolding $F:\mathcal{P}_{n,n_{0}}\times\mathcal{M}\to\mathcal{M}$ which is the restriction of the original family.
Lemma 19.
The domains $\mathcal{P}_{n,n_{0}}$ are pairwise disjoint for $n$ large enough and the diameter goes to zero.
Proof.
By Proposition 4 the curve $b_{n,n_{0}}$ is the graph of the function $b$ and by the angle estimate in Lemma 13, there exists a uniform constant $K>0$ such that
$\frac{1}{K}\frac{1}{n\left(\lambda^{\theta}\mu^{2}\right)^{n}}\leq\left|\left\{t\left|\right.(t,b(t))\in\mathcal{A}_{n}\right\}\right|\leq K\frac{1}{n\left(\lambda^{\theta}\mu^{2}\right)^{n}}.$ (6.2)
Moreover the proof of Proposition 3 gives $\text{dist}\left(f_{n,n_{0}},f_{n,n_{0}+1}\right)$ is proportional to 1/n{1}/{n}. The disjointness follows from this estimates, the fact that $\lambda^{\theta}\mu^{2}\geq\left(\lambda^{2\theta}\mu^{3}\right)^{1/2}>1$, see (2.16) and the fact that $\mathcal{A}_{n}$ are pairwise disjoint. ∎
The next proposition ensures that the family restricted to $\mathcal{P}_{n,n_{0}}$ is again an unfolding. This allows an inductive procedure.
Proposition 6.
For $n$ large enough, the map $f_{n,n_{0}}$ has a strong homoclinic tangency and the restriction $F:\mathcal{P}_{n,n_{0}}\times\mathcal{M}\to\mathcal{M}$ can be reparametrized to become an unfolding. Moreover each map in $\mathcal{P}_{n,n_{0}}$ has a sink of period $n+N$.
Proof.
The proof is given in Lemma 17. ∎
Inductively we are going to construct parameters with multiple sinks of higher and higher periods and a strong homoclinic tangency. Let $\mathfrak{N}^{1}=\left\{(n,n_{0})\left|\right.n_{0}\in\mathfrak{N}_{n}\right\}\subset{\mathbb{N}}^{2}$ be the set of labels of the Newhouse boxes of first generation $\mathcal{P}_{n,n_{0}}$. As inductive hypothesis assume that there exist sets $\mathfrak{N}^{k}\in\left({\mathbb{N}}^{2}\right)^{k}$ such that, each $\mathfrak{N}^{k}$ is the set of labels of the Newhouse boxes of generation $k$. Moreover the natural projections $\mathfrak{N}^{1}\leftarrow\mathfrak{N}^{2}\leftarrow\dots\leftarrow\mathfrak{N}^{g}$ with $\mathfrak{N}^{k}\in\left({\mathbb{N}}^{2}\right)^{k}$ are countable to 11 and correspond to inclusion of Newhouse boxes of successive generations. They satisfy the following inductive hypothesis.
For $\underline{n}=\left\{(n^{(k)},n_{0}^{(k)})\right\}_{k=1}^{g}\in\mathfrak{N}^{g}$, there exist sets $\mathcal{P}^{k}_{n^{(k)},n_{0}^{(k)}}\subset\mathcal{P}$, where $n^{(k)}$ labels the sink and $n_{0}^{(k)}$ labels the secondary tangency, such that
  • $\text{diam}\left(\mathcal{P}^{k}_{n^{(k)},n_{0}^{(k)}}\right)\leq\frac{1}{k}$,
  • $\mathcal{P}^{k+1}_{n^{(k+1)},n_{0}^{(k+1)}}\subset\mathcal{P}^{k}_{n^{(k)},n_{0}^{(k)}}$ for $k=1,\dots,g-1$,
  • there exists a map $f^{k}_{n^{(k)},n_{0}^{(k)}}\in\mathcal{P}^{k}_{n^{(k)},n_{0}^{(k)}}$ which has a strong homoclinic tangency,
  • the restriction $F:{\mathcal{P}^{k}_{n^{(k)},n_{0}^{(k)}}}\times\mathcal{M}\to\mathcal{M}$ can be reparametrized to become and unfolding of $f^{k}_{n^{(k)},n_{0}^{(k)}}$,
  • every map in $\mathcal{P}^{k}_{n^{(k)},n_{0}^{(k)}}$ has at least $k$ sinks of different periods.
By induction, using Proposition 6, we get an infinite sequence of sets $\mathfrak{N}^{k}$.
Figure 8: Newhouse boxes
Definition 6.
The sets $\mathcal{P}^{g}_{n^{(g)},n_{0}^{(g)}}$ are called Newhouse boxes of generation $g$, see Figure 8, and
$NH=\bigcap_{g}\bigcup_{\underline{n}\in\mathfrak{N}^{g}}\mathcal{P}^{g}_{n^{(g)},n_{0}^{(g)}}.$
The set $NH$, consisting of parameters for which the corresponding maps have infinitely many sinks, accumulates on the curve of the original tangency. The inductive construction of these parameters implies that the set $NH$ accumulates on all tangency curves given by $b_{n,n_{0}}$. We have the following lemma.
Lemma 20.
$\overline{NH}\supset[-t_{0},t_{0}]\times\left\{0\right\}.$
In particular,
$\overline{NH}\cap\mathcal{P}_{n,n_{0}}\supset\text{\rm graph}(b_{n,n_{0}}).$ (6.3)
Proof.
Given $(t,0)$, by Proposition 3, for every $n$ large enough, $a_{n}$ has a secondary tangency at $(t_{n},a_{n}(t_{n}))$ with $|t-t_{n}|=O\left(\frac{1}{n}\right)$. Hence, there exists a sequence of Newhouse boxes in $\mathfrak{N_{1}}$ accumulating at $(t,0)$. By construction, each box in $\mathfrak{N_{1}}$ contains points of $NH$. The lemma follows. ∎
Given any family $F$ of diffeomorphisms, we define the Newhouse set $NH_{F}$ as the set of parameters having infinitely many sinks. The upper Minkowski dimension is denoted by $MD$.
Theorem 3.
Let $F:\mathcal{P}\times\mathcal{M}\to\mathcal{M}$ be an unfolding then
    $NH\subset NH_{F}$, every map in $NH$ has infinitely many sinks,
    NHNH is homeomorphic to $\mathbb{R}\setminus{\mathbb{Q}}$,
    $MD(NH)\geq\frac{2}{3}$.
Proof.
The inductive construction, using Proposition 6, implies that all maps in $NH$ have infinitely many sinks. From Lemma 19 we know that
$\bigcup_{\underline{n}\in\mathfrak{N}^{g}}\mathcal{P}^{g}_{n^{(g)},n_{0}^{(g)}},$
consists of countably many disjoint boxes. Each box $P^{g}_{n^{(g)},n_{0}^{(g)}}$ contains countably many pairwise disjoint boxes of the next generation. Hence, the nested intersection $NH$ is homeomorphic to $\mathbb{R}\setminus{\mathbb{Q}}$, see [22].
For the last property, let $(t^{*},0)$ such that ${\log{\lambda(t^{*},0)^{-1}}}/{\log\mu(t^{*},0)}$ is the maximum of ${\log{\lambda(t,0)^{-1}}}/{\log\mu(t,0)}$. Consider a sequence of first generation Newhouse boxes $\mathcal{P}_{n,n_{0}}\in\mathfrak{N_{1}}$ accumulating at $(t^{*},0)$. This is possible because of Lemma 20. Choose $\epsilon>0$ and let $n$ be maximal such that $\epsilon\leq{\epsilon_{0}}/{\mu(t^{*},a^{*})^{2n}}$. Because (6.2), (6.3) and the fact that the vertical size of $\mathcal{P}_{n,n_{0}}$ is ${\epsilon_{0}}/{\mu(t^{*},a^{*})^{2n}}$, we need at least ${K}/{\lambda(t^{*},a^{*})^{\theta n}}$ balls of radius $\epsilon$ to cover $\overline{NH}\cap\mathcal{P}_{n,n_{0}}$. As consequence
$MD(NH)\geq\frac{\theta}{2}\max_{NH}\left(\frac{\log\frac{1}{\lambda}}{\log\mu}\right),$
and $MD(NH)\geq{2}/{3}$, where we used (2.17). ∎
Remark 12.
Observe that the estimate for the upper Minkowski dimension is not sharp. Other dimension estimates for maps with infinitely many sinks were obtained in [10, 40, 45].
7 Dynamical Partitions, Directed Graphs and Strips
In this section, we introduce the objects and the induction step (see Proposition 7) that will allow to build later the main player of this work: the lamination and the invariant Cantor set.
7.1 Dynamical Partitions
We give now the definition of a dynamical partition associated to a map. The reader can refer to the Figure 9. The definition models the combinatorics as described in Subsection 4.1.
Figure 9: $(1,N,M)$ dynamical partition of $f$
Recall the rescaling done in Subsection 2.1.1 while introducing our families of unfoldings and let $Q$ be the domain of linearization.
We denote by $W^{u}_{\text{loc}}(p)$ a connected component of $W^{u}(p)\cap Q$ and by $W^{s}_{\text{loc}}(p)$ a connected component of $W^{s}(p)\cap Q$.
Definition 7.
Let $\mathcal{M}$ be a $m$-dimensional manifold. A map $f:\mathcal{M}\to\mathcal{M}$ has a $(1,N,M)$ dynamical partition
$\mathfrak{D}=\left\{Q,C^{0},C^{1},\dots,C^{N-1},H^{0},H^{1},\dots,H^{M-1}\right\},$
where $N$ and $M$ are positive natural numbers, $Q,C^{i},Hj$ are open sets for all $i\in\left\{0,\dots,N-1\right\}$ and $j\in\left\{0,\dots,M-1\right\}$ if the following holds.
  1. 1.
    The elements of $\mathfrak{D}$ are pairwise disjoints.
  2. 2.
    The map $f$ has a saddle point $p\in Q$.
  3. 3.
    The $C$-sets satisfy the followings:
    • $C^{0}\subset f(Q)\setminus Q$,
    • $f:C^{k}\to C^{k+1}$, for all $k\in\left\{0,1,\dots,N-2\right\}$ is a diffeomorphism,
    • $f:C^{N-1}\to f\left(C^{N-1}\right)=C^{N}$ is a diffeomorphism,
    • $f\left(C^{N-1}\right)\subset Q\setminus f(Q)$,
    • $C^{0}\cap W^{u}_{\text{loc}}(p)\neq\emptyset$, where $W^{u}_{\text{loc}}(p)$ is a connected component of $W^{u}(p)\cap Q$,
    • $f\left(C^{N-1}\right)\cap W^{s}_{\text{loc}}(p)\neq\emptyset$.
  4. 4.
    The $H$-sets satisfy the followings:
    • $H^{0}\subset f(Q)\setminus Q$,
    • $f:H^{k}\to H^{k+1}$, for all $k\in\left\{0,1,\dots,M-2\right\}$ is a diffeomorphism,
    • $f:H^{M-1}\to f\left(H^{M-1}\right)=H^{M}$ is a diffeomorphism,
    • $f\left(H^{M-1}\right)\subset Q\setminus f(Q)$,
    • $H^{0}\cap W^{u}_{\text{loc}}(p)\neq\emptyset$, where $W^{u}_{\text{loc}}(p)$ is a connected component of $W^{u}(p)\cap Q$,
    • $f\left(H^{M-1}\right)\cap W^{s}_{\text{loc}}(p)\neq\emptyset$,
    • $f^{M}\left(H_{0}\cap W^{u}_{\text{loc}}(p)\right)$ intersects transversally $W^{u}_{\text{loc}}(p)$ in the point $q_{2}$.
Definition 8.
Let $\mathcal{M}$ be a $m$-dimensional manifold, $f:\mathcal{M}\to\mathcal{M}$ a map and let $\mathfrak{D}$ be the $(1,N,M)$ dynamical partition of $f$. We define the mesh of $\mathfrak{D}$ as
$\text{mesh}\left(\mathfrak{D}\right)=\max_{U\in\mathfrak{D}}\left(\text{diam(U)}\right),$
where diam(U) is the diameter of $U$.
Moreover we denote by $\mathcal{D}$ the union of elements in $\mathfrak{D}$. Namely,
$\mathcal{D}=\bigcup_{U\in\mathfrak{D}}U.$
7.2 Directed Graphs
Fixed two positive natural numbers, $N$ and $M$, we define the $(1,N,M)$ directed graph. The reader can refer to Figure 10. As before this models the combinatorics as described in Subsection 4.1.
Definition 9.
Given two positive natural numbers, $N$ and $M$, a directed graph $X$ is called the $(1,N,M)$ directed graph if it has $1+N+M$ vertices, $q,c^{0},\dots,c^{N-1},h^{0},\dots,h{M-1}$ and unique edges:
$\overrightarrow{qq}$
$\overrightarrow{qh^{0}}$ $\overrightarrow{h^{k}h^{k+1}}\text{ for all }k\in\left\{1,\dots,M-2\right\}$ $\overrightarrow{h^{M-1}q}$
$\overrightarrow{qc^{0}}$ $\overrightarrow{c^{k}c^{k+1}}\text{ for all }k\in\left\{1,\dots,N-2\right\}$ $\overrightarrow{c^{N-1}q}$
The sub graph formed by $q$ is called the $q$-loop and is denote by $\lambda^{q}$.
The sub graph formed by $q,h^{0},\dots,h^{M-1}$ is called the $h$-loop and is denote by $\lambda^{h}$.
The sub graph formed by $q,c0,\dots,c^{M-1}$ is called the $c$-loop and is denote by $\lambda^{c}$.
Remark 13.
Observe that $q$ has three ingoing and outgoing edges. All other vertices have exactly one ingoing and one outgoing edge.
Let $X$ be the $(1,N,M)$ directed graph. A loop $\lambda$ is a path in $X$ which starts and ends in $q$. The loops $\lambda^{q}$, $\lambda^{h}$ and $\lambda^{c}$ are called the prime loops of $X$, respectively $q$-loop, $h$-loop and $c$-loop.
Observe that a loop $\lambda$ is a concatenation of prime loops $\gamma_{i}$, namely
$\lambda=\gamma_{u}\ast\cdots\ast\gamma_{1}.$
The winding of $\lambda$ is the word
$\omega(\lambda)\in\left\{q,h,c\right\}^{n},$
such that $\gamma_{i}=\lambda_{\omega(\lambda)(i)}$ for all $i\in\left\{1,\dots,u\right\}$.
Figure 10: $(1,N,M)$ directed graph
Graph Morphisms.
Given $N,M,\tilde{N},\tilde{M}$ positive natural numbers, consider $the (1,N~,M~)(1,\tilde{N},\tilde{M}) directed graph X~\tilde{X}$ and the graph morphism
$w:\tilde{X}\mapsto X.$
Observe that, because $q$ is the only vertex in $X$ which has an edge from itself to itself then $w(\tilde{q})=q$ and that the image of every loop in $\tilde{X}$ is a loop in $X$. In particular, the graph morphism $w$ is completely determined by the windings of the primary loops,
[…]
Projections.
In the following lemma we introduce the projection of $D$ on $X$.
Lemma 21.
Let $f$ be a map. Consider $\mathfrak{D}$ the $(1,N,M)$ dynamical partition of $f$ and $X$ the $(1,N,M)$ directed graph $X$. There exists a map $\pi:\mathcal{D}\mapsto X$ satisfying the following properties:
  1. 1.
    for all vertices $v\in X$, $\pi^{-1}(v)\in\mathfrak{D}$,
  2. 2.
    if $x\in\mathcal{D}$ and $f(x)\in\mathcal{D}$ $\implies$ there exists an edge $\overrightarrow{\pi(x)\pi\left(f(x)\right)}$,
  3. 3.
    if there is an edge $\overrightarrow{v_{1}v_{2}}$ $\implies$ there exists $x\in\pi^{-1}(v_{1})$ such that $f(x)\in\pi^{-1}(v_{2})$.
The map $\pi$ is called the projection of $\mathcal{D}$ on $X$.
7.3 Strips
We introduce strips in the plane that will enable us to later construct, inductively, the object of primary interest to us: the lamination. Observe that the strip are an abstract generalization of the $\mathcal{B}_{n}$ strips built earlier to contain the curves of secondary tangencies.
Definition 10.
Let $r>0$, $r\in\mathbb{N}$ and let $F$ be an unfolding. An $r$-tangency strip $S$ is an open region in $t_{0},t_{0}]\times[-a_{0},a_{0}]$ satisfying the following:
  • $S$ is bounded by the graph of two functions $s_{l},s_{u}:[-t_{0},t_{0}]\to\mathbb{R}$ such that, for all $t\in[-t_{0},t_{0}]$,
    $\left|s_{l}(t)-s_{u}(t)\right|\leq C\rho^{r(1+\alpha)-1},$
    where $0<\rho<1$, $\rho<\max_{t}\left\{\frac{1}{\mu(t)}\right\}$, $0<\alpha<1$ and $C$ is a positive constant.
  • $S$ contains $r$ global curves $\left\{b_{k}\right\}_{k=1}^{r}$, whcih are graph of functions $b_{k}:[-t_{0},t_{0}]\mapsto\mathbb{R}$ for all $k$, such that the following holds:
    • (i)
      $\forall$ $k\in\left\{1,\dots,r\right\}$ and for all $t\in[-t_{0},t_{0}]$, $F_{t,b_{k}(t)}$ is a map with a strong homoclinic tangency,
    • (ii)
      $\forall$ $k\in\left\{1,\dots,r\right\}$ there exists a neighborhood of $b_{k}$, $U_{k}$, such that $F_{|U_{k}}$ is an unfolding of $F_{t,b_{k}(t)}$,
    • (iii)
      $\forall$ $k,k^{\prime}\in\left\{1,\dots,r\right\}$, $U_{k}\cap U_{k^{\prime}}=\emptyset$,
    • (iv)
      $\forall$ $k\in\left\{1,\dots,r\right\}$ there exist positive numbers $N_{k},M_{k}$ such that, for all $f\in U_{k}$, the map $f$ has a $(1,N_{k},M_{k})$ dynamical partition $\mathfrak{D}_{k}(f)$ which depends continuously on the parameters.
For all $k\in\left\{1,\dots,r\right\}$, the pair $\left\{b_{k},U_{k}\right\}$ is called a $(1,N_{k},M_{k})$ descendant of $S$.
We start by defining $r$-tangency strips, with $r$ a positive real number; see Figure 11.
Figure 11: A $r$-tangency strip
7.4 The induction step
In this subsection we present the induction step that will allow later to build the lamination and the invariant Cantor set.
Proposition 7.
Let $S$ be a $r$-tangency strip and let $\left\{b,U\right\}$ be a $(1,N,M)$ descendant of $S$. Then there exist $n\in\mathbb{N}$, $n>>r$ large enough and a countable sequence of regions $\left\{S_{j}\right\}_{j=1}^{\infty}\subset U\subset S$ satisfying the following:
  • (i)
    $\forall j$, $S_{j}$ is a $(n+j)$-tangency strip,
  • (ii)
    $\forall j,j^{\prime}$ with $j\neq j^{\prime}$, $S_{j}\cap S_{j}^{\prime}=\emptyset$.
Moreover, for all $j$ and for all $k\in\left\{1,\dots,n+j\right\}$ let $\left\{b_{k}^{(j)},U_{k}^{(j)}\right\}$ be a $\left(1,N_{k}^{(j)},M_{k}^{(j)}\right)$ descendant of $S_{j}$. Then the following holds:
  • (iii)
    $\exists C>0$ such that $\forall t\in\left[-t_{0},t_{0}\right]$,
    $\frac{1}{C}\frac{1}{\mu(t)^{n+j}}\leq\left|b_{k}^{(j)}(t)-b(t)\right|\leq\frac{C}{\mu(t)^{n+j}},$
  • (iv)
    $\exists\tilde{C}>0$ such that
    $\frac{1}{\tilde{C}}\frac{n+j}{\mu(t)^{n+1+j}}\frac{\partial\mu}{\partial t}\leq\left|\frac{db_{k}^{(j)}}{dt}-\frac{db}{dt}\right|\leq\tilde{C}\frac{n+j}{\mu(t)^{n+1+j}}\frac{\partial\mu}{\partial t},$ (7.1)
  • (v)
    $\forall f\in U_{k}^{(j)}\subset U$, let $\mathfrak{D}^{(j)}_{k}(f)$ be the $\left(1,N_{k}^{j},M_{k}^{j}\right)$ dynamical partition of $f$ and let $\mathfrak{D}(f)$ be the $\left(1,N,M\right)$ dynamical partition of $f$. Then $\mathfrak{D}^{(j)}_{k}(f)$ is a refinement of $\mathfrak{D}(f)$ and
    $\text{mesh}\left(\mathfrak{D}^{(j)}_{k}(f)\right)<\left(\frac{1}{k}\right)\text{mesh}\left(\mathfrak{D}(f)\right).$
Furthermore, if $X$ is the $\left(1,N,M\right)$ directed graph, $X_{k}^{j}$ is the $\left(1,N_{k}^{j},M_{k}^{j}\right)$ directed graph, $i:\mathcal{D}_{k}^{(j)}\mapsto\mathcal{D}$ is the inclusion, $\pi:\mathcal{D}\mapsto X$ is the projection of $\mathcal{D}$ on $X$ and $\pi_{k}^{j}:\mathcal{D}_{k}^{(j)}\mapsto X_{k}^{(j)}$ is the projection of $\mathcal{D}_{k}^{(j)}$ on $X_{k}^{(j)}$, there exists a unique graph morphism $w:X_{k}^{(j)}\mapsto X$ such that
$\pi\circ i_{k}^{(j)}=w\circ\pi_{k}^{(j)}.$
Proof.
Without loss of generality, we can assume that the $r$-strip $S$ is delimitated by straight lines. In particular we can make a reparametrization to ensure that the curve $b$ corresponds to a straight line in the $a$-direction.. For all $j$, consider the strips $\mathcal{B}_{n+j}$ as defined in (4.1) and take $S_{j}=\mathcal{B}_{n+j}$. Observe that for $n$ large enough $S_{j}\subset U\subset S$ and by construction $S_{j}\cap S_{j}^{\prime}=\emptyset$ for all $j\neq j^{\prime}$. Apply Section $4$ and Section $5$ to ensure that $\mathcal{B}_{n+j}$ is indeed a $n+j$- strip.
The descendants are the pairs $\left\{b_{k}^{(j)},U_{k}^{(j)}\right\}$ with $k\in\left\{1,\dots,n\right\}$ where $b_{k}^{(j)}$ is the curve of secondary tangencies as constructed in Proposition 4 and $U_{k}^{(j)}$ is a neighborhood of $b_{k}^{(j)}$ as built in Lemma 17. In particular the family $F_{|U_{k}^{(j)}}$ is an unfolding of the homoclinic tangency $b_{k}^{(j)}$.
Given a descendent of $S_{j}$, $\left\{b_{k}^{(j)},U_{k}^{(j)}\right\}$. Observe that $a_{n+j},b_{k}^{(j)}\in S_{j}=\mathcal{B}_{n+j}$ and the diameter of $\mathcal{B}_{n+j}$ is smaller than the distance between $a_{n}$ and the previous tangency $b$ that is indeed of then order of $1/\mu^{n+j}$, see 2.14. This proves $(iii)$. Condition $(iv)$ on the derivative comes from 5.4 and 5.6. Condition $(v)$ and the following statement are consequence of the construction. ∎
Remark 14.
Observe that, using the notation as in the previous proposition and by using the construction in Subsection 4.1, there exist $n_{k}^{(j)},\underline{\ell}_{k}^{(j)},\overline{\ell}_{k}^{(j)},\theta_{k}^{(j)},n_{0,k}^{(j)}$ such that
$\underline{w}\left(\omega_{k}^{(j)}\right)=\left(q,{q}^{\underline{\ell}_{k}^{(j)}}{c}{q}^{\theta_{k}^{(j)}n_{k}^{(j)}}{c}{q}^{n_{k}^{(j)}}{c}{q}^{n_{0,k}^{(j)}}{h}{q}^{\overline{\ell}_{k}^{(j)}},{q}^{\underline{\ell}_{k}^{(j)}}{h}{q}^{\overline{\ell}_{k}^{(j)}}\right).$
8 Finite Collet-Eckmann Condition
In this section we set up the induction step needed to prove later the existence of maps with a Collet-Eckmann point. More in details we prove the existence of maps having a point satisfying the finite Collet-Eckmann condition. Roughly speaking a point $x$ is a $T$-Collet-Eckmann (T-CE) point if there are vectors $v$ in the tangent space at $x$ such that the derivative of the map along the orbit of $x$ in the direction of $v$ grows at the least exponentially fast up to time $T$. Below is the formal definition.
Definition 11.
Let $f:\mathcal{M}\to\mathcal{M}$ be a map and let $x\in\mathcal{M}$. Given constants $\varphi>0$, $0<C<1$, $\rho>1$, we say that $x$ is a $T$-Collet-Eckmann (T-CE) point if the following holds. Denote by $K_{\varphi}(x)$ the cone at $x$ of angle $\varphi$}, and by $\text{T}_{x}$ the tangent space at $x$, then for all vectors $v\in K_{\varphi}\subset\text{T}_{x}$,
$\left|Df^{s}_{x}v\right|\geq C\rho^{s}\left|v\right|\text{, \hskip 13.942pt }\forall\text{ }0\leq s\leq T.$
From now on we are going to assume an extra condition on $\theta$, namely we ask that,
$\lambda^{\theta}\mu^{1+\alpha}>\lambda^{\alpha},$ (8.1)
where $\alpha$ is defined in (4.17). By (2.17) and (4.17), there exists $\theta\in(\theta_{0},\theta_{1})$ satisfying (8.1), see Appendix.
Take constants $\varphi>0$, $0<C<1$ and $\rho>1$ such that,
$\text{ (o) }\cos\varphi>\frac{\rho}{\mu}$
$\text{ (oo) }1<\rho<\lambda^{\theta}\mu^{2}<\mu.$
We introduce now some notation. Let $x$ be a point. If $\Gamma$ is a curve, we denote by $\text{d}_{\text{v}}\left(x,\Gamma\right)$ the vertical distance between a point $x$ and the curve $\Gamma$. The angle between two vectors is denoted by $\angle$.
Let $S$ be a tangency strip and let $\left\{b,U\right\}$ be a $(1,N,M)$ descendant of $S$. For all $f\in U$, let $\mathfrak{D}(f)$ be the $\left(1,N,M\right)$ dynamical partition of $f$. Given $T\in\left\{1,\dots,N-1\right\}$, one can define a vector field $\mathfrak{X}\left(f^{T}\right)$ on $C^{N-T}$ as
$Df^{T}(x)v(x)=e_{1}\text{ for all }x\in C^{N-T}.$
Observe that, since $f^{T}:C^{N-T}\to C^{N}$ is a diffeomorphism, then $\mathfrak{X}\left(f^{T}\right)$ is well defined.
We are now ready to present the induction step needed to prove the existence of laminations of maps with a Collet–Eckmann point. We use the notation introduced above.
Proposition 8.
Let $S$ be a tangency strip and let $\left\{b,U\right\}$ be a $(1,N,M)$ descendant of $S$. Suppose that for all $f\in U$, there exist a positive constant $K$ and $T\in\left\{1,\dots,N-1\right\}$ such that the following holds,
  • $\text{IH}_{Q}:$
    $f_{|Q}$ is a linear map, i.e.
    $f_{|Q}=\left(\begin{matrix}\lambda&0\\ 0&\mu\end{matrix}\right).$
  • $\text{IH}_{H^{0}}:$
    The family $\left\{f^{i}_{|H^{0}}\right\}_{i=0}^{M}$ is a compact ${\mathcal{C}}^{2}$}-family.
  • $\text{IH}_{C^{0}}:$
    • (i)
      There exists a curve $\Gamma$ in $C^{0}$ which is a graph over the $x$-axis and
      $Df_{x}^{N-T}e_{2}\in\text{span}\left(\mathfrak{X}\left(f^{N-T}\right)\right)\iff x\in\Gamma.$
    • (ii)
      For all $x\in C^{0}$,
      $\frac{1}{K}{{d}_{\text{v}}}\left(x,\Gamma\right)\leq\angle\left(Df_{x}^{N-T}e_{2},\mathfrak{X}\left(f^{N-T}\right)\right)\leq Kd_{\text{v}}\left(x,\Gamma\right).$
    • (iii)
      The family $\left\{f^{i}_{|C^{0}}\right\}_{i=0}^{N-T}$ is a compact ${\mathcal{C}}^{2}$}-family.
  • $\text{IH}_{C^{N-T}}:$
    • (i)
      For all $x\in C^{N-T}$ and for all $v\in K_{\varphi}(x)\subset\text{T}_{x}$,
      $\left|Df_{x}^{s}v\right|\geq C\rho^{s}\left|v\right|,\text{ }\forall\text{ }0\leq s\leq T.$
    • (ii)
      For all $x\in C^{N-T}$ and for all $v\in K_{\varphi}(x)\subset\text{T}_{x}$,
      $Df_{x}^{T}v\in K_{\varphi}\left(f^{T}(x)\right)\subset\text{T}_{f^{T}(x)}.$
    • (iii)
      The family $\left\{f^{i}_{|C^{N-T}}\right\}_{i=0}^{T}$ is a compact ${\mathcal{C}}^{2}$}-family.
[…]
  • $\text{IH}_{\tilde{Q}}$:
    $f_{|\tilde{Q}}$ is a linear map, i.e.
    $f_{|\tilde{Q}}=\left(\begin{matrix}\lambda&0\\ 0&\mu\end{matrix}\right).$
  • $\text{IH}_{\tilde{H}^{0}}:$
    The family $\left\{f^{i}_{|\tilde{H}^{0}}\right\}_{i=0}^{\tilde{M}}$ is a compact ${\mathcal{C}}^{2}$}-family.
  • $\text{IH}_{\tilde{C}^{0}}:$
    • (i)
      There exists a curve $\tilde{\Gamma}$ in $\tilde{C}^{0}$ which is a graph over the $x$-axis and
      $Df^{\tilde{N}-\tilde{T}}(x)e_{2}\in\text{span}\left({\mathfrak{X}}\left(f^{\tilde{N}-\tilde{T}}\right)\right)\iff x\in\tilde{\Gamma}$.
    • (ii)
      For all $x\in\tilde{C}^{0}$,
[…]
    • (iii)
      The family $\left\{f^{i}_{|\tilde{C}^{0}}\right\}_{i=0}^{\tilde{N}-\tilde{T}}$ is a compact ${\mathcal{C}}^{2}$}-family.
  • $\text{IH}_{\tilde{C}^{\tilde{N}-\tilde{T}}}$:
    • (i)
      For all $x\in\tilde{C}^{\tilde{N}-\tilde{T}}$ and for all $v\in K_{\varphi}(x)\subset\text{T}_{x}$,
      $\left|Df_{x}^{s}v\right|\geq C\rho^{s}\left|v\right|\text{ }\forall\text{ }0\leq s\leq\tilde{T}$.
    • (ii)
      For all $x\in\tilde{C}^{\tilde{N}-\tilde{T}}$ and for all $v\in K_{\varphi}(x)\subset\text{T}_{x}$,
      $Df_{x}^{\tilde{T}}v\in K_{\varphi}\left(f^{\tilde{T}}(x)\right)\subset\text{T}_{f^{\tilde{T}}(x)}.$
    • (iii)
      The family $\left\{f^{i}_{|\tilde{C}^{\tilde{N}-\tilde{T}}}\right\}_{i=0}^{\tilde{T}}$ is a compact ${\mathcal{C}}^{2}$}-family.
Proof.
[…]
$\tilde{T}=T+n+N-T+T+n_{0}+M+\overline{\ell},$
and for all $\tilde{f}\in\tilde{U}$ let $\tilde{K}$ be a positive constant assuring that $\text{IH}_{\tilde{H}^{0}}$, $\text{IH}_{\tilde{C}^{0}}$-(iii) and $\text{IH}_{\tilde{C}^{\tilde{N}-\tilde{T}}}-$(iii) hold. The existence of such a constant is guaranteed by the initial conditions on the family. Observe that $\text{IH}_{\tilde{Q}}$ is trivially satisfies.
Define $\tilde{\Gamma}=f^{-\overline{\ell}}\left(\Gamma\right)$, then $\text{IH}_{\tilde{C}^{0}}$-(i) and $\text{IH}_{\tilde{C}^{0}}$-(ii) are satisfied by construction.
The rest of the proof is devoted to prove $\text{IH}_{\tilde{C}^{\tilde{N}-\tilde{T}}}-$(i) and $\text{IH}_{\tilde{C}^{\tilde{N}-\tilde{T}}}-$(ii). This will be achieved by steps. Each step corresponds to a time defining $\tilde{T}$. In order to lighten the notation, we denote $\tilde{C}=\tilde{C}^{\tilde{N}-\tilde{T}}$. Let $x\in\tilde{C}$ and let $v\in K_{\varphi}(x)\subset\text{T}_{x}$. Observe that $\tilde{C}\subset C^{N-T}$.
  • Step 1:
    We begin by proving $\text{IH}_{\tilde{C}^{\tilde{N}-\tilde{T}}}-$(i) from time 0 up to time $T$. By $\text{IH}_{C^{N-T}}$-(i),
    • $\text{IH}_{\tilde{C}}-$(i)P1:
      $\left|Df_{x}^{s}v\right|\geq C\rho^{s}\left|v\right|\text{ }\forall\text{ }0\leq s\leq T.$
    Moreover, by $\text{IH}_{C^{N-T}}$-(ii)
    $Df_{x}^{T}v\in K_{\varphi}\left(f^{T}(x)\right)\subset\text{T}_{f^{T}(x)}.$
    In particular, if we denote by
    $v^{(1)}=Df^{T}_{x}v=\left(\begin{matrix}v^{(1)}_{1}\\ v^{(1)}_{2}\end{matrix}\right)$,
    it follows that,
    $v^{(1)}_{2}>\cos\varphi\left|v^{(1)}\right|\text{ and }\frac{v^{(1)}_{1}}{v^{(1)}_{2}}<\tan\varphi.$ (8.3)
  • Step 2:
    We begin by proving $\text{IH}_{\tilde{C}^{\tilde{N}-\tilde{T}}}-$(i) from time $T+1$ up to time $T+n$. Indeed, by $\text{IH}_{Q}$ and $\text{IH}_{\tilde{C}}-$(i)P1, for all $1\leq k\leq n$, we get,
    • $\text{IH}_{\tilde{C}}-$(i)P2:
      $\left|Df^{T+k}_{x}v\right|$ $=$ $\left|\left(\begin{matrix}\lambda^{k}&0\\ 0&\mu^{k}\end{matrix}\right)\left(\begin{matrix}v^{(1)}_{1}\\ v^{(1)}_{2}\end{matrix}\right)\right|>\mu^{k}v^{(1)}_{2}\geq\mu^{k}\cos\varphi\left|v^{(1)}\right|$
      $=$ $\mu^{k}\cos\varphi\left|Df^{T}_{x}v\right|\geq\mu^{k}\cos\varphi C\rho^{T}\left|v\right|\geq C\rho^{T+k}\left|v\right|.$
    Observe that in the last inequality of $\text{IH}_{\tilde{C}}-$(i)P2 we used that
    $\mu^{k}\cos\varphi\geq\rho^{k},\text{ }\forall 1\leq k\leq n.$
    This follow from the condition imposed in (8)-(o). In particular, $\text{IH}_{\tilde{C}}-$(i)P2 proves
    $\left|Df^{T+n}_{x}v\right|\geq\mu^{n}\cos\varphi C\rho^{T}\left|v\right|.$ (8.4)
    Denote by
    $v^{(2)}=Df^{T+n}_{x}v=\left(\begin{matrix}v^{(2)}_{1}\\ v^{(2)}_{2}\end{matrix}\right)$.
    By $\text{IH}_{Q}$ and by (8.3),
    $\frac{v^{(2)}_{1}}{v^{(2)}_{2}}=\left(\frac{\lambda}{\mu}\right)^{n}\frac{v^{(1)}_{1}}{v^{(1)}_{2}}<\left(\frac{\lambda}{\mu}\right)^{n}\tan\varphi=\left(\frac{\lambda}{\mu}\right)^{n}\varphi.$ (8.5)
  • Step 3:
    We start proving $\text{IH}_{\tilde{C}^{\tilde{N}-\tilde{T}}}-$(i) from time $T+n+1$ up to time T+n+NTT+n+N-T. Indeed, by $\text{IH}_{C^{0}}$-(iii) and (8.4), for all $1\leq k\leq N-T$, we get,
    • $\text{IH}_{\tilde{C}}-$(i)P3:
      $\left|Df^{T+n+k}_{x}v\right|\geq\frac{1}{K}\left|Df^{T+n}_{x}v\right|>\frac{1}{K}\mu^{n}\cos\varphi C\rho^{T}\left|v\right|\geq C\rho^{T+n+k}\left|v\right|.$
    Observe that in the last inequality of $\text{IH}_{\tilde{C}}-$(i)P3 we used that
    $\frac{1}{K}\left(\frac{\mu}{\rho}\right)^{n}\cos\varphi\geq\rho^{k},\text{ }\forall 1\leq k\leq N-T$
    which is equivalent to ask
    $\frac{1}{K}\left(\frac{\mu}{\rho}\right)^{n}\cos\varphi\geq\rho^{N-T}.$
    This follow from the condition imposed in (8)-(oo) and taking $n$ large enough. In particular, $\text{IH}_{\tilde{C}}-$(i)P3 proves
    $\left|Df^{T+n+N-T}_{x}v\right|\geq\frac{C}{K}cos\varphi\mu^{n}\ \rho^{T}\left|v\right|.$ (8.6)
    Moreover by $\text{IH}_{C^{0}}$-(ii) and Lemma 7,
    $\angle\left(Df_{f^{T+n}(x)}^{N-T}e_{2},\mathfrak{X}\left(f^{T+n+N-T}\right)\right)\geq\frac{1}{K}{\text{d}_{\text{v}}}\left(f^{T+n}(x),\Gamma\right)\geq\frac{1}{K^{2}}\left(\lambda^{\theta}\mu\right)^{n},$
    and by (8.5),
    $\tan\left(v^{(2)},e_{2}\right)\leq\left(\frac{\lambda}{\mu}\right)^{n}\tan\varphi.$
    The last two inequalities together with $\text{IH}_{C^{0}}$-(iii), gives that, for $n$ large enough,
    $\angle\left(Df_{x}^{T+n+N-T}v,\mathfrak{X}\left(f^{T+n+N-T}\right)\right)\geq\frac{1}{2K^{2}}\left(\lambda^{\theta}\mu\right)^{n}.$ (8.7)
  • Step 4:
    We start proving $\text{IH}_{\tilde{C}^{\tilde{N}-\tilde{T}}}-$(i) from time T+nNT+1T+nN-T+1 up to time T+n+NT+TT+n+N-T+T. Indeed, by $\text{IH}_{C^{N-T}}$-(iii) and (8.6), for all $1\leq k\leq T$, we get,
    • $\text{IH}_{\tilde{C}}-$(i)P4:
      $\left|Df^{T+n+N-T+k}_{x}v\right|\geq\frac{1}{K}\left|Df^{T+n+N-T}_{x}v\right|\geq\frac{C}{K^{2}}\cos\varphi\mu^{n}\rho^{T}\left|v\right|\geq C\rho^{T+n+N-T+k}\left|v\right|.$
    Observe that in the last inequality of $\text{IH}_{\tilde{C}}-$(i)P4 we used that
    $\frac{1}{K^{2}}\cos\varphi\left(\frac{\mu}{\rho}\right)^{n}\rho^{T}\geq\rho^{N+k},\text{ }\forall 1\leq k\leq T$
    which is equivalent to ask
    $\frac{1}{K^{2}}\cos\varphi\left(\frac{\mu}{\rho}\right)^{n}\geq\rho^{N}.$
    This follow from the condition imposed in (8)-(oo) and taking $n$ large enough. In particular, $\text{IH}_{\tilde{C}}-$(i)P4 proves
    $\left|Df^{T+n+N}_{x}v\right|\geq\frac{C}{K^{2}}cos\varphi\mu^{n}\ \rho^{T}\left|v\right|.$ (8.8)
    Equation (8.7), $\text{IH}_{C}$-(iii) and the definition of the vector field $\mathfrak{X}$ give that, for $n$ large enough,
    $\angle\left(Df_{x}^{T+n+N}v,e_{1}\right)\geq\frac{1}{2K^{3}}\left(\lambda^{\theta}\mu\right)^{n}.$ (8.9)
    In particular, if we denote by
    $v^{(4)}=Df^{T+n+N}_{x}v=\left(\begin{matrix}v^{(4)}_{1}\\ v^{(4)}_{2}\end{matrix}\right)$,
    then, by (8.9),
    $\frac{v^{(4)}_{2}}{\left|v^{(4)}\right|}\geq\sin\frac{1}{2K^{3}}\left(\lambda^{\theta}\mu\right)^{n}\geq\frac{1}{4K^{3}}\left(\lambda^{\theta}\mu\right)^{n}$ (8.10)
  • Step 5:
    We start proving $\text{IH}_{\tilde{C}^{\tilde{N}-\tilde{T}}}-$(i) from time $T+nN+1$ up to time $T+n+N+n_{0}$. Indeed, by $\text{IH}_{Q}$, (8.10) and (8.8), for all $1\leq k\leq n_{0}$, we get,
    • $\text{IH}_{\tilde{C}}-$(i)P5:
      $\left|Df^{T+n+N+k}_{x}v\right|$ $\geq$ $\mu^{k}v^{(4)}_{2}\geq\mu^{k}\frac{1}{4K^{3}}\left(\lambda^{\theta}\mu\right)^{n}\left|Df^{T+n+N}_{x}v\right|$
      $\geq$ $\frac{C}{4K^{5}}\cos\varphi\mu^{k+n}\rho^{T}\left(\lambda^{\theta}\mu\right)^{n}\left|v\right|\geq C\rho^{T+n+N+k}\left|v\right|.$
    Observe that in the last inequality of $\text{IH}_{\tilde{C}}-$(i)P5 we used that
    $\frac{1}{4K^{5}}\cos\varphi\left(\frac{\mu}{\rho}\right)^{n+k}\left(\lambda^{\theta}\mu\right)^{n}\geq\rho^{N},\text{ }\forall 1\leq k\leq n_{0}$
    which is equivalent to ask
    $\frac{1}{4K^{5}}\cos\varphi\left(\frac{\mu}{\rho}\right)\left(\lambda^{\theta}\mu^{2}\right)^{n}\geq\rho^{n+N}.$
    This follow from the condition imposed in (8)-(oo) and taking $n$ large enough. In particular, $\text{IH}_{\tilde{C}}-$(i)P5 proves
    $\left|Df^{T+n+N+n_{0}}_{x}v\right|\geq\frac{C}{4K^{5}}\cos\varphi\mu^{n_{0}}\rho^{T}\left(\lambda^{\theta}\mu^{2}\right)^{n}\left|v\right|.$ (8.11)
    Moreover, by (8.9),
    $\tan\angle\left(Df_{x}^{T+n+N}v,e_{2}\right)\leq\frac{2K^{3}}{\left(\lambda^{\theta}\mu\right)^{n}}$
    which together with (8.1) and by taking $n$ large enough imply that,
    $\tan\angle\left(Df_{x}^{T+n+N+n_{0}}v,e_{2}\right)\leq\left(\frac{\lambda}{\mu}\right)^{n_{0}}\frac{2K^{3}}{\left(\lambda^{\theta}\mu\right)^{n}}<<1.$ (8.12)
  • Step 6:
    We start proving $\text{IH}_{\tilde{C}^{\tilde{N}-\tilde{T}}}-$(i) from time $T+nN+n_{0}+1$ up to time $T+n+N+n_{0}+M$. Indeed, by $\text{IH}_{H}$ and (8.11), for all $1\leq k\leq M$, we get,
    • $\text{IH}_{\tilde{C}}-$(i)P6:
      $\left|Df^{T+n+N+n_{0}+k}_{x}v\right|$ $\geq$ $\frac{1}{K}\left|Df^{T+n+N+n_{0}}_{x}v\right|$
      $\geq$ $\frac{C}{4K^{6}}\cos\varphi\mu^{n_{0}}\rho^{T}\left(\lambda^{\theta}\mu^{2}\right)^{n}\left|v\right|$
      $\geq$ $C\rho^{T+n+N+n_{0}+k}\left|v\right|.$
    Observe that in the last inequality of $\text{IH}_{\tilde{C}}-$(i)P6 we used that
    $\frac{1}{4K^{6}}\cos\varphi\left(\frac{\mu}{\rho}\right)^{n+n_{0}}\left(\lambda^{\theta}\mu\right)^{n}\geq\rho^{N+k},\text{ }\forall 1\leq k\leq M$
    which is equivalent to ask
    $\frac{1}{4K^{6}}\cos\varphi\left(\frac{\mu}{\rho}\right)^{n_{0}}\left(\frac{\lambda^{\theta}\mu^{2}}{\rho}\right)^{n}\geq\rho^{N+M}.$
    This follow from the condition imposed in (8)-(oo) and taking $n$ large enough. In particular, $\text{IH}_{\tilde{C}}-$(i)P6 proves
    $\left|Df^{T+n+N+n_{0}+M}_{x}v\right|\geq\frac{C}{4K^{6}}\cos\varphi\mu^{n_{0}}\rho^{T}\left(\lambda^{\theta}\mu^{2}\right)^{n}\left|v\right|.$ (8.13)
    Observe that, there exists $\beta>0$ such that
    $\angle\left(Df^{M}_{q_{2}}e_{2},e_{1}\right)>\beta$
    and by (8.12),
    $\text{d}_{\text{v}}\left(f^{T+n+N+n_{0}}(x),q_{2}\right)<<1.$
    The previous two inequalities together with $\text{IH}_{H}$ gives
    $\angle\left(Df^{T+n+N+n_{0}+M}_{x}v,e_{1}\right)>\frac{\beta}{2}.$
    In particular, if we denote by
    $v^{(6)}=Df^{T+n+N+n_{0}+M}_{x}v=\left(\begin{matrix}v^{(6)}_{1}\\ v^{(6)}_{2}\end{matrix}\right)$,
    it follows that
    $\frac{v^{(6)}_{2}}{\left|v^{(6)}\right|}\geq\sin\frac{\beta}{2}.$ (8.14)
  • Step 7:
    We start proving $\text{IH}_{\tilde{C}^{\tilde{N}-\tilde{T}}}-$(i) from time $T+nN+n_{0}+M+1$ up to time $T+n+N+n_{0}+M+\overline{\ell}$. Indeed, by $\text{IH}_{Q}$, (8.14) and (8.13), for all $1\leq k\leq\overline{\ell}$, we get,
    • $\text{IH}_{\tilde{C}}-$(i)P7:
      $\left|Df^{T+n+N+n_{0}+M+k}_{x}v\right|$ $\geq$ $\mu^{k}\sin\frac{\beta}{2}\left|Df^{T+n+N+n_{0}+M}_{x}v\right|$
      $\geq$ $\frac{C}{4K^{6}}\cos\varphi\sin\frac{\beta}{2}\mu^{n_{0}+k}\rho^{T}\left(\lambda^{\theta}\mu^{2}\right)^{n}\left|v\right|$
      $\geq$ $C\rho^{T+n+N+n_{0}+M+k}\left|v\right|.$
    Observe that in the last inequality of $\text{IH}_{\tilde{C}}-$(i)P7 we used that
    $\frac{1}{4K^{6}}\cos\varphi\sin\frac{\beta}{2}\left(\frac{\mu}{\rho}\right)^{n_{0}+k}\left(\frac{\lambda^{\theta}\mu^{2}}{\rho}\right)^{n}\geq\rho^{N+M}.$
    This follow from the condition imposed in (8)-(oo) and taking $n$ large enough. Denote by
    $v^{(7)}=Df^{T+n+N+n_{0}+M+\overline{\ell}}_{x}v=\left(\begin{matrix}v^{(7)}_{1}\\ v^{(7)}_{2}\end{matrix}\right)$,
    then, by $\text{IH}_{Q}$ and by (8.14), we get
    $\frac{v^{(7)}_{1}}{v^{(7)}_{2}}=\left(\frac{\lambda}{\mu}\right)^{\overline{\ell}}\frac{v^{(6)}_{1}}{v^{(6)}_{2}}\leq\left(\frac{\lambda}{\mu}\right)^{\overline{\ell}}\cot\frac{\beta}{2}<\tan\varphi,$
    with the last inequality achieved by taking $\overline{\ell}$ large enough.
In conclusion, the inequality above proves $\text{IH}_{\tilde{C}}-$(ii) and $\text{IH}_{\tilde{C}}-$(i) is achieved by putting together $\text{IH}_{\tilde{C}}-$(i)P1, $\text{IH}_{\tilde{C}}-$(i)P2, $\text{IH}_{\tilde{C}}-$(i)P3, $\text{IH}_{\tilde{C}}-$(i)P4, $\text{IH}_{\tilde{C}}-$(i)P5, $\text{IH}_{\tilde{C}}-$(i)P6 and $\text{IH}_{\tilde{C}}-$(i)P7.
9 The main objects: the lamination and the invariant Cantor set
In this section we build a lamination in parameter space whose maps have an invariant Cantor set.
9.1 The lamination
Consider the $1$-tangency strip $S$ given by $S=[-t_{0},t_{0}]\times[-a_{0},a_{0}]$ with tangency line $b=\left\{a=0\right\}$. Take $U=S$. In Subsection 2.1, we defined a family $F$ which is an unfolding of the tangency $b$ and each map $f\in S$ has a $(1,N_{1},M_{1})$ dynamical partition with $N_{1}=N$ and $M_{1}=M$, where $N$ and $M$ are defined in Definition 2, property $(f10)$.
Denote by
$\mathfrak{S}_{1}=S$
and by
$\mathfrak{S}_{2}=\left\{S_{j,2}\right\}_{j=1}^{\infty}$
where each $S_{j,2}$ is a $n_{2}+j$-tangency strip contained in $S$ whose existence is guaranteed by Proposition 7. By applying Proposition 7 inductively, for all $g\geq 3$ one can define a sequence of tangency strips
$\mathfrak{S}_{g}=\left\{S_{j,g}\right\}_{j=1}^{\infty}$
satisfying the following properties:
  • $\forall j,j^{\prime}$ with $j\neq j^{\prime}$, $S_{j,g}\cap S_{j^{\prime},g}=\emptyset$,
  • $\forall S_{j,g}\in\mathfrak{S}_{g}$ there exists a unique $S_{i,g-1}\in\mathfrak{S}_{g-1}$ with $S_{j,g}\subset S_{i,g-1}$.
We are now ready to define the lamination
$\mathfrak{L}=\bigcap_{g=1}^{\infty}\bigcup_{j=1}^{\infty}S_{j,g}.$ (9.1)
A leave of the lamination is defined as
$\ell=\bigcap_{g=1}^{\infty}S_{g}$ (9.2)
where, for all $g$, $S_{g}\in\mathfrak{S}_{g}$ and $S_{g}\subset S_{g-1}$. Observe that, by Equation (7.1), $\mathfrak{L}$ is a ${\mathcal{C}}^{1}$} lamination, i.e. each leave is the graph of a ${\mathcal{C}}^{1}$} function.
The following proposition proves Theorem A-(v).
Proposition 9.
Every leaves of $\mathfrak{L}$ is the graph of a smooth function $\ell:[-t_{0},t_{0}]\mapsto[-a_{0},a_{0}]$. Morever if $t\in[-t_{0},t_{0}]$, the closure of the set $L_{t}=\left\{(t,a)\in\mathfrak{L}\right\}$ is a Cantor set.
Proof.
The proof follows from the construction of the lamination, and Proposition 7 and Proposition 4. The main point is that each strip contains leaves corresponding to maps with a new homoclinic tangency and these leaves are graphs of global curves. ∎
9.2 The invariant Cantor set
The following discussion is consequence of Proposition 7.
Given a leave of the lamination $\ell$ and a map $f\in\ell$, one can associate a sequence $\left\{\left(b_{g},U_{g},\mathfrak{D}_{g}\right)\right\}_{g=1}^{\infty}$ and a sequence of numbers $\left\{\left(1,N_{g},M_{g}\right)\right\}_{g=1}^{\infty}$ such that, for all $g$, $(b_{g},U_{g})$ is the $(1,N_{g},M_{g})$ descendant of $S_{g-1}$ and $\mathfrak{D}_{g}$ is the $(1,N_{g},M_{g})$ dynamical partition of $f$. The sequence $\left\{\left(1,N_{g},M_{g}\right)\right\}_{g=1}^{\infty}$ is called the type of $\ell$.
Moreover for each $g$, $\mathfrak{D}_{g}$ is a refinement of $\mathfrak{D}_{g-1}$, with mesh going to zero when $g$ goes to infinity. We can now define the set associated to $f$
$A(f)=\bigcap_{g=1}^{\infty}\mathcal{D}_{g}=\bigcap_{g=1}^{\infty}\bigcup_{D_{g}\in\mathfrak{D}_{g}}{D}_{g}$
where $\mathcal{D}_{g}$ is the union of the elements of $\mathfrak{D}_{g}$. This set is going to play a fundamental role in our research. Let us begin by studying its properties. Following Proposition 7, for all $g$, let $X_{g}$ be the $(1,N_{g},M_{g})$ directed graph and let $w_{g}:X_{g+1}\to X_{g}$ be the projection and $w_{g^{\prime}g}:X_{g}\to X_{g^{\prime}}$ be
$w_{g^{\prime}g}=w_{g^{\prime}}\circ\cdots w_{g-1}\circ w_{g}.$
9.3 Properties of the set AA
In this subsection we collect properties of the set $A$. In particular, the following lemma proves Theorem A-(i).
Lemma 22.
Let $\ell$ be a leave of the lamination $\mathfrak{L}$ and let $f\in\ell$. Then the set $A(f)$ is a Cantor set.
Proof.
The set $A(f)$ is compact because it is the intersection of a nested sequence of compact sets and it is totally disconnected because the $\text{mesh}(\mathfrak{D}_{g})$ goes to zero when $g$ goes to infinity.
It is left to prove that $A(f)$ is perfect. Let us start by observing that, by Remark 14, each set $D_{g}\in\mathfrak{D}_{g}$ contains at least two sets $D_{g+1},D^{\prime}_{g+1}\in\mathfrak{D}_{g+1}$.
Now take $x\in A(f)$, then for every $g$ there exists $D_{g}\in\mathfrak{D}_{g}$ such that $x\in D_{g}$. Fix $\epsilon>0$, since $\text{mesh}(\mathfrak{D}_{g})$ goes to zero, there exists $g$ with $\text{diam}(D_{g})<\epsilon$. Let $D_{g+1}\in\mathfrak{D}_{g+1}$ with $x\in D_{g+1}\subset D_{g}$. By the previous observation, there exists $D^{\prime}_{g+1}\in\mathfrak{D}_{g+1}$ such that $D^{\prime}_{g+1}\subset D_{g}$ and $D_{g+1}\cap D^{\prime}_{g+1}=\emptyset$. Take $y\in A(f)\cap D^{\prime}_{g+1}$, then $y\neq x$ and $\text{dist}(x,y)\leq\text{dist}(D_{g})\leq\epsilon$. The proof is now complete. ∎
In the following lemma we describe the possible limit behavior of points in $A$. We denote by $\omega(x)$, the $\omega$-limit set of the point $x$. In particular, it implies that the set $A$ is non-hyperbolic.
Recall that $q_{2}$ is the point with a transversal homoclinic intersection. Let
$O_{2}=\left\{f^{t}(q_{2})\left|\right.t\in\mathbb{Z}\right\}\cup\left\{p\right\}.$
The proof of the following lemmas relies on the directed graphs introduced before. We will need the following notation. Denote the corresponding prime loops in $X_{g}$ respectively by $\lambda^{q}_{g}$, $\lambda^{h}_{g}$, and $\lambda^{c}_{g}$. We will use frequently the following observation. If a loop in $X_{g}$ consists purely of $h$- and $q$-loops of $X_{g}$ then $w_{g^{\prime}g}(\lambda)$ consists purely of $h$- and $q$-loops of $X_{g^{\prime}}$.
Lemma 23.
Let $\ell$ be a leave of the lamination $\mathfrak{L}$ and let $f\in\ell$. Each $x\in A(f)$ satisfies one of the following:
  • $\omega(x)=\left\{p\right\}$,
  • $\omega(x)=O_{2}$,
  • $\omega(x)=A(f)$.
Proof.
Let $x\in A(f)$, $x\neq p$, in particular $x=\bigcap_{g}D_{g}$ where, for all $g$, $D_{g}\in\mathfrak{D}_{g}$ and $D_{g}\subset D_{g+1}$. Fix $g$ and denote by $x_{g}=\pi_{g}(D_{g})\in X_{g}$ where $\pi_{g}$ is the projection of $D_{g}$ on $X_{g}$. Observe that $w_{g-1}(x_{g})=x_{g-1}$. We claim that, if for infinitely many $g$,
$w_{g-1}\left(\overrightarrow{x_{g}q_{g}}\right)\supset\lambda^{c}_{g-1}$
then $\omega(x)=A(f)$. Indeed, take an open set $U\subset A(f)$ and let $g_{0}$ be such that there exists $\tilde{D}_{g_{0}}\in\mathfrak{D}_{g_{0}}$ with $\tilde{D}_{g_{0}}\subset U$ and
$w_{g_{0}+1}\left(\overrightarrow{x_{g_{0}+2}q_{g_{0}+2}}\right)\supset\lambda^{c}_{g_{0}+1}.$
Observe,
$w_{g_{0}}(\lambda^{c}_{g_{0}+1})=X_{g_{0}}.$
Hence, there exists $t$ such that $f^{t}\left(D_{g_{0}+2}\right)\subset\tilde{D}_{g_{0}}$ and in particular $f^{t}(x)\in\tilde{D}_{g_{0}}\subset U$.
Now consider the opposite situation when there exists $g_{0}\in\mathbb{N}$ such that $w_{g-1}\left(\overrightarrow{x_{g}q_{g}}\right)\subset X_{g-1}$ does not contain $\lambda^{c}_{g-1}$ for every $g\geq g_{0}$. We will prove in this case that
$\omega(x)=O_{2}.$
The proof relies on the following claim. Let $g_{0}\leq g^{\prime}<g$. There exists $t_{g^{\prime}}\geq 0$ such that for every $g>g^{\prime}$
$w_{g^{\prime}g}(\overrightarrow{y_{g}q_{g}})=\lambda_{g^{\prime}}^{q}\cup\lambda_{g^{\prime}}^{h},$
where $y=f^{t_{g^{\prime}}}(x)$. In other words, the loop $w_{g^{\prime}g}(\overrightarrow{y_{g}q_{g}})$ is a concatenation of multiple times the $q$-loop and $h$-loop of $X_{g^{\prime}}$.
The proof of this Claim is by induction. First observe
$w_{g^{\prime}}(\overrightarrow{x_{g^{\prime}+1}q_{g^{\prime}+1}})=\gamma\lambda,$
where $\gamma=w_{g^{\prime}}(\overrightarrow{x_{g^{\prime}+1}y_{g^{\prime}+1}})\subset\lambda^{c}_{g^{\prime}}$ is a path of length $t_{g^{\prime}}$ and $y=f^{t_{g^{\prime}}}(x)$. Moreover, $\lambda$ is a loop which is a concatenation of multiple $\lambda_{g^{\prime}}^{q}$ and one $\lambda_{g^{\prime}}^{h}$.
Suppose the Claim is proved for $g>g^{\prime}$. Observe,
$w_{g}(\overrightarrow{y_{g+1}q_{g+1}})=\overrightarrow{y_{g}q_{g}}\lambda,$
where $\lambda$ is a loop which is a concatenation of multiple $\lambda_{g}^{q}$ and one $\lambda_{g}^{h}$. This is true because $w_{g}(\overrightarrow{y_{g+1}q_{g+1}})$ cannot contain the $c$-loop of $X_{g}$, we are in the second case. The induction assumption implies that
$w_{g^{\prime}g}(\overrightarrow{y_{g}q_{g}})=\lambda_{g^{\prime}}^{q}\cup\lambda_{g^{\prime}}^{h}$
Moreover, using that projections of $q$- and $h$-loops are concatenations of $q$- and $h$-loops, we get
$w_{g^{\prime}g}(\lambda)=\lambda_{g^{\prime}}^{q}\cup\lambda_{g^{\prime}}^{h}.$
This finishes the proof of the Claim.
The Claim implies that for all $g^{\prime}\geq g_{0}$,
$\text{Orbit}(f^{t_{g}^{\prime}}(x))\subset Q_{g^{\prime}}\cup\bigcup_{j=1}^{M_{g^{\prime}}-1}H_{g^{\prime}}(j).$
Hence,
$\omega(x)\subset\bigcap_{g\geq g_{0}}\left(Q_{g}\cup\bigcup_{j=1}^{M_{g}-1}H_{g}(j)\right)$
where for all $g$ and for all $j$, $Q_{g}$ and $H_{g}(j)$ are the sets in $\mathfrak{D}_{g}$ as described in Definition 7. In order to complete the proof one should establish that
$\bigcap_{g}\left(Q_{g}\cup\bigcup_{j=1}^{M_{g}-1}H_{g}(j)\right)=O_{2}.$
One inclusion is trivial. Let us concentrate on the other one and for this purpose let us take a point
$x\in\bigcap_{g}\left(Q_{g}\cup\bigcup_{j=1}^{M_{g}-1}H_{g}(j)\right).$
If $x=p$ then $x\in O_{2}$. Assume $x\neq p$, then from a certain moment on
$x\in\bigcap_{g\geq\tilde{g}}\bigcup_{j=1}^{M_{g}-1}H_{g}(j).$
Here we are using the fact that if $x_{g}\neq q_{g}$, then the same holds for all next points in the sequence defining the point $x$. Say, for $g\geq\tilde{g}$
$x\in H_{g}(j_{x}).$
In particular, there exists a number $t_{g}$ such that
$f^{t_{g}}(q_{2})\in H_{g}(j_{x}).$
Because, $f^{t}(q_{2})\to p$ when $|t|\to\infty$ we have that $t_{g}$ is bounded. Observe also that $f^{t_{g}}(q_{2})$ is close to $x$ because the diameter of $H_{g}(j_{x})$ goes to zero when $g$ goes to infinity. It follows that $f^{t}(q_{2})=x$ form some $t\in\mathbb{Z}$. Hence, $x\in O_{2}$ and the proof is complete. ∎
A natural question at this point is: does a point with dense orbit exist? The answer is yes and the proof is the following lemma which also proves Theorem A-(ii).
Lemma 24.
Let $\ell$ be a leave of the lamination $\mathfrak{L}$ and let $f\in\ell$. There exists a point $c\in A(f)$ such that $\omega(c)=A(f).$
Proof.
Fix $C_{1}\in\mathfrak{D}_{1}$ and consider $c_{1}=\pi_{1}(C_{1})\in X_{1}$ where $\pi_{1}$ is the projection of $\mathcal{D}_{1}$ on $X_{1}$. We want to prove that for infinitely many $g$ there is an element $c_{g}$ in the $c$-loop of $X_{g}$ such that
$w_{g-1}\left(\overrightarrow{c_{g}q_{g}}\right)\supset\lambda^{c}_{g-1}$
and $w_{g-1}(c_{g})=c_{g-1}$.
Suppose we have constructed $c_{1},\dots,c_{g-1}$. Consider the unique path $\gamma=\overrightarrow{c_{g-1}q_{g-1}}$ contained in the $c$-loop of $X_{g-1}$. Take the first vertex $c_{g}$ in the $c$-loop of $X_{g}$ with the property that $w_{g-1}(c_{g})=c_{g-1}$. Observe that
$w_{g-1}\left(\overrightarrow{c_{g}q_{g}}\right)=\gamma\circ\lambda$
where $\lambda$ is a loop and, by Remark 14 $\lambda$ contains all primary loops of $X_{g-1}$. In particular $w_{g-1}\left(\overrightarrow{c_{g}q_{g}}\right)\supset\lambda^{c}_{g-1}$.
For all $g$, consider the set of $\mathfrak{D}_{g}$ defined by $\pi_{g}(C_{g})=c_{g}$, where $\pi_{g}$ is the projection of $D_{g}$ on $X_{g}$. Our selected point is
$c=\bigcap_{g}C_{g}\in A(f).$
Following the proof of Lemma 23 one gets that $\omega(c)=A(f)$. ∎
The following lemma proves Theorem A-(iii).
Lemma 25.
Let $\ell$ be a leave of the lamination $\mathfrak{L}$ and let $f\in\ell$. There exists a point $b\in A(f)$ such that $\omega(b)=O_{2}.$
Proof.
Observe, for every $x_{g-1}\in\lambda^{c}_{g-1}\subset X_{g-1}$ there exists a unique $x_{g}\in\lambda^{c}_{g}\subset X_{g}$ such that
$w_{g-1}(\overrightarrow{x_{g}q_{g}})=\overrightarrow{x_{g-1}q_{g-1}}\lambda,$
where $\lambda$ is a concatenation of multiple $\lambda^{q}_{g-1}$ and only one $\lambda^{h}_{g-1}$.
Choose $x_{1}\in\lambda^{c}_{1}$, $x_{1}\neq q_{1}$, and use the above observation to construct the sequence $x_{g}$, $g\geq 1$. Let $D_{g}\in\mathfrak{D}_{g}$ such that
$x_{g}=\pi_{g}(D_{g})$
and $x=\bigcap_{g}D_{g}$.
Given $g_{0}$, let $t_{g_{0}}$ be such that the point $y=f^{t_{g_{0}}}(x)$ satisfies
$w_{g_{0}}(\overrightarrow{y_{g_{0}+1}q_{g_{0}+1}})=\lambda,$
where $\lambda$ is a concatenation of multiple $\lambda^{q}_{g_{0}}$ and only one $\lambda^{h}_{g_{0}}$. Now we claim that for all $g_{0}<g$
$w_{g_{0}g}(\overrightarrow{y_{g}q_{g}})=\lambda,$
where $\lambda$ is a concatenation of multiple $\lambda^{q}_{g_{0}}$ and $\lambda^{h}_{g_{0}}$.
The proof of the Claim is by induction. It holds for $g=g_{0}+1$. By construction we have
$w_{g}(\overrightarrow{y_{g+1}q_{g+1}})=\overrightarrow{y_{g}q_{g}}\lambda,$
where $\lambda$ is a concatenation of multiple $\lambda^{q}_{g}$ and one $\lambda^{h}_{g}$. Then
$w_{g_{0}g}(\overrightarrow{y_{g+1}q_{g+1}})=w_{g_{0}g}(\overrightarrow{y_{g}q_{g}})w_{g_{0}g}(\lambda).$
The induction hypothesis implies that $w_{g_{0}g}(\overrightarrow{y_{g}q_{g}})$ is a concatenation of multiple $\lambda^{q}_{g_{0}}$ and $\lambda^{h}_{g_{0}}$. As used before we have $w_{g_{0}g}(\lambda)$ is a concatenation of multiple $\lambda^{q}_{g_{0}}$ and $\lambda^{h}_{g_{0}}$. This finishes the proof of the Claim.
The Claim implies for all $g\geq 1$
$\omega(x)\subset Q_{g}\cup\bigcup_{j=1}^{M_{g}-1}H_{g}(j)$
and
$\omega(x)\cap Q_{g}\neq\emptyset,$
and for all $j\leq M_{g}-1$
$\omega(x)\cap H_{g}(j)\neq\emptyset.$
This means that $\omega(x)=O_{2}$ where we use that
$\bigcap_{g}\left(Q_{g}\cup\bigcup_{j=1}^{M_{g}-1}H_{g}(j)\right)=O_{2}.$
In the next proposition we prove that $A$ is an invariant set.
Proposition 10.
Let $\ell$ be a leave of the lamination $\mathfrak{L}$ and let $f\in\ell$. Then the map $f:A(f)\mapsto A(f)$ is an homeomorphism.
Proof.
Since $f$ is injective, its restriction to $A$ is injective as well. We only need to prove that $f(A(f))=A(f)$. Observe that for all $x$, $f(\omega(x))=\omega(x)$. In particular, when $x=c$, $f(\omega(c))=\omega(c)$ and by Lemma 24 we obtain $f(A(f))=A(f)$. The proof is complete. ∎
10 Topological Equivalence along the Leaves
The aim of this section is to prove that two maps which are in the same leave of the lamination are conjugate on their respective Cantor sets, see Theorem 4. In order to do that, we need to build a topological model for the invariant Cantor sets.
10.1 Inverse limits
Take a leave $\ell$ of the lamination $\mathfrak{L}$ as constructed in (9.1) and (9.2). By the construction following Proposition 7 we know that for all $g$ there is a $(1,N_{g},M_{g})$-directed graph and a graph morphism $w_{g}:X_{g+1}\mapsto X_{g}$. We can then define the inverse limit as
$X=\underleftarrow{\lim}X_{g}$
with the corresponding topology.
A sequence $\left\{\underline{x}_{n}\right\}_{n\in\mathbb{N}}=\left\{\left(x_{g,n}\right)\right\}_{n\in\mathbb{N}}\subset X$ converges to a point $\underline{x}=\left(x_{g}\right)\in X$, namely
$\left\{\underline{x}_{n}\right\}_{n\in\mathbb{N}}\mapsto\underline{x}\iff\forall g\text{ }\exists n_{0}\in\mathbb{N}\text{ s.t }\forall n\geq n_{0}\text{, }x_{g,n}=x_{g}.$
Lemma 26.
Let $\underline{x}\in X$, $\underline{x}=(x_{g})$. If there exist $g_{0}$ such that $x_{g_{0}}=q_{g_{0}}$ then for all $g\leq g_{0}$, $x_{g}=q_{g}$.
Proof.
Since for all $g$, $w_{g-1}(q_{g})=q_{g-1}$, then $x_{g_{0}-1}=w_{g_{0}-1}(x_{g_{0}})=w_{g_{0}-1}(q_{g_{0}})=q_{g_{0}-1}$. The same can be repeated for all $g\leq g_{0}$. ∎
In the following proposition we prove the existence of an homeomorphism on the inverse limit defined by the inverse graphs.
Proposition 11.
There exists a unique homeomorphism $\phi:X\to X$ with the following property. For all sequences $\underline{x}=(x_{g})\in X$, if $\phi(\underline{x})=\underline{y}=(y_{g})\in X$ then $\overrightarrow{x_{g}y_{g}}$ is an edge in $X_{g}$ for all $g$.
Proof.
Let $\underline{x}=(x_{g})\in X$. By Lemma 26, either $\underline{x}=\underline{q}=(q_{g})$ or there exists $g_{0}$ such that, for all $g\geq g_{0}$, $x_{g}\neq q_{g}$. For the case when $\underline{x}=\underline{q}$, define $\phi(\underline{x})=\underline{q}$. In the other case define $\phi(\underline{x})=\underline{y}=(y_{g})$ where $y_{g}=q_{g}$ for all $g<g_{0}$ and $y_{g}$ is the vertex in the unique outgoing edge starting in $x_{g}$. Since $w_{g}(x_{g+1})=x_{g}$ and since $x_{g}$ has a unique outgoing edge, if $w_{g}(\overrightarrow{x_{g+1}y_{g+1}})=\overrightarrow{x_{g}z}$, then $\overrightarrow{x_{g}z}=\overrightarrow{x_{g}y_{g}}$ and $w_{g}(y_{g+1})=y_{g}$. Hence $\phi(\underline{x})=\underline{y}=(y_{g})\in X$.
It is left to prove that $\phi$ is an homeomorphism. For the injectivity, take $\underline{x}=(x_{g}),\underline{\tilde{x}}=(\tilde{x}_{g})\in X$ with $\underline{x}\neq\underline{\tilde{x}}$. We start proving that if $\underline{x}\neq\underline{q}$ then $\phi(\underline{x})\neq\underline{q}$. If not, $\phi(\underline{x})=\underline{y}=(y_{g})=\underline{q}=(q_{g})$ and for all $g$, $y_{g}=q_{g}$. Since $\underline{x}\neq\underline{q}$, there exists $g_{0}$ such that, for all $g\geq g_{0}$ $x_{g}\neq q_{g}$. In particular, for all $g\geq g_{0}$, $x_{g}$ either is the precursor of $q_{g}$ in the $c$-loop or $x_{g}$ is the precursor of $q_{g}$ in the $h$-loop. Denote this loop by $\lambda_{g}$. Take $g\geq g_{0}+1$, then there exists $z_{g}\in\lambda_{g}\setminus q_{g}$ such that $\overrightarrow{z_{g}q_{g}}\subset w_{g-1}^{-1}(q_{g-1})\cap\lambda_{g}$. Observe that $x_{g}$ cannot be contained in $\overrightarrow{z_{g}q_{g}}$, otherwise $w_{g-1}(x_{g})=q_{g-1}=x_{g-1}$. This contradicts the fact that $x_{g-1}\neq q_{g-1}$. Hence, if $\underline{x}\neq\underline{q}$, then $\phi(\underline{x})\neq\underline{q}$.
[…]
In order to prove that $\phi$ is a surjection, let $\underline{y}=(y_{g})\in X$. If $\underline{y}=\underline{q}$, then $\phi(\underline{q})=\underline{y}$. Suppose $\underline{y}\neq\underline{q}$. By Lemma 26, there exists $g_{0}$ such that, for all $g\geq g_{0}$, $y_{g}\neq q_{g}$ and $y_{g}$ has a unique precursor $x_{g}$. If $w_{g}(\overrightarrow{x_{g+1}y_{g+1}})=\overrightarrow{z_{g}y_{g}}$, by the uniqueness of the predecessor $z_{g}=x_{g}$ and $w_{g}(x_{g+1})=x_{g}$. Hence, the sequence $\underline{x}=(x_{g})\in X$ and $\phi(\underline{x})=\underline{y}$.
We are now going to prove that $\phi$ is continuous.
Take a sequence $\underline{x}_{n}\in X$ converging to an element $\underline{x}\in X$.
Case 1. If $\underline{x}\neq\underline{q}$ then, by Lemma 26 there exists $g_{0}$ such that for all $g\geq g_{0}$ $x_{g}\neq q_{g}$ and $\phi(x_{g})=y_{g}$ is the unique successor of $x_{g}$. Moreover, since $\underline{x}_{n}\to\underline{x}$, there exists $n_{0}$ such that, for all $n\geq n_{0}$, $x_{g,n}=x_{g}$. This says that for all $n\geq n_{0}$, $y_{g}$ is also the successor of $x_{n,g}$ and $\phi(x_{g,n})=y_{g,n}=y_{g}$. Hence, $\phi(\underline{x}_{n})\to\underline{y}=\phi(\underline{x})$.
Case 2. If $\underline{x}=\underline{q}$, then for all $g$, $x_{g}=q_{g}$. Choose $g$. Since $\underline{x}_{n}\to\underline{x}$, there exists $n_{0}$ such that, for all $n\geq n_{0}$, $x_{g+1,n}=q_{g+1}=x_{g+1}$. Let $y_{g+1,n}$ be the successor of $x_{g+1,n}=q_{g+1}$. Then $y_{g+1,n}\in Q_{g+1}\bigcup C_{g+1}(1)\bigcup H_{g+1}(1)\subset w_{g}^{-1}(q_{g})$. This implies that $y_{g,n}=w_{g+1}(y_{g+1,n})=q_{g}$. Hence, for all $n\geq n_{0}$, $y_{g,n}=q_{g}$ and $\phi(x_{g,n})=q_{g}$. In conlcusion $\phi(\underline{x})\to\underline{q}$. The construction shows that the homeomorphism is unique. ∎
As for the directed graphs, if for all $g$, $\mathfrak{D}_{g}$ is the $(1,N_{g},M_{g})$ dynamical partition of a map $f\in\ell$, $\mathcal{D}_{g}$ is the union of the elements of $\mathfrak{D}_{g}$ and $i_{g}:\mathcal{D}_{g+1}\to\mathcal{D}_{g}$ is the inclusion, we can define the inverse limit as
$\mathcal{D}=\underleftarrow{\lim}\mathcal{D}_{g}.$
10.2 Diagram
In this section we build a diagram structure as in Figure 12. This is the the core of the proof of Theorem 4.
Figure 12: Diagram Structure
Orserve that, for all $g$, $\pi_{g+1}\circ w_{g}=i_{g}\circ\pi_{g}$. Moreover, for all $g$, let $j_{g}:A\to\mathcal{D}_{g}$ be the inclusion and let $j:A\to\mathcal{D}$. By construction the map $j$ is an homeomorphism. Similarly define the projection $\pi:\mathcal{D}\to X$.
Lemma 27.
The map $\pi:\mathcal{D}\to X$ is an homeomorphism.
Proof.
Observe that $\pi$ is continuous by construction and since $\mathcal{D}$ is compact it is enough to show that $\pi$ is injective and surjective.
Let $\underline{x}=(x_{g})\in X$ and for all $g$, let $\pi^{-1}_{g}(x_{g})=U_{g}\subset\mathcal{D}_{g}$. Since for all $g$, $\pi_{g+1}\circ w_{g}=i_{g}\circ\pi_{g}$, it follows that $\dots\subset U_{g+1}\subset U_{g}$. Let ${a}\in\bigcap U_{g}$. Then $(a)\in\mathcal{D}$ and $\pi((a))=\underline{x}$. This proves that $\pi$ is a surjective map.
Moreover, since $\text{mesh}(U_{g})$ tends to zero, then $\left\{a\right\}=\bigcap U_{g}$. Hence $\pi^{-1}(\underline{x})=(a)$.
The proof is complete. ∎
We are now ready to define an homeomorphism $h:A\to X$ as
$h=\pi\circ j.$
Lemma 28.
The following diagram commutes,
[…]
Namely $h\circ f=\phi\circ h$.
Proof.
Let $a\in A$. If $a=p$, then $h(f((p))=h(p)=q$ and $\phi(h(p))=\phi(q)=q$.
Assume $a\neq p$, $a=\bigcap_{g}U_{g}$ where for all $g$, $U_{g}$ is a set in $\mathcal{D}_{g}$. Let $x_{g}=\pi_{g}(U_{g})$. By definition $h(a)=\underline{x}=(x_{g})$. Since $a\neq p$, there exists $g_{0}$ such that, for all $g\geq g_{0}$, $U_{g}\neq Q_{g}$ and $x_{g}\neq q_{g}$. In particular, $x_{g}$ has a unique successor $y_{g}$. Let $V_{g}=\pi^{-1}(y_{g})$. Since $\overrightarrow{x_{g}y_{g}}$ is an edge, $f(U_{g})\subset V_{g}$. Hence $\phi(h(a))=\phi(\underline{x})=(y_{g})=f(h(a))$. ∎
The following theorem proves Theorem A-(vi).
Theorem 4.
If $f$ and $\tilde{f}$ are in the same leave $\ell$ of the lamination $\mathfrak{L}$ then they are conjugate on their respective Cantor sets.
Proof.
The dynamical partitions of both maps $f$ and $\tilde{f}$ give rise to the same $(1,N_{g},M_{g})$-directed graph and graph morphism $w_{g}:X_{g+1}\mapsto X_{g}$, for all $g\geq 1$. According to Lemma 28 both $A(f)$ and A(f~)A($\tilde${f}) are conjugate to the corresponding $X$. ∎
Remark 15.
Probably, the Cantor sets of maps on different leaves are not conjugated.
10.3 Codimension-1 Stability
In this section we prove that the leaves of the lamination constructed above are the loci of infinitely renormalizable maps in the following sense.
Recall, from the proof of Lemma 13, there exists a smooth function
$m_{1}:S_{g}\mapsto\mathbb{R}^{2}$
such that
  • $m_{1}\in W^{u}_{\text{loc}}(p)\cap C_{g}(0)$,
  • $m_{7}=F^{3N_{g}+(1+\theta_{g})n_{g}+n_{0,g}}(m_{1})\in W^{u}(p)$
and $W^{u}(p)$ has horizontal tangency at $m_{7}$.
Definition 12.
Let $\ell$ be a leave of type $\left\{\left(1,N_{g},M_{g}\right)\right\}_{g=1}^{\infty}$. A map $f\in S_{g}$ is renormalizable of type $(1,N_{g+1},M_{g+1})$ if $m_{7}(f)\in H_{g}(0)(f)$. Moreover, $f$ is infinitely renormalizable of type $\left\{\left(1,N_{g},M_{g}\right)\right\}_{g=1}^{\infty}$ if it is renormalizable of type $(1,N_{g+1},M_{g+1})$ for all $g\geq 1$.
Definition 13.
Let $f$ be an infinitely renormalizable map of type $\left\{\left(1,N_{g},M_{g}\right)\right\}_{g=1}^{\infty}$. The Cantor set $A(f)$ associated to $f$ is codimension-$k$ stable if there exists a codimension-$k$ manifold $W$ in parameter space containing $f$ such that every map in $W$ is infinitely renormalizable of type $\left\{\left(1,N_{g},M_{g}\right)\right\}_{g=1}^{\infty}$ and there does not exists a manifold of codimension k1k-1 with the same property. The manifold $W$ is called the manifold of stability.
The following theorem proves Theorem A-(vii).
Theorem 5.
Let $\ell$ be a leave of type $\left\{\left(1,N_{g},M_{g}\right)\right\}_{g=1}^{\infty}$. Every map $f\in\ell$ is infinitely renormalizable of type $\left\{\left(1,N_{g},M_{g}\right)\right\}_{g=1}^{\infty}$ and $A(f)$ is codimension-$1$ stable. Moreover, $\ell$ is the manifold of stability.
Proof.
[…]
11 Invariant Measures
In this section we prove that each set $A$ has a unique invariant probability measure. In particular, the following theorem proves Theorem A-(iv).
Theorem 6.
For each function $f\in\ell$, the set $A(f)$ has a unique invariant probability measure. This invariant measure is the point mass on the fixed point in $A(f)$ .
The following corollary proves Theorem A-(i).
Corollary 1.
The set $A$ is not hyperbolic.
Proof.
If $A$ is hyperbolic, then $A$ is equivalent to a subshift of finite type. Since $A$ has a dense orbit, this subset of finite type, is irreducible, i.e. it has a continuum of ergodic invariant measures. This contradicts Theorem 6. (Another reason is that $A$ has only one periodic orbit.) ∎
The proof of Theorem 6 requires some preparation. Let $X$ be the $(1,N,M)$ directed graph. We denote by $\delta^{q}$ the measure on $X$ which assigns unit mass to the vertex of the $q$-loop, $\lambda^{q}$. Similarly, we denote by $\delta^{c}$ the measure on $X$ which assigns unit mass to each vertex of the $c$-loop, $\lambda^{c}$ and by $\delta^{h}$ the measure which assigns unit mass to each vertex of the $h$-loop, $\lambda^{h}$. In follows that $\delta^{q}$ is a probability measure, $\delta^{c}(X)=N$ and $\delta^{h}(X)=M$.
Consider the span of the measures defined above,
$\mathcal{M}=\text{span}\left\{\delta^{q},\delta^{c},\delta^{h}\right\}.$
Observe that $\mathcal{M}$ consists of invariant measures in the sense that, if $\mu\in\mathcal{M}$, then for all $x\in X\setminus\{q,c_{0},h_{0}\}$,
$\mu(x)=\sum_{\overrightarrow{yx}}\mu(y).$
Similarly, for all $g$, if $X_{g}$ is the $(1,N_{g},M_{g})$ directed graph, we denote the span of the measures $\delta_{g}^{q},\delta_{g}^{c},\delta_{g}^{h}$ by $\mathcal{M}_{g}.$ In particular, $\left\{\delta_{g}^{q},\delta_{g}^{c},\delta_{g}^{h}\right\}$ is called the standard basis for $\mathcal{M}_{g}$.
Moreover the map $w_{g}:X_{q+1}\to X_{g}$, induces, by pushing forward the measure, a linear map
$w_{g*}:\mathfrak{M}_{g+1}\to\mathfrak{M}_{g}.$
As in Remark 14, the winding of $w_{g}$ is given by
$\underline{w}\left(w_{g}\right)=\left(q,{q}^{\underline{\ell}_{g}}{c}{q}^{\theta_{g}n_{g}}{c}{q}^{n_{g}}{c}{q}^{n_{0,g}}{h}{q}^{\overline{\ell}_{g}},{q}^{\underline{\ell}_{g}}{h}{q}^{\overline{\ell}_{g}}\right).$
Then the matrix for $w_{g*}$, in the standard basis of $\mathfrak{M}_{g}$ and $\mathfrak{M}_{g+1}$ is
$W_{g}=\left(\begin{matrix}1&\underline{\ell}_{g}+\theta_{g}n_{g}+n_{g}+n_{0,g}+\overline{\ell}_{g}&\underline{\ell}_{g}+\overline{\ell}_{g}\\ 0&3&0\\ 0&1&1\end{matrix}\right)$
The simplex $\mathcal{P}_{g}\subset\mathfrak{M}_{g}$ consisting of probability measures is
$\mathcal{P}_{g}=\{\mu=x\delta^{q}_{g}+y\delta^{c}_{g}+z\delta^{h}_{g}|x,y,z\geq 0\text{ and }\mu(X_{g})=1\}.$
Observe,
$W_{g}:\mathcal{P}_{g+1}\to\mathcal{P}_{g}.$
The inverse limit defines probability measures on $X=\underleftarrow{\lim}X_{g}$,
$\mathcal{P}=\underleftarrow{\lim}\mathcal{P}_{g}.$
Let $\mathcal{I}$ be the set of invariant probability measures of $f:X\to X$. If $\mu\in\mathcal{I}$ denote the restriction of the measure to the $\varsigma$-algebra of $X_{g}$ by
$\mu_{g}=w_{g,\infty*}(\mu)\in\mathcal{P}_{g}.$
Observe, $w_{g*}(\mu_{g+1})=\mu_{g}$. Hence, $\mu=(\mu_{g})\in\mathcal{P}$. By construction, the map
$\mathcal{I}\ni\mu\mapsto(\mu_{g})\in\mathcal{P}$
is a bijection. For more details on this construction see LABEL:GM.
Denote the vertices of $\mathcal{P}_{g}$ by
$\underline{\delta}^{c}_{g}=\frac{1}{N_{g}},\underline{\delta}^{h}_{g}=\frac{1}{M_{g}},\underline{\delta}^{q}_{g}=\delta^{q}_{g}.$
Observe,
$\mathcal{P}_{g}=\text{hull}\{\underline{\delta}^{c}_{g},\underline{\delta}^{h}_{g},\underline{\delta}^{q}_{g}\}\subset\mathfrak{M}_{g}.$
Define the norm $|.|$ on $\mathfrak{M}_{g}$ by
$\mu=\max\{|x|,|y|,|z|\},$
where
$\mu=x\underline{\delta}^{c}_{g}+y\underline{\delta}^{h}_{g}+z\underline{\delta}^{q}_{g}.$
Observe, if $\mu=x\underline{\delta}^{c}_{g}+y\underline{\delta}^{h}_{g}+z\underline{\delta}^{q}_{g}$ then
$\mu(\lambda^{q}_{g})=\frac{x}{N_{g}}+\frac{y}{M_{g}}+z.$
$\mu(\lambda^{c}_{g}\setminus\lambda^{q}_{g})=x\cdot\frac{N_{g}-1}{N_{g}}.$
$\mu(\lambda^{h}_{g}\setminus\lambda^{q}_{g})=y\cdot\frac{M_{g}-1}{M_{g}}.$
The map $W_{g}:\mathcal{P}_{g+1}\to\mathcal{P}_{g}.$ contracts towards $\underline{\delta}^{q}_{g}$. Namely,
Lemma 29.
If $N_{g+1},M_{g+1}\geq 32\max\{N_{g},M_{g}\}$ and $N_{g},M_{g}\geq 3$ then for all $\mu\in\mathcal{P}_{g+1}$
$|W_{g}\mu-\underline{\delta}^{q}_{g}|\leq\frac{1}{2}|\mu-\underline{\delta}^{q}_{g+1}|.$
Proof.
Observe, $W_{g}\underline{\delta}^{q}_{g+1}=\underline{\delta}^{q}_{g}$. Hence, it suffices to prove the estimate for $\mu=\underline{\delta}^{c}_{g}$ and $\mu=\underline{\delta}^{h}_{g}$. We will only present the estimate for $\mu=\underline{\delta}^{c}_{g}$, the other is similar.
Observe,
$W_{g}\underline{\delta}^{c}_{g+1}(\lambda^{q}_{g})=\frac{\underline{\ell}_{g}+\theta_{g}n_{g}+n_{g}+n_{0,g}+\overline{\ell}_{g}}{N_{g+1}}$
Because,
$N_{g+1}=\underline{\ell}_{g}+N_{g}+\theta_{g}n_{g}+N_{g}+n_{g}+N_{g}+n_{0,g}+M_{g}+\overline{\ell}_{g}$
we get
$W_{g}\underline{\delta}^{c}_{g+1}(\lambda^{q}_{g})=\frac{N_{g+1}-3N_{g}-M_{g}}{N_{g+1}}\geq\frac{7}{8}$
when $N_{g+1}\geq 32\max\{N_{g},M_{g}\}$. Hence,
$W_{g}\underline{\delta}^{c}_{g+1}(\lambda^{c}_{g}\setminus\lambda^{q}_{g}),W_{g}\underline{\delta}^{c}_{g+1}(\lambda^{h}_{g}\setminus\lambda^{q}_{g})\leq\frac{1}{8}.$
Use the coordinates of
$W_{g}\underline{\delta}^{c}_{g+1}=x\underline{\delta}^{c}_{g}+y\underline{\delta}^{h}_{g}+z\underline{\delta}^{q}_{g}.$
Recall, $x,y,z\geq 0$ and $x+y+z=1$. Then
$x=\frac{N_{g}}{N_{g}-1}W_{g}\underline{\delta}^{c}_{g+1}(\lambda^{c}_{g}\setminus\lambda^{q}_{g})\leq\frac{3}{16},$
when $N_{g}\geq 3$. Moreover,
$y=\frac{M_{g}}{M_{g}-1}W_{g}\underline{\delta}^{c}_{g+1}(\lambda^{h}_{g}\setminus\lambda^{q}_{g})\leq\frac{3}{16},$
when $M_{g}\geq 3$. Because $x+y+z=1$ we get
$z\geq\frac{5}{8}.$
The coordinates of $W_{g}\underline{\delta}^{c}_{g+1}-\underline{\delta}^{q}_{g}$ are (x,y,z1)(x,y,z-1). Hence,
$|W_{g}\underline{\delta}^{c}_{g+1}-\underline{\delta}^{q}_{g}|\leq\frac{1}{2}.$
Proof of Theorem 6: Let $\mu\in\mathcal{P}$, say $\mu=(\mu_{g})$. Recall, this measure corresponds to an invariant probability of $f$. We will show that $\mu_{g}=\underline{\delta}^{q}_{g}$, for all $g\geq 1$. The previous lemma implies that for all $s\geq 1$
$|\mu_{g}-\underline{\delta}^{q}_{g}|\leq\frac{1}{2^{s}}|\mu_{g+s}-\underline{\delta}^{q}_{g+s}|\leq\frac{1}{2^{s}}.$
Hence, $\mu_{g}=\underline{\delta}^{q}_{g}$, for all $g\geq 1$. This means that the only invariant measure of $f:A\to A$ is the unit mass on its fixed point.
12 Properties of the Lamination: Collet-Eckmann and Newhouse points
In this section, we prove the existence of a sub-lamination whose leaves have an interesting property. In particular we prove Theorem B and Theorem C.
The invariant non-hyperbolic Cantor set, associated to each map in the leaves of this sublamination have a Collet-Eckmann point, which is dense in the Cantor set. Let us start by giving the formal definition of a Collet-Eckmann point.
Definition 14.
Let $f:\mathcal{M}\to\mathcal{M}$ be a map and let $x\in\mathcal{M}$. We say that $x$ is a Collet-Eckmann (CE) point if there exists a vector $v$, in the tangent space at $x$, such that
$\left|Df^{s}_{x}v\right|\geq C\rho^{s}\left|v\right|\text{, \hskip 13.942pt }\forall\text{ }s\geq 0$
where $C$ and $\rho$ are two constants, $C>0$ and $\rho>1$.
The following Theorem proves Theorem B.
Theorem 7.
There exists a sub-lamination $\mathfrak{L}_{CE}\subset\mathfrak{L}$ satisfying the following property. For each function $f\in\ell\subset\mathfrak{L}_{CE}$, the invariant non-hyperbolic Cantor set $A(f)$ contains a Collet-Eckmann point $c$ such that $\omega(c)=A(f)$. Moreover the closure $\overline{\mathfrak{L}_{CE}}=\overline{\mathfrak{L}}$.
Proof.
The proof goes by induction on the generations $g$ defining the lamination. The induction step is contained in Proposition 8. Using the notation in that proposition we prove the base of induction, i.e., we prove that the conditions are verified at generation $g=1$.
Consider the $1$-tangency strip $S$ given by $S=[-t_{0},t_{0}]\times[-a_{0},a_{0}]$ with tangency line $b=\left\{a=0\right\}$. Take $U=S$. In Subsection 2.1, we defined a family $F$ which is an unfolding of the tangency $b$ and each map $f\in S$ has a $(1,N_{1},M_{1})$ dynamical partition with $N_{1}=N$ and $M_{1}=M$, where $N$ and $M$ are defined in Definition 2, property $(f10)$. Define $T=0$ and for all $f\in U$ let $K$ be a positive constant assuring that $\text{IH}_{H^{0}}$, $\text{IH}_{C^{0}}$-(iii) and $\text{IH}_{C^{N}}$-(iii) hold. The existence of such a constant is guaranteed by the initial conditions on the family. Observe that $\text{IH}_{Q}$, $\text{IH}_{H^{0}}$, $\text{IH}_{C^{0}}$-(i) and $\text{IH}_{C^{0}}$-(ii) are consequence of the definition of our family, while $\text{IH}_{C^{N}}$ is guaranteed since $T=0$. As for $\text{IH}_{C^{0}}$-(ii), the lower bound comes from the fact that $b$ is a non-degenerate tangency, while the upper bound is a consequence of the fact that the family is made of ${\mathcal{C}}^{2}$} diffeomorphisms. This concludes the base of induction. As said above, the induction step is Proposition 8. We can now draw our conclusions. Let
$c=\bigcap_{g}C^{N_{g}-T_{g}}\in A(f).$
Since $T_{g}$ goes to infinity, $c$ is a Collet-Eckmann point. The proof of the density is the same as the proof of Lemma 24. ∎
Even more surprisingly, there exists a sub-lamination of $\mathfrak{L}_{CE}$ whose leaves contains a map with infinitely many sinks of higher and higher order accumulating at the orbit of the Collet-Eckmann point. The following Theorem proves Theorem C.
Theorem 8.
There exists a sub-lamination $\mathfrak{L}_{N}\subset\mathfrak{L}_{CE}\subset\mathfrak{L}$ satisfying the following property. Each leave $\ell\subset\mathfrak{L}_{N}$ contains a map $f_{\ell}$ with infinitely many attracting periodic points, $p_{k}$ of arbitrarily high period, such that,
$\bigcap_{n}\overline{\bigcup_{k\geq n}\text{Orb}(p_{k})}=A(f_{\ell}).$
Moreover if $NC=\left\{f_{\ell}\right\}_{f_{\in}\ell\subset\mathfrak{L}_{N}}$, then the closure $\overline{{NC}}=\overline{\mathfrak{L}_{CE}}=\overline{\mathfrak{L}}$.
Proof.
The theorem is consequence of Theorem 7 and Theorem 3. ∎
13 Appendix
We prove that there exists a $\theta$ satisfying (8.1).
Lemma 30.
There exists $\theta\in\left(\theta_{0},\theta_{1}\right)$ such that
$\lambda^{\theta}\mu^{1+\alpha}>\lambda^{\alpha}.$
Proof.
By (4.17), there exists $K>1$ such that
$\alpha=\frac{\log K}{\log\mu}+2\theta\frac{\log\lambda}{\log\mu}+3.$
We need to prove that
$\alpha\left(\log\mu-\log\lambda\right)+\theta\log\lambda+\log\mu>0.$
We substitute the value of $\alpha$ and we get
$\log K\left(1-\frac{\log\lambda}{\log\mu}\right)-3\log\lambda+4\log\mu>\theta\left(2\frac{(\log\lambda)^{2}}{\log\mu}-3\log\lambda\right).$
The inequality is satisfied if one substitute the value of $\theta_{0}=-4\log\mu/3\log\lambda$. Hence, for a value $\theta$ larger then $\theta_{0}$ and sufficiently close to it, the inequality is still satisfied. The lemma is then proved. ∎

### v3

@@
-Stability of Non-Hyperbolic Sets
+arXiv:2511.12926v3[math.DS] 01 Mar 2026
 Abstract
-In families of unfoldings we find codimension-one laminations consisting of maps exibiting an invariant, non-hyperbolic Cantor set containing a Collet–Eckmann point whose orbit is dense in the set.
-The topology of this Cantor set remains unchanged along the leaves of the lamination, ensuring codimension-one stability.
-Furthermore, within this lamination we identify a sublamination whose leaves contain maps with infinitely many sinks accumulating on the non-hyperbolic Cantor set containing the Collet–Eckmann point.
-These results provide a new perspective on the stability of chaotic non-hyperbolic dynamics in two-dimensional systems and reveal surprising coexistence phenomena that highlight the richness and complexity of such dynamics.
+In two-dimensional unfoldings of homoclinic tangencies, the parameter space contains codimension-1 laminations whose leaves consist of maps with invariant non-hyperbolic Cantor sets.
+These Cantor sets are wild both in the sense of Hofbauer–Keller and in the sense of Newhouse, and they contain Collet–Eckmann points with dense orbits.
+Hence, wildness and non-uniform chaotic hyperbolicity can coexist on a single invariant set, while persisting along codimension-1 manifolds.
+In addition, each leaf of the lamination contains a map with infinitely many sinks accumulating on the Cantor set containing the Collet-Eckmann point.
+This surprising symbiosis of contraction and expansion may not, in fact, be pathological.
+Along the way, we introduce a generalized renormalization scheme for two-dimensional systems.
 1 Introduction
 This paper addresses two central questions in dynamical systems: the existence and stability of dynamically relevant sets.
-Given a dynamical system $(\mathcal{M},f)$ where $\mathcal{M}$ is a manifold and $f$ an evolution acting on it, one aims to identify the regions where the dynamics effectively take place—that is, the sets on which most orbits spend most of their time.
+Given a dynamical system $(\mathcal{M},f)$ where $\mathcal{M}$ is a manifold and $f$ a map acting on it, one aims to identify the regions where the dynamics effectively take place, that is, invariant sets capturing the long-term behaviour of typical orbits.
 Establishing the existence of such sets can already be a challenging task.
 Once their existence is established, one naturally seeks to understand whether these sets, together with their dynamical properties, persist under perturbations of the system.
-Are these sets isolated?
-Are their properties robust and observable across nearby systems?
+Are these properties robust and observable across nearby systems?
 In other words, are they stable?
-Let us now make the notion of stability more precise.
-Given two dynamical systems $(M,f)$ and $(M,g)$ with relevant sets $A(f)$ and $A(g)$, we say that they are conjugate if there exists a homeomorphism $h:A(f)\to A(g)$ such that $h\circ f=g\circ h.$ One can think of this homeomorphism as a change of coordinates.
+We now introduce a notion of stability tailored to this setting.
+Given two dynamical systems $(\mathcal{M},f)$ and $(\mathcal{M},g)$ with relevant sets $A(f)$ and $A(g)$, we say that they are conjugate if there exists a homeomorphism $h:A(f)\to A(g)$ such that $h\circ f=g\circ h.$ One can think of this homeomorphism as a change of coordinates.
 For instance, if $A(f)$ and $A(g)$ are Cantor sets, the conjugacy map “rearranges” the pieces of one Cantor set to produce the other.
-Roughly speaking, this means that the two sets share the same topological structure, they are topologically equivalent.
-Within the space of dynamical systems, one can thus consider the class of maps whose relevant sets are topologically equivalent in this sense.
-The geometry of this class provides information about the stability of the corresponding sets.
-We say that the set $A(f)$ is codimension-$k$ stable if the class of systems conjugate to it forms a codimension-$k$ submanifold of the space of dynamical systems.
-This manifold is referred to as the manifold of stability.
-The sets exhibiting the strongest form of stability are sinks or closed attracting periodic orbits.
-Indeed, such sets are codimension-0 stable: their topology does not change under small perturbations of the system.
-More interesting examples arise in one-dimensional dynamics, such as $\mathcal{C}^{3}$ circle maps.
-In this class of systems, the topology of the relevant set (in this case, the entire circle) persists on codimension-one manifolds.
-The set is therefore codimension-$1$ stable.
-This phenomenon follows from one of the most profound results in one-dimensional dynamics, namely Herman’s Theorem, see [28]; for more modern expositions, see [20] and [30].
-The stability of sets in one-dimensional dynamics is by now largely understood, and no major surprises are expected.
-Obtaining this complete picture required several decades of work and the efforts of many leading mathematicians.
-Although the main questions have been settled, a number of subtle and intriguing problems remain open.
-In contrast, the situation in two dimensions is far richer and less complete.
-We now turn to the family of systems at the center of our study, namely the Hénon family (see Section 3 for a precise definition).
-In fact, in this paper we consider a broader class of families, of which the Hénon family is only one example.
-Specifically, we study unfoldings of maps with a homoclinic tangency and a transversal intersection.
+In this sense, the two sets share the same topological structure and are therefore topologically equivalent.
+Within the space of dynamical systems, one can therefore consider the class of maps whose relevant sets are topologically equivalent.
+The structure of this class encodes the stability of the corresponding sets.
+We say that the set $A(f)$ is codimension-$k$ stable if the set of systems whose relevant sets are conjugate forms a codimension-$k$ submanifold of the space of dynamical systems.
+This submanifold is called the manifold of stability.
+The sets exhibiting the strongest form of stability are closed attracting periodic orbits and, more generally, hyperbolic invariant sets.
+Indeed, such sets are codimension-0 stable, meaning that their topology does not change under small perturbations of the system.
+More subtle phenomena already arise in one-dimensional dynamics.
+For instance, in the class of $\mathcal{C}^{3}$ circle maps, the relevant invariant set is the entire circle, but its topological structure is no longer stable under arbitrary perturbations.
+Instead, it persists only along codimension-1 manifolds in parameter space, and is therefore codimension-1 stable.
+This phenomenon follows from one of the most profound results in one-dimensional dynamics, namely Herman’s Theorem, see [40], and for more modern expositions [26, 43].
+The stability of invariant sets in one-dimensional dynamics is by now largely understood, and no fundamentally new phenomena are expected.
+That said, a number of subtle and interesting questions remain open, mostly concerning finer properties of the dynamics and their dependence on parameters.
+In contrast, the situation in two dimensions is much richer and far from completely understood.
+Several mechanisms of instability may interact in nontrivial ways, making questions of existence and stability more delicate.
+We now turn to the family of systems at the center of our study, namely unfoldings of maps with a homoclinic tangency and a transversal intersection.
+Such unfoldings provide a natural setting in which complex invariant sets and coexistence phenomena arise.
 The precise definition is given in Subsection 2.1.
-Which relevant sets are known to exist in such systems and what do we know about their stability?
-As expected, finitely many sinks can exist, and these are codimension-0 stable.
+A central example is provided by the Hénon family (see Section 3 for a precise definition).
+What relevant invariant sets arise in this context, and what is known about their stability?
+As expected, a finite number of attracting periodic orbits (sinks) may coexist, and these are codimension-0 stable.
 Indeed, sinks are hyperbolic sets.
-The existence of infinitely many sinks has also been established (see [33]), together with some preliminary results concerning their stability (see [23]).
-More intricate dynamics have likewise been shown to occur.
-In particular, the existence of a period-doubling Cantor attractor—an invariant, non-hyperbolic, minimal set carrying a measure with zero Lyapunov exponent—has been proved (see [14, 31]).
+More surprisingly, the existence of infinitely many sinks has also been established (see [48]), together with initial results concerning their stability (see [35]).
+Beyond periodic behaviour, non periodic dynamics have likewise been shown to occur.
+In particular, the existence of a period-doubling Cantor attractor, an invariant, non-hyperbolic, minimal set carrying a measure with zero Lyapunov exponent, has been proved (see [17, 33, 45]).
 This set is known to be codimension-1 stable.
-Furthermore, the existence of two period-doubling Cantor attractors has been demonstrated, and these are codimension-2 stable (see [36]).
-Even more complex invariant sets, known as strange attractors, have been shown to exist in this family.
-Strange attractors are invariant, non-hyperbolic, and chaotic sets (they carry a measure having positive Lyapunov exponent).
-The seminal work establishing their existence is [3], where a remarkably intricate and ingenious proof was developed.
-Similar constructions adapted to more general systems and further studies were later given in [7, 4, 6, 27, 32, 43, 44].
-Coexistence of these sets is established in [5].
-What makes these sets chaotic is the existence of a point satisfying the Collet–Eckmann condition, which, roughly speaking, asserts that the derivative of the map grows at least exponentially fast along the orbit of this point (see Definition 14 for a precise formulation).
+In addition, maps exhibiting two period-doubling Cantor attractors have been constructed, and these are codimension-2 stable (see [51]).
+Even more intricate invariant sets, known as strange attractors, have been shown to exist in Hénon-like families.
+Strange attractors are invariant, non-hyperbolic, and chaotic sets, in the sense that they support an invariant measure with positive Lyapunov exponent.
+Their existence was first established in the seminal work [4], which introduced a new and highly nontrivial construction.
+Similar constructions and further developments were later given in [5, 7, 8, 39, 47, 58, 59], and the coexistence of multiple strange attractors was established in [6].
+In many constructions, the analysis of chaotic dynamics relies on the identification of a point satisfying the Collet–Eckmann condition.
+Roughly speaking, this condition asserts that the derivative of the map grows exponentially fast along the orbit of this point (see Definition 12.0.1 for a precise formulation).
 A natural question then arises: what can be said about the stability of strange attractors?
-Their chaotic nature makes this question extremely challenging.
-It is known that, in parameter space, there exists a positive Lebesgue measure set of maps exhibiting such attractors; however, their topology changes as the parameter varies.
-So far, we lack the tools to clearly address this problem.
-The present paper aims to provide a partial answer to this question.
-A new construction and new tools are given.
-Furthermore, while trying to answer this question, we found new fascinating phenomena, namely the existence of maps with infinitely many sinks accumulating on a set containing a Collet-Eckmann point.
-We now summarize our main results, which together describe the existence and stability of non-hyperbolic Cantor sets in families of unfoldings.”
+Because of their chaotic nature, this question is particularly challenging.
+It is known that, in parameter space, there exists a positive Lebesgue measure set of maps exhibiting such attractors.
+However, the structure of these attractors typically changes as the parameter varies.
+At present, we lack general tools to determine whether, and in what sense, such chaotic sets can be stable.
+The present paper provides a first answer to this question for a class of non-hyperbolic chaotic invariant sets.
+Our approach is to identify a precise notion of stability for dynamics that lie well beyond the uniformly hyperbolic setting.
+Before summarizing our main results, we explain the viewpoint underlying this work.
+The invariant sets we construct display several types of dynamical behaviour that are commonly regarded as wild.
+First, they are statistically wild in the sense of Hofbauer and Keller, as stated in Theorem A-(iv), in that a generic orbit in the set is dense, while from a statistical point of view the behavior is degenerate, meaning that typical orbits accumulate on a very singular measure, as described for instance in [15, 41, 44].
+Second, they are wild in the sense of Newhouse, see Theorem A-(viii), in the classical terminology of [12, 48, 50]: arbitrarily small perturbations exhibit infinitely many sinks.
+Moreover, all constructed Cantor sets contain a Collet–Eckmann point with dense orbit, see Theorem B.
+Furthermore, we show that these sets persist along the leaves of a codimension-1 lamination.
+In summary, wild and non-hyperbolic chaotic behavior can coexist on a single invariant set while persisting along codimension-1 manifolds.
+In addition, Theorem C states that each leaf of the lamination contains a map with infinitely many sinks accumulating on the Cantor set containing the Collet-Eckmann point.
+This surprising symbiosis of contraction and expansion might not even be pathological, and motivates the Conjecture stated below.
+The main contribution of this paper is to show that all these phenomena can coexist on a single invariant Cantor set and, moreover, persist along codimension-one laminations in parameter space.
+Theorems A, B, and C make this precise by progressively incorporating these behaviours and establishing a unified notion of stability for such highly non-hyperbolic dynamics.
+We now summarize the main results of the paper.
 Theorem A.
-Let $\mathcal{M}$ be a two-dimensional manifold and let $F:\left([-t_{0},t_{0}]\times[-a_{0},a_{0}]\right)\times\mathcal{M}\to\mathcal{M}$ be a family of unfolding.
-There exists a codimension-$1$ lamination $\mathfrak{L}$ in parameter space such that, each leave $\ell\subset\mathfrak{L}$ and each map $f\in\ell$ have a set $A(f)$ satisfying the following.
+Let $\mathcal{M}$ be a two-dimensional manifold and let $F:\left([-t_{0},t_{0}]\times[-a_{0},a_{0}]\right)\times\mathcal{M}\to\mathcal{M}$ be an unfolding with $t_{0},t_{0}]\times\left\{0\right\}$ as homoclinic tangency locus of a saddle point.
+There exists a codimension-$1$ lamination $\mathfrak{L}$ in parameter space such that, each leaf $\ell\subset\mathfrak{L}$ and each map $f\in\ell$ have a set $A(f)$ satisfying the following.
   • (i)
-    The set $A(f)$ in an invariant, non-trivial, non-hyperbolic Cantor set.
+    The set $A(f)$ in an invariant, non-hyperbolic Cantor set.
   • (ii)
@@
   • (iii)
-    The set $A(f)$ contains a point whose limit set coincides with the orbit of a transversal homoclinic intersection point.
+    The set $A(f)$ contains a saddle point together with a corresponding transversal homoclinic intersection.
   • (iv)
@@
   • (v)
-    Every leaves of $\mathfrak{L}$ is the graph of a smooth function $\ell:[-t_{0},t_{0}]\mapsto[-a_{0},a_{0}]$.
+    Every leaf of $\mathfrak{L}$ is the graph of a smooth function $\ell:[-t_{0},t_{0}]\to[-a_{0},a_{0}]$.
 Moreover if $t\in[-t_{0},t_{0}]$, the closure of the set $L_{t}=\left\{(t,a)\in\mathfrak{L}\right\}$ is a Cantor set.
   • (vi)
-    If $f$ and $\tilde{f}$ are in the same leave $\ell$ of the lamination $\mathfrak{L}$ then they are conjugate on their respective Cantor sets.
+    If $f$ and $\tilde{f}$ are in the same leaf $\ell$ of the lamination $\mathfrak{L}$ then they are conjugate on their respective Cantor sets $A(f)$ and A(f~)A($\tilde${f}).
   • (vii)
     The set $A(f)$ is codimension-$1$ stable and $\ell$ is the smooth manifold of stability.
-Theorem $A$ is the summary of several results proved along the paper. In order: Lemma 22, Corollary 1, Lemma 24, Lemma 25, Theorem 6, Proposition 9, Theorem 4, Theorem 5.
+  • (viii)
+    Every open neighborhood of $f\in\mathfrak{L}$ contains maps with infinitely many sinks.
+Theorem A is the summary of several results proved in the paper.
+In order: Lemma 9.3.1, Corollary 11.0.2, Lemma 9.3.4, Lemma 9.3.5, Theorem 11.0.1, Proposition 9.1.3, Theorem 10.2.3, Theorem 10.3.3, Remark 9.1.5.
 Theorem A establishes the existence of an invariant, non-hyperbolic Cantor set that is codimension-1 stable.
-In particular, any two maps lying on the same leaf of the lamination are conjugate on their corresponding Cantor sets.
-In other words, along each leaf of the lamination, the corresponding Cantor sets are topologically conjugate, and their structure remains invariant.
-These results are obtained through the introduction of dynamical partitions and the associated directed graphs.
-The inverse limit of these graphs is homeomorphic to the Cantor set.
-This combinatorial structure provides effective control over both the topological dynamics along the leaves and the measure-theoretical properties of the Cantor set.
-These inverse limits and directed graphs were introduced in [21].
-In the present work, we adapt these ideas to the setting of families of maps exhibiting homoclinic tangencies.
-To formally establish the codimension-one stability of the Cantor set, we introduce the notion of infinitely renormalizable maps (see Subsection 10.3).
-For simplicity, we refer to this process as renormalization, although it is, strictly speaking, a pre-renormalization.
-In particular, we study the first-return maps to the sets generating the dynamical partition, without performing any rescaling.
-Since our main focus lies on topological properties, this pre-renormalization suffices for our purposes.
-To the best of our knowledge, this represents the first attempt to formulate a generalized renormalization framework with two branches in the setting of two-dimensional systems.
-A formulation of renormalization for two-dimensional systems using only one branch has been explored in [8, 14, 16, 17, 36], but our construction differs in that it employs two branches and more closely resembles generalized renormalization as used for one-dimensional systems.
-Let us go back to analyze the statements in Theorem A.
-In particular, one should compare these results with those of [14], where the period-doubling Cantor attractor was shown to be codimension-1 stable.
-In that case, the topology of the Cantor set also persists along a codimension-1 leave; however, the constructed leave is very short.
-Numerical experiments suggest that the period-doubling locus should, in fact, form a long global curve, but proving this remains a challenging open problem.
-In contrast, our approach allows for the construction of global curves that are genuine manifolds of stability.
-We next strengthen these results by showing the existence of a sub-lamination such that the corresponding non-hyperbolic invariant Cantor sets contain a Collet–Eckmann point with a dense orbit in the Cantor set.
-These sets, which exhibit chaotic dynamics, are likewise codimension-1 stable, i.e.
-their topology is preserved along the leaves of the lamination.
-This result represents an initial step toward the study of stability of strange attractors.
-To the best of our knowledge, no prior results are available in this direction.
-The proof of the theorem relies on the dynamical structure described above, together with a detailed analytical control along the tangency loci.
+In particular, any two maps lying on the same leaf of the lamination are conjugate on their corresponding Cantor sets, so that the dynamical structure is preserved along the leaf.
+This stability is striking given the wildness of these Cantor sets.
+Indeed, they are Newhouse-unstable and support only a trivial invariant probability measure, namely the Dirac mass at a saddle point.
+In this sense, Theorem A already reveals a first form of stability for dynamics with wild properties.
+These results should be compared with those of [17], where the period-doubling Cantor attractor was shown to be codimension-1 stable.
+In that case, the topology of the Cantor set likewise persists along a codimension-1 leaf; however, the constructed leaf is very short.
+Numerical experiments suggest that the period-doubling locus should in fact form a long global curve, but establishing this rigorously remains a challenging open problem.
+By contrast, our approach exploits the unstable expansion at the saddle point to construct a lamination in parameter space whose leaves have uniform size and act as manifolds of stability.
+The underlying idea behind the proof of Theorem A is a renormalization scheme, which can be understood as an iterative procedure based on induced maps.
+More precisely, one studies the dynamics by restricting the map to carefully chosen regions of phase space and considering the map induced by the first return to these regions.
+These induced maps describe the behaviour of the system at progressively smaller scales.
+Renormalization is defined by iterating this construction: starting from the original map, one passes to an induced map on a suitable domain, rescales it, and then repeats the same induced-map procedure.
+This iteration thus defines a renormalization operator acting on the space of maps.
+The regions selected at successive steps form a nested sequence that progressively refines the phase space.
+This nested structure gives rise to what are known as dynamical partitions, which encode the combinatorial and geometric organization of the dynamics across scales.
+In our study, we focus primarily on the analysis of these dynamical partitions and rely less on the associated renormalization operator itself.
+The combinatorics of our renormalization scheme is intrinsically unbounded.
+In such unbounded settings, a central difficulty is to control the geometry of the partitions and to obtain uniform estimates on how the relevant quantities vary with the parameters.
+This difficulty is overcome in Proposition 4.2.2 and its proof.
+Roughly speaking, this result provides quantitative control on how certain distinguished points (critical values) move with the parameters, and how the corresponding dynamical partitions are shaped.
+While the proof of Proposition 4.2.2 is very technical, it contains the core of our analysis.
+The combinatorics of the renormalization scheme and the ideas underlying Proposition 4.2.2 are inspired by [4].
+A fundamental difference, however, is that we only need to control a linearly growing number of critical points, whereas in [4] an exponentially growing number of such points must be handled.
+This reduced combinatorial complexity allows the dynamics to be organized in a more rigid and transparent way, which enables the constructions described below.
+The lamination described in Theorem A consists of infinitely renormalizable maps (see Subsection 10.3 for a formal definition).
+As explained above, the renormalization procedure associated with such maps produces a sequence of dynamical partitions at finer and finer scales.
+The intersection of these nested partitions is an invariant Cantor set.
+Rather than studying the dynamics on the invariant Cantor set directly, we encode it using a combinatorial model based on the dynamical partitions.
+To this end, we associate a directed graph to each level of the renormalization procedure.
+The vertices of the graph correspond to the elements of the dynamical partition at that level, and the edges record how these elements are mapped onto one another by the dynamics.
+In this way, each graph provides a finite combinatorial description of the dynamics at a given scale.
+The inverse limit of these graphs carries a natural dynamics and is a Cantor set.
+This dynamical system is conjugate to the dynamics induced by the original map on its invariant Cantor set.
+As a result, this combinatorial description allows effective control of the topological, geometric, and measure-theoretical properties of the Cantor sets.
+Sequences of directed graphs of this type were introduced in [32].
+In the present work, we extend this construction to the setting of families of maps exhibiting homoclinic tangencies.
+We would like to stress at this point that our approach represents the first successful attempt to formulate a generalized renormalization framework with two branches in the setting of two-dimensional systems with positive entropy.
+Renormalization schemes for two-dimensional systems using only one branch have been explored in [1, 9, 17, 20, 21, 22, 23, 27, 28, 29, 30, 31, 33, 45, 46, 51].
+Our construction differs in that it employs two branches and, in this respect, is closer in spirit to the generalized renormalization frameworks developed for one-dimensional systems.
+We next extend the list of properties of the Cantor sets described in Theorem A.
+In particular, we show that these non-hyperbolic invariant Cantor sets exhibit chaotic dynamics, in the sense that they contain a Collet–Eckmann point with dense orbit.
+Moreover, these sets remain codimension-1 stable, and their Collet–Eckmann points persist continuously along the leaves of the lamination.
+This shows that strong chaotic behaviour, manifested by exponential growth along a dense orbit, is compatible with a robust topological structure in parameter space.
+Theorem B thus represents a first step toward the study of the stability of chaotic non-hyperbolic invariant sets.
+At the present stage, however, since our analysis controls only a linearly growing number of critical points, it would be premature to formulate a general conjecture concerning the stability of strange attractors.
 Theorem B.
 There exists a sub-lamination $\mathfrak{L}_{CE}\subset\mathfrak{L}$ satisfying the following property.
-For each function $f\in\ell\subset\mathfrak{L}_{CE}$, the invariant non-hyperbolic Cantor set $A(f)$ contains a Collet-Eckmann point $c$ such that $\omega(c)=A(f)$.
-Moreover the closure $\overline{\mathfrak{L}_{CE}}=\overline{\mathfrak{L}}$.
-Theorem B is proved in Theorem 7.
+For each leaf $\ell\subset\mathfrak{L}_{CE}$ and for each map $f\in\ell$, the invariant non-hyperbolic Cantor set $A(f)$ contains a Collet-Eckmann point $c_{f}$ satisfying the following properties.
+  • The omega-limit set $\omega(c_{f})=A(f)$.
+  • The point $c_{f}$ depends continuously on $f\in\mathfrak{L}_{CE}$.
+[…]
+Moreover, every leaf of $\mathfrak{L}_{CE}$ is the graph of a smooth function $\ell:[-t_{0},t_{0}]\mapsto[-a_{0},a_{0}]$ and if $t\in[-t_{0},t_{0}]$, the closure of the set $L^{CE}_{t}=\left\{(t,a)\in\mathfrak{L}_{CE}\right\}$ is a Cantor set.
+Theorem B is a reformulation of Theorem 12.0.2.
 In pursuing the stability of chaotic non-hyperbolic sets, we encounter a striking coexistence phenomenon that further illustrates the richness of the dynamics in these systems.
-We prove the existence of a sub-lamination in which each leaf contains a map with infinitely many sinks of increasing period, accumulating on a non-hyperbolic Cantor set that contains a Collet–Eckmann point.
-This coexistence of infinitely many sinks with chaotic dynamics is both unusual and unexpected.
+We prove that each leaf of the lamination contains a map with infinitely many sinks of increasing period.
+Moreover these periodic sinks accumulate on the non-hyperbolic Cantor set that contains a Collet–Eckmann point.
+The coexistence of infinitely many sinks with chaotic dynamics is surprising.
 The formal statement of the theorem is given below.
@@
 There exists a sub-lamination $\mathfrak{L}_{N}\subset\mathfrak{L}_{CE}\subset\mathfrak{L}$ satisfying the following property.
-Each leave $\ell\subset\mathfrak{L}_{N}$ contains a map $f_{\ell}$ with infinitely many attracting periodic points, $p_{k}$ of arbitrarily high period, such that,
-$\bigcap_{n}\overline{\bigcup_{k\geq n}\text{Orb}(p_{k})}=A(f_{\ell}).$
-Moreover if $NC=\left\{f_{\ell}\right\}_{f_{\in}\ell\subset\mathfrak{L}_{N}}$, then the closure $\overline{{NC}}=\overline{\mathfrak{L}_{CE}}=\overline{\mathfrak{L}}$.
-Theorem C is proved in Theorem 8.
-The proof relies on an intermediate result, Theorem 3, which establishes the existence of maps with infinitely many sinks in families of unfoldings.
-Maps exhibiting the Newhouse phenomenon have been constructed in various contexts.
-In particular, numerous examples of Baire [^1] A Baire set in a locally compact Hausdorff space X X is a countable intersection of open and dense subsets.
-sets of maps with infinitely many sinks in the space of dynamical systems are known; see, for instance, [1, 9, 13, 11, 15, 18, 19, 24, 25, 29, 33, 34, 41, 42].
-In the classical Newhouse construction, the coexistence of infinitely many sinks is produced by the persistence of homoclinic tangencies.
-The underlying mechanism relies on the intersection of stable and unstable Cantor sets whose thicknesses exceed one, ensuring that tangencies survive under perturbation.
-Our approach is fundamentally different.
-To prove Theorem C, we avoid using the persistence of tangencies altogether.
-Instead, we introduce a quantitative condition (2.2) that involves only the eigenvalues of a saddle point.
-This spectral condition replaces the geometric thickness criterion of Newhouse and provides an explicit, verifiable rule for when the construction can be iterated.
-It is worth noting that, in the Hénon family, the set of maps satisfying the thickness condition and the set satisfying condition (2.2) intersect, but neither is contained in the other.
-The proof proceeds inductively: starting from a map with a sink and a homoclinic tangency, we show that the sink persists under small parameter changes, while the tangency gives rise to a new tangency and an additional sink.
-Iterating this mechanism yields infinitely many sinks accumulating on a non-hyperbolic Cantor set containing a Collet–Eckmann point.
-The most significant difference between our approach and Newhouse’s original method [33] lies in the creation of the secondary tangency.
-In Newhouse’s construction, the persistence of tangencies guarantees that every nearby map possesses a tangency somewhere, but this dependence on the map is highly discontinuous, making the analysis delicate.
-Our method—inspired by [3, 5]—replaces this mechanism with the use of critical points and binding points, thereby avoiding any dependence on the persistence of tangencies.
-Thus, our eigenvalue-based and constructive method offers an alternative route to the Newhouse phenomenon and clarifies how coexistence of chaos and stability can occur.
+Each leaf $\ell\subset\mathfrak{L}_{N}$ contains a map $f_{\ell}$ with infinitely many attracting periodic points $p_{g}$, of arbitrarily high period, such that,
+$\bigcap_{n}\overline{\bigcup_{g\geq n}\text{Orb}(p_{g})}=A(f_{\ell}).$
+Moreover, every leaf of $\mathfrak{L}_{N}$ is the graph of a smooth function $\ell:[-t_{0},t_{0}]\mapsto[-a_{0},a_{0}]$ and if $t\in[-t_{0},t_{0}]$, the closure of the set $L^{N}_{t}=\left\{(t,a)\in\mathfrak{L}_{N}\right\}$ is a Cantor set.
+Theorem C is a reformulation of Theorem 12.0.3.
+For a map with infinitely many sinks, it is natural to expect that each sink arises from a local return map which, after rescaling, is Hénon-like.
+The creation of further sinks at smaller scales therefore requires that these return maps be parameter unstable, as stability under parameter variations would prevent the appearance of new sinks.
+Since parameter instability is often associated with dynamical instability of the underlying system, this leads to the following conjecture.
+Conjecture.
+Let $f$ be a diffeomorphism of a compact two-dimensional manifold with infinitely many attracting periodic points $p_{g}$, of arbitrarily high period. If the set
+$A=\bigcap_{n}\overline{\bigcup_{g\geq n}\text{Orb}(p_{g})},$
+is transitive then it contains a Collet-Eckmann point with dense orbit.
+The proof of Theorem C relies on an intermediate result, Theorem 6.2.7, which establishes the existence of maps with infinitely many sinks in unfoldings.
+The coexistence of infinitely many attracting periodic orbits is known as the Newhouse phenomenon, first discovered by Newhouse in [48].
+Since its discovery, the Newhouse phenomenon has been observed in a wide range of settings.
+In particular, many constructions yield Baire sets of maps with infinitely many sinks in the space of dynamical systems; see, for instance, [2, 10, 13, 16, 18, 24, 25, 36, 37, 42, 48, 49, 56, 57].
+These results demonstrate that the phenomenon is abundant from a topological point of view, although its detailed dynamical structure often remains poorly understood.
+In Newhouse’s original construction, the coexistence of infinitely many sinks is produced by the persistence of homoclinic tangencies.
+The key mechanism relies on the interaction of stable and unstable Cantor sets whose thicknesses exceed one, which guarantees that tangencies persist under perturbations of the system.
+This persistence, in turn, allows the repeated creation of sinks.
+Our proof of Theorem C follows a different, more explicit strategy.
+The argument proceeds inductively.
+Starting from a map with a sink and a homoclinic tangency, we show that the sink persists under small parameter variations, while the tangency gives rise to a new, explicitly constructed tangency, called secondary tangency, and an additional sink.
+The construction of the secondary tangency is intrinsically related to the renormalization scheme.
+Repeating this procedure produces infinitely many sinks of increasing period, which accumulate on a non-hyperbolic Cantor set containing a Collet–Eckmann point.
+The main difference between our approach and Newhouse’s original method [48] lies precisely in the treatment of the secondary tangency.
+In the classical construction, the persistence of tangencies guarantees that nearby maps exhibit tangencies somewhere, but the location of these tangencies depends highly discontinuously on the map, making a detailed analysis difficult.
+By contrast, our method, inspired by [4, 6], replaces persistence arguments with an explicit construction of the next tangency at each step.
+This constructive approach provides an alternative route to the Newhouse phenomenon and clarifies how the coexistence of chaotic dynamics and stable periodic behaviour can arise within the same system.
+Moreover, the explicit nature of the construction allows for a precise geometric analysis in both parameter space and phase space.
+To the best of our knowledge, this is the first time that the asymptotic dynamics of Newhouse sinks accumulating on a chaotic non-hyperbolic set is described in such detail.
 Acknowledgements.
 The authors express their gratitude to Michael Benedicks for his continued support throughout the preparation of this paper and for the many invaluable discussions at both the inspirational and technical levels.
+The authors are also grateful to Michael Lyubich for shedding light on aspects of our results that we had previously overlooked, and for his valuable insights connecting the different parts of this work.
 The second author was partially supported by the starting grant SFB 67575 and by the VR grant 67578, Renormalization in dynamics.
@@
 To define the families of maps that will concern us, we first need a suitable linearization theorem. The following well-known result, due to Sternberg (see [54]), provides exactly this.
-Theorem 1.
+Theorem 2.0.1.
 Given $\left(\mu,\lambda\right)\in\mathbb{R}^{2}$, there exists $k\in{\mathbb{N}}$ such that the following holds.
@@
 If
-$\lambda\neq\mu^{k_{1}}\text{ and }\mu\neq\lambda^{k_{2}}$ (2.1)
+$\lambda\neq\mu^{k_{1}}\text{ and }\mu\neq\lambda^{k_{2}}$ (2.0.2)
 for $k=(k_{1},k_{2})\in{\mathbb{N}}^{2}$ and $2\leq|k|=k_{1}+k_{2}\leq k$, then $f$ is ${\mathcal{C}}^{4}$} linearizable.
-Definition 1.
+Definition 2.0.3.
 Let $\mathcal{M}$ be an two dimensional $\mathcal{C}$$^{\infty}$} manifold and let $f:\mathcal{M}\to\mathcal{M}$ be a diffeomorphism with saddle point $p\in\mathcal{M}$.
-We say that $p$ satisfies the ${\mathcal{C}}^{4}$} non-resonance condition if (2.1) holds.
-Theorem 2.
+We say that $p$ satisfies the ${\mathcal{C}}^{4}$} non-resonance condition if (2.0.2) holds.
+Theorem 2.0.4.
 Let $\mathcal{M}$ be a two dimensional $\mathcal{C}$$^{\infty}$} manifold and let $f:\mathcal{M}\to\mathcal{M}$ be a diffeomorphism with a saddle point $p\in\mathcal{M}$ which satisfies the ${\mathcal{C}}^{4}$} non-resonance condition.
 Let $0\in\mathcal{P}\subset\mathbb{R}^{n}$ and let $F:\mathcal{P}\times\mathcal{M}\to\mathcal{M}$ be a $\mathcal{C}$$^{\infty}$} family with $F_{0}=f$.
-Then, there exists a neighborhood $U$ of $p$ and a neighborhood $V$ of 0 such that, for every $t\in V$, $F_{t}$ has a saddle point $p_{t}\in U$ satisfying the ${\mathcal{C}}^{4}$} non-resonance condition.
+Then, there exists a neighborhood $U$ of $p$ and a neighborhood $V$ of $0\in\mathcal{P}$ such that, for every $t\in V$, $F_{t}$ has a saddle point $p_{t}\in U$ satisfying the ${\mathcal{C}}^{4}$} non-resonance condition.
 Moreover $p_{t}$ is ${\mathcal{C}}^{4}$} linearizable in the neighborhood $U$ and the linearization depends ${\mathcal{C}}^{4}$} on the parameters.
-The proofs of Theorem 1 and Theorem 2 can be found in [12, 26].
-2.1 Families of unfoldings
+The proofs of Theorem 2.0.1 and Theorem 2.0.4 can be found in [14, 38].
+2.1 Unfoldings
 In the sequel we define a map with a strong homoclinic tangency by listing several conditions.
@@
 Moreover to implement the construction, a condition on the eigenvalues is also required, see $(f2)$.
-The basic properties of a map with a strong homoclinic tangency are summarized in Figure 1.
-Definition 2.
+The basic properties of a map with a strong homoclinic tangency are illustrated in Figure 1.
+Definition 2.1.1.
 Let $\mathcal{M}$ be a two dimensional $\mathcal{C}$$^{\infty}$} manifold and let $f:\mathcal{M}\to\mathcal{M}$ be a local diffeomorphism satisfying the following conditions:
@@
 Figure 1: A map with a strong homoclinic tangency
-Remark 1.
+Remark 2.1.2.
 The conditions defining a map with a strong homoclinic tangency are natural, except $(f2)$.
@@
 It plays a crucial role in many fundamental places towards the paper.
-Remark 2.
-Observe that all conditions are open in the space of maps with an homoclinic and transversal tangency. Also, except for $(f2)$, all conditions are dense.
-Remark 3.
+Remark 2.1.3.
+Observe that all conditions are open in the space of maps with an homoclinic and transversal tangency.
+Remark 2.1.4.
 If the unstable eigenvalue is negative, $\mu<-1$, then $(f6)$, $(f7)$, and $(f8)$ are redundant.
-Following [35], we construct now a family of unfolding of a map $f$ with a homoclinic tangency and a transversal intersection.
+Following [50], we consider now an unfolding of a map $f$ with a homoclinic tangency and a transversal intersection.
 Let $\mathcal{P}=[-r,r]\times[-r,r]$ with $r>0$.
@@
     let $\mu_{\text{max}}=\max_{(t,a)}|\mu(t,a)|$, $\lambda_{\text{max}}=\max_{(t,a)}|\lambda(t,a)|$ and assume
-    $\lambda_{\text{max}}\mu_{\text{max}}^{3}<1,$ (2.2)
+    $\lambda_{\text{max}}\mu_{\text{max}}^{3}<1,$ (2.1.5)
   • $(F4)$
@@
     there exists a ${\mathcal{C}}^{2}$} function $r,r]^{2}\ni(t,a)\mapsto q_{2}(t,a)\in W^{u}(p(t,a))\cap W^{s}(p(t,a))$ such that $q_{2}(t,a)$ is a transversal homoclinic intersection of $F_{t,a}$.
-Remark 4.
+Remark 2.1.6.
 Without loss of generality, we may assume that
@@
 Observe that, by $(F3)$ and by continuity, for $t_{0}$ small enough,
-$0<\frac{\log\mu_{\text{max}}}{\log{\frac{1}{\lambda_{\text{max}}}}}<\frac{3}{2}\frac{\log\mu_{\text{min}}}{\log{\frac{1}{\lambda_{\text{min}}}}}<\frac{1}{2}.$ (2.3)
+$0<\frac{\log\mu_{\text{max}}}{\log{\frac{1}{\lambda_{\text{max}}}}}<\frac{3}{2}\frac{\log\mu_{\text{min}}}{\log{\frac{1}{\lambda_{\text{min}}}}}<\frac{1}{2}.$ (2.1.7)
 2.1.1 Rescaled Families
-According to Theorem 2 we may make a change of coordinates to ensure that the family $F$ is ${\mathcal{C}}^{4}$} and for all $(t,a)\in[-r_{0},r_{0}]\times[-r_{0},r_{0}]$ with $0<r_{0}<r$ and by rescaling we can assume that $F_{t,a}$ is linear on the ball $2,2]^{2}$, namely
-$F_{t,a}=\left(\begin{matrix}\lambda(t,a)&0\\ 0&\mu(t,a)\\ \end{matrix}\right).$ (2.4)
+According to Theorem 2.0.4 we may make a change of coordinates to ensure that the family $F$ is ${\mathcal{C}}^{4}$} and for all $(t,a)\in[-t_{0},t_{0}]\times[-a_{0},a_{0}]$ with $0<t_{0},a_{0}<r$ and by rescaling we can assume that $F_{t,a}$ is linear on the ball $2,2]^{2}$, namely
+$F_{t,a}=\left(\begin{matrix}\lambda(t,a)&0\\ 0&\mu(t,a)\\ \end{matrix}\right).$ (2.1.8)
 Moreover the saddle point $p(t,a)=(0,0)$ and the local stable and unstable manifolds satisfy:
@@
 $(X_{t,a}(x,y),Y_{t,a}(x,y))=F^{N}_{t,a}(x,y).$
-The following lemma says that $F^{N}_{t,a}(x,y)$ produces an unfolding in the sense of [35]. The formal definition of unfolding is given in Definition 4 below.
-Lemma 1.
+The following lemma says that $F^{N}_{t,a}(x,y)$ produces an unfolding in the sense of [50]. The formal definition of unfoldings is given in Definition 2.1.14 below.
+Lemma 2.1.9.
 There exist $x_{0},a_{0}>0$, a ${\mathcal{C}}^{2}$} function $c:[-x_{0},x_{0}]\times[-t_{0},t_{0}]\times[-a_{0},a_{0}]\to\mathbb{R}$ and a positive constant $Q$ such that
-$\frac{\partial Y_{t,a}}{\partial y}\left(x,c(x,t,a)\right)=0,$ (2.5)
+$\frac{\partial Y_{t,a}}{\partial y}\left(x,c(x,t,a)\right)=0,$ (2.1.10)
 and
 $\frac{\partial^{2}Y_{t,a}}{\partial y^{2}}\left(x,c(x,t,a)\right)\geq Q.$
-Moreover
-$|c(x,t,a)-c(0,t,a)|=O\left(|x|\right).$ (2.6)
+Moreover,
+$|c(x,t,a)-c(0,t,a)|=O\left(|x|\right).$ (2.1.11)
 Proof.
@@
 ∎
-Remark 5.
+Remark 2.1.12.
 Without lose of generality, by a smooth change of coordinates in the $y$ direction, we may assume that $c(0,t,a)=1$.
-Definition 3.
+Definition 2.1.13.
 Let $\left(t,a\right)\in[-t_{0},t_{0}]\times[-a_{0},a_{0}]$. We call the point
@@
 This is another instance where, from a technical point of view, the notion of critical point is not that important.
-Definition 4.
-A family $F_{t,a}$ is called an unfolding of $f$ if it can be reparametrized such that
+A similar definition of critical point in a more general setting is given in [22].
+Definition 2.1.14.
+A family $F_{t,a}$ is called an unfoldings of $f$ if it can be reparametrized such that
   • $(P1)$
@@
 Figure 2: Unfolding
-Remark 6.
-Without lose of generality by a suitable coordinate change we may assume that if $F$ is an unfolding then $z_{y}(t,a)=a$, the primary critical value is at height $a$ and the primary critical point $c(t,a)=(0,1)$, see Figure 2.
-Remark 7.
-A generic two-dimensional family trough f can locally be reparametrized to become an unfolding.
-Remark 8.
+Remark 2.1.15.
+Without lose of generality by a suitable coordinate change we may assume that if $F$ is an unfolding then $z_{y}(t,a)=a$, the primary critical value is at height $a$, $z_{x}(t,a)=1$ and the primary critical point $c_{t,a}=(0,1)$, see Figure 2.
+Remark 2.1.16.
+A generic two-dimensional family trough f can locally be reparametrized to become an unfoldings.
+Remark 2.1.17.
 The results and the techniques presented in this paper can be generalized to higher dimensional families, both in parameter and phase space.
@@
 Moreover, because
-$W_{n}(x)=\mu^{-n}W\left(\lambda^{n}x\right),$ (2.7)
+$W_{n}(x)=\mu^{-n}W\left(\lambda^{n}x\right),$ (2.2.1)
 then
-$\frac{1}{2\mu}\mu^{-n}\leq|W_{n}|\leq 2\mu^{-n},$ (2.8)
-$\left|\frac{\partial W_{n}}{\partial x}\right|=O\left(\frac{\lambda}{\mu}\right)^{n},$ (2.9)
+$\frac{1}{2\mu}\mu^{-n}\leq|W_{n}|\leq 2\mu^{-n},$ (2.2.2)
+$\left|\frac{\partial W_{n}}{\partial x}\right|=O\left(\frac{\lambda}{\mu}\right)^{n},$ (2.2.3)
 and
-$\left|\frac{\partial^{2}W_{n}}{\partial x^{2}}\right|=O\left(\frac{\lambda^{2}}{\mu}\right)^{n}.$ (2.10)
+$\left|\frac{\partial^{2}W_{n}}{\partial x^{2}}\right|=O\left(\frac{\lambda^{2}}{\mu}\right)^{n}.$ (2.2.4)
 We estimate now the speed of the stable manifold $W_{n}$ as the graph of a function in the phase space.
-Lemma 2.
+Lemma 2.2.5.
 $\frac{\partial W_{n}}{\partial t}=-\frac{n}{\mu}\frac{\partial\mu}{\partial t}W_{n}+\left(\frac{\lambda}{\mu}\right)^{n}\frac{n}{\lambda}\frac{\partial\lambda}{\partial t}\frac{\partial W}{\partial x}x+\frac{1}{\mu^{n}}\frac{\partial W}{\partial t}=O\left(\frac{n}{\mu^{n}}\right),$
@@
 We denote by $W_{n}$ the manifold corresponding to $(t,a)$ and by $W_{n}+\Delta W_{n}$ the manifold to $(t,a+\Delta a)$.
-Then, since the maps $F^{n}_{t,a}$ are linear, by differentiating (2.7) we obtain
+Then, since the maps $F^{n}_{t,a}$ are linear, by differentiating (2.2.1) we obtain
 $\left(\mu+\frac{\partial\mu}{\partial a}\Delta a\right)^{n}\left(W_{n}+\Delta W_{n}\right)(x)=\mu^{n}W_{n}(x)+\frac{\partial W}{\partial x}\left(\left(\left(\lambda+\frac{\partial\lambda}{\partial a}\Delta a\right)^{n}-\lambda^{n}\right)x\right)+\frac{\partial W}{\partial a}\Delta a.$
@@
 In the next lemma we build a curve $a_{n}$ of points, in the parameter space, whose corresponding critical values are mapped after $n$ steps into $\Gamma_{t,a}$.
-Lemma 3.
+Lemma 2.2.6.
 For $n$ large enough, there exists a ${\mathcal{C}}^{2}$} function $a_{n}:[-t_{0},t_{0}]\to(0,a_{0}]$ such that
@@
 Moreover
-$\frac{da_{n}}{dt}=-n\frac{\partial\mu}{\partial t}\frac{1}{\mu^{n+1}}\left[1+O\left(\lambda^{n}\right)\right].$ (2.11)
+$\frac{da_{n}}{dt}=-n\frac{\partial\mu}{\partial t}\frac{1}{\mu^{n+1}}\left[1+O\left(\lambda^{n}\right)\right].$ (2.2.7)
 Proof.
@@
 $\Gamma=\left\{\left(x,c(x,t,a),t,a\right)|\left(x,t,a\right)\in[-x_{0},x_{0}]\times[-t_{0},t_{0}]\times[-a_{0},a_{0}]\right\}.$
-Then $\Gamma$ is a ${\mathcal{C}}^{2}$} codimension $1$ manifold transversal to $W^{u}(0)$. For $n\geq 0$ let
-$\Gamma_{n}=\left\{\left(x,y,t,a\right)|F^{n}_{t,a}\left(x,y\right)\in\Gamma\right\},$ (2.12)
+Then $\Gamma$ is a ${\mathcal{C}}^{2}$} codimension-11 manifold transversal to $W^{u}(0)$. For $n\geq 0$ let
+$\Gamma_{n}=\left\{\left(x,y,t,a\right)|F^{n}_{t,a}\left(x,y\right)\in\Gamma\right\},$ (2.2.8)
 and the limit of the $\Gamma_{n}$:s as $n\to\infty$ is,
@@
 where $z_{t,a}=\left(z_{x}(t,a),z_{y}(t,a)\right)$. Observe that
-$\frac{\partial z_{y}}{\partial a}=1\text{ and }z_{y}(t,0)=0.$ (2.13)
+$\frac{\partial z_{y}}{\partial a}=1\text{ and }z_{y}(t,0)=0.$ (2.2.9)
 Let $Z=\text{Image}(z)$.
-Because of (2.13), $Z$ is a manifold transversal to $\Gamma_{\infty}$.
+Because of (2.2.9), $Z$ is a manifold transversal to $\Gamma_{\infty}$.
 Hence, there exists $n_{1}>0$ such that, for all $n\geq n_{1}$, $Z$ is transversal to $\Gamma_{n}$.
@@
 $A_{n}=z^{-1}\left(\Gamma_{n}\right)$
-is a ${\mathcal{C}}^{2}$} codimension $1$ manifold.
+is a ${\mathcal{C}}^{2}$} codimension-11 manifold.
 We define $a_{n}:t\mapsto a_{n}(t)$ as a function whose graph is $A_{n}$.
-Observe that, by Lemma 1 and Remark 5, the ${\mathcal{C}}^{2}$} function $c$ satisfies, $c(0,t,a)=1$.
+Observe that, by Lemma 2.1.9 and Remark 2.1.12, the ${\mathcal{C}}^{2}$} function $c$ satisfies, $c(0,t,a)=1$.
 Hence $\partial c/\partial t=O(x)$ and $\partial c/\partial a=O(x)$.
 Recall that
-$\mu^{n}a_{n}=c\left(\lambda^{n}z_{x},t,a_{n}\right).$ (2.14)
-By differentiating (2.14) and using $\partial c/\partial a,\partial c/\partial t=O(x)$ we have
+$\mu^{n}a_{n}=c\left(\lambda^{n}z_{x},t,a_{n}\right).$ (2.2.10)
+By differentiating (2.2.10) and using $\partial c/\partial a,\partial c/\partial t=O(x)$ we have
 $\mu^{n}\frac{da_{n}}{dt}+na_{n}\mu^{n-1}\frac{\partial\mu}{\partial t}=O\left(\lambda^{n}\right).$
 The lemma follows. ∎
-A proof similar to that of Lemma 2.7 gives the following.
-Lemma 4.
-Let $\Gamma_{n}$ be as in (2.12), then
+A proof similar to that of Lemma 2.2.1 gives the following.
+Lemma 2.2.11.
+Let $\Gamma_{n}$ be as in (2.2.8), then
 $\frac{\partial\Gamma_{n}}{\partial t}=-\frac{n}{\mu}\frac{\partial\mu}{\partial t}\Gamma_{n}+\left(\frac{\lambda}{\mu}\right)^{n}\frac{n}{\lambda}\frac{\partial\lambda}{\partial t}\frac{\partial\Gamma}{\partial x}x+\frac{1}{\mu^{n}}\frac{\partial\Gamma}{\partial t}=O\left(\frac{n}{\mu^{n}}\right),$
 $\frac{\partial\Gamma_{n}}{\partial a}=-\frac{n}{\mu}\frac{\partial\mu}{\partial a}\Gamma_{n}+\left(\frac{\lambda}{\mu}\right)^{n}\frac{n}{\lambda}\frac{\partial\lambda}{\partial a}\frac{\partial\Gamma}{\partial x}x+\frac{1}{\mu^{n}}\frac{\partial\Gamma}{\partial a}=O\left(\frac{n}{\mu^{n}}\right).$
-Exact form of the map $F^{N}$.
-For families of unfoldings the following holds.
-Lemma 5.
+Form of the map $F^{N}$.
+For unfoldings the following holds.
+Lemma 2.2.12.
 There exist $x^{\prime}_{0}<x_{0}$, $a^{\prime}_{0}<a_{0}$, $b>0$ and $Q>0$, such that, for all $(t,a)\in[-t_{0},t_{0}]\times[-a^{\prime}_{0},a^{\prime}_{0}]$ and for every $(x,y)\in\Gamma_{t,a}$ with $|x|<|x^{\prime}_{0}|$ the following holds.
-There exist matrices $A_{x,y}$, $B_{x,y}\neq 0$ and $C_{x,y}$ such that $F^{N}_{t,a}$ in coordinates centered in $(x,y)$ and $F^{N}_{t,a}(x,y)$ has the form
-$F^{N}_{t,a}\left(\begin{matrix}\Delta x\\ \Delta y\end{matrix}\right)$ $=$ $\left(\begin{matrix}A_{x,y}&B_{x,y}\\ C_{x,y}&0\end{matrix}\right)\left(\begin{matrix}\Delta x\\ \Delta y\end{matrix}\right)$ (2.15)
+There exist ${\mathcal{C}}^{3}$} functions $A_{x,y}$, $B_{x,y}\neq 0$ and $C_{x,y}$ such that $F^{N}_{t,a}$ in coordinates centered in $(x,y)$ and $F^{N}_{t,a}(x,y)$ has the form
+$F^{N}_{t,a}\left(\begin{matrix}\Delta x\\ \Delta y\end{matrix}\right)$ $=$ $\left(\begin{matrix}A_{x,y}&B_{x,y}\\ C_{x,y}&0\end{matrix}\right)\left(\begin{matrix}\Delta x\\ \Delta y\end{matrix}\right)$ (2.2.13)
 $+$ $\left(\begin{matrix}O_{1,1}\Delta x^{2}+O_{1,2}\Delta x\Delta y+O_{1,3}\Delta y^{2}\\ Q_{x,y}\Delta y^{2}+O_{2,1}\Delta x^{2}+O_{2,2}\Delta x\Delta y+O_{2,3}\Delta y^{3}\end{matrix}\right),$
-where the matrices $A_{x,y}$, $C_{x,y}$ and the vector $B_{x,y}$ are ${\mathcal{C}}^{2}$} dependent on $x$ and $y$, $Q_{x,y}>Q$, $|B_{x,y}|>b>0$ and the vector valued functions $O_{i,j}$ are ${\mathcal{C}}^{2}$} dependent on $x$ and $y$ and they are uniformly bounded.
+where $Q_{x,y}>Q$, $|B_{x,y}|>b>0$ and the ${\mathcal{C}}^{2}$} functions $O_{i,j}$ are uniformly bounded.
 Proof.
 The lemma gives the Taylor expansions of $F^{N}_{t,a}$ when $(x,y)\in\Gamma_{t,a}$.
-It follows immediately from Lemma 1 where we state that a vertical curve trough $(x,y)$ in $\Gamma_{t,a}$ is mapped to a curve with a non degenerate horizontal tangency.
-In particular $\partial Y_{t,a}/\partial y=0$ by (2.5) and the horizontal tangency,
+It follows immediately from Lemma 2.1.9 where we state that a vertical curve trough $(x,y)$ in $\Gamma_{t,a}$ is mapped to a curve with a non degenerate horizontal tangency.
+In particular $\partial Y_{t,a}/\partial y=0$ by (2.1.10) and the horizontal tangency,
 $DF^{N}_{t,a}(x,y)\left(\begin{matrix}0\\ 1\end{matrix}\right)=\left(\begin{matrix}B_{x,y}\\ 0\end{matrix}\right)$
@@
 Let $t\in[-t_{0},t_{0}]$.
-Because $F^{N}_{t,0}(q_{3}(t))=q_{1}(t)$ is a non degenerate homoclinic tangency, we know that the vector $B_{q_{3}(t)}\neq 0$ and $Q_{q_{3}(t)}>0$.
+Because $F^{N}_{t,0}(q_{3}(t))=q_{1}(t)$ is a non degenerate homoclinic tangency, we know that $B_{q_{3}(t)}\neq 0$ and $Q_{q_{3}(t)}>0$.
 By taking $|a|<|a^{\prime}_{0}|$, $|x|<|x^{\prime}_{0}|$ small enough, the lower bounds on $Q_{x,y}$ and $B_{x,y}$ follow.
@@
 2.3 Choice of $\theta$
-In this subsection, we introduce a small number $\theta$ that is going to play a fundamental role in the construction of our combinatorics.
+In this subsection, we introduce a number $\theta$ that is going to play a fundamental role in the construction of our combinatorics.
 The condition $\lambda_{\text{max}}\mu_{\text{max}}^{3}<1$, see $(F3)$, allows us to choose $\theta\in(0,\frac{1}{2})$ such that
-$1<\lambda_{\text{min}}^{2\theta}\mu_{\text{min}}^{3}\text{ and }\lambda_{\text{max}}^{3\theta}\mu^{4}_{\text{max}}<1.$ (2.16)
+$1<\lambda_{\text{min}}^{2\theta}\mu_{\text{min}}^{3}\text{ and }\lambda_{\text{max}}^{3\theta}\mu^{4}_{\text{max}}<1.$ (2.3.1)
 We choose any $\theta$ satisfying
-$0<\theta_{0}=\frac{4}{3}\frac{\log\mu_{\text{max}}}{\log{\frac{1}{\lambda_{\text{max}}}}}<\theta<\frac{3}{2}\frac{\log\mu_{\text{min}}}{\log{\frac{1}{\lambda_{\text{min}}}}}=\theta_{1}<\frac{1}{2},$ (2.17)
-where we used $(F3)$, the initial condition $\lambda_{\text{max}}\mu_{\text{max}}^{3}<1$ and (2.3).
-Moreover, by restricting the parameter domain, we choose $\theta$ such that
-$\left(\frac{\theta}{\lambda_{\text{max}}}\frac{\partial\lambda}{\partial t}+\frac{1}{\mu_{\text{max}}}\frac{\partial\mu}{\partial t}\right)>0.$ (2.18)
+$0<\theta_{0}=\frac{4}{3}\frac{\log\mu_{\text{max}}}{\log{\frac{1}{\lambda_{\text{max}}}}}<\theta<\frac{3}{2}\frac{\log\mu_{\text{min}}}{\log{\frac{1}{\lambda_{\text{min}}}}}=\theta_{1}<\frac{1}{2},$ (2.3.2)
+where we used $(F3)$, the initial condition $\lambda_{\text{max}}\mu_{\text{max}}^{3}<1$ and (2.1.7).
+Moreover, by adjusting the parameter domain, we can choose $\theta$ such that
+$\left|\frac{\theta}{\lambda}\frac{\partial\lambda}{\partial t}+\frac{1}{\mu}\frac{\partial\mu}{\partial t}\right|>w>0.$ (2.3.3)
 3 The Hénon family
@@
 a two parameter family. In this section we are going to prove that the Hénon family is an unfolding of a map with a strong homoclinic tangency and a transversal intersection.
-Proposition 1.
+Proposition 3.0.1.
 The real Hénon family is an unfolding.
@@
 Hence, there exists an analytic curve $b\mapsto a(b)$ with $a(0)=2$ such that the parameter $(a(b),b)$ corresponds to a Hénon map with an homoclinic tangency of the saddle point $p(b)$ which is a continuation of p(0)=(2,2)p(0)=(-2,-2).
-For all $b$ positive and small enough $F_{a(b),b}$ has a strong homoclinic tangency and the Hénon family $F_{a,b}$ restricted to a small neighborhood of $(a(b),b)$ is an unfolding.
+For all $b$ positive and small enough $F_{a(b),b}$ has a strong homoclinic tangency and the Hénon family $F_{a,b}$ restricted to a small neighborhood of $(a(b),b)$ is an unfoldings.
 ∎
 4 Existence of secondary tangencies
-In this section we prove the existence of secondary tangencies in a certain strip of parameters that we introduce below.
-Remember the choice of $\theta$ in (2.16) and (2.17).
-Denote by $C$ the entry of the matrix in Lemma 5 corresponding to the critical point $c$, see Definition 3.
-Let $\alpha\in(0,1)$ and let $n_{0}=[n\alpha]+O(1)$.
+Given an unfolding we create in this section curves corresponding to maps with a so-called secondary tangency.
+These curves are graphs over the $t$-axes accumulating on the primary tangency curve at $a=0$.
+Moreover they are fully contained in a well defined strip in parameter space which is defined below.
+The restriction of the initial unfolding to a neighborhood of any of these curves is again an unfolding of the new secondary tangency.
+Remember the choice of $\theta$ in (2.3.1) and (2.3.2).
+Denote by $C$ the entry of the matrix in Lemma 2.2.12 corresponding to the critical point $c$, see Definition 2.1.13.
+Observe that $C$, $c$, $\lambda$ and $\mu$ all depend on the parameters $(t,a)$.
+For simplicity of notation we omit this dependence.
+Let
+$n_{0}^{*}=\left[\frac{\log\left(\lambda^{2\theta}\mu^{3}\right)}{2\log\mu}n\right].$ (4.0.1)
 Choose $\epsilon>0$, $T$ a positive large constant and define, for large $n$,
-$\mathcal{B}_{n}=\left\{(t,a)\in[-t_{0},t_{0}]\times[-a_{0},a_{0}]\left|\right.\left(\frac{\epsilon}{\mu^{\frac{n_{0}}{2}}}-C\right)\lambda^{\theta n}\leq(a-a_{n}(t))\leq\frac{T}{\mu^{n+\frac{n_{0}}{2}}}\right\}.$ (4.1)
-The strip $\mathcal{B}_{n}$ is built to contain the curves of tangencies. The exact choice of $n_{0}$ is made in (4.16).
-Denote by $E=C-\frac{\epsilon}{\mu^{{n_{0}}/{2}}}$.
-Observe that in this definition of the strip, $\mu$ and $\lambda$ are taken in the parameter $(t,a_{n}(t))$.
-The following remark points out the distortion of the eigenvalues when changing the parameters in $\mathcal{B}_{n}$.
-Remark 9.
-[…]
-$\left(\frac{\mu(t,a_{n}(t))}{\mu(\tilde{t},\tilde{a})}\right)^{n}=1+O\left(n\lambda(t,a_{n}(t))^{\theta n}\right),\left(\frac{\lambda(t,a_{n}(t))}{\lambda(\tilde{t},\tilde{a})}\right)^{n}=1+O\left(n\lambda(t,a_{n}(t))^{\theta n}\right)$
-where we used the fact that $\partial\mu/\partial t$ and $\partial\mu/\partial a$ are bounded. Moreover if $(t,a)\in\mathcal{B}_{n}$, then
-$\frac{\mu(t,a_{n}(t))^{n}(a-a_{n}(t))}{\lambda(t,a)^{\theta n}\mu(t,a)^{n}}\geq{E},$
+$\mathcal{B}_{n}=\left\{(t,a)\in[-t_{0},t_{0}]\times[-a_{0},a_{0}]\left|\right.\left(\frac{\epsilon}{\mu^{\frac{n_{0}^{*}}{2}}}-C\right)\lambda^{\theta n}\leq(a-a_{n}(t))\leq\frac{T}{\mu^{n+\frac{n_{0}^{*}}{2}}}\right\}.$ (4.0.2)
+The strip $\mathcal{B}_{n}$ is built to contain the curves of secondary tangencies. The exact choice of $n_{0}$ is made in (4.1.46). Denote by $E=C-\frac{\epsilon}{\mu^{{n_{0}^{*}}/{2}}}$.
+Remark 4.0.3.
+Observe that, if $(t,a)\in\mathcal{B}_{n}$, then
+$\left(\frac{\mu(t,a)}{\mu(t,a_{n}(t))}\right)^{n}=1+O\left(n(a-a_{n}(t))\right),$
 and
-${\mu(t,a)^{n+\frac{n_{0}}{2}}(a-a_{n}(t))}\leq{T}.$
+$\left(\frac{\lambda(t,a)}{\lambda(t,a_{n}(t))}\right)^{n}=1+O\left(n(a-a_{n}(t))\right).$
 4.1 The combinatorics
-Fix $(t,a)\in\mathcal{B}_{n}$ and in the sequel when there is no possibility of confusion, we will suppress this choice in the notation, for example, $z=z_{t,a}$, $c=c_{t,a}$ (see Definition 3), $\lambda=\lambda(t,a)$ and $\mu=\mu(t,a)$.
+Fix $(t,a)\in\mathcal{B}_{n}$ and in the sequel when there is no possibility of confusion, we will suppress this choice in the notation, for example, $z=z_{t,a}$, $c=c_{t,a}$ (see Definition 2.1.13), $\lambda=\lambda(t,a)$ and $\mu=\mu(t,a)$.
 The first loop: $N+\theta n+N$ iterates.
@@
 $c^{\prime}\in F^{\theta n}_{t,a}\left(W^{u}_{\text{loc}}(z)\right)\cap\Gamma,$
-and a local unstable manifold $W^{u}_{\text{loc}}(c^{\prime})$ of diameter $0<L^{\frac{1}{2}}\lambda^{\frac{\theta n}{2}}$ where is a large constant that will be chosen later.
+and a local unstable manifold $W^{u}_{\text{loc}}(c^{\prime})$ of vertical size $L^{\frac{1}{2}}\lambda^{\frac{\theta n}{2}}$ where $L$ is a large constant that will be chosen later.
+The curve $W^{u}_{\text{\rm loc}}(c^{\prime})$, is a graph over the $y$-axes.
+In the following lemma we give an expression for the curve $W^{u}_{\text{\rm loc}}(c^{\prime})$.
+Lemma 4.1.1.
+In a coordinate system centered in $c^{\prime}$ the ${\mathcal{C}}^{4}$} curve $W^{u}_{\text{\rm loc}}(c^{\prime})$ is given by
+$\Delta x=O\left(\left(\frac{\lambda^{2}}{\mu}\right)^{\frac{\theta n}{2}}\right)\Delta y+O\left(\left(\frac{\lambda^{2}}{\mu}\right)^{\frac{\theta n}{2}}\right)(\Delta y)^{2},$
+where $0\leq\Delta y<L^{\frac{1}{2}}\lambda^{\frac{\theta n}{2}}$, and $\Delta x=O\left(\left({\lambda^{3}}/{\mu}\right)^{\frac{\theta n}{2}}\right)$.
+Proof.
+Use coordinates centered at the critical point $c^{\prime}$ of the parameter $(t,a)$ and let $(\Delta x_{c^{\prime}},\Delta y_{c^{\prime}})\in W^{u}_{\text{loc}}(c^{\prime})$.
+Observe that $\Delta x_{c^{\prime}}$ is a function of $\Delta y_{c^{\prime}}$.
+First we are going to prove that
+$\Delta x_{c^{\prime}}=O\left(\left(\frac{\lambda^{2}}{\mu}\right)^{\frac{\theta n}{2}}\right),$ (4.1.2)
+$\frac{d\Delta x_{c^{\prime}}}{d\Delta y_{c^{\prime}}}=O\left(\left(\frac{\lambda^{2}}{\mu}\right)^{\frac{\theta n}{2}}\right),$ (4.1.3)
+and
+$\frac{d^{2}\Delta x_{c^{\prime}}}{d\left(\Delta y_{c^{\prime}}\right)^{2}}=O\left(\left(\frac{\lambda^{2}}{\mu}\right)^{\frac{\theta n}{2}}\right).$ (4.1.4)
+Because the first coordinate of $z$ is equal to $1$, see Remark 2.1.15 , we have that
+$c^{\prime}_{y}=1+O\left(\lambda^{\theta n}\right)$ (4.1.5)
+and since $(t,a)\in\mathcal{B}_{n}$, then
+$z_{y}\leq a_{n}(t)+\frac{T}{\mu^{n+n_{0}^{*}/2}}\leq\frac{1}{\mu^{n}}\left(1+O\left(\lambda^{n}\right)\right)+\frac{T}{\mu^{n+n_{0}^{*}/2}},$
+where we used (2.2.10). Let $m$ be the highest point of $W^{u}_{\text{loc}}(c^{\prime})$. Then
+$m_{y}=1+\left(L\lambda^{\theta n}\right)^{1/2}+O\left(\lambda^{\theta n}\right)=1+O\left(\lambda^{\theta n/2}\right).$
+Let $Y$ be the vertical distance from $F^{-\theta n}(m)$ and $z$ and $X$ the horizontal one. Then, since $(t,a)\in\mathcal{B}_{n}$,
+$Y$ $=$ $\frac{1}{\mu^{\theta n}}\left(1+O\left(\lambda^{\theta n/2}\right)\right)-z_{y}$ (4.1.6)
+$\leq$ $\frac{1}{\mu^{\theta n}}\left(1+O\left(\lambda^{\theta n/2}\right)\right)-\left[\frac{1}{\mu^{n}}\left(1+O\left(\lambda^{\theta n}\right)\right)-E\lambda^{\theta n}\right]=O\left(\frac{1}{\mu^{\theta n}}\right),$
+and since $W^{u}_{\text{loc}}(z)$ is essentially a parabola, $X\leq O\left({1}/{\mu^{\theta n/2}}\right)$.
+From this and from the fact that $z_{x}=1$, see Remark 2.1.15, formula (4.1.2) follows.
+Observe that, the estimate on $X$ also gives that
+$c^{\prime}_{x}=\lambda^{\theta n}+O\left(\left(\frac{\lambda^{2}}{\mu}\right)^{\frac{\theta n}{2}}\right).$ (4.1.7)
+To prove (4.1.3), use coordinates $\left(\Delta x_{z},\Delta y_{z}\right)\in W^{u}_{\text{\rm loc}}(z)$ centered around $z$ and coordinates $\left(\Delta x_{F^{\theta n}(z)},\Delta y_{F^{\theta n}(z)}\right)\in W^{u}_{\text{\rm loc}}(c^{\prime})$ centered around $F^{\theta n}(z)$, i.e.
+$F^{\theta n}\left(\Delta x_{z},\Delta y_{z}\right)=\left(\Delta x_{F^{\theta n}(z)},\Delta y_{F^{\theta n}(z)}\right)$.
+Let $(u,v)$ be a tangent vector to $W^{u}_{\text{loc}}(z)$ at the point $\left(\Delta x_{z},\Delta y_{z}\right)$.
+Then, using the same estimates as in (4.1.6), there exist uniform constants $K_{1},K_{2}$ such that
+${K_{1}^{-1}}{\mu^{-\theta n}}\leq\Delta y_{z}\leq K_{1}{\mu^{-\theta n}},$ (4.1.8)
+and because of the quadratic behavior of $W^{u}_{\text{loc}}(z)$,
+$K_{2}^{-1}{\mu^{-\frac{\theta n}{2}}}\leq|\Delta x_{z}|\leq K_{2}{\mu^{-\frac{\theta n}{2}}}.$
+In particular
+$K_{2}^{-1}{\mu^{-\frac{\theta n}{2}}}|u|\leq|v|\leq K_{2}{\mu^{-\frac{\theta n}{2}}}|u|.$
+Observe that
+$\frac{d\Delta x_{c^{\prime}}}{d\Delta y_{c^{\prime}}}=O\left(\frac{\lambda^{\theta n}}{\mu^{\theta n}}\frac{u}{v}\right).$
+The estimate for the slope follows. This estimate also give a better bound for $\Delta x$, namely $\Delta x=O\left(\left({\lambda^{3}}/{\mu}\right)^{\frac{\theta n}{2}}\right)$.
+We are left to prove (4.1.4). Since $W^{u}_{\text{\rm loc}}(z)$ is the image of a straight line segment, there is a non zero uniformly bounded ${\mathcal{C}}^{2}$} function $R$ such that
+$\Delta x_{z}=R\left(\Delta y_{z}\right)\left(\Delta y_{z}\right)^{\frac{1}{2}},$
+and by (4.1.8), there exists a uniform constant $K_{1}$, such that
+${K_{1}^{-1}}{\mu^{-\theta n}}\leq\Delta y_{z}.$
+In particular,
+$\frac{d^{2}\Delta x_{z}}{d\left(\Delta y_{z}\right)^{2}}=O\left(\left(\Delta y_{z}\right)^{-\frac{3}{2}}\right)=O\left(\mu^{\frac{3}{2}\theta n}\right).$
+By applying the linear map $F^{\theta n}$ we have
+$\frac{d^{2}\Delta x_{c^{\prime}}}{d\left(\Delta y_{c^{\prime}}\right)^{2}}=\frac{d^{2}\Delta x_{F^{\theta n}(z)}}{d\left(\Delta y_{F^{\theta n}(z)}\right)^{2}}=O\left(\mu^{\frac{3}{2}\theta n}\left(\frac{\lambda}{\mu^{2}}\right)^{\theta n}\right)=O\left(\left(\frac{\lambda^{2}}{\mu}\right)^{\frac{\theta n}{2}}\right).$
+Finally, using (4.1.3) and (4.1.4) the curve $W^{u}_{\text{\rm loc}}(c^{\prime})$ in coordinates centered around the point $c^{\prime}$ is essentially a straight line given by
+$\Delta x_{c^{\prime}}=O\left(\left(\frac{\lambda^{2}}{\mu}\right)^{\frac{\theta n}{2}}\right)\Delta y_{c^{\prime}}+O\left(\left(\frac{\lambda^{2}}{\mu}\right)^{\frac{\theta n}{2}}\right)(\Delta y_{c^{\prime}})^{2}.$ (4.1.9)
 Let $z^{(1)}=(z_{x}^{(1)},z_{y}^{(1)})$ be the lowest point of $F^{N}\left(W^{u}_{\text{loc}}(c^{\prime})\right)$, see Figure 3.
-Figure 3: $N+\theta n$ iterates
-Lemma 6.
-For $n$ large enough and for all $(t,a)\in\mathcal{B}_{n}$,
-$C\lambda^{\theta n}+O\left(\lambda^{2\theta n}\right)\leq|z_{y}^{(1)}(t,a)-z_{y}(t,a)|\leq\left(C+{QL}\right)\lambda^{\theta n}+O\left(\lambda^{2\theta n}\right),$
-and the tangent space $T_{z^{(1)}}W^{u}_{\text{loc}}(z^{(1)})\subset\mathbb{R}^{m-1}\times\left\{0\right\}$ is horizontal.
-Moreover, the limit $\lim_{n\to\infty}{|z_{y}^{(1)}(t,a)-z_{y}(t,a)|}/{\lambda^{\theta n}}$ exists.
+Figure 3: $N+\theta n+N$ iterates
+Lemma 4.1.10.
+For all $(t,a)\in\mathcal{B}_{n}$,
+$z_{y}^{(1)}(t,a)-z_{y}(t,a)=C\lambda^{\theta n}+O\left(\left(\frac{\lambda^{2}}{\mu}\right)^{\frac{\theta n}{2}}\right).$
+In particular, the limit $\lim_{n\to\infty}{|z_{y}^{(1)}(t,a)-z_{y}(t,a)|}/{\lambda^{\theta n}}=C$. Moreover if $(x,y)\in W^{u}_{\text{loc}}(z^{(1)})$,
+$C\lambda^{\theta n}+O\left(\left(\frac{\lambda^{2}}{\mu}\right)^{\frac{\theta n}{2}}\right)\leq y-z_{y}(t,a)\leq(C+QL)\lambda^{\theta n}+O\left(\left(\frac{\lambda^{2}}{\mu}\right)^{\frac{\theta n}{2}}\right).$
+Proof.
+We need to determine the the point $c^{(1)}$ in $W^{u}_{\text{loc}}(c^{\prime})$ which is mapped to $z^{(1)}$.
+Use coordinates centered at $c^{\prime}$ of the parameter $(t,a)$ and denote by $v^{\prime}=F^{N}(c^{\prime})$.
+Let $(\Delta x_{c^{\prime}},\Delta y_{c^{\prime}})\in W^{u}_{\text{loc}}(c^{\prime})$ with $(\Delta x_{v^{\prime}},\Delta y_{v^{\prime}})\in F_{t,a}^{N}\left(W^{u}_{\text{loc}}(c^{\prime})\right)$ centered in $v^{\prime}$.
+By Lemma 2.2.12 we have
+$\Delta y_{v^{\prime}}$ $=$ $C_{c^{\prime}}\Delta x_{c^{\prime}}+Q_{c^{\prime}}(\Delta y_{c^{\prime}})^{2}+O\left(\Delta x_{c^{\prime}}^{2}+\Delta x_{c^{\prime}}\Delta y_{c^{\prime}}+\Delta y_{c^{\prime}}^{3}\right).$ (4.1.11)
+The point $c^{(1)}$ is given by
+$0=\frac{d\Delta y_{v^{\prime}}}{d\Delta y_{c^{\prime}}}.$
+By taking the derivate of (4.1.11), we obtain
+$0=\frac{d\Delta y_{v^{\prime}}}{d\Delta y_{c^{\prime}}}$ $=$ $C_{c^{\prime}}\frac{d\Delta x_{c^{\prime}}}{d\Delta y_{c^{\prime}}}+2Q_{c^{\prime}}\Delta y_{c^{\prime}}+O\left(2\Delta x_{c^{\prime}}\frac{d\Delta x_{c^{\prime}}}{d\Delta y_{c^{\prime}}}+\frac{d\Delta x_{c^{\prime}}}{d\Delta y_{c^{\prime}}}\Delta y_{c^{\prime}}+\Delta x_{c^{\prime}}+3\Delta y_{c^{\prime}}^{2}\right)$
+$=$ $C_{c^{\prime}}\frac{d\Delta x_{c^{\prime}}}{d\Delta y_{c^{\prime}}}\left(1+o(1)\right)+2Q_{c^{\prime}}\Delta y_{c^{\prime}}\left(1+o(1)\right)+O\left(\Delta x_{c^{\prime}}\right),$
+where we used Lemma 4.1.1. Using (4.1.2), (4.1.3) we get
+$\Delta y_{c^{\prime}}=O\left(\left(\frac{\lambda^{2}}{\mu}\right)^{\frac{\theta n}{2}}\right).$ (4.1.12)
+Use now the point $c^{(1)}$ with coordinates centered at $c$.
+Let $c^{(1)}=(\Delta x_{c},\Delta y_{c})\in W^{u}_{\text{loc}}(c^{\prime})$ with $z^{(1)}=(\Delta x_{z},\Delta y_{z})\in F_{t,a}^{N}\left(W^{u}_{\text{loc}}(c^{\prime})\right)$ centered in $z$.
+Observe that $\Delta x_{c}=\Delta x_{c^{\prime}}+c^{\prime}_{x}$ and $\Delta y_{c}=\Delta y_{c^{\prime}}+c^{\prime}_{y}-1$.
+Hence, by (4.1.2), (4.1.7) and (4.1.5),
+$\Delta x_{c}=\lambda^{\theta n}+O\left(\left(\frac{\lambda^{2}}{\mu}\right)^{\frac{\theta n}{2}}\right)\text{ and }\Delta y_{c}=O\left(\lambda^{{\theta n}}\right).$ (4.1.13)
+Moreover, by Lemma 2.2.12 we have
+$\Delta y_{z}$ $=$ $C\Delta x_{c}+Q(\Delta y_{c})^{2}+O\left(\Delta x_{c}^{2}+\Delta x_{c}\Delta y_{c}+\Delta y_{c}^{3}\right).$ (4.1.14)
+Using (4.1.13) in (4.1.14) we have
+$z_{y}^{(1)}(t,a)-z_{y}(t,a)=\Delta y_{z}=C\lambda^{\theta n}+O\left(\left(\frac{\lambda^{2}}{\mu}\right)^{\frac{\theta n}{2}}\right).$
+If $(x,y)\in W^{u}_{\text{loc}}(z^{(1)})$, using again (4.1.14), (4.1.13) and the fact that $0\leq|\Delta y_{c}|\leq(L\lambda^{\theta n})^{1/2}+O\left(\lambda^{\theta n}\right)$, we get
+$C\lambda^{\theta n}+O\left(\left(\frac{\lambda^{2}}{\mu}\right)^{\frac{\theta n}{2}}\right)\leq y-z_{y}(t,a)\leq(C+QL)\lambda^{\theta n}+O\left(\left(\frac{\lambda^{2}}{\mu}\right)^{\frac{\theta n}{2}}\right).$
+Observe that, in the last two inequalities, for the estimate of the order terms we used $(F3)$. ∎
+The curve $W^{u}_{\text{\rm loc}}(z^{(1)})$ is a graph over the $x$-axes. In the following lemma we study the expression for the curve $W^{u}_{\text{\rm loc}}(z^{(1)})$.
+Lemma 4.1.15.
+In a coordinate system centered in $z^{(1)}$ the ${\mathcal{C}}^{4}$} curve $W^{u}_{\text{\rm loc}}(z^{(1)})$ is given by
+$\Delta y=\frac{Q}{B^{2}}\left(1+O\left(\lambda^{\theta n}\right)\right)\Delta x^{2}+O\left(\Delta x^{3}\right),$
+where $0\leq\Delta y<L\lambda^{\theta n}$, and $\Delta x=O\left(\lambda^{\frac{\theta n}{2}}\right)$.
+Proof.
+Denote by $v^{\prime}=(v^{\prime}_{x},v^{\prime}_{y})=F^{N}(c^{\prime})$ and consider coordinates $(\Delta x_{c^{\prime}},\Delta y_{c^{\prime}})\in W^{u}_{\text{loc}}(c^{\prime})$ centered at $c^{\prime}$ with $(\Delta x_{v^{\prime}},\Delta y_{v^{\prime}})\in F_{t,a}^{N}\left(W^{u}_{\text{loc}}(c^{\prime})\right)$ centered in $v^{\prime}$.
+Consider also coordinates $(\Delta x_{c^{(1)}},\Delta y_{c^{(1)}})\in W^{u}_{\text{loc}}(c^{\prime})$ centered at $c^{(1)}=F^{-N}\left(z^{(1)}\right)$ with $(\Delta x_{z^{(1)}},\Delta y_{z^{(1)}})\in\left(W^{u}_{\text{loc}}(z^{(1)})\right)$ centered at $z^{(1)}=(z^{(1)}_{x},z^{(1)}_{y})$.
+Denote by $s=\left({\lambda^{2}}/{\mu}\right)^{\frac{\theta n}{2}}$ and observe that by Lemma 4.1.1,
+$\Delta x_{c^{\prime}}$ $=$ $O\left(s\right)\Delta y_{c^{\prime}}\left[1+\Delta y_{c^{\prime}}\right],$ (4.1.16)
+and by construction,
+$\Delta x_{v^{\prime}}$ $=$ $z^{(1)}_{x}-v^{\prime}_{x}+\Delta x_{z^{(1)}}=\Delta x_{z^{(1)}}+O\left(s\right),$ (4.1.17)
+$\Delta y_{v^{\prime}}$ $=$ $z^{(1)}_{y}-v^{\prime}_{y}+\Delta y_{z^{(1)}}=\Delta y_{z^{(1)}}+O\left(s^{2}\right),$ (4.1.18)
+where we also used (4.1.12). Applyng in order (4.1.17), Lemma 2.2.12 and (4.1.16) we get the following series of equalities,
+$\Delta x_{z^{(1)}}$ $=$ $\Delta x_{v^{\prime}}+O\left(s\right)$
+$=$ $A_{c^{\prime}}\Delta x_{c^{\prime}}+B_{c^{\prime}}\Delta y_{c^{\prime}}+O\left(\Delta x_{c^{\prime}}^{2}+\Delta x_{c^{\prime}}\Delta y_{c^{\prime}}+\Delta y_{c^{\prime}}^{2}\right)+O\left(s\right)$
+$=$ $B_{c^{\prime}}\Delta y_{c^{\prime}}\left[1+O\left(s\right)+O\left(\Delta y_{c^{\prime}}\right)\right]+O(s).$
 In particular,
-$|z_{y}^{(1)}(t,a_{n}(t))-z_{y}(t,a_{n}(t))|=C\lambda^{\theta n}+O\left(\lambda^{2\theta n}\right).$ (4.2)
-Proof.
-Use coordinates centered at the critical point $c$ of the parameter $(t,a)$ and let $(\Delta x,\Delta y)\in W^{u}_{\text{loc}}(c^{\prime})$ with $(\Delta\tilde{x},\Delta\tilde{y})\in F_{t,a}^{N}\left(W^{u}_{\text{loc}}(c^{\prime})\right)$ centered in $z$.
-By Lemma 5 we have
-$\Delta\tilde{y}$ $=$ $C\Delta x+Q(\Delta y)^{2}+O\left(|\Delta x|^{2}+|\Delta x||\Delta y|+|\Delta y|^{3}\right).$
-Because the first coordinate of $q_{1}$ is equal to $1$, see $(F4)$ and the following normalization, we have that $\Delta x=\lambda^{\theta n}+O\left(\lambda^{2\theta n}\right)$ and since $0\leq|\Delta y|\leq L^{\frac{1}{2}}\lambda^{\frac{\theta n}{2}}$,
-$C\lambda^{\theta n}+O\left(\lambda^{2\theta n}\right)\leq\Delta\tilde{y} \displaystyle\leq$ $\left(C+{QL}\right)\lambda^{\theta n}+O\left(\lambda^{2\theta n}\right).$
-The first claim of the lemma is then proved. The fact that, in the parameter $(t,a_{n}(t))$, $\Delta y$ coincides with $c^{\prime}$, ensures that 4.2 is also true.
-For the second claim of the lemma, use the coordinates centered in $c^{\prime}$.
-Let $(\Delta x,\Delta y)\in W^{u}_{\text{loc}}(c^{\prime})$ with $(\Delta\tilde{x},\Delta\tilde{y})=F_{t,a}^{N}\left((\Delta x,\Delta y)\right)$ written in coordinates centered in $F^{N}(c^{\prime})$.
-Consider $W^{u}_{\text{loc}}(c^{\prime})$ as a graph on the $y$-axis and take a tangent vector $(\Delta u,\Delta v)\in T_{c^{\prime}}W^{u}_{\text{loc}}(c^{\prime})$.
-Define the slope-vector at $c^{\prime}$ as $s=\Delta u/\Delta v$.
-Then
-$s=O\left(\left(\frac{\lambda^{2}}{\mu}\right)^{\frac{\theta n}{2}}\right).$ (4.3)
-This estimate is obtained as follows.
-Let $(z_{x}+\Delta z_{x},z_{y}+\Delta z_{y})\in W^{u}_{\text{loc}}(z)$ such that $F^{\theta n}\left(z_{x}+\Delta z_{x},z_{y}+\Delta z_{y}\right)=c^{\prime}$ and $(\Delta u_{0},\Delta v_{0})$ be a tangent vector to $W^{u}_{\text{loc}}(z)$ at the point $\left(z_{x}+\Delta z_{x},z_{y}+\Delta z_{y}\right)$.
-Then there exist uniform constants $K_{1},K_{2}$ such that
-${K_{1}^{-1}}{\mu^{-\theta n}}\leq\Delta z_{y}\leq K_{1}{\mu^{-\theta n}},$
-and because of the quadratic behavior of $W^{u}_{\text{loc}}(z)$,
-$K_{2}^{-1}{\mu^{-\frac{\theta n}{2}}}\leq|\Delta z_{x}|\leq K_{2}{\mu^{-\frac{\theta n}{2}}}.$
-In particular
-$K_{2}^{-1}{\mu^{-\frac{\theta n}{2}}}|\Delta u_{0}|\leq|\Delta v_{0}|\leq K_{2}{\mu^{-\frac{\theta n}{2}}}|\Delta u_{0}|.$
-Observe that
-$s=O\left(\frac{\lambda^{\theta n}}{\mu^{\theta n}}\frac{\Delta u_{0}}{\Delta v_{0}}\right).$
-The estimate for the slope follows.
-Use now coordinates $\left(\Delta x^{\prime},\Delta y^{\prime}\right)$ centered around $z$ with $F^{\theta n}\left(\Delta x^{\prime},\Delta y^{\prime}\right)=\left(\Delta x,\Delta y\right)$.
-Then there is a non zero smooth function $R$ such that
-$|\Delta x^{\prime}|=R\left(\Delta y^{\prime}\right)\left(\Delta y^{\prime}\right)^{\frac{1}{2}},$
-and
-$\Delta y^{\prime}=O\left(\frac{1}{\mu^{\theta n}}\right).$
-In particular,
-$\frac{d^{2}|\Delta x^{\prime}|}{d\left(\Delta y^{\prime}\right)^{2}}=O\left(\left(\Delta y^{\prime}\right)^{-\frac{3}{2}}\right)=O\left(\mu^{\frac{3}{2}\theta n}\right).$
-This implies that
-$\frac{d^{2}|\Delta x|}{d\left(\Delta y\right)^{2}}=O\left(\mu^{\frac{3}{2}\theta n}\frac{\lambda^{\theta n}}{\mu^{2\theta n}}\right)=O\left(\left(\frac{\lambda^{2}}{\mu}\right)^{\frac{\theta n}{2}}\right).$
-Hence,
-$\Delta x=s\Delta y+O\left(\left(\frac{\lambda^{2}}{\mu}\right)^{\frac{\theta n}{2}}\right)(\Delta y)^{2}.$ (4.4)
-From (4.4) and (4.3) we have that $|\Delta x|=O\left(s\Delta y\right)=o\left(\Delta y\right)$. From Lemma 5,
-$\Delta\tilde{x}$ $=$ $B\Delta y+O_{x}\left(\Delta y\right)(\Delta y)^{2}$
-$\Delta\tilde{y}$ $=$ $s\Delta y+Q(\Delta y)^{2}+O_{y}\left(\Delta y\right)(\Delta y)^{3},$
-where $B,Q\neq 0$, $O_{x}$ and $O_{y}$ are bounded differentiable functions in $\Delta y$. After differentiating one obtains
-$\frac{d\tilde{y}}{d\tilde{x}}=\frac{2Q}{B}\Delta y+O\left(\left(\frac{\lambda^{2}}{\mu}\right)^{\frac{\theta n}{2}}\right),$ (4.5)
-where we used (4.3). As consequence $z^{(1)}$ which is the point where ${d\tilde{y}}/{d\tilde{x}}=0$ corresponds to
-$\Delta y=O\left(\left(\frac{\lambda^{2}}{\mu}\right)^{\frac{\theta n}{2}}\right)\leq L^{\frac{1}{2}}\lambda^{\frac{\theta n}{2}}.$
-Indeed $W^{u}_{\text{loc}}(z^{(1)})$ has a horizontal tangency. ∎
+$\Delta y_{c^{\prime}}$ $=$ $\frac{\Delta x_{z^{(1)}}+O(s)+O\left(\Delta y_{c^{\prime}}^{2}\right)}{B_{c^{\prime}}\left(1+O(s)\right)}$ (4.1.19)
+$=$ $\frac{\Delta x_{z^{(1)}}}{B_{c^{\prime}}\left(1+O(s)\right)}+O\left(\lambda^{{\theta n}}\right),$ (4.1.20)
+where we used that $\Delta y_{c^{\prime}}=O\left(\lambda^{\frac{\theta n}{2}}\right)$, see Lemma 4.1.1. Applying in order (4.1.18), Lemma 2.2.12, (4.1.16) and (4.1.19) we get
+$\Delta y_{z^{(1)}}$ $=$ $\Delta y_{v^{\prime}}+O\left(s^{2}\right)$
+$=$ $C_{c^{\prime}}\Delta x_{c^{\prime}}+Q_{c^{\prime}}\Delta y_{c^{\prime}}^{2}+O\left(\Delta x_{c^{\prime}}^{2}+\Delta x_{c^{\prime}}\Delta y_{c^{\prime}}+\Delta y_{c^{\prime}}^{3}\right)+O\left(s^{2}\right)$
+$=$ $C_{c^{\prime}}O(s)\Delta y_{c^{\prime}}\left(1+O\left(\lambda^{\frac{\theta n}{2}}\right)\right)+Q_{c^{\prime}}\Delta y_{c^{\prime}}^{2}+O\left(s\Delta y_{c^{\prime}}^{2}+\Delta y_{c^{\prime}}^{3}\right)+O\left(s^{2}\right)$
+$=$ $\left[O\left(s^{2}\right)+O\left(s\lambda^{{\theta n}}\right)+O\left(\lambda^{{2\theta n}}\right)\right]+\left[O\left(s\right)+O\left(\lambda^{{\theta n}}\right)\right]\Delta x_{z^{(1)}}$
+$+$ $\left[\frac{Q_{c^{\prime}}\left(1+O(s)\right)}{B_{c^{\prime}}^{2}}+O\left(\lambda^{{\theta n}}\right)\right]\Delta x_{z^{(1)}}^{2}+\left[O\left(1\right)\right]\Delta x_{z^{(1)}}^{3}.$
+Observe, that this polynomial expression has a special form, the constant and linear term are zero. This is because $z^{(}1)$ is the lowest point of the curve $W^{u}_{\text{loc}}(z^{(1)})$. Hence,
+$\Delta y_{z^{(1)}}$ $=$ $\frac{Q_{c^{\prime}}}{B_{c^{\prime}}^{2}}\left(1+O\left(\lambda^{{\theta n}}\right)\right)\Delta x_{z^{(1)}}^{2}+O\left(\Delta x_{z^{(1)}}^{3}\right)$
+$=$ $\frac{Q}{B^{2}}\left(1+O\left(\lambda^{{\theta n}}\right)\right)\Delta x_{z^{(1)}}^{2}+O\left(\Delta x_{z^{(1)}}^{3}\right),$
+where we used that, by construction, $c$ and $c^{\prime}$ are at distance to each other proportional to $\lambda^{{\theta n}}$. ∎
 Figure 4: $2N+n$ iterates
 The second loop: $N+\theta n+N+n+N$ iterates.
-Let $z^{(2)}=(z_{x}^{(2)},z_{y}^{(2)})=F_{t,a}^{n}(z^{(1)})$ and take a local unstable manifold $W^{u}_{\text{loc}}(z^{(2)})$ of vertical size $L\left(\lambda^{\theta}\mu\right)^{n}$ with $L$ a large constant which is going to be chosen later.
-Observe that $z^{(2)}$ is located close to $c$, see Figure 4.
+Let $(t,a)\in\mathcal{B}_{n}$ and let $z^{(2)}=(z_{x}^{(2)},z_{y}^{(2)})=F_{t,a}^{n}(z^{(1)})$ and take a local unstable manifold $W^{u}_{\text{loc}}(z^{(2)})$ of vertical size $L\left(\lambda^{\theta}\mu\right)^{n}$ with $L$ a large constant which is going to be chosen later, see Figure 4.
 The following lemma gives information on the size and the position of $W^{u}_{\text{\rm loc}}(z^{(2)})$.
-In particular Lemma 7 states how much $W^{u}_{\text{loc}}(z^{(2)})$ is shifted up with respect to the critical point $c$, see Figure 4.
-Its proof follows from Lemma 6 and from the fact that we are in the domain of linearization.
-Lemma 7.
-For $n$ large enough and $(x,y)\in W^{u}_{\text{\rm loc}}$, if $(t,a)\in\mathcal{B}_{n}$,
-$(C-E)\left(\lambda^{\theta}\mu\right)^{n}+O\left(\mu^{n}\lambda^{2\theta n}\right)$ $\leq$
-$y-\mu_{0}^{n}z_{n,y}+\mu^{n}(a-a_{n}(t))$ $\leq$
-$\frac{T}{\mu^{n_{0}/2}}+(C+QL)\left(\lambda^{\theta}\mu\right)^{n}+O\left(\mu^{n}\lambda^{2\theta n}\right),$
-where $z_{n}=z(t,a_{n}(t))$, $\mu_{0}=\mu(t,a_{n}(t))$ and $\mu=\mu(t,a)$. Moreover
-$x=O\left(\lambda^{n}\right).$
-In particular, when $a=a_{n}(t)$,
-$y-\mu_{0}^{n}z_{n,y}=C\left(\lambda^{\theta}\mu\right)^{n}+O\left(\mu^{n}\lambda^{2\theta n}\right).$
-The curves $W^{u}_{\text{\rm loc}}(z^{(1)})$, $W^{u}_{\text{\rm loc}}(z^{(2)})$ and $W^{u}_{\text{\rm loc}}(z^{(3)})$ (which will be introduced later) are graphs over smooth curves in the $x$-plane.
-We can describe these curves in coordinates centered at $z^{(i)}$.
-In particular, for each component of $W^{u}_{\text{\rm loc}}(z^{(i)})\setminus z^{(i)}$, there exists a smooth function $w$ such that if $(\Delta x,\Delta y)\in W^{u}_{\text{\rm loc}}(z^{(i)})$ then $\Delta y=w\left(|\Delta x|\right)$.
-In the following lemma we study the curvature of $W^{u}_{\text{\rm loc}}(z^{(2)})$.
-Lemma 8.
-For $n\geq 1$ large enough, in a coordinate system centered in $z^{(2)}$ the ${\mathcal{C}}^{4}$} curve $W^{u}_{\text{\rm loc}}(z^{(2)})$ is given by
-$\Delta y=Q_{2}\left(\frac{\mu}{\lambda^{2}}\right)^{n}|\Delta x|^{2}+O\left(|\Delta x|^{3}\left(\frac{\mu}{\lambda^{3}}\right)^{n}\right),$
-where $(\Delta x,\Delta y)\in W^{u}_{\text{\rm loc}}(z^{(2)})$, $0\leq\Delta y<L\left(\lambda^{\theta}\mu\right)^{n}$, $|\Delta x|=O\left(\lambda^{2+\theta}\right)^{\frac{n}{2}}$ and the constant $Q_{2}>0$ is a uniformly bounded and bounded away from zero.
+Lemma 4.1.21.
+For all $(t,a)\in\mathcal{B}_{n}$,
+$z_{y}^{(2)}(t,a)-\mu^{n}_{0}z_{n,y}=C\left(\lambda^{\theta}\mu\right)^{n}+\left(\mu^{n}+O(n)\right)\left(a-a_{n}(t)\right)+O\left(\left({\lambda^{\theta}}\mu^{1-\frac{\theta}{2}}\right)^{n}\right),$
+where $z_{n}=z(t,a_{n}(t))$ and $\mu_{0}=\mu(t,a_{n}(t))$. Moreover if $(x,y)\in W^{u}_{\text{loc}}(z^{(2)})$,
+$C\left(\lambda^{\theta}\mu\right)^{n}+\left(\mu^{n}+O(n)\right)\left(a-a_{n}(t)\right)+O\left(\left({\lambda^{\theta}}\mu^{1-\frac{\theta}{2}}\right)^{n}\right)$
+$\leq y-\mu_{0}^{n}z_{n,y}\leq$
+$(C+QL)\left(\lambda^{\theta}\mu\right)^{n}+\left(\mu^{n}+O(n)\right)\left(a-a_{n}(t)\right)+O\left(\left({\lambda^{\theta}}\mu^{1-\frac{\theta}{2}}\right)^{n}\right),$
+and
+$x=\lambda^{n}+o\left(\lambda^{n}\right).$
 Proof.
-Consider the $W^{u}_{\text{loc}}(z^{(1)})$ of vertical size $L\lambda^{\theta n}$ and let $B\in T_{z^{(1)}}W^{u}_{\text{loc}}(z^{(1)})$ be a unit vector.
-Then in coordinate centered in $z^{(1)}$ the ${\mathcal{C}}^{4}$} curve $W^{u}_{\text{loc}}(z^{(1)})$ is given by
-$\Delta y$ $=$ $Q_{1}|\Delta x|^{2}+O(|\Delta x|^{3}).$
-where $Q_{1}$ is a uniform constant, see (4.5).
-Given the fact that $B$ has a non zero first coordinate, see $(F5)$, then the linear map $F^{n}_{t,a}$ turns the curve $W^{u}_{\text{loc}}(z^{(1)})$ into a curve of the form as stated in the lemma.
-Namely, recall that the curves $W^{u}_{\text{loc}}(z^{(1)})$ and $W^{u}_{\text{loc}}(z^{(2)})$ are graphs of functions over smooth curves in the $x$-plane.
-The curve $W^{u}_{\text{loc}}(z^{(1)})$ is parametrized by $|\Delta x|$ and by iteration by the linear map we obtain that the curve $W^{u}_{\text{loc}}(z^{(2)})$ is parametrized by $\lambda^{n}|\Delta x|$.
-Observe also that for $n$ large enough and for all $(t,a)\in\mathcal{B}_{n}$, the highest vertical point of $W^{u}_{\text{loc}}(z^{(2)})$ is still in the domain of linearization.
-The factor $\left({\mu}/{\lambda^{2}}\right)^{n}$ comes from the change of curvature by $n$ linear iterates.
+From Lemma 4.1.10, (2.2.9), (2.2.10), Remark 4.0.3 and from the fact that we are in the domain of linearization,
+$z_{y}^{(2)}(t,a)-\mu^{n}_{0}z_{n,y}$ $=$ $z_{y}^{(2)}(t,a)-\mu^{n}z_{y}+\mu^{n}z_{y}-\mu^{n}_{0}z_{n,y}$
+$=$ $C\left(\lambda^{\theta}\mu\right)^{n}+O\left(\left({\lambda^{\theta}}\mu^{1-\frac{\theta}{2}}\right)^{n}\right)+z_{n,y}\left(\mu^{n}-\mu^{n}_{0}\right)+\mu^{n}(a-a_{n}(t))$
+$=$ $C\left(\lambda^{\theta}\mu\right)^{n}+O\left(\left({\lambda^{\theta}}\mu^{1-\frac{\theta}{2}}\right)^{n}\right)+O\left(\left(\frac{\mu}{\mu_{0}}\right)^{n}-1\right)+\mu^{n}(a-a_{n}(t))$
+$=$ $C\left(\lambda^{\theta}\mu\right)^{n}+\left(\mu^{n}+O(n)\right)\left(a-a_{n}(t)\right)+O\left(\left({\lambda^{\theta}}\mu^{1-\frac{\theta}{2}}\right)^{n}\right).$
+Repeat the same estimates using the second inequalities from Lemma 4.1.10 to get the second inequalities in this lemma.
+For the last one, it is enough to use again Lemma 4.1.10 and the fact that we are in the linearization domain.
+∎
+The curve $W^{u}_{\text{\rm loc}}(z^{(2)})$, is a graph over the $y$-axes. In the following lemma we give an expression for the curve $W^{u}_{\text{\rm loc}}(z^{(2)})$.
+Lemma 4.1.22.
+In a coordinate system centered in $z^{(2)}$ the ${\mathcal{C}}^{4}$} curve $W^{u}_{\text{\rm loc}}(z^{(2)})$ is given by
+$\Delta y=\frac{Q}{B^{2}}\left(1+O\left(\lambda^{\theta n}\right)\right)\left(\frac{\mu}{\lambda^{2}}\right)^{n}\Delta x^{2}+O\left(\Delta x^{3}\left(\frac{\mu}{\lambda^{3}}\right)^{n}\right),$
+where $0\leq\Delta y<L\left(\lambda^{\theta}\mu\right)^{n}$ and $\Delta x=O\left(\lambda^{2+\theta}\right)^{\frac{n}{2}}$.
+Proof.
+The lemma follows from the fact that the curve $W^{u}_{\text{loc}}(z^{(1)})$ is a graph over the $x$-axes and $W^{u}_{\text{loc}}(z^{(2)})$ is obtained by applying the linear map $F^{n}_{t,a}$ to the inequality defining the shape of $W^{u}_{\text{loc}}(z^{(1)})$ in Lemma 4.1.15.
 ∎
 Let $z^{(3)}=(z_{x}^{(3)},z_{y}^{(3)})$ be the lowest point of $F^{N}_{t,a}\left(W^{u}_{\text{loc}}(z^{(2)})\right)$.
-In the next lemma we prove that the new critical value $z^{(3)}$ is far from the critical value $z$ and from now on they will have independent behavior.
-In particular, while $z$ is contained in the basin of the sink, the new critical value $z^{(3)}$ it is outside of the basin and it will be involved in the creation of the new tangency.
-Lemma 9.
+At this moment we have three critical values, $z$, $z^{(1)}$ and $z^{(3)}$.
+The critical value $z^{(1)}$ is too closely connected to $z$ to use it to create a secondary tangency.
+In the next lemma we prove that the new critical value $z^{(3)}$ is far enough and independent enough from the critical value $z$ that we can use it to create the secondary tangency.
+This independent behavior of $z^{(3)}$ is the key of our constructions.
+Lemma 4.1.23.
 For all $(t,a)\in\mathcal{B}_{n}$,
-$0\leq Q(C-E)^{2}\left(\lambda^{\theta}\mu\right)^{2n}+O\left(\left(\lambda^{\theta}\mu\right)^{3n}\right)$
-$\leq z_{y}^{(3)}(t,a)-z_{y}(t,a)\leq$
-$Q(C+QL)^{2}\left(\lambda^{\theta}\mu\right)^{2n}+\frac{QT^{2}}{\mu^{n_{0}}}+O\left(\frac{\left(\lambda^{\theta}\mu\right)^{n}}{\mu^{n_{0}/2}}\right)+O\left(\left(\lambda^{\theta}\mu\right)^{3n}\right),$
-Moreover $T_{z^{(3)}}W^{u}_{\text{loc}}(z^{(3)})\subset\mathbb{R}^{m-1}\times\left\{0\right\}$ is horizontal. In particular,
-$z_{y}^{(3)}(t,a_{n}(t))-z_{y}(t,a_{n}(t))=QC^{2}\left(\lambda^{\theta}\mu\right)^{2n}+O\left(\left(\lambda^{\theta}\mu\right)^{3n}\right).$
+$z_{y}^{(3)}(t,a)-z_{y}(t,a)=QH^{2}\left[1+O\left(H\right)\right].$
+where $H=C\left(\lambda^{\theta}\mu\right)^{n}+\left(\mu^{n}+O(n)\right)\left(a-a_{n}(t)\right)$. In particular,
+$z_{y}^{(3)}(t,a_{n}(t))-z_{y}(t,a_{n}(t))=QC^{2}\left(\lambda^{\theta}\mu\right)^{2n}\left[1+O\left(\left(\lambda^{\theta}\mu\right)^{n}\right)\right].$
 Proof.
-Use coordinates centered at $c$ and let $(\Delta x,\Delta y)\in W^{u}_{\text{loc}}(z^{(2)})=F_{t,a}^{n}\left(W^{u}_{\text{loc}}(z^{(1)})\right)$. Observe that
-$\Delta y$ $=$ $y-\mu_{0}^{n}z_{n,y}+\mu_{0}^{n}z_{n,y}-1$ (4.6)
-$=$ $y-\mu_{0}^{n}z_{n,y}+O\left(\lambda(t,a_{n}(t))|^{n}\right),$
-where we used that $F^{n}_{t,a_{n}(t)}(z_{n})\in\Gamma$ which implies that
-$|\mu_{0}^{n}z_{n,y}-1|=O\left(\left|F^{n}_{t,a_{n}(t)}(z_{n})_{x}\right|\right)=O\left(\lambda(t,a_{n}(t))^{n}\right).$
-From Lemma 7 and for all $(t,a)\in\mathcal{B}_{n}$ we get
-$(C-E)\left(\lambda^{\theta}\mu\right)^{n}+O\left(\lambda(t,a_{n}(t))^{n}\right)$ $\leq\Delta y\leq$ (4.7)
-$\frac{T}{\mu^{n_{0}/2}}+(C+QL)\left(\lambda^{\theta}\mu\right)^{n}+O\left(\lambda(t,a_{n}(t))^{n}\right),$ (4.8)
-$|\Delta x|=O\left(\lambda^{n}\right).$ (4.9)
-[…]
-$\Delta\tilde{y}=O\left(\lambda^{n}\right)+Q\left(\Delta y\right)^{2}+O\left(\left(\lambda^{\theta}\mu\right)^{3n}\right)=Q\left(\Delta y\right)^{2}+O\left(\left(\lambda^{\theta}\mu\right)^{3n}\right),$
-[…]
-In the following lemma we study the curvature of $W^{u}_{\text{\rm loc}}(z^{(3)})$.
-One can observe that the curvature of $W^{u}_{\text{\rm loc}}(z^{(3)})$ has grown considerably compared to that of $W^{u}_{\text{\rm loc}}(z)$.
-Lemma 10.
-Let $(t,a)\in\mathcal{B}_{n}$ then $W^{u}_{\text{loc}}(z^{(3)}(t,a))$ is the graph of a function and its curvature satisfies
-$\text{\rm curv}\left(W^{u}_{\text{loc}}(z^{(3)}(t,a))\right)\geq T\left(\frac{\mu^{4}}{\lambda^{2-3\theta}}\right)^{n},$
-with $T>0$ a uniform constant and ${\mu^{4}}/{\lambda^{2-3\theta}}>1$.
+Use coordinates $(\Delta x_{c},\Delta y_{c})\in W^{u}_{\text{loc}}(z^{(2)})$ centered at $c$.
+We are interested in the special point $m_{5}=(\Delta x_{c},\Delta y_{c})\in W^{u}_{\text{loc}}(z^{(2)})$ such that $F^{N}(m_{5})=z^{(3)}$.
+In particular, if $F^{N}(m_{5})=(\Delta x_{z},\Delta y_{z})\in W^{u}_{\text{loc}}(z^{(3)})$ centered at $z$.
+Along the proof we will need as well coordinates $(\Delta x_{z^{(2)}},\Delta y_{z^{(2)}})\in W^{u}_{\text{loc}}(z^{(2)})$ centered at $z^{(2)}$.
+By construction and by Lemma 4.1.22 we get
+$\Delta x_{c}=z^{(2)}_{x}-m_{5,x}+\Delta x_{z^{(2)}}=O\left(\left(\lambda^{2+\theta}\right)^{\frac{n}{2}}\right)+\Delta x_{z^{(2)}},$ (4.1.24)
+and
+$\Delta y_{c}=z_{y}^{(2)}(t,a)-\mu^{n}_{0}z_{n,y}+\Delta y_{z^{(2)}}=H+\Delta y_{z^{(2)}}.$ (4.1.25)
+Observe that, by Lemma 4.1.21 and by the definition of the strip $\mathcal{B}_{n}$, $H>\left(C-E\right)\left(\lambda^{\theta}\mu\right)^{n}+O\left(\left({\lambda^{\theta}}\mu^{1-\frac{\theta}{2}}\right)^{n}\right)>O\left(\lambda^{2+\theta}\right)^{\frac{n}{2}}$.
+Use in order, Lemma 2.2.12, (4.1.24), (4.1.25), (2.3.1), (2.3.2) and the fact that $\Delta x=O\left(\lambda^{2+\theta}\right)^{\frac{n}{2}}$ (see Lemma 4.1.22) to get
+$\Delta y_{z}$ $=$ $C\Delta x_{c}+Q\Delta y_{c}^{2}+O\left(\Delta x_{c}^{2}+\Delta x_{c}\Delta y_{c}+\Delta y_{c}^{3}\right)$
+$=$ $C\Delta x_{z^{(2)}}+Q\Delta y_{z^{(2)}}^{2}+2QH\Delta y_{z^{(2)}}+QH^{2}+O\left(\lambda^{2+\theta}\right)^{\frac{n}{2}}$
+$+$ $O\left(\Delta x_{z^{(2)}}^{2}+\Delta x_{z^{(2)}}H+\Delta x_{z^{(2)}}\Delta y_{z^{(2)}}+\Delta y_{z^{(2)}}^{3}+\Delta y_{z^{(2)}}^{2}H+\Delta y_{z^{(2)}}H^{2}+H^{3}\right)$
+$=$ $C\Delta x_{z^{(2)}}\left[1+O(H)\right]+Q\Delta y_{z^{(2)}}^{2}\left[1+O(H)\right]+2QH\Delta y_{z^{(2)}}\left[1+O(H)\right]$
+$+$ $QH^{2}\left[1+O(H)\right].$
+Observe that the special point $m_{5}$ we are looking for will satisfy the equation $d\Delta y_{z}/d\Delta x_{z^{(2)}}=0$. Hence, by differentiating (4.1),
+$0=\frac{d\Delta y_{z}}{d\Delta x_{z^{(2)}}}$ $=$ $C+2Q\Delta y_{z^{(2)}}\frac{d\Delta y_{z^{(2)}}}{d\Delta x_{z^{(2)}}}+2QH\frac{d\Delta y_{z^{(2)}}}{d\Delta x_{z^{(2)}}}$
+$+$ $O\left(\Delta x_{z^{(2)}}+\Delta y_{z^{(2)}}+\Delta x_{z^{(2)}}\frac{d\Delta y_{z^{(2)}}}{\Delta x_{z^{(2)}}}+\Delta y_{z^{(2)}}^{2}\frac{d\Delta y_{z^{(2)}}}{d\Delta x_{z^{(2)}}}\right$.
+$+$ $\left.\Delta y_{z^{(2)}}\frac{d\Delta y_{z^{(2)}}}{d\Delta x_{z^{(2)}}}H+\frac{d\Delta y_{z^{(2)}}}{d\Delta x_{z^{(2)}}}H^{2}+H\right).$
+By Lemma 4.1.22,
+$\Delta y_{z^{(2)}}$ $=$ $\frac{Q}{B^{2}}\left(1+O\left(\lambda^{\theta n}\right)\right)\left(\frac{\mu}{\lambda^{2}}\right)^{n}\Delta x_{z^{(2)}}^{2}+O\left(\Delta x_{z^{(2)}}^{3}\left(\frac{\mu}{\lambda^{3}}\right)^{n}\right)$ (4.1.29)
+$=$ $\frac{Q}{B^{2}}\left(\frac{\mu}{\lambda^{2}}\right)^{n}\Delta x_{z^{(2)}}^{2}\left[1+O\left(\lambda^{\frac{\theta n}{2}}\right)\right]$
+and
+$\frac{d\Delta y_{z^{(2)}}}{d\Delta x_{z^{(2)}}}$ $=$ $2\frac{Q}{B^{2}}\left(1+O\left(\lambda^{\theta n}\right)\right)\left(\frac{\mu}{\lambda^{2}}\right)^{n}\Delta x_{z^{(2)}}+O\left(\Delta x_{z^{(2)}}^{2}\left(\frac{\mu}{\lambda^{3}}\right)^{n}\right)$
+$=$ $2\frac{Q}{B^{2}}\left(\frac{\mu}{\lambda^{2}}\right)^{n}\Delta x_{z^{(2)}}\left[1+O\left(\lambda^{\frac{\theta n}{2}}\right)\right].$
+Using the last two inequalities in (4.1) we get
+$0=\frac{d\Delta y_{z}}{d\Delta x_{z^{(2)}}}$ $=$ $2\frac{Q^{3}}{B^{4}}\left(\frac{\mu}{\lambda^{2}}\right)^{2n}\Delta x_{z^{(2)}}^{3}\left[1+O\left(H\right)\right]+4\frac{Q^{2}}{B^{2}}H\left(\frac{\mu}{\lambda^{2}}\right)^{n}\Delta x_{z^{(2)}}\left[1+O\left(H\right)\right]$ (4.1.30)
+$+$ $C\left[1+O(H)\right]$
+where we also use Lemma 4.1.22 and (2.3.1). By (4.1.30) it follows that
+$\Delta x_{z^{(2)}}=O\left(\left(\frac{\lambda^{2-\theta}}{\mu^{2}}\right)^{n}\right)$ (4.1.31)
+and by (4.1.29),
+$\Delta y_{z^{(2)}}=O\left(\left(\frac{\lambda^{2-2\theta}}{\mu^{3}}\right)^{n}\right).$ (4.1.32)
+Use the last two inequalities in (4.1) to get
+$\Delta y_{z}=QH^{2}\left[1+O(H)\right],$
+where we also used the lower bound for $H$, (2.3.1) and (2.3.2). The lemma is proved. Moreover by using (4.1.31) in (4.1.30) we get an exact estimate for $\Delta x_{z^{(2)}}$. Namely,
+$0$ $=$ $4\frac{Q^{2}}{B^{2}}H\left(\frac{\mu}{\lambda^{2}}\right)^{n}\Delta x_{z^{(2)}}\left[1+O\left(\left(\frac{\mu}{\lambda^{2}}\right)^{n}\frac{\Delta x_{z^{(2)}}^{2}}{H}\right)+O(H)\right]+C\left[1+O(H)\right]$
+$=$ $4\frac{Q^{2}}{B^{2}}H\left(\frac{\mu}{\lambda^{2}}\right)^{2n}\Delta x_{z^{(2)}}\left[1+O(H)\right]+C\left[1+O(H)\right]$
+and
+$\Delta x_{z^{(2)}}=-\frac{CB^{2}}{4Q^{2}H}\left(\frac{\lambda^{2}}{\mu}\right)^{n}\left[1+O(H)\right].$ (4.1.33)
+The curve $W^{u}_{\text{\rm loc}}(z^{(3)})$, is a graph over the $y$-axes. In the following lemma we give an expression for the curve $W^{u}_{\text{\rm loc}}(z^{(3)})$.
+Lemma 4.1.34.
+In a coordinate system centered in $z^{(3)}$ the ${\mathcal{C}}^{4}$} curve $W^{u}_{\text{\rm loc}}(z^{(3)})$ is given by
+$\Delta y=\frac{8Q^{4}}{C^{2}B^{4}}{H^{3}}\left(\frac{\mu}{\lambda^{2}}\right)^{n}\Delta x^{2}+O\left(\Delta x^{3}\right),$
+where $H=C\left(\lambda^{\theta}\mu\right)^{n}+\left(\mu^{n}+O(n)\right)\left(a-a_{n}(t)\right)$.
 Proof.
-Use coordinates centered at $z^{(2)}$ and let $m_{5}=(\Delta x,\Delta y)\in W^{u}_{\text{\rm loc}}(z^{(2)})$ be the preimage of $z^{(3)}$ under $N$ iterates, i.e.
-$F^{N}_{t,a}(\Delta x,\Delta y)=z^{(3)}$.
-Then $0\leq\Delta y\leq L\left(\lambda^{\theta}\mu\right)^{n}$.
-Let $v=\left(\begin{matrix}v_{1}\\ v_{2}\end{matrix}\right)$ be the tangent vector at $m_{5}$ to $W^{u}_{\text{loc}}(z^{(2)})$.
-We use Lemma 8 and we obtain that
-$|v_{2}|=\left[2Q_{2}\left(\frac{\mu}{\lambda^{2}}\right)^{n}|\Delta x|+O\left(|\Delta x|^{2}\left(\frac{\mu}{\lambda^{3}}\right)^{n}\right)\right]|v_{1}|.$ (4.10)
-By Lemma 1 we have that ${\partial^{2}Y}/{\partial y^{2}}$ is bounded away from zero in a neighbourhood of $c$. As a consequence, by Lemma 5 we have
-$DF^{N}_{t,a}(m_{5})=\left(\begin{matrix}A&B\\ C&D\end{matrix}\right),$
-where $A$, $B$, $C$ are matrices and $D$ is a number. Observe that, from Lemma 1, there exists a uniform constant $d^{\prime}>0$ such that
-$D=d^{\prime}\left(m_{5,y}-1\right)+O(m_{5,x}),$
-and by Lemma 7, $m_{5,x}=O\left(\lambda^{n}\right)$.
-By using again the Lemma 7 and the bound on $m_{5,x}$, by taking $\epsilon_{1}$ small enough, we get that $D\geq d\left(\lambda^{\theta}\mu\right)^{n}$ where $d>0$ is a uniform constant.
-Because $F^{N}_{t,a}(m_{5})$ is the minimum of the curve $W^{u}_{\text{loc}}(z^{(3)})$, then
-$Cv_{1}+Dv_{2}=0.$ (4.11)
-Use coordinates $\left(\Delta x^{\prime},\Delta y^{\prime}\right)$ along $W^{u}_{\text{loc}}\left(z^{(1)}\right)$ centered around $z^{(1)}$ such that $F^{n}\left(\Delta x^{\prime},\Delta y^{\prime}\right)=\left(\Delta x,\Delta y\right)$.
-Then, from (4.7), $\Delta y^{\prime}=O\left(\lambda^{\theta n}\right)$ and by the quadratic behavior, $|\Delta x^{\prime}|=O\left(\lambda^{\frac{\theta}{2}n}\right)$.
-This implies that
-$|\Delta x|=O\left(\lambda^{\left(1+\frac{\theta}{2}\right)n}\right).$ (4.12)
-By (4.10), (4.11), Lemma 8, the fact that $D\geq d\left(\lambda^{\theta}\mu\right)^{n}$ and that $C$ is a uniform constant, we have
-$2dQ_{2}|\Delta x|+O\left(|\Delta x|^{2}\left(\frac{1}{\lambda}\right)^{n}\right)=O\left(\left(\frac{\lambda^{2-\theta}}{\mu^{2}}\right)^{n}\right).$
-Moreover by (4.12), $|\Delta x|\left(\frac{1}{\lambda}\right)^{n}=O\left(\lambda^{\frac{\theta}{2}n}\right)$ and because $\lambda^{\frac{\theta}{2}n}<1$ we obtain
-$|\Delta x|=O\left(\left(\frac{\lambda^{2-\theta}}{\mu^{2}}\right)^{n}\right).$ (4.13)
-Consider the projection of the curve $W^{u}_{\text{loc}}(z^{(2)})$ on the the $(x,y)$-plane.
-The curvature of this projected curve is proportional to the curvature of the curve $W^{u}_{\text{loc}}(z^{(2)})$.
-Because $F^{n}$ is linear, we can estimate the curvature of the projection of the curve $W^{u}_{\text{loc}}(z^{(2)})$ to the $(x,y)$-plane by calculating how the linear map, which preserves the $(x,y)$-plane, changes the curvature of $W^{u}_{\text{loc}}(z^{(1)})$ on the $(x,y)$-plane.
-We denote the $x$-coordinate by $x$.
-We calculate now the second derivatives using again Lemma 8 and the fact that the curve $W^{u}_{\text{loc}}(z^{(2)})$ is in $n$ ${\mathcal{C}}^{2}$} exponentially close to the projected curve to the $xy$-plane.
-We get
-$\frac{d^{2}y}{dx^{2}}\geq\frac{3}{2}Q_{2}\left(\frac{\mu}{\lambda^{2}}\right)^{n}+O\left(|\Delta x|\left(\frac{\mu}{\lambda^{3}}\right)^{n}\right)\geq Q_{2}\left(\frac{\mu}{\lambda^{2}}\right)^{n},$ (4.14)
-where we used (4.13). Observe that $W^{u}_{\text{loc}}(m_{5})$ is a graph over an $x$-direction with second derivative at $m_{5}$ given by (4.14). By (4.10) and (4.13), the slope satisfies
-$\frac{dy}{dx}=O\left(\left(\frac{1}{\lambda^{\theta}\mu}\right)^{n}\right).$ (4.15)
-Consider now the same curve as a graph over the $y$-axis. Then, by (4.14), (4.15) and using $d^{2}x/dy^{2}=-1/\left(dy/dx\right)^{3}d^{2}y/dx^{2}$, we have
-$\frac{d^{2}x}{dy^{2}}\geq C\left(\frac{\mu^{4}}{\lambda^{2-3\theta}}\right)^{n}.$
-The map $F^{N}_{t,a}$ will preserve this order of curvature.
-As final remark observe that, by (2.16), ${\mu^{4}}/{\lambda^{2-3\theta}}=\left({\mu}/{\lambda^{2-\theta}}\right){\mu^{3}}{\lambda^{2\theta}}>1$.
-∎
+In order to be able to use Lemma 2.2.12, use coordinates $(\Delta x_{c},\Delta y_{c})\in W^{u}_{\text{loc}}(z^{(2)})$ centered at $c$ and coordinates $(\Delta x_{z},\Delta y_{z})\in W^{u}_{\text{loc}}(z^{(3)})$ centered at $z$.
+By Lemma 2.2.12, we have
+$\Delta x_{z}$ $=$ $F^{N}\left(\Delta x_{c}\right)=A\Delta x_{c}+B\Delta y_{c}+O\left(\Delta x_{c}^{2}+\Delta x_{c}\Delta y_{c}+\Delta y_{c}^{2}\right)$ (4.1.35)
+$\Delta y_{z}$ $=$ $F^{N}\left(\Delta y_{c}\right)=C\Delta x_{c}+Q\Delta y_{c}^{2}+O\left(\Delta x_{c}^{2}+\Delta x_{c}\Delta y_{c}+\Delta y_{c}^{3}\right).$ (4.1.36)
+For the aim of proving the lemma we are interested in coordinates $(\Delta x_{z^{(2)}},\Delta y_{z^{(2)}})\in W^{u}_{\text{loc}}(z^{(2)})$ centered at $z^{(2)}$ and coordinates $(\Delta x_{z^{(3)}},\Delta y_{z^{(3)}})\in W^{u}_{\text{loc}}(z^{(3)})$ centered at $z^{(3)}$.
+By construction, by Lemma 4.1.21, Lemma 4.1.23 and Lemma 4.1.22 we have,
+$\Delta x_{z}$ $=$ $\Delta x_{z^{(3)}}+1-{z_{x}^{(3)}}=\Delta x_{z^{(3)}}+O\left(H\right),$
+$\Delta y_{z}$ $=$ $\Delta y_{z^{(3)}}+{z_{y}^{(3)}}-z_{y}=\Delta x_{z^{(3)}}+O\left(H^{2}\right),$
+$\Delta x_{c}$ $=$ $\Delta x_{z^{(2)}}+O\left(\lambda^{n}\right),$
+$\Delta y_{c}$ $=$ $\Delta y_{z^{(2)}}+H=\frac{Q}{B^{2}}\left(1+O\left(\lambda^{\theta n}\right)\right)\left(\frac{\mu}{\lambda^{2}}\right)^{n}\Delta x_{z^{(2)}}^{2}+H+O\left(\left(\lambda^{\frac{3}{2}\theta}\mu\right)^{n}\right).$
+Using the equalities above in (4.1.35), (4.1.36) and after several order cancellations, we get
+$\Delta x_{z^{(3)}}$ $=$ $H\left[B+O(1)\right]+A\Delta x_{z^{(2)}}\left[1+O(H)\right]+\frac{Q}{B}\left(\frac{\mu}{\lambda^{2}}\right)^{n}\Delta x_{z^{(2)}}^{2}\left[1+O(H)\right]$ (4.1.37)
+$+$ $O\left(\left(\frac{\mu}{\lambda^{2}}\right)^{n}\Delta x_{z^{(2)}}^{3}+\left(\frac{\mu^{2}}{\lambda^{4}}\right)^{n}\Delta x_{z^{(2)}}^{4}\right)=f\left(\Delta x_{z^{(2)}}\right),$ (4.1.38)
+$\Delta y_{z^{(3)}}$ $=$ $H^{2}\left[Q+O(1)\right]+C\Delta x_{z^{(2)}}\left[1+O(H)\right]+2H\frac{Q^{2}}{B^{2}}\left(\frac{\mu}{\lambda^{2}}\right)^{n}\Delta x_{z^{(2)}}^{2}\left[1+O(H)\right]$ (4.1.39)
+$+$ $O\left(\left(\frac{\mu}{\lambda^{2}}\right)^{n}\Delta x_{z^{(2)}}^{3}+\left(\frac{\mu^{2}}{\lambda^{4}}\right)^{n}H\Delta x_{z^{(2)}}^{4}+\left(\frac{\mu^{3}}{\lambda^{6}}\right)^{n}\Delta x_{z^{(2)}}^{6}\right)=g\left(\Delta x_{z^{(2)}}\right).$ (4.1.40)
+By Lemma 4.1.22,
+$\Delta y_{z^{(3)}}=\frac{C_{2}}{2}\Delta x_{z^{(3)}}^{2}+O\left(\Delta x_{z^{(3)}}^{3}\right)$ (4.1.41)
+with $C_{2}\neq 0$. We need then to calculate the coefficient $C_{2}$. Observe that
+$C_{2}=\frac{d^{2}\Delta y_{z^{(3)}}}{d\Delta x_{z^{(3)}}^{2}}$ $=$ $g^{\prime\prime}\left(\Delta x_{z^{(2)}}\right)\left(\frac{d\Delta x_{z^{(2)}}}{d\Delta x_{z^{(3)}}}\right)^{2}+g^{\prime}\left(\Delta x_{z^{(2)}}\right)\frac{d^{2}\Delta x_{z^{(2)}}}{d\Delta x_{z^{(3)}}^{2}}$ (4.1.42)
+$=$ $g^{\prime\prime}\left(\Delta x_{z^{(2)}}\right)\left(\frac{d\Delta x_{z^{(2)}}}{d\Delta x_{z^{(3)}}}\right)^{2},$
+where we used that $z^{(3)}$ is the lowest point of the $W^{u}_{\text{\rm loc}}(z^{(3)})$ and hence $g^{\prime}\left(\Delta x_{z^{(2)}}\right)=0$. By (4.1.39),
+$g^{\prime}=\frac{d\Delta y_{z^{(3)}}}{d\Delta x_{z^{(2)}}}$ $=$ $C\left[1+O(H)\right]+4H\frac{Q^{2}}{B^{2}}\left(\frac{\mu}{\lambda^{2}}\right)^{n}\Delta x_{z^{(2)}}\left[1+O(H)\right]$
+$+$ $O\left(\left(\frac{\mu}{\lambda^{2}}\right)^{n}\Delta x_{z^{(2)}}^{2}+\left(\frac{\mu^{2}}{\lambda^{4}}\right)^{n}H\Delta x_{z^{(2)}}^{3}+\left(\frac{\mu^{3}}{\lambda^{6}}\right)^{n}\Delta x_{z^{(2)}}^{5}\right),$
+and
+$g^{\prime\prime}=\frac{d^{2}\Delta y_{z^{(3)}}}{d\Delta x_{z^{(2)}}^{2}}$ $=$ $4H\frac{Q^{2}}{B^{2}}\left(\frac{\mu}{\lambda^{2}}\right)^{n}\left[1+O(H)\right]$
+$+$ $O\left(\left(\frac{\mu}{\lambda^{2}}\right)^{n}\Delta x_{z^{(2)}}+\left(\frac{\mu^{2}}{\lambda^{4}}\right)^{n}H\Delta x_{z^{(2)}}^{2}+\left(\frac{\mu^{3}}{\lambda^{6}}\right)^{n}\Delta x_{z^{(2)}}^{4}\right).$
+By the previous inequality and by (4.1.31) after cleaning the order terms we get
+$g^{\prime\prime}\left(\Delta x_{z^{(2)}}\right)=4H\frac{Q^{2}}{B^{2}}\left(\frac{\mu}{\lambda^{2}}\right)^{n}\left[1+O(H)\right].$ (4.1.43)
+Observe that
+$\frac{d\Delta x_{z^{(2)}}}{d\Delta x_{z^{(3)}}}=\frac{1}{f^{\prime}\left(\Delta x_{z^{(2)}}\right)}$ (4.1.44)
+and by (4.1.37),
+$f^{\prime}=\frac{d\Delta x_{z^{(3)}}}{d\Delta x_{z^{(2)}}}$ $=$ $A\left[1+O(H)\right]+2\frac{Q}{B}\left(\frac{\mu}{\lambda^{2}}\right)^{n}\Delta x_{z^{(2)}}\left[1+O(H)\right]$
+$+$ $O\left(\left(\frac{\mu}{\lambda^{2}}\right)^{n}\Delta x_{z^{(2)}}^{2}+\left(\frac{\mu^{2}}{\lambda^{4}}\right)^{n}\Delta x_{z^{(2)}}^{3}\right).$
+In particular, by (4.1.33)
+$f^{\prime}\left(\Delta x_{z^{(2)}}\right)=-\frac{CB}{2Q}\frac{1}{H}\left[1+O(H)\right],$
+and by (4.1.44)
+$\frac{d\Delta x_{z^{(2)}}}{d\Delta x_{z^{(3)}}}=-\frac{2Q}{CB}{H}\left[1+O(H)\right].$ (4.1.45)
+Use (4.1.43) and (4.1.44) in (4.1.42) to get
+$C_{2}$ $=$ $\frac{16Q^{4}}{C^{2}B^{4}}{H^{3}}\left(\frac{\mu}{\lambda^{2}}\right)^{n}\left[1+O(H)\right],$
+and by (4.1.41),
+$\Delta y_{z^{(3)}}=\frac{8Q^{4}}{C^{2}B^{4}}{H^{3}}\left(\frac{\mu}{\lambda^{2}}\right)^{n}\Delta x_{z^{(3)}}^{2}+O\left(\Delta x_{z^{(3)}}^{3}\right).$
 The full cycle: $N+\theta n+N+n+N+n_{0}$ iterates.
-We are now ready to chose $n_{0}$. Let
-$S_{n}=\left\{(x,y)\in[-2,2]^{m}|y\geq W_{n}(x)\right\}.$
-Observe that $z\in S_{n}\setminus S_{n-1}$. Fix $t$ and let $n_{0}$ be the maximal $n\in{\mathbb{N}}$ such that
-$W^{u}_{\text{loc}}(z^{(3)}(t,a_{n}(t)))\subset S_{n-n_{0}}.$ (4.16)
-Observe that such a choice of $n_{0}$ gives information on the position of $W^{u}_{\text{loc}}(z^{(3)}(t,a_{n}(t)))$ with respect to the pull-backs of the stable manifold, see Figure 5.
+For all $(t,a)\in\mathcal{B}_{n}$, let $n_{0}$ be the maximal $n\in{\mathbb{N}}$ such that
+$z^{(3)}\subset\left\{(x,y)\in[-2,2]^{2}|y\geq W_{n-n_{0}}(x)\right\}.$ (4.1.46)
+Observe that the funtion $(t,a)\mapsto n_{0}$ gives information on the position of $W^{u}_{\text{loc}}(z^{(3)}(t,a_{n}(t)))$ with respect to the pull-backs of the stable manifold, see Figure 5.
 A more precise estimate on the size of $n_{0}$ comes from the following lemma.
-Let $\alpha={\log\left(\lambda^{2\theta}\mu^{3}\right)}/{\log\mu}$.
-By (2.17), $\alpha\in\left(0,{1}\right)$.
-Lemma 11.
-The integer $n_{0}$ satisfies the following:
-$n_{0}=[n\alpha]+O(1).$
+Let
+$\alpha=\frac{\log\left(\lambda^{2\theta}\mu^{3}\right)}{\log\mu}$ (4.1.47)
+and recall the definition of $n_{0}^{*}$ in (4.0.1). By (2.3.2), $\alpha\in\left(0,{1}\right)$.
+Lemma 4.1.48.
+For every $(t,a_{n}(t))\in\mathcal{B}_{n}$, the integer $n_{0}$ satisfies the following:
+$n_{0}=[n\alpha]+O(1)>n_{0}^{*}.$
+In particular there exists a uniform positive constant $K$ such tha
+$\frac{1}{K}\left(\lambda^{2\theta}\mu^{3}\right)^{n}\leq\mu^{n_{0}}\leq K\left(\lambda^{2\theta}\mu^{3}\right)^{n}.$ (4.1.49)
 Proof.
-Fiz $t$ as in the definition of $n_{0}$. From Lemma 9, (2.8) and the fact that $z\in S_{n}\setminus S_{n-1}$, there exist uniform constants $K_{1},K_{2}>0$ such that
-$z_{y}^{(3)}(t,a_{n}(t))=K_{1}\left(\frac{1}{\mu^{n}}+\left(\lambda^{\theta}\mu\right)^{2n}\right).$
-Moreover the definition of $n_{0}$ implies that
+Let $(t,a_{n}(t))\in\mathcal{B}_{n}$. From Lemma 4.1.23, (2.2.2) and the fact that $W_{n+1}(1)<z_{y}<W_{n}(1)$, there exists a uniform constant $K_{1}>0$ such that
+$\frac{1}{K_{1}}\left(\frac{1}{\mu^{n}}+\left(\lambda^{\theta}\mu\right)^{2n}\right)\leq z_{y}^{(3)}\leq K_{1}\left(\frac{1}{\mu^{n}}+\left(\lambda^{\theta}\mu\right)^{2n}\right).$
+Moreover the definition of $n_{0}$ implies the existence of a uniform constant $K_{2}>0$ such that
 $\frac{1}{K_{2}}\frac{1}{\mu^{n-n_{0}}}\leq z_{y}^{(3)}\leq{K_{2}}\frac{1}{\mu^{n-n_{0}}}.$
 The two previous inequalities imply the existence of a positive constant $K$ such that
-$\frac{1}{K}\leq\frac{1}{\mu^{n_{0}}}\left(1+\left(\lambda^{2\theta}\mu^{3}\right)^{n}\right)\leq K.$ (4.17)
-The lemma follows from (2.16). ∎
-Remark 10.
-Observe that by combining (4.17) and (2.16), one gets
-$\lambda^{2\theta}\mu^{2}<\frac{1}{\mu^{\alpha}}.$ (4.18)
-This simplifies the formulas in Lemmas 7 and 9.
+$\frac{1}{K}\leq\frac{1}{\mu^{n_{0}}}\left(1+\left(\lambda^{2\theta}\mu^{3}\right)^{n}\right)\leq K.$
+The lemma follows from (2.3.1). ∎
+Remark 4.1.50.
+Observe that by combining (4.1.47) and (2.3.1), one gets
+$\lambda^{2\theta}\mu^{2}<\frac{1}{\mu^{\alpha}}.$ (4.1.51)
 4.2 The secondary tangencies
-In this subsection we construct secondary tangencies along the curve $a_{n}$. We start by giving the formal definition of secondary tangency.
-Definition 5.
+In this subsection we locate secondary tangencies along the curve $a_{n}$. We start by giving the formal definition of secondary tangency.
+Definition 4.2.1.
 A tangency between $W^{u}_{\text{loc}}(z^{(3)})$ and $W_{n-n_{0}}$ is called a secondary tangency of type $n_{0}$.
@@
 Comparing this speed with the one of the pullbacks of the stable manifold at $q_{2}$ allows finding the new tangencies.
-We take a vector at a point $m_{1}\in W^{u}_{\text{\rm loc}}\left(c\right)$ and calculate its variation along its orbit at crucial moments, step by step, until we get a vector at the point $F^{3N+\theta n+n}(m_{1})\in W^{u}_{\text{\rm loc}}(z^{(3)})$.
-The steps will be defined precisely in the proof of the proposition.
+We take a point $m_{1}\in W^{u}_{\text{\rm loc}}\left(c\right)$ and consider its orbit at crucial moments, step by step, until we get to the point $F^{3N+\theta n+n}(m_{1})\in W^{u}_{\text{\rm loc}}(z^{(3)})$.
+In particular we are interested in the derivative of these points with respect to $m_{1},t$ and $a$.
+The crucial moments are the ones defined by the combinatorics as presented in the previous section.
 The reader can refer to Figure 4.
-A crucial application of Proposition 2 is the speed of $W^{u}_{\text{\rm loc}}(z^{(3)})$ when varying the parameters along the curve $a_{n}$.
-The dominating terms in the estimates for the partial derivatives will cancel in the calculation of this speed; see (4.48).
-This is the reason why we cannot suppress more terms in the Taylor development.
-This cancellation is also crucial in the proof of Proposition 4.
-Proposition 2.
-For $n\geq 1$ large enough and for $(t,a)\in\mathcal{B}_{n}$,
-$D(C-E)\left(\lambda^{\theta}\mu\right)^{n}\left\{\frac{n}{\mu}\frac{\partial\mu}{\partial t}+n\left(\lambda^{\theta}\mu\right)^{n}\left[C\frac{\theta}{\lambda}\frac{\partial\lambda}{\partial t}+\frac{(C-E)}{\mu}\frac{\partial\mu}{\partial t}\right]\right\}+O\left(\left(\lambda^{\theta}\mu\right)^{2n}\right)$
+The reader might be surprised by the smaller order terms we need to keep track of in Proposition 4.2.2.
+Unfortunately the leading terms will cancel and the speed of $W^{u}_{\text{\rm loc}}(z^{(3)})$ when varying the parameters along the curve $a_{n}$ will be determined by smaller order terms, see (4.2.33).
+Smaller order terms also turned out to be crucial in the proof of Proposition 5.3.1.
+Proposition 4.2.2.
+For $n\geq 1$ large enough and $(t,a)\in\mathcal{B}_{n}$,
+$DS_{C-E}\left\{\frac{n}{\mu}\frac{\partial\mu}{\partial t}+n\left(\lambda^{\theta}\mu\right)^{n}\left[C\frac{\theta}{\lambda}\frac{\partial\lambda}{\partial t}+\frac{(C-E)}{\mu}\frac{\partial\mu}{\partial t}\right]+O\left(\left(\lambda^{\theta}\mu\right)^{n}\right)\right\}$
 $\leq\frac{\partial z_{y}^{(3)}}{\partial t}\leq$
-$\frac{DT}{\mu^{n_{0}/2}}\left\{\frac{n}{\mu}\frac{\partial\mu}{\partial t}+\frac{n}{\mu}\frac{\partial\mu}{\partial t}\frac{T}{\mu^{n_{0}/2}}\right\}+O\left(\frac{1}{\mu^{n_{0}}}\right),$
+$\frac{DT}{\mu^{n_{0}^{*}/2}}\left\{\frac{n}{\mu}\frac{\partial\mu}{\partial t}+\frac{n}{\mu}\frac{\partial\mu}{\partial t}\frac{T}{\mu^{n_{0}^{*}/2}}+O\left(\frac{1}{\mu^{n_{0}^{*}/2}}\right)\right\},$
 and
-$D(C-E)\left(\lambda^{\theta}\mu^{2}\right)^{n}+1+D(C-E)\left(\lambda^{\theta}\mu\right)^{n}\left\{\frac{n}{\mu}\frac{\partial\mu}{\partial a}+n\left(\lambda^{\theta}\mu\right)^{n}\left[C\frac{\theta}{\lambda}\frac{\partial\lambda}{\partial a}+\frac{(C-E)}{\mu}\frac{\partial\mu}{\partial a}\right]\right\}$
-$+O\left(\left(\lambda^{\theta}\mu\right)^{2n}\right)\leq\frac{\partial z_{y}^{(3)}}{\partial a}\leq$
-$DT\mu^{n-n_{0}/2}+1+\frac{DT}{\mu^{n_{0}/2}}\left\{\frac{n}{\mu}\frac{\partial\mu}{\partial a}+\frac{n}{\mu}\frac{\partial\mu}{\partial a}\frac{T}{\mu^{n_{0}/2}}\right\}+O\left(\frac{1}{\mu^{n_{0}}}\right).$
-Moreover $\lambda^{\theta}\mu^{2}>1$, CE,D,T>0C-E,D,T>0 and they converge to a non zero limit as $n\to\infty$.
-In particular, when $a=a_{n}$,
-$\frac{\partial z_{y}^{(3)}}{\partial t}=DC\left(\lambda^{\theta}\mu\right)^{n}\left\{\frac{n}{\mu}\frac{\partial\mu}{\partial t}+Cn\left(\lambda^{\theta}\mu\right)^{n}\left[\frac{\theta}{\lambda}\frac{\partial\lambda}{\partial t}+\frac{1}{\mu}\frac{\partial\mu}{\partial t}\right]\right\}+O\left(\left(\lambda^{\theta}\mu\right)^{2n}\right),$
+$DS_{C-E}\left\{\mu^{n}+\frac{1}{DS_{C_{E}}}+\frac{n}{\mu}\frac{\partial\mu}{\partial a}+n\left(\lambda^{\theta}\mu\right)^{n}\left[C\frac{\theta}{\lambda}\frac{\partial\lambda}{\partial a}+\frac{(C-E)}{\mu}\frac{\partial\mu}{\partial a}\right]+O\left(\left(\lambda^{\theta}\mu\right)^{n}\right)\right\}$
+$\leq\frac{\partial z_{y}^{(3)}}{\partial a}\leq$
+$DT\mu^{n-n_{0}^{*}/2}+1+\frac{DT}{\mu^{n_{0}^{*}/2}}\left\{\frac{n}{\mu}\frac{\partial\mu}{\partial a}+\frac{n}{\mu}\frac{\partial\mu}{\partial a}\frac{T}{\mu^{n_{0}^{*}/2}}+O\left(\frac{1}{\mu^{n_{0}^{*}/2}}\right)\right\},$
+where $S_{C-E}=(C-E)\left(\lambda^{\theta}\mu\right)^{n}+O\left(\left(\lambda^{\theta}\mu^{1-\frac{\theta}{2}}\right)^{n}\right)$.
+Moreover C,CE,DC,C-E,D are positive smooth functions in $t,a$ and depending also on $n$.
+These functions converge to a non zero function in $t$ as $n\to\infty$.
+The constant $T$ is the same as in the definition of $\mathcal{B}_{n}$.
+In particular, when $a=a_{n}(t)$,
+$\frac{\partial z_{y}^{(3)}}{\partial t}=DS_{C}\left\{\frac{n}{\mu}\frac{\partial\mu}{\partial t}+Cn\left(\lambda^{\theta}\mu\right)^{n}\left[\frac{\theta}{\lambda}\frac{\partial\lambda}{\partial t}+\frac{1}{\mu}\frac{\partial\mu}{\partial t}\right]+O\left(\left(\lambda^{\theta}\mu\right)^{n}\right)\right\},$
 and
-$\frac{\partial z_{y}^{(3)}}{\partial a}$ $=$ $DC\left(\lambda^{\theta}\mu^{2}\right)^{n}+1+DC\left(\lambda^{\theta}\mu\right)^{n}\left\{\frac{n}{\mu}\frac{\partial\mu}{\partial a}+Cn\left(\lambda^{\theta}\mu\right)^{n}\left[\frac{\theta}{\lambda}\frac{\partial\lambda}{\partial a}+\frac{1}{\mu}\frac{\partial\mu}{\partial a}\right]\right\}$
-$+$ $O\left(\left(\lambda^{\theta}\mu\right)^{2n}\right).$
+$\frac{\partial z_{y}^{(3)}}{\partial a}$ $=$ $DS_{C}\left\{\mu^{n}+\frac{1}{DS_{C}}+\frac{n}{\mu}\frac{\partial\mu}{\partial a}+Cn\left(\lambda^{\theta}\mu\right)^{n}\left[\frac{\theta}{\lambda}\frac{\partial\lambda}{\partial a}+\frac{1}{\mu}\frac{\partial\mu}{\partial a}\right]+O\left(\left(\lambda^{\theta}\mu\right)^{n}\right)\right\},$
+where $S_{C}=C\left(\lambda^{\theta}\mu\right)^{n}+O\left(\left(\lambda^{\theta}\mu^{1-\frac{\theta}{2}}\right)^{n}\right)$.
 Proof.
-Choose $t_{1}\in[-t_{0},t_{0}]$.
-Observe that, by Lemma 5, the definition of unfolding, by using coordinates $(x,y,t,a)$ centered around $(0,1,t_{1},0)$ in the domain and coordinates centered around $F_{t_{1},0}^{N}(0,1)$ we get the following expression for $F_{t,a}^{N}(x,y)$,
-$F_{t,a}^{N}(x,y)$ $=$ $\left(\begin{matrix}Ax+By+tE+aF\\ Cx+Qy^{2}+a\left[1+x\varphi_{x}+y^{2}\varphi_{y}\right]\end{matrix}\right),$ (4.19)
-where the uniformly bounded coefficients satisfy
-    $A(x,y,t,a)$ is a $(n-1)\times(n-1)$ matrix valued ${\mathcal{C}}^{3}$} function,
-    $B(x,y,t,a),E(x,y,t,a),F(x,y,t,a)$ are n1n-1 vector valued ${\mathcal{C}}^{3}$} functions,
-    $C(x,y,t,a)$ is a $1\times(n-1)$ matrix valued ${\mathcal{C}}^{3}$} function,
-    $Q(x,y,t,a),\varphi_{x}(x,y,t,a),\varphi_{y}(x,y,t,a)$ are ${\mathcal{C}}^{3}$} functions.
-Moreover $B,Q\neq 0$.
-The $x$ component of $F_{t,a}^{N}(x,y)$ is a general expression and the fact that $B\neq 0$ follows from $(f5)$.
-For the $y$ component of $F_{t,a}^{N}(x,y)$ observe that the linear terms in $t$ and $y$ are absent because in $a=0$, the point $F^{N}_{t,0}(0,1)$ is a non-degenerate homoclinic tangency, see Remark 6.
-The $a$-dependence of the second component of $F^{N}_{t,a}$ follows from Remark 6.
+Choose $t_{1}\in(-t_{0},t_{0})$.
+By Lemma 2.2.12 and by Remark 2.1.15, we get the following expression for $F_{t,a}^{N}(x,y)$, in coordinates $(x,y,t,a)$ centered around $(0,1,t_{1},0)=(c,t_{1},0)$ and coordinates centered around $z=F_{t_{1},0}^{N}(0,1)$,
+$F_{t,a}^{N}(x,y)$ $=$ $\left(\begin{matrix}Ax+By+tE+aF\\ Cx+Qy^{2}+a\left[1+x\varphi_{x}+y^{2}\varphi_{y}\right]\end{matrix}\right),$ (4.2.3)
+where $A(x,y,t,a),B(x,y,t,a),C(x,y,t,a),E(x,y,t,a),F(x,y,t,a),\varphi_{x}(x,y,t,a)$ are ${\mathcal{C}}^{3}$} functions and $Q(x,y,t,a),\varphi_{y}(x,y,t,a)$ are ${\mathcal{C}}^{2}$} function over $B,Q\neq 0$.
+The fact that $B\neq 0$ follows from $(f4)$.
+For the $y$ component of $F_{t,a}^{N}(x,y)$ observe that the linear terms in $t$ and $y$ are absent because in $a=0$, the point $F^{N}_{t,0}(0,1)$ is a non-degenerate homoclinic tangency, see Remark 2.1.15.
+The $a$-dependence of the second component of $F^{N}_{t,a}$ follows from Remark 2.1.15.
 Recall that $c^{\prime}\in F^{\theta n}_{t,a}\left(W^{u}_{\text{loc}}(z)\right)\cap\Gamma$, $z^{(1)}$ is defined as the lowest point of $F^{N}_{t,a}\left(W^{u}_{\text{loc}}(c^{\prime})\right)$, $z^{(2)}$ as the lowest point of $F^{n}_{t,a}\left(W^{u}_{\text{loc}}(z^{(1)})\right)$ and $z^{(3)}$ as the lowest point of $F^{N}_{t,a}\left(W^{u}_{\text{loc}}(z^{(2)})\right)$.
@@
     $m_{6}\in W^{u}_{\text{loc}}(z^{(3)})$ sufficiently close to $z^{(3)}$.
-    $m_{5}=(m_{5,x},m_{5,y})\in W^{u}_{\text{loc}}(z^{(2)})$, $F_{t,a}^{-N}(m_{6})=m_{5}$. By Lemma 7, Remark 9 and Remark 10,
-    $(C-E)\left(\lambda^{\theta}\mu\right)^{n}+O\left(\mu^{n}\lambda^{2\theta n}\right)$ $\leq$
-    $|m_{5,y}-c_{n,y}|$ $\leq$ (4.20)
-    $\frac{T}{\mu^{n_{0}/2}}+O\left(\mu^{n}\lambda^{\theta n}\right),$
-    and by (4.13)
-    $|m_{5,x}-z^{(2)}_{x}|=O\left(\left(\frac{\lambda^{2-\theta}}{\mu^{2}}\right)^{n}\right).$ (4.21)
-    In particular, when $a=a_{n}(t)$,
-    $|m_{5,y}-c_{n,y}|=C\left(\lambda^{\theta}\mu\right)^{n}+O\left(\mu^{n}\lambda^{2\theta n}\right).$ (4.22)
-    $m_{4}=(m_{4,x},m_{4,y})\in W^{u}_{\text{loc}}(z^{(1)})$, $F_{t,a}^{-n}(m_{5})=m_{4}$. By Lemma 6 and the fact that $z_{n,y}$ can be estimated by $\mu^{-n}$, see Remark 9,
-    $|m_{4,y}|=O\left(\frac{1}{\mu^{n}}\right),$ (4.23)
-    and by (4.21), using the linear map backward and $(F5)$ (which ensures that the vector $m_{4,x}-z^{(1)}_{x}$ has a dominating first component) we get,
-    $|m_{4,x}-z^{(1)}_{x}|=O\left(\left(\frac{\lambda^{1-\theta}}{\mu^{2}}\right)^{n}\right).$ (4.24)
-    $m_{3}=(m_{3,x},m_{3,y})\in W^{u}_{\text{loc}}(c^{\prime})$, $F_{t,a}^{-N}(m_{4})=m_{3}$. By Lemma 6 (see Figure 3) and (4.24),
-    $|m_{3}-c|$ $=$ $O\left({\lambda^{\theta{n}}}\right),$ (4.25)
-    $|m_{3}-c^{\prime}|$ $=$ $O\left(\left(\frac{\lambda^{1-\theta}}{\mu^{2}}\right)^{n}\right).$ (4.26)
+    $m_{5}=(m_{5,x},m_{5,y})\in W^{u}_{\text{loc}}(z^{(2)})$, $F_{t,a}^{-N}(m_{6})=m_{5}$. By Lemma 4.1.21 and (4.0.1),
+    $(C-E)\left(\lambda^{\theta}\mu\right)^{n}+O\left(\left(\lambda^{\theta}\mu^{1-\frac{\theta}{2}}\right)^{n}\right)$ $\leq$
+    $m_{5,y}-c_{n,y}$ $\leq$ (4.2.4)
+    $\frac{T}{\mu^{n_{0}^{*}/2}}+O\left(\frac{n}{\mu^{n+{n_{0}^{*}}/{2}}}\right),$
+    and by (4.1.31)
+    $m_{5,x}-z^{(2)}_{x}=O\left(\left(\frac{\lambda^{2-\theta}}{\mu^{2}}\right)^{n}\right).$ (4.2.5)
+    In particular, when $a=a_{n}(t)$ by Lemma 4.1.21 and (4.1.32),
+    $m_{5,y}-c_{n,y}=C\left(\lambda^{\theta}\mu\right)^{n}+O\left(\left(\lambda^{\theta}\mu^{1-\frac{\theta}{2}}\right)^{n}\right).$ (4.2.6)
+    $m_{4}=(m_{4,x},m_{4,y})\in W^{u}_{\text{loc}}(z^{(1)})$, $F_{t,a}^{-n}(m_{5})=m_{4}$. By Lemma 4.1.10 and the fact that $z_{n,y}$ can be estimated by $\mu^{-n}$, see Remark 4.0.3,
+    $m_{4,y}=O\left(\frac{1}{\mu^{n}}\right),$ (4.2.7)
+    and by (4.2.5), using the linear map backward we get,
+    $m_{4,x}-z^{(1)}_{x}=O\left(\left(\frac{\lambda^{1-\theta}}{\mu^{2}}\right)^{n}\right).$ (4.2.8)
+    $m_{3}=(m_{3,x},m_{3,y})\in W^{u}_{\text{loc}}(c^{\prime})$, $F_{t,a}^{-N}(m_{4})=m_{3}$. By (4.1.7) and (4.2.8),
+    $m_{3}-c$ $=$ $O\left({\lambda^{\theta{n}}}\right),$ (4.2.9)
+    $m_{3}-c^{\prime}$ $=$ $O\left(\left(\frac{\lambda^{1-\theta}}{\mu^{2}}\right)^{n}\right).$ (4.2.10)
     $m_{2}=(m_{2,x},m_{2,y})\in W^{u}_{\text{loc}}(z)$, $F_{t,a}^{-\theta n}(m_{3})=m_{2}$. Therefore, there exists a uniform constant $K_{2}>0$ such that
-    $\frac{1}{K_{2}}\frac{1}{\mu^{\theta{n}}}\leq|m_{2,y}|\leq K_{2}\frac{1}{\mu^{\theta{n}}}.$ (4.27)
-    $m_{1}=(m_{1,x},m_{1,y})\in W^{u}_{\text{loc}}(c)$, $F_{t,a}^{-N}(m_{2})=m_{1}$. By (4.27) and because of the quadratic behavior, there exists a uniform constant $K_{1}>0$ such that
-    $\frac{1}{K_{1}}\frac{1}{\mu^{\theta\frac{n}{2}}}\leq|m_{1,y}-c_{y}|\leq K_{1}\frac{1}{\mu^{\theta\frac{n}{2}}}.$ (4.28)
+    $\frac{1}{K_{2}}\frac{1}{\mu^{\theta{n}}}\leq m_{2,y}\leq K_{2}\frac{1}{\mu^{\theta{n}}}.$ (4.2.11)
+    $m_{1}=(m_{1,x},m_{1,y})\in W^{u}_{\text{loc}}(c)$, $F_{t,a}^{-N}(m_{2})=m_{1}$. By (4.2.11), (2.2.10) and Remark 4.0.3,
+    $\frac{1}{K_{2}}\frac{1}{\mu^{\theta{n}}}\leq m_{2,y}-z_{y}\leq K_{2}\frac{1}{\mu^{\theta{n}}}.$
+    and because of the quadratic behavior, see Lemma 4.1.15,
+    $\frac{1}{K_{2}}\frac{1}{\mu^{\theta{n}/2}}\leq m_{2,x}-1\leq K_{2}\frac{1}{\mu^{\theta{n}/2}}$
+    where we adjusted the value of the constant $K_{2}$. Hence, there exists a uniform constant $K_{1}>0$ such that
+    $\frac{1}{K_{1}}\frac{1}{\mu^{\theta\frac{n}{2}}}\leq m_{1,y}-c_{y}\leq K_{1}\frac{1}{\mu^{\theta\frac{n}{2}}}.$ (4.2.12)
 Let us recall that, $m_{2}=F^{N}(m_{1})$, $m_{3}=F^{\theta n}(m_{2})$, $m_{4}=F^{N}(m_{3})$, $m_{5}=F^{n}(m_{4})$, and $m_{6}=F^{N}(m_{5})$.
-Moreover, $t_{1}\in[-t_{0},t_{0}]$ was chosen arbitrary.
-Take $(t_{1}+t,a)\in\mathcal{B}_{n}$ and consider the ${\mathcal{C}}^{4}$} map
+Moreover, $t_{1}\in[-t_{0},t_{0}]$ was chosen arbitrarly.
+Take $(t_{1}+\Delta t,a)\in\mathcal{B}_{n}$ and consider the ${\mathcal{C}}^{4}$} map
 $(\Delta t,\Delta a,\Delta y)\longmapsto F^{3N+(\theta+1)n}_{t+\Delta t,a+\Delta a}{(m_{1}+(0,\Delta y))}=m_{6}+(\Delta m_{6,x},\Delta m_{6,y}).$
 We are interested in the partial derivatives of $m_{6}$. Observe that for $i=1,3,5$,
-$dm_{i+1}$ $=$ $DF^{N}_{t,a}\left(m_{i}\right)dm_{i}+\frac{\partial F^{N}_{t,a}\left(m_{i}\right)}{\partial t}dt+\frac{\partial F^{N}_{t,a}\left(m_{i}\right)}{\partial a}da,$ (4.29)
+$dm_{i+1}$ $=$ $DF^{N}_{t,a}\left(m_{i}\right)dm_{i}+\frac{\partial F^{N}_{t,a}\left(m_{i}\right)}{\partial t}dt+\frac{\partial F^{N}_{t,a}\left(m_{i}\right)}{\partial a}da,$ (4.2.13)
 for $i=4$,
-$dm_{i+1}$ $=$ $DF^{n}_{t,a}\left(m_{i}\right)dm_{i}+\frac{\partial F^{n}_{t,a}\left(m_{i}\right)}{\partial t}dt+\frac{\partial F^{n}_{t,a}\left(m_{i}\right)}{\partial a}da,$ (4.30)
+$dm_{i+1}$ $=$ $DF^{n}_{t,a}\left(m_{i}\right)dm_{i}+\frac{\partial F^{n}_{t,a}\left(m_{i}\right)}{\partial t}dt+\frac{\partial F^{n}_{t,a}\left(m_{i}\right)}{\partial a}da,$ (4.2.14)
 and for $i=2$,
-$dm_{i+1}$ $=$ $DF^{\theta n}_{t,a}\left(m_{i}\right)dm_{i}+\frac{\partial F^{\theta n}_{t,a}\left(m_{i}\right)}{\partial t}dt+\frac{\partial F^{\theta n}_{t,a}\left(m_{i}\right)}{\partial a}da.$ (4.31)
-All partial derivatives in (4.29) are uniformly bounded. However we will need more careful estimate for $\Delta m_{i+1,y}$. Namely, for $i=1,3,5$,
-$DF^{N}_{t,a}\left(m_{i}\right)$ $=$ $\left(\begin{matrix}A_{i}&B_{i}\\ C_{i}&O\left(|m_{i,x}|\right)+D_{i}(m_{i,y}-1)\end{matrix}\right),$
-which follows from (4.19). Similarly from (4.19) one obtains
+$dm_{i+1}$ $=$ $DF^{\theta n}_{t,a}\left(m_{i}\right)dm_{i}+\frac{\partial F^{\theta n}_{t,a}\left(m_{i}\right)}{\partial t}dt+\frac{\partial F^{\theta n}_{t,a}\left(m_{i}\right)}{\partial a}da.$ (4.2.15)
+All partial derivatives in (4.2.13) are uniformly bounded. However we will need more careful estimate for $\Delta m_{i+1,y}$. Namely, for $i=1,3,5$,
+$DF^{N}_{t,a}\left(m_{i}\right)$ $=$ $\left(\begin{matrix}A_{i}&B_{i}\\ C_{i}&O\left(m_{i,x}\right)+D_{i}(m_{i,y}-1)\end{matrix}\right),$
+which follows from (4.2.3). Similarly from (4.2.3) one obtains
 $\frac{\partial m_{i+1,y}}{\partial t}$ $=$ $\frac{\partial C}{\partial t}m_{i,x}+\frac{\partial Q}{\partial t}\left(m_{i,x}-1\right)^{2}+am_{i,x}\frac{\partial\varphi_{x}}{\partial t}+a\left(m_{i,y}-1\right)^{2}\frac{\partial\varphi_{y}}{\partial t}$
-$=$ $d_{i}\left|m_{i,y}-1\right|^{2}+O\left(|m_{i,x}|\right),$
-where $d_{i}\neq 0$ is uniformly bounded away from zero. By differentiation of (4.19) with respect of $\Delta a$ we obtain
-$\frac{\partial m_{i+1,y}}{\partial a}$ $=$ $1+O\left(\left|m_{i,y}-1\right|^{2}\right)+O\left(|m_{i,x}|\right).$
+$=$ $d_{i}\left(m_{i,y}-1\right)^{2}+O\left(m_{i,x}\right),$
+where $d_{i}\neq 0$ is uniformly bounded away from zero. By differentiation of (4.2.3) with respect to $\Delta a$ we obtain
+$\frac{\partial m_{i+1,y}}{\partial a}$ $=$ $1+O\left(\left(m_{i,y}-1\right)^{2}\right)+O\left(m_{i,x}\right).$
 For $i=3$ we refine the estimate
-$DF^{N}_{t,a}\left(m_{3}\right)$ $=$ $\left(\begin{matrix}A_{3}&B_{3}\\ C_{3}&O\left(|m_{3}-c^{\prime}|\right)\end{matrix}\right),$
+$DF^{N}_{t,a}\left(m_{3}\right)$ $=$ $\left(\begin{matrix}A_{3}&B_{3}\\ C_{3}&O\left(m_{3}-c^{\prime}\right)\end{matrix}\right),$
 where we estimate $DF^{N}(m_{3})$ with $DF^{N}(c^{\prime})$ evaluated at a point at distance $|m_{3}-c^{\prime}|$. For $i=1,5$ we get in an analogous way
-$dm_{i+1}$ $=$ $\left(\begin{matrix}A_{i}&B_{i}\\ C_{i}&O\left(|m_{i,x}|\right)+D_{i}(m_{i,y}-1)\end{matrix}\right)dm_{i}$
-$+$ $\left(\begin{matrix}O\left(dt\right)+O\left(da\right)\\ \left[d_{i}|m_{i,y}-1|^{2}+O\left(|m_{i,x}|\right)\right]dt+\left[1+O\left(|m_{i,y}-1|^{2}+|m_{i,x}|\right)\right]da\end{matrix}\right),$
+$dm_{i+1}$ $=$ $\left(\begin{matrix}A_{i}&B_{i}\\ C_{i}&O\left(m_{i,x}\right)+D_{i}(m_{i,y}-1)\end{matrix}\right)dm_{i}$
+$+$ $\left(\begin{matrix}O\left(dt\right)+O\left(da\right)\\ \left[d_{i}(m_{i,y}-1)^{2}+O\left(m_{i,x}\right)\right]dt+\left[1+O\left((m_{i,y}-1)^{2}+m_{i,x}\right)\right]da\end{matrix}\right),$
 and
-$dm_{4}$ $=$ $\left(\begin{matrix}A_{3}&B_{3}\\ C_{3}&O\left(|m_{3}-c^{\prime}|\right)\end{matrix}\right)dm_{3}$
-$+$ $\left(\begin{matrix}O\left(dt\right)+O\left(da\right)\\ \left[d_{3}|m_{3,y}-1|^{2}+O\left(|m_{3,x}|\right)\right]dt+\left[1+O\left(|m_{3,y}-1|^{2}+|m_{3,x}|\right)\right]da\end{matrix}\right),$
+$dm_{4}$ $=$ $\left(\begin{matrix}A_{3}&B_{3}\\ C_{3}&O\left(m_{3}-c^{\prime}\right)\end{matrix}\right)dm_{3}$
+$+$ $\left(\begin{matrix}O\left(dt\right)+O\left(da\right)\\ \left[d_{3}(m_{3,y}-1)^{2}+O\left(m_{3,x}\right)\right]dt+\left[1+O\left((m_{3,y}-1)^{2}+m_{3,x}\right)\right]da\end{matrix}\right),$
 where $D_{i},B_{i},C_{i}\neq 0$ (because $q_{1}$ is a non degenerate tangency in general direction and $DF^{N}(0,1)$ is non singular). Moreover, by construction,
-$C_{3}=C(1+\lambda^{\theta_{n}}).$ (4.34)
-From $dm_{4}$ to $dm_{5}$ we use the linear map $F^{n}$. From (4.30) and using that $m_{5,y}=\mu^{n}m_{4,y}$ we get
-$dm_{5}$ $=$ $\left(\begin{matrix}\lambda^{n}&0\\ 0&\mu^{n}\end{matrix}\right)dm_{4}+\left(\begin{matrix}X_{4}\frac{n\lambda^{n}}{\lambda}\left[\frac{\partial\lambda}{\partial t}dt+\frac{\partial\lambda}{\partial a}da\right]\\ m_{5,y}\frac{n}{\mu}\left[\frac{\partial\mu}{\partial t}dt+\frac{\partial\mu}{\partial a}da\right]\end{matrix}\right),$ (4.35)
+$C_{3}=C(1+O\left(\lambda^{\theta_{n}}\right)).$ (4.2.18)
+From $dm_{4}$ to $dm_{5}$ we use the linear map $F^{n}$. From (4.2.14) and using that $m_{5,y}=\mu^{n}m_{4,y}$ we get
+$dm_{5}$ $=$ $\left(\begin{matrix}\lambda^{n}&0\\ 0&\mu^{n}\end{matrix}\right)dm_{4}+\left(\begin{matrix}X_{4}\frac{n\lambda^{n}}{\lambda}\left[\frac{\partial\lambda}{\partial t}dt+\frac{\partial\lambda}{\partial a}da\right]\\ m_{5,y}\frac{n}{\mu}\left[\frac{\partial\mu}{\partial t}dt+\frac{\partial\mu}{\partial a}da\right]\end{matrix}\right),$ (4.2.19)
 with $0\neq X_{4}=m_{4,x}\approx 1$.
-A similar formula holds going from $dm_{2}$ to $dm_{3}$ where we use the linear map $F^{\theta n}$, (4.31) and the fact that $m_{3,y}=\mu^{\theta n}m_{2,y}$.
+A similar formula holds going from $dm_{2}$ to $dm_{3}$ where we use the linear map $F^{\theta n}$, (4.2.15) and the fact that $m_{3,y}=\mu^{\theta n}m_{2,y}$.
 Namely,
-$dm_{3}$ $=$ $\left(\begin{matrix}\lambda^{\theta n}&0\\ 0&\mu^{\theta n}\end{matrix}\right)dm_{2}+\left(\begin{matrix}X_{2}\frac{\theta n\lambda^{\theta n}}{\lambda}\left[\frac{\partial\lambda}{\partial t}dt+\frac{\partial\lambda}{\partial a}da\right]\\ m_{3,y}\frac{\theta n}{\mu}\left[\frac{\partial\mu}{\partial t}dt+\frac{\partial\mu}{\partial a}da\right]\end{matrix}\right)$ (4.36)
+$dm_{3}$ $=$ $\left(\begin{matrix}\lambda^{\theta n}&0\\ 0&\mu^{\theta n}\end{matrix}\right)dm_{2}+\left(\begin{matrix}X_{2}\frac{\theta n\lambda^{\theta n}}{\lambda}\left[\frac{\partial\lambda}{\partial t}dt+\frac{\partial\lambda}{\partial a}da\right]\\ m_{3,y}\frac{\theta n}{\mu}\left[\frac{\partial\mu}{\partial t}dt+\frac{\partial\mu}{\partial a}da\right]\end{matrix}\right),$ (4.2.20)
 where
-$0\neq X_{2}=m_{2,x}\approx\left(\begin{matrix}1&0&\ldots&0\end{matrix}\right)$ (4.37)
-Observe that $m_{1,x}=0$ and recall that $\left(dm_{1,x},dm_{1,y}\right)=\left(0,dy\right)$. By (4.2) and by (4.28),
+$0\neq X_{2}=m_{2,x}\approx 1.$ (4.2.21)
+Observe that $m_{1,x}=0$ and recall that $\left(dm_{1,x},dm_{1,y}\right)=\left(0,dy\right)$. By (4.2) and by (4.2.12),
 $\left(\begin{matrix}dm_{2,x}\\ dm_{2,y}\end{matrix}\right)=\left(\begin{matrix}A_{1}&B_{1}\\ C_{1}&D_{1}\tilde{K}_{1}{\mu^{\frac{-\theta{n}}{2}}}\end{matrix}\right)\left(\begin{matrix}0\\ dy\end{matrix}\right)+\left(\begin{matrix}O\left(dt\right)+O\left(da\right)\\ \left[d_{1}\tilde{K}_{1}^{2}{\mu^{{-\theta{n}}}}\right]dt+\left[1+O\left({\mu^{{-\theta{n}}}}\right)\right]da\end{matrix}\right),$
-where $\tilde{K}_{1}$ is the constant which gives an equality in (4.28). As a consequence
+where $\tilde{K}_{1}$ is the constant which gives an equality in (4.2.12). As a consequence
 $dm_{2,x}$ $=$ $O\left(dt\right)+O\left(da\right)+O\left(dy\right),$
 $dm_{2,y}$ $=$ $\frac{\tilde{d}_{1}}{\mu^{\theta{n}}}dt+\left[1+O\left(\frac{1}{\mu^{\theta{n}}}\right)\right]da+\frac{\tilde{D}_{1}}{\mu^{\theta\frac{n}{2}}}dy,$
-where $\tilde{d}_{1}=d_{1}\tilde{K}_{1}^{2}$ and $\tilde{D}_{1}=D_{1}\tilde{K}_{1}$. By using the fact that $F^{\theta n}_{t,a}$ is linear and using (4.36),
+where $\tilde{d}_{1}=d_{1}\tilde{K}_{1}^{2}$ and $\tilde{D}_{1}=D_{1}\tilde{K}_{1}$. By using the fact that $F^{\theta n}_{t,a}$ is linear and using (4.2.20),
 $\left(\begin{matrix}dm_{3,x}\\ dm_{3,y}\end{matrix}\right)$ $=$ $\left(\begin{matrix}\lambda^{\theta n}&0\\ 0&\mu^{\theta n}\end{matrix}\right)\left(\begin{matrix}dm_{2,x}\\ dm_{2,y}\end{matrix}\right)+\left(\begin{matrix}X_{2}\frac{\theta n\lambda^{\theta n}}{\lambda}\left[\frac{\partial\lambda}{\partial t}dt+\frac{\partial\lambda}{\partial a}da\right]\\ \left(\left(m_{3,y}-c_{n,y}\right)+c_{n,y}\right)\frac{\theta n}{\mu}\left[\frac{\partial\mu}{\partial t}dt+\frac{\partial\mu}{\partial a}da\right]\end{matrix}\right),$
@@
 $+$ $\left[\mu^{\theta n}+\frac{\theta n}{\mu}\frac{\partial\mu}{\partial a}+O\left(1\right)\right]da+\tilde{D}_{1}{\mu^{\theta\frac{n}{2}}}dy,$
-where $X_{2}\neq 0$ is pointing in the direction of $e_{1}$, see Figure 3. By (4.2) and (4.25) (observe that $\left\{|m_{3,y}-c_{n,y}|,|m_{3,x}|\right\}=O\left(\lambda^{\theta n}\right)$), we get
+where $X_{2}\neq 0$. By (4.2) and (4.2.9) (observe that $m_{3,y}-c_{n,y},m_{3,x}=O\left(\lambda^{\theta n}\right)$), we get
 $\left(\begin{matrix}dm_{4,x}\\ dm_{4,y}\end{matrix}\right)=\left(\begin{matrix}A_{3}&B_{3}\\ C_{3}&O\left(\left(\frac{\lambda^{1-\theta}}{\mu^{2}}\right)^{n}\right)\end{matrix}\right)\left(\begin{matrix}dm_{3,x}\\ dm_{3,y}\end{matrix}\right)+\left(\begin{matrix}O\left(dt\right)+O\left(da\right)\\ O\left(\lambda^{\theta n}\right)dt+\left[1+O\left(\lambda^{\theta n}\right)\right]da\end{matrix}\right).$
-As consequence, using also (4.37) and (4.34),
+As consequence, using also (4.2.21) and (4.2.18),
 $dm_{4,x}$ $=$ $\left[B_{3}\frac{\theta n}{\mu}\frac{\partial\mu}{\partial t}+O\left(1\right)\right]dt$
@@
 $+$ $\left[1+C\frac{\theta n\lambda^{\theta n}}{\lambda}\frac{\partial\lambda}{\partial a}+O\left(\lambda^{\theta n}\right)\right]da$
-$+$ $O\left(\lambda|^{\theta{n}}dy\right).$
-By (4.35) we get
+$+$ $O\left(\lambda^{\theta{n}}dy\right).$
+By (4.2.19) we get
 $dm_{5,x}$ $=$ $\left[B_{3}\frac{\theta n\lambda^{n}}{\mu}\frac{\partial\mu}{\partial t}+X_{4}\frac{n\lambda^{n}}{\lambda}\frac{\partial\lambda}{\partial t}+O\left(\lambda^{n}\right)\right]dt$
@@
 $+$ $\left[\mu^{n}+\frac{n}{\mu}\frac{\partial\mu}{\partial a}+n\left(\lambda^{\theta}\mu\right)^{n}\left[C\frac{\theta}{\lambda}\frac{\partial\lambda}{\partial a}+\frac{(C-E)}{\mu}\frac{\partial\mu}{\partial a}\right]+O\left(\left(\lambda^{\theta}\mu\right)^{n}\right)\right]da$
-$+$ $O\left(\left(\lambda|^{\theta}\mu\right)^{{n}}dy\right),$
-where we used that $m_{5,y}=(m_{5,y}-c_{n,y})+c_{n,y}$, the fact that the point $c$ is at height 11 and (-4.2). Similarly,
-$dm_{5,y}$ $\leq$ $\left[\frac{n}{\mu}\frac{\partial\mu}{\partial t}+\frac{n}{\mu}\frac{\partial\mu}{\partial t}\frac{T}{\mu^{n_{0}/2}}+n\left(\lambda^{\theta}\mu\right)^{n}C\frac{\theta}{\lambda}\frac{\partial\lambda}{\partial t}+O\left(\left(\lambda^{\theta}\mu\right)^{n}\right)\right]dt$
-$+$ $\left[\mu^{n}+\frac{n}{\mu}\frac{\partial\mu}{\partial a}+\frac{n}{\mu}\frac{\partial\mu}{\partial a}\frac{T}{\mu^{n_{0}/2}}+n\left(\lambda^{\theta}\mu\right)^{n}C\frac{\theta}{\lambda}\frac{\partial\lambda}{\partial a}+O\left(\left(\lambda^{\theta}\mu\right)^{n}\right)\right]da$
-$+$ $O\left(\left(\lambda|^{\theta}\mu\right)^{{n}}dy\right).$
-In particular, when $a=a_{n}$, using the same equalities as before and 4.22,
+$+$ $O\left(\left(\lambda^{\theta}\mu\right)^{{n}}dy\right),$
+where we used that $m_{5,y}=(m_{5,y}-c_{n,y})+c_{n,y}$, the fact that the point $c$ is at height $1$ and (-4.2). Similarly,
+$dm_{5,y}$ $\leq$ $\left[\frac{n}{\mu}\frac{\partial\mu}{\partial t}+\frac{n}{\mu}\frac{\partial\mu}{\partial t}\frac{T}{\mu^{n_{0}*/2}}+n\left(\lambda^{\theta}\mu\right)^{n}C\frac{\theta}{\lambda}\frac{\partial\lambda}{\partial t}+O\left(\left(\lambda^{\theta}\mu\right)^{n}\right)\right]dt$
+$+$ $\left[\mu^{n}+\frac{n}{\mu}\frac{\partial\mu}{\partial a}+\frac{n}{\mu}\frac{\partial\mu}{\partial a}\frac{T}{\mu^{n_{0}^{*}/2}}+n\left(\lambda^{\theta}\mu\right)^{n}C\frac{\theta}{\lambda}\frac{\partial\lambda}{\partial a}+O\left(\left(\lambda^{\theta}\mu\right)^{n}\right)\right]da$
+$+$ $O\left(\left(\lambda^{\theta}\mu\right)^{{n}}dy\right).$
+In particular, when $a=a_{n}$, using the same equalities as before and 4.2.6,
 $dm_{5,y}$ $=$ $\left[\frac{n}{\mu}\frac{\partial\mu}{\partial t}+n\left(\lambda^{\theta}\mu\right)^{n}C\left[\frac{\theta}{\lambda}\frac{\partial\lambda}{\partial t}+\frac{1}{\mu}\frac{\partial\mu}{\partial t}\right]+O\left(\left(\lambda^{\theta}\mu\right)^{n}\right)\right]dt$
 $+$ $\left[\mu^{n}+\frac{n}{\mu}\frac{\partial\mu}{\partial a}+n\left(\lambda^{\theta}\mu\right)^{n}C\left[\frac{\theta}{\lambda}\frac{\partial\lambda}{\partial a}+\frac{1}{\mu}\frac{\partial\mu}{\partial a}\right]+O\left(\left(\lambda^{\theta}\mu\right)^{n}\right)\right]da$
-$+$ $O\left(\left(\lambda|^{\theta}\mu\right)^{{n}}dy\right),$
-By (4.2), the lower bound in (-4.2) and (4.21) we get
-$\left(\begin{matrix}dm_{6,x}\\ dm_{6,y}\end{matrix}\right)$ $\geq$ $\left(\begin{matrix}A_{5}&B_{5}\\ C_{5}&O\left(\lambda^{n}\right)+D_{5}(C-E)\left(\lambda|^{\theta}\mu\right)^{{n}}\end{matrix}\right)\left(\begin{matrix}dm_{5,x}\\ dm_{5,y}\end{matrix}\right)$
-$+$ $\left(\begin{matrix}O\left(dt\right)+O\left(da\right)\\ \left[d_{5}(C-E)^{2}\left(\lambda|^{\theta}\mu\right)^{{2n}}+O\left(\lambda^{n}\right)\right]dt+\left[1+O\left(\left(\lambda|^{\theta}\mu\right)^{{2n}}\right)\right]da\end{matrix}\right).$
-As consequence,
+$+$ $O\left(\left(\lambda^{\theta}\mu\right)^{{n}}dy\right),$
+By (4.2), the lower bound in (-4.2) and (4.2.5) we get
+$\left(\begin{matrix}dm_{6,x}\\ dm_{6,y}\end{matrix}\right)$ $\geq$ $\left(\begin{matrix}A_{5}&B_{5}\\ C_{5}&D_{5}S_{C-E}\end{matrix}\right)\left(\begin{matrix}dm_{5,x}\\ dm_{5,y}\end{matrix}\right)$
+$+$ $\left(\begin{matrix}O\left(dt\right)+O\left(da\right)\\ \left[d_{5}S_{C-E}^{2}\right]dt+\left[1+O\left(S_{C-E}^{2}\right)\right]da\end{matrix}\right),$
+where $S_{C-E}=(C-E)\left(\lambda^{\theta}\mu\right)^{n}+O\left(\left(\lambda^{\theta}\mu^{1-\frac{\theta}{2}}\right)^{n}\right)$. As consequence,
 $dm_{6,x}$ $\geq$ $\left[B_{5}\frac{n}{\mu}\frac{\partial\mu}{\partial t}+O\left(1\right)\right]dt+\left[B_{5}\mu^{n}+B_{5}\frac{n}{\mu}\frac{\partial\mu}{\partial a}+O\left(1\right)\right]da$
-$+$ $O\left(\left(\lambda|^{\theta}\mu\right)^{{n}}dy\right),$
-$dm_{6,y}$ $\geq$ $\left[D_{5}(C-E)\left(\lambda^{\theta}\mu\right)^{n}\frac{n}{\mu}\frac{\partial\mu}{\partial t}+D_{5}(C-E)n\left(\lambda^{\theta}\mu\right)^{2n}\left[C\frac{\theta}{\lambda}\frac{\partial\lambda}{\partial t}+\frac{(C-E)}{\mu}\frac{\partial\mu}{\partial t}\right]\right$.
-$+$ $\left.O\left(\left(\lambda^{\theta}\mu\right)^{2n}\right)\right]dt$
-$+$ $\left[D_{5}(C-E)\left(\lambda^{\theta}\mu^{2}\right)^{n}+1+D_{5}(C-E)\left(\lambda^{\theta}\mu\right)^{n}\frac{n}{\mu}\frac{\partial\mu}{\partial a}\right$.
-$+$ $\left.D_{5}(C-E)n\left(\lambda^{\theta}\mu\right)^{2n}\left[C\frac{\theta}{\lambda}\frac{\partial\lambda}{\partial a}+\frac{(C-E)}{\mu}\frac{\partial\mu}{\partial a}\right]\right$.
-$+$ $\left.O\left(\left(\lambda^{\theta}\mu\right)^{2n}\right)\right]da+O\left(\left(\lambda|^{\theta}\mu\right)^{{2n}}\right)dy,$
-where $D_{5},B_{5},(C-E)\neq 0$ and $\lambda^{\theta}\mu^{2}>1$ (see (2.16)). Similarly, By (4.2), the upper bound in (-4.2) and (4.21) we get
-$\left(\begin{matrix}dm_{6,x}\\ dm_{6,y}\end{matrix}\right)$ $\leq$ $\left(\begin{matrix}A_{5}&B_{5}\\ C_{5}&{D_{5}T}/{\mu^{n_{0}/2}}+O\left(\lambda^{\theta}\mu\right)^{n}\end{matrix}\right)\left(\begin{matrix}dm_{5,x}\\ dm_{5,y}\end{matrix}\right)$
-$+$ $\left(\begin{matrix}O\left(dt\right)+O\left(da\right)\\ \left[{d_{5}T^{2}}/{\mu^{n_{0}}}+O\left(\lambda^{\theta}\mu\right)^{2n}\right]dt+\left[1+O\left({1}/{\mu^{n_{0}}}\right)\right]da\end{matrix}\right)$
+$+$ $O\left(\left(\lambda^{\theta}\mu\right)^{{n}}dy\right),$
+$dm_{6,y}$ $\geq$ $\left[D_{5}S_{C-E}\frac{n}{\mu}\frac{\partial\mu}{\partial t}+D_{5}S_{C-E}n\left(\lambda^{\theta}\mu\right)^{n}\left[C\frac{\theta}{\lambda}\frac{\partial\lambda}{\partial t}+\frac{(C-E)}{\mu}\frac{\partial\mu}{\partial t}\right]\right$.
+$+$ $\left.O\left(S_{C-E}\left(\lambda^{\theta}\mu\right)^{n}\right)\right]dt$
+$+$ $\left[D_{5}S_{C-E}\mu^{n}+1+D_{5}S_{C-E}\frac{n}{\mu}\frac{\partial\mu}{\partial a}+D_{5}S_{C-E}n\left(\lambda^{\theta}\mu\right)^{n}\left[C\frac{\theta}{\lambda}\frac{\partial\lambda}{\partial a}+\frac{(C-E)}{\mu}\frac{\partial\mu}{\partial a}\right]\right$.
+$+$ $\left.O\left(S_{C-E}\left(\lambda^{\theta}\mu\right)^{n}\right)\right]da+O\left(\left(\lambda^{\theta}\mu\right)^{{2n}}\right)dy$ (4.2.25)
+where $D_{5},B_{5},(C-E)\neq 0$ and $S_{C-E}\mu^{n}>1$ (see (2.3.1)). Similarly, By (4.2), the upper bound in (-4.2) and (4.2.5) we get
+$\left(\begin{matrix}dm_{6,x}\\ dm_{6,y}\end{matrix}\right)$ $\leq$ $\left(\begin{matrix}A_{5}&B_{5}\\ C_{5}&{D_{5}T}/{\mu^{n_{0}^{*}/2}}+O\left({n}/{\mu^{n+n_{0}^{*}/2}}\right)\end{matrix}\right)\left(\begin{matrix}dm_{5,x}\\ dm_{5,y}\end{matrix}\right)$
+$+$ $\left(\begin{matrix}O\left(dt\right)+O\left(da\right)\\ \left[{d_{5}T^{2}}/{\mu^{n_{0}^{*}}}+O\left({n}/{\mu^{n+n_{0}^{*}}}\right)\right]dt+\left[1+O\left({1}/{\mu^{n_{0}^{*}}}\right)\right]da\end{matrix}\right)$
 As consequence, if $(t,a)\in\mathcal{B}_{n}$
 $dm_{6,x}$ $\leq$ $\left[B_{5}\frac{n}{\mu}\frac{\partial\mu}{\partial t}+O\left(1\right)\right]dt+\left[B_{5}\mu^{n}+B_{5}\frac{n}{\mu}\frac{\partial\mu}{\partial a}+O\left(1\right)\right]da$
-$+$ $O\left(\left(\lambda|^{\theta}\mu\right)^{{n}}dy\right),$
-$dm_{6,y}$ $\leq$ $\left[\frac{D_{5}T}{\mu^{n_{0}/2}}\frac{n}{\mu}\frac{\partial\mu}{\partial t}+\frac{n}{\mu}\frac{\partial\mu}{\partial t}\frac{D_{5}T^{2}}{\mu^{n_{0}}}+O\left(\frac{1}{\mu^{n_{0}}}\right)\right]dt$ (4.43)
-$+$ $\left[D_{5}T\mu^{n-n_{0}/2}+1+\frac{D_{5}T}{\mu^{n_{0}/2}}\frac{n}{\mu}\frac{\partial\mu}{\partial a}+\frac{n}{\mu}\frac{\partial\mu}{\partial a}\frac{D_{5}T^{2}}{\mu^{n_{0}}}+O\left(\frac{1}{\mu^{n_{0}}}\right)\right]da$
-$+$ $O\left(\left(\lambda|^{\theta}\mu\right)^{{2n}}\right)dy,$
-where $D_{5},B_{5},T\neq 0$ and $\mu^{n-n_{0}/2}>1$ (see (2.16)).
-In particular, when $a=a_{n}$, using the same equalities as before, (4.22) and (4.2) we get,
+$+$ $O\left(\left(\lambda^{\theta}\mu\right)^{{n}}dy\right),$
+$dm_{6,y}$ $\leq$ $\left[\frac{D_{5}T}{\mu^{n_{0}^{*}/2}}\frac{n}{\mu}\frac{\partial\mu}{\partial t}+\frac{n}{\mu}\frac{\partial\mu}{\partial t}\frac{D_{5}T^{2}}{\mu^{n_{0}^{*}}}+O\left(\frac{1}{\mu^{n_{0}^{*}}}\right)\right]dt$ (4.2.27)
+$+$ $\left[D_{5}T\mu^{n-n_{0}^{*}/2}+1+\frac{D_{5}T}{\mu^{n_{0}^{*}/2}}\frac{n}{\mu}\frac{\partial\mu}{\partial a}+\frac{n}{\mu}\frac{\partial\mu}{\partial a}\frac{D_{5}T^{2}}{\mu^{n_{0}^{*}}}+O\left(\frac{1}{\mu^{n_{0}^{*}}}\right)\right]da$
+$+$ $O\left(\left(\lambda^{\theta}\mu\right)^{{2n}}\right)dy,$
+where $D_{5},B_{5},T\neq 0$ and $\mu^{n-n_{0}^{*}/2}>1$ (see (4.0.1) and (2.3.1)).
+In particular, when $a=a_{n}$, using the same equalities as before, (4.2.6) and (4.2) we get,
 $dm_{6,x}$ $=$ $\left[B_{5}\frac{n}{\mu}\frac{\partial\mu}{\partial t}+O\left(1\right)\right]dt+\left[B_{5}\mu^{n}+B_{5}\frac{n}{\mu}\frac{\partial\mu}{\partial a}+O\left(1\right)\right]da$
-$+$ $O\left(\left(\lambda|^{\theta}\mu\right)^{{n}}dy\right),$
+$+$ $O\left(\left(\lambda^{\theta}\mu\right)^{{n}}dy\right),$
 and
-$dm_{6,y}$ $=$ $\left[D_{5}C\left(\lambda^{\theta}\mu\right)^{n}\frac{n}{\mu}\frac{\partial\mu}{\partial t}+D_{5}Cn\left(\lambda^{\theta}\mu\right)^{2n}C\left[\frac{\theta}{\lambda}\frac{\partial\lambda}{\partial t}+\frac{1}{\mu}\frac{\partial\mu}{\partial t}\right]\right$.
-$+$ $\left.O\left(\left(\lambda^{\theta}\mu\right)^{2n}\right)\right]dt$
-$+$ $\left[D_{5}C\left(\lambda^{\theta}\mu^{2}\right)^{n}+1+D_{5}C\left(\lambda^{\theta}\mu\right)^{n}\frac{n}{\mu}\frac{\partial\mu}{\partial a}\right$.
-$+$ $\left.D_{5}Cn\left(\lambda^{\theta}\mu\right)^{2n}C\left[\frac{\theta}{\lambda}\frac{\partial\lambda}{\partial a}+\frac{1}{\mu}\frac{\partial\mu}{\partial a}\right]\right$.
-$+$ $\left.O\left(\left(\lambda^{\theta}\mu\right)^{2n}\right)\right]da+O\left(\left(\lambda|^{\theta}\mu\right)^{{2n}}\right)dy,$
-where, by (2.18)
-$\left[\frac{\theta}{\lambda}\frac{\partial\lambda}{\partial t}+\frac{1}{\mu}\frac{\partial\mu}{\partial t}\right]>0.$
-Refer to Figure 4.
+$dm_{6,y}$ $=$ $\left[D_{5}S_{C}\frac{n}{\mu}\frac{\partial\mu}{\partial t}+D_{5}S_{C}Cn\left(\lambda^{\theta}\mu\right)^{n}\left[\frac{\theta}{\lambda}\frac{\partial\lambda}{\partial t}+\frac{1}{\mu}\frac{\partial\mu}{\partial t}\right]\right$.
+$+$ $\left.O\left(S_{C}\left(\lambda^{\theta}\mu\right)^{n}\right)\right]dt$
+$+$ $\left[D_{5}S_{C}\mu^{n}+1+D_{5}S_{C}\frac{n}{\mu}\frac{\partial\mu}{\partial a}+D_{5}S_{C}Cn\left(\lambda^{\theta}\mu\right)^{n}\left[\frac{\theta}{\lambda}\frac{\partial\lambda}{\partial a}+\frac{1}{\mu}\frac{\partial\mu}{\partial a}\right]\right$.
+$+$ $\left.O\left(S_{C}\left(\lambda^{\theta}\mu\right)^{n}\right)\right]da+O\left(\left(\lambda^{\theta}\mu\right)^{{2n}}\right)dy$ (4.2.31)
+where $S_{C}=C\left(\lambda^{\theta}\mu\right)^{n}+O\left(\left(\lambda^{\theta}\mu^{1-\frac{\theta}{2}}\right)^{n}\right)$, and by (2.3.3)
+$\left|\frac{\theta}{\lambda}\frac{\partial\lambda}{\partial t}+\frac{1}{\mu}\frac{\partial\mu}{\partial t}\right|>w>0.$
 The formula for $m_{6,y}$ holds in general for all $\Delta t,\Delta a$ and $\Delta y$.
@@
 The result is achieved by comparing the rate of speed of $W^{u}_{\text{loc}}(z^{(3)})$ and $W_{n-n_{0}}$ when changing parameters in the phase space and it relies on the cancellation of the main term.
-Proposition 3.
-For all $t\in\left(-t_{0},t_{0}\right)$ and for $n$ large enough there exists $t_{n}\in\left(-t_{0},t_{0}\right)$ such that the following holds:
-    $|t-t_{n}|=O\left(\frac{1}{n}\right)$,
-    $F_{t_{n},a_{n}(t_{n})}$ has a secondary tangency $q_{1,n,n_{0}}$.
+Proposition 4.2.32.
+For all $t\in\left(-t_{0},t_{0}\right)$ and for $n$ large enough there exists $t_{n}\in\left(-t_{0},t_{0}\right)$ such $t-t_{n}=O\left({1}/{n}\right)$ and $F_{t_{n},a_{n}(t_{n})}$ has a secondary tangency.
 Proof.
-Choose $t^{*}\in\left(-t_{0},t_{0}\right)$. Use the notation of Proposition 2 and notice that, by the choice of $\theta$, see (2.18),
-$DC^{2}\left[\frac{\theta}{\lambda}\frac{\partial\lambda}{\partial t}+\frac{1}{\mu}\frac{\partial\mu}{\partial t}\right]=v>0$
-Take $n$ large enough and a point $(t,a_{n}(t))\in A_{n}$ near $(t^{*},a_{n}(t^{*}))\in A_{n}$. By Proposition 2 and Lemma 3
-$z^{(3)}_{y}\left(t,a_{n}(t)\right)$ $=$ $z^{(3)}_{y}\left(t^{*},a_{n}(t^{*})\right)+\int_{0}^{(t-t^{*})}\left[\frac{\partial z^{(3)}}{\partial t}+\frac{\partial z^{(3)}}{\partial a}\frac{da_{n}}{dt}\right]dt=z^{(3)}_{y}\left(t^{*},a_{n}(t^{*})\right)$ (4.48)
-$+$ $\left\{DC\left(\lambda^{\theta}\mu\right)^{n}\frac{n}{\mu}\frac{\partial\mu}{\partial t}+DC^{2}n\left(\lambda^{\theta}\mu\right)^{2n}\left[\frac{\theta}{\lambda}\frac{\partial\lambda}{\partial t}+\frac{1}{\mu}\frac{\partial\mu}{\partial t}\right]\right$.
-- $\left.DC\left(\lambda^{\theta}\mu\right)^{n}\frac{n}{\mu}\frac{\partial\mu}{\partial t}+O\left(\left(\lambda^{\theta}\mu\right)^{2n}\right)\right\}(t-t^{*})$
-$=$ $z^{(3)}_{y}\left(t^{*},a_{n}(t^{*})\right)+n\left(\lambda^{\theta}\mu\right)^{2n}\frac{v}{2}(t-t^{*}),$
-where we used that $\lambda^{2\theta}\mu^{3}>1$ and $\lambda^{\theta}\mu^{2}>1$ see (2.16).
-Observe that we have a cancellation of the dominant terms in the partial derivatives obtained by combining Proposition 2 and Lemma 3.
-Let $n^{*}_{0}$ be such that $W^{u}_{\text{loc}}\left(z^{(3)}\left(t^{*},a_{n}(t^{*})\right)\right)\subset S_{n-n_{0}^{*}}$, then, by (2.8),
-$z^{(3)}_{y}\left(t^{*},a_{n}(t^{*})\right)\geq\frac{2}{\mu}\frac{1}{\mu^{n-n^{*}_{0}}}.$ (4.49)
-By (4.48) and (4.49)
-$z^{(3)}_{y}\left(t,a_{n}(t)\right)\geq\frac{2}{\mu}\frac{1}{\mu^{n-n^{*}_{0}}}+n\left(\lambda^{\theta}\mu\right)^{2n}\frac{v}{2}(t-t^{*}).$ (4.50)
-Choose $\kappa\geq 2$. Then by (2.8), in a neighborhood of $(z^{(3)}\left(t^{*},a_{n}(t^{*})\right),t^{*},a_{n}(t^{*}))$
-$\max W_{n-n_{0}^{*}-\kappa}$ $\leq$ $\frac{2}{\mu^{n-n^{*}_{0}-\kappa}}.$
-From the previous inequality, (4.50), we get that $W^{u}_{\text{loc}}(z^{(3)}(t,a_{n}(t)))$ is above $W_{n-n_{0}^{*}-\kappa}$ if
-$n\left(\lambda^{\theta}\mu\right)^{2n}\frac{v}{2}(t-t^{*})\geq\frac{2}{\mu^{n-n^{*}_{0}}}\left[\mu^{\kappa}-\frac{2}{\mu}\right]\geq 2C^{\prime}\left(\lambda^{\theta}\mu\right)^{2n}\left[\mu^{\kappa}-\frac{2}{\mu}\right],$
-where we also used (4.17) and $C^{\prime}$ is a uniform constant. As consequence, if
-$t-t^{*}\geq\frac{4C^{\prime}}{n}\frac{\mu^{\kappa}}{v},$
-then, $W^{u}_{\text{loc}}(z^{(3)}(t,a_{n}(t)))$ is above $W_{n-n_{0}^{*}-\kappa}$.
-Because $W^{u}_{\text{loc}}(z^{(3)}(t^{*},a_{n}(t^{*})))$ contains a point below $W_{n-n_{0}^{*}-1}$, there exists a parameter between $(t,a_{n}(t))$ and $(t^{*},a_{n}(t^{*}))$ for which a secondary homoclinic tangency of type $n_{0}=n_{0}^{*}+\kappa$ occurs and $|t^{*}-t|=O\left(\frac{1}{n}\right)$.
-Moreover by Lemma 2, (2.8), Lemma 11 and the definition of $\alpha$ we get
-$\frac{\partial W_{n-n_{0}}}{\partial t}(q_{1,n,n_{0}})=-2n\left(\lambda^{\theta}\mu\right)^{2n}\mu^{\kappa-1}\left(1-\alpha-\frac{\kappa}{n}\right)\frac{\partial\mu}{\partial t}+O\left(\left(\lambda^{\theta}\mu\right)^{2n}\right).$ (4.51)
-Observe that, from $(F6)$, the value of $\alpha$ and the lower bound for $\theta$, see (2.17), we have
-$\left(1-\alpha-\frac{\kappa}{n}\right)\frac{\partial\mu}{\partial t}>0.$ (4.52)
+Choose $t\in\left(-t_{0},t_{0}\right)$. Use the notation of Proposition 4.2.2 and notice that, by the choice of $\theta$, see (2.3.3),
+$\left|\frac{\theta}{\lambda}\frac{\partial\lambda}{\partial t}+\frac{1}{\mu}\frac{\partial\mu}{\partial t}\right|>w>0.$
+Suppose that,
+$DC^{2}\left[\frac{\theta}{\lambda}\frac{\partial\lambda}{\partial t}+\frac{1}{\mu}\frac{\partial\mu}{\partial t}\right]=v>0,$
+the negative case is handled similarly by reversing the time. Take $n$ large enough and a point $(s,a_{n}(s))\in A_{n}$ near $(t,a_{n}(t))\in A_{n}$. By Proposition 4.2.2 and Lemma 2.2.6
+$z^{(3)}_{y}\left(s,a_{n}(s)\right)$ $=$ $z^{(3)}_{y}\left(t,a_{n}(t)\right)+\int_{0}^{(s-t)}\left[\frac{\partial z^{(3)}}{\partial t}+\frac{\partial z^{(3)}}{\partial a}\frac{da_{n}}{dt}\right]dt=z^{(3)}_{y}\left(t,a_{n}(t)\right)$ (4.2.33)
+$+$ $DS_{C}\left\{\frac{n}{\mu}\frac{\partial\mu}{\partial t}+Cn\left(\lambda^{\theta}\mu\right)^{n}\left[\frac{\theta}{\lambda}\frac{\partial\lambda}{\partial t}+\frac{1}{\mu}\frac{\partial\mu}{\partial t}\right]-\frac{n}{\mu}\frac{\partial\mu}{\partial t}+O\left(\left(\lambda^{\theta}\mu\right)^{n}\right)\right\}(s-t)$
+$=$ $z^{(3)}_{y}\left(t,a_{n}(t)\right)+\left\{DC^{2}n\left(\lambda^{\theta}\mu\right)^{2n}\left[\frac{\theta}{\lambda}\frac{\partial\lambda}{\partial t}+\frac{1}{\mu}\frac{\partial\mu}{\partial t}\right]+O\left(\left(\lambda^{\theta}\mu\right)^{2n}\right)\right\}(s-t)$
+$\geq$ $z^{(3)}_{y}\left(t,a_{n}(t)\right)+n\left(\lambda^{\theta}\mu\right)^{2n}\frac{v}{2}(s-t),$
+where we used that $\lambda^{2\theta}\mu^{3}>1$ and $\lambda^{\theta}\mu^{2}>1$ see (2.3.1).
+Observe that we have a cancellation of the dominant terms in the partial derivatives obtained by combining Proposition 4.2.2 and Lemma 2.2.6.
+Let $n_{0}=n_{0}(t,a_{n}(t))$ be as in (4.1.46).
+Hence, by (2.2.2),
+$z^{(3)}_{y}\left(t,a_{n}(t)\right)\geq\frac{2}{\mu}\frac{1}{\mu^{n-n_{0}}}.$ (4.2.34)
+By (4.2.33) and (4.2.34)
+$z^{(3)}_{y}\left(s,a_{n}(s)\right)\geq\frac{2}{\mu}\frac{1}{\mu^{n-n_{0}}}+n\left(\lambda^{\theta}\mu\right)^{2n}\frac{v}{2}(s-t).$ (4.2.35)
+Choose $\kappa\geq 2$ such that $\left[\mu^{\kappa}-{2}/{\mu}\right]\geq v/4$. Then by (2.2.2), in a neighborhood of $(z^{(3)}\left(t,a_{n}(t)\right),t,a_{n}(t))$
+$\max W_{n-n_{0}-\kappa}$ $\leq$ $\frac{2}{\mu^{n-n_{0}-\kappa}}.$
+From the previous inequality, (4.2.35), we get that $W^{u}_{\text{loc}}(z^{(3)}(s,a_{n}(s)))$ is above $W_{n-n_{0}^{*}-\kappa}$ if
+$n\left(\lambda^{\theta}\mu\right)^{2n}\frac{v}{2}(s-t)\geq\frac{2}{\mu^{n-n_{0}}}\left[\mu^{\kappa}-\frac{2}{\mu}\right]\geq 2C^{\prime}\left(\lambda^{\theta}\mu\right)^{2n}\left[\mu^{\kappa}-\frac{2}{\mu}\right]\geq C^{\prime}\left(\lambda^{\theta}\mu\right)^{2n}\frac{v}{2},$
+where we also used (4.1.49) and $C^{\prime}$ is a uniform constant. As consequence, if
+$s-t\geq\frac{C^{\prime}}{n},$
+then, $W^{u}_{\text{loc}}(z^{(3)}(s,a_{n}(s)))$ is above $W_{n-n_{0}-\kappa}$.
+Because $W^{u}_{\text{loc}}(z^{(3)}(t,a_{n}(t)))$ contains a point below $W_{n-n_{0}-1}$, there exists a parameter between $(s,a_{n}(s))$ and $(t,a_{n}(t))$ for which a secondary homoclinic tangency of type $n_{0}+1$ occurs and $t-s=O\left({1}/{n}\right)$.
+∎
 Figure 5: Position of $z^{(3)}$
 The next lemma estimates the curvature of any secondary tangency point of type $n_{0}$ in $\mathcal{B}_{n}$ satisfies the following estimate.
-Lemma 12.
+Lemma 4.2.36.
 Let $(t,a)\in\mathcal{B}_{n}$ be such that $W^{u}_{\text{\rm loc}}(z^{(3)}(t,a))$ has a secondary tangency $q_{1,n,n_{0}}$ of type $n_{0}$, then $q_{1,n,n_{0}}$ is non degenerate.
@@
 Proof.
-Use coordinates centered at $z^{(2)}$ and let $(\Delta x,\Delta y)\in W^{u}_{\text{loc}}(z^{(2)})$ such that $F^{N}_{t,a}(\Delta x,\Delta y)=q_{1,n,n_{0}}$.
-Then $0\leq\Delta y\leq L\left(\lambda^{\theta}\mu\right)^{n}$.
-Let $v=\left(\begin{matrix}v_{1}\\ v_{2}\end{matrix}\right)$ be the tangent vector at $(\Delta x,\Delta y)$ to $W^{u}_{\text{loc}}(z^{(2)})$.
-We use Lemma 8 and we obtain that
-$|v_{2}|=\left[2Q_{2}\left(\frac{\mu}{\lambda^{2}}\right)^{n}|\Delta x|+O\left(|\Delta x|^{2}\left(\frac{\mu}{\lambda^{3}}\right)^{n}\right)\right]|v_{1}|.$ (4.53)
-By Lemma 1 we have that ${\partial^{2}Y}/{\partial y^{2}}$ is bounded away from zero in a neighborhood of $c$.
-As consequence, by Lemma 7, Lemma 7 and Lemma 5 we have $DF^{N}_{t,a}(\Delta x,\Delta y)=\left(\begin{matrix}A&B\\ C&D\end{matrix}\right)$ where $D\geq d\left(\lambda^{\theta}\mu\right)^{n}$ and $d>0$ is a uniform constant.
-Because $F^{N}_{t,a}(\Delta x,\Delta y)$ is a tangency at $q_{1,n,n_{0}}$, then
-$Cv_{1}+Dv_{2}=O\left(\left(\frac{\lambda}{\mu}\right)^{n-n_{0}}\left(|v_{1}|+|v_{2}|\right)\right),$ (4.54)
-where we used (2.9). The proof of the lemma is completed by following exactly that of Lemma 10. ∎
+Use coordinates centered at $z^{(3)}$ and let $(\Delta x_{z^{(3)}},\Delta y_{z^{(3)}})\in W^{u}_{\text{loc}}(z^{(3)})$ corresponding to the tangency point $q_{1,n,n_{0}}$. Then by (2.2.3),
+$\frac{d\Delta y_{z^{(3)}}}{d\Delta x_{z^{(3)}}}=O\left(\left(\frac{\lambda}{\mu}\right)^{n-n_{0}}\right)$
+and by differentiating the equation in Lemma 4.1.34,
+$\frac{d\Delta y_{z^{(3)}}}{d\Delta x_{z^{(3)}}}=O\left(\left(\frac{\lambda}{\mu}\right)^{n-n_{0}}\right)=\frac{16Q^{4}}{C^{2}B^{4}}{H^{3}}\left(\frac{\mu}{\lambda^{2}}\right)^{n}\Delta x_{z^{(3)}}+O\left(\Delta x_{z^{(3)}}^{2}\right).$ (4.2.37)
+In particular,
+${\Delta x_{z^{(3)}}}=O\left(\left(\frac{\lambda}{\mu}\right)^{n-n_{0}}\left(\frac{\lambda^{2}}{\mu}\right)^{n}\frac{1}{H^{3}}\right).$
+The curvature at the tangency point is found by differentiating (4.2.37). Hence,
+$\frac{d^{2}\Delta y_{z^{(3)}}}{d\Delta x_{z^{(3)}}^{2}}$ $=$ $\frac{16Q^{4}}{C^{2}B^{4}}{H^{3}}\left(\frac{\mu}{\lambda^{2}}\right)^{n}+O\left(\Delta x_{z^{(3)}}\right)$
+$\geq$ $\frac{16Q^{4}}{C^{2}B^{4}}{H^{3}}\left(\frac{\mu}{\lambda^{2}}\right)^{n}\left[1-O\left(\left(\frac{\lambda}{\mu}\right)^{n-n_{0}}\left(\frac{\lambda^{4}}{\mu^{2}}\right)^{n}\frac{1}{H^{6}}\right)\right]$
+$\geq$ $\frac{8Q^{4}}{C^{2}B^{4}}{H^{3}}\left(\frac{\mu}{\lambda^{2}}\right)^{n}.$
+where we used the expression of $H$ given in Lemma 4.1.23 and the lower bound of the strip $\mathcal{B}_{n}$. The lower bound for $H$ gives the statement of the Lemma. ∎
 5 Curves of secondary tangencies
-In the previous section, we proved the existence of tangency points on the curve $A_{n}$.
-In the sequel, we extend these points to create global curves.
-All points along these curves have a secondary tangency; refer to Figure 6.
-In particular, we prove that the curves of secondary tangencies are actually long uniform curves that never leave the strip $\mathcal{B}_{n}$.
+In the previous section, we proved the existence of maps with secondary tangency points on the curve $a_{n}$.
+In the sequel, we show that secondary tangencies persists along curves of uniform size.
+In particular these curves are graphs of functions defined on all of the $t$ domain, see Figure 6.
+Moreover, we prove that the curves of secondary tangencies are actually long uniform curves that are actually contained in the strip $\mathcal{B}_{n}$.
 Figure 6: Curves of Secondary Tangencies
-5.1 Continuation of secondary tangencies: the implicit function theorem
+5.1 Persistence of secondary tangencies
 The next lemma is the first step in proving the existence of local tangency curves.
-In particular, we prove that each parameter in $\mathcal{B}_{n}$ corresponding to a map having a secondary tangency of type $n_{0}$ can be locally enlarged to create a curve.
+In particular, we prove that each parameter in $\mathcal{B}_{n}$ corresponding to a map having a secondary tangency of type $n_{0}$ is contained in a local curve.
 Each point on this curve corresponds to a map with a $n_{0}$ secondary tangency.
-This result is achieved by implementing and using the implicit function theorem.
-Lemma 13.
-Let $(t^{*},a^{*})\in{\mathcal{B}_{n}}$ be a parameter corresponding to a map with a secondary tangency of type $n_{0}$. Then there exist $\epsilon>0$ and a smooth curve
-$b_{n,n_{0}}:[t^{*}-\epsilon,t^{*}+\epsilon]\mapsto\mathbb{R}$
-such that, for all $t\in[t^{*}-\epsilon,t^{*}+\epsilon]$ the parameter $(t,b_{n,n_{0}}(t))$ corresponds to a map with a secondary tangency of type $n_{0}$. Moreover
-$\frac{(C-E)}{T}\frac{n}{\mu^{n+1}}\frac{\partial\mu}{\partial t}+O\left(n\lambda^{\theta n}\right)\leq\frac{db}{dt}\leq-\frac{T}{(C-E)}\frac{n}{\mu^{n+1}}\frac{\partial\mu}{\partial t}+O\left(\frac{n}{\mu^{n_{0}/2}}\right).$
-In particular, there exists a positive constant $v$ such that,
-$\frac{db}{dt}=\frac{da_{n}}{dt}+v{n\lambda^{\theta n}}+O\left(\lambda^{\theta n}\right)$
-for all $t\in[\tilde{t}^{*}-\epsilon,\tilde{t}^{*}+\epsilon]$, with $(\tilde{t}^{*},b_{n,n_{0}}(\tilde{t}^{*}))\in a_{n}$.
+Lemma 5.1.1.
+Let $(t_{n_{0}},a_{n_{0}})\in{\mathcal{B}_{n}}$ be a parameter corresponding to a map with a secondary tangency of type $n_{0}$. Then there exist $\epsilon>0$ and a smooth curve
+$b_{n,n_{0}}:[t_{n_{0}}-\epsilon,t_{n_{0}}+\epsilon]\mapsto\mathbb{R}$
+such that, for all $t\in[t_{n_{0}}-\epsilon,t_{n_{0}}+\epsilon]$ the parameter $(t,b_{n,n_{0}}(t))$ corresponds to a map with a secondary tangency of type $n_{0}$.
+Moreover, the slopes of the curves $b_{n,n_{0}}$ tend exponentially fast to zero, namely
+$\left|\frac{db_{n,n_{0}}}{dt}\right|=O\left(\frac{n}{\left(\lambda^{\theta}\mu^{2}\right)^{n}\mu^{n_{0}^{*}/2}}\right).$
+In particular,
+$\frac{db_{n,n_{0}}}{dt}=\frac{da_{n}}{dt}+{n\lambda^{\theta n}}C\left[\frac{\theta}{\lambda}\frac{\partial\lambda}{\partial t}+\frac{1}{\mu}\frac{\partial\mu}{\partial t}\right]+O\left(\lambda^{\theta n}\right),$
+for all $t\in[t_{n_{0}}-\epsilon,t_{n_{0}}+\epsilon]$, with $(t,b_{n,n_{0}}(t))\in a_{n}$.
 Proof.
-Let $(t^{*},a^{*})\in\mathcal{B}_{n}\cap A_{n}$ be a parameter corresponding to a map with a secondary tangency of type $n_{0}$, $q_{1,n,n_{0}}$.
-Its existence is proved in Proposition 3.
-We construct a local function whose graph is contained in a neighborhood of $t^{*}$.
+First we consider the case when the given point $(t_{n_{0}},a_{n_{0}})\in\mathcal{B}_{n}$ is on the curve $a_{n}$, i.e.
+$a_{n_{0}}=a_{n}(t_{n_{0}})$.
+In particular this point corresponds to a map with a secondary tangency of type $n_{0}$, $q_{1,n,n_{0}}$.
+We construct a local function whose graph contains the point $(t_{n_{0}},a_{n_{0}})$.
 Let $m_{1}\in W^{u}_{\text{loc}}(c)$ such that $F_{t,a}^{3N+(1+\theta)n}(m_{1})=q_{1,n,n_{0}}$.
-To describe the perturbation of $W^{u}_{\text{loc}}(q_{1,n,n_{0}})$ we define the following function by choosing coordinates centered in the image at $q_{1,n,n_{0}}$.
-Take some $\epsilon>0$ and consider the ${\mathcal{C}}^{4}$} function $(\tilde{x},\tilde{y}):(-\epsilon,\epsilon)^{3}\mapsto\mathbb{R}^{m-1}\times\mathbb{R}$ defined by
-$\left(\tilde{x}(\Delta y,\Delta t,\Delta a),\tilde{y}(\Delta y,\Delta t,\Delta a)\right)=F_{t+\Delta t,a+\Delta a}^{3N+(1+\theta)n}\left(m_{1}+(0,\Delta y)\right),$
-which describes $W^{u}_{\text{loc}}\left(F_{t+\Delta t,a+\Delta a}^{3N+(1+\theta)n}(m_{1})\right)$.
-Observe that the manifolds $W_{n-n_{0}}$ of $F_{t+\Delta t,a+\Delta a}$ are described locally, near $q_{1,n,n_{0}}$, as the graph of a ${\mathcal{C}}^{4}$} function
+To describe the perturbation of $W^{u}_{\text{loc}}(q_{1,n,n_{0}})$ we define the following function by choosing coordinates centered in $q_{1,n,n_{0}}$.
+Take some $\epsilon>0$ and consider the ${\mathcal{C}}^{4}$} function $(\tilde{x},\tilde{y}):(-\epsilon,\epsilon)^{3}\mapsto\mathbb{R}\times\mathbb{R}$ defined by
+$\left(\tilde{x}(\Delta y,\Delta t,\Delta a),\tilde{y}(\Delta y,\Delta t,\Delta a)\right)=F_{t_{n_{0}}+\Delta t,a_{n_{0}}+\Delta a}^{3N+(1+\theta)n}\left(m_{1}+(0,\Delta y)\right),$
+which describes $W^{u}_{\text{loc}}\left(F_{t_{n_{0}}+\Delta t,a_{n_{0}}+\Delta a}^{3N+(1+\theta)n}(m_{1})\right)$.
+Observe that the manifolds $W_{n-n_{0}}$ of $F_{t_{n_{0}}+\Delta t,a_{n_{0}}+\Delta a}$ are described locally, near $q_{1,n,n_{0}}$, as the graph of a ${\mathcal{C}}^{4}$} function
 $w_{n-n_{0}}:(\Delta x,\Delta t,\Delta a)\longmapsto w_{n-n_{0}}(\Delta x,\Delta t,\Delta a)\in\mathbb{R}.$
@@
 Observe that $\Psi^{-1}(0)$ describes locally the perturbation of the secondary tangency $q_{1,n,n_{0}}$ and that
-$D\Psi({0,0,0})=\left(\begin{matrix}0&\left(\lambda^{\theta}\mu\right)^{n}\Psi_{1,2}&\left(\lambda^{\theta}\mu^{2}\right)^{n}\Psi_{1,3}\\ \left(\frac{\mu^{3}}{\lambda^{2-2\theta}}\right)^{n}\Psi_{2,1}&\Psi_{2,2}&\Psi_{2,3}\end{matrix}\right).$ (5.1)
-In particular, by (4.2), (2.9), (4.2), (4.17) (in the estimation of the order term)
-$\Psi_{1,2}$ $=$ $\frac{1}{\left(\lambda^{\theta}\mu\right)^{n}}\left[\frac{\partial m_{6,y}}{\partial t}-\frac{\partial W_{n-n_{0}}}{\partial x}\frac{\partial m_{6,x}}{\partial t}-\frac{\partial W_{n-n_{0}}}{\partial t}\right]$ (5.2)
-$=$ $DC\frac{n}{\mu}\frac{\partial\mu}{\partial t}+DC^{2}n\left(\lambda^{\theta}\mu\right)^{n}\left[\frac{\theta}{\lambda}\frac{\partial\lambda}{\partial t}+\frac{1}{\mu}\frac{\partial\mu}{\partial t}\right]$
-$+$ $O\left(\left(\lambda^{\theta}\mu\right)^{n}\right)+O\left(n\left(\frac{\lambda}{\mu}\right)^{n-n_{0}}\frac{1}{\left(\lambda^{\theta}\mu\right)^{n}}\right)-\frac{1}{\left(\lambda^{\theta}\mu\right)^{n}}\frac{\partial W_{n-n_{0}}}{\partial t}$
-$=$ $DC\frac{n}{\mu}\frac{\partial\mu}{\partial t}+n\left(\lambda^{\theta}\mu\right)^{n}\left\{{DC^{2}}\left[\frac{\theta}{\lambda}\frac{\partial\lambda}{\partial t}+\frac{1}{\mu}\frac{\partial\mu}{\partial t}\right]-\frac{1}{n\left(\lambda^{\theta}\mu\right)^{2n}}\frac{\partial W_{n-n_{0}}}{\partial t}\right\}$
-$+$ $O\left(\left(\lambda^{\theta}\mu\right)^{n}\right)=DC\frac{n}{\mu}\frac{\partial\mu}{\partial t}+n\left(\lambda^{\theta}\mu\right)^{n}v+O\left(\left(\lambda^{\theta}\mu\right)^{n}\right)$
-$=$ $DC\frac{n}{\mu}\frac{\partial\mu}{\partial t}+O\left(n\left(\lambda^{\theta}\mu\right)^{n}\right),$
-and by (4.2), (2.9), (4.2), Lemma 2 and (4.17) (in the estimation of the order term)
-$\Psi_{1,3}$ $=$ $\frac{1}{\left(\lambda^{\theta}\mu^{2}\right)^{n}}\left[\frac{\partial m_{6,y}}{\partial a}-\frac{\partial W_{n-n_{0}}}{\partial x}\frac{\partial m_{6,x}}{\partial a}-\frac{\partial W_{n-n_{0}}}{\partial a}\right]$ (5.3)
-$=$ $\frac{1}{\left(\lambda^{\theta}\mu^{2}\right)^{n}}\left\{DC\left(\lambda^{\theta}\mu^{2}\right)^{n}+1+DC\left(\lambda^{\theta}\mu\right)^{n}\frac{n}{\mu}\frac{\partial\mu}{\partial a}\right$.
-$+$ $\left.DC^{2}n\left(\lambda^{\theta}\mu\right)^{2n}\left[\frac{\theta}{\lambda}\frac{\partial\lambda}{\partial a}+\frac{1}{\mu}\frac{\partial\mu}{\partial a}\right]+O\left(\left(\lambda^{\theta}\mu\right)^{2n}\right)\right$.
-$+$ $\left.O\left(\mu^{n}\left(\frac{\lambda}{\mu}\right)^{n-n_{0}}\right)+O\left(\frac{n-n_{0}}{\mu^{n-n_{0}}}\right)\right\}$
-$=$ $DC\left[1+O\left({\left({\lambda}^{1+\theta-\alpha}{\mu}\right)^{n}}\right)\right].$
-Moreover $\Psi_{1,2},\Psi_{1,3}>0$ and $\Psi_{2,1}\neq 0$, see Lemma 12, (4.2) and (2.10).
-Because $D\Psi({0,0,0})$ is onto we get that the set of secondary tangencies $\Psi^{-1}(0)$ is locally the graph of a ${\mathcal{C}}^{2}$} function $b$.
+$D\Psi({0,0,0})=\left(\begin{matrix}0&\left(\lambda^{\theta}\mu\right)^{n}\Psi_{1,2}&\left(\lambda^{\theta}\mu^{2}\right)^{n}\Psi_{1,3}\\ \left(\frac{\mu^{3}}{\lambda^{2-2\theta}}\right)^{n}\Psi_{2,1}&\Psi_{2,2}&\Psi_{2,3}\end{matrix}\right).$ (5.1.2)
+In particular, by (4.2), (2.2.3), (4.2), Lemma 2.2.5, (4.1.49) (in the estimation of the order term)
+$\Psi_{1,2}$ $=$ $\frac{1}{\left(\lambda^{\theta}\mu\right)^{n}}\left[\frac{\partial m_{6,y}}{\partial t}-\frac{\partial W_{n-n_{0}}}{\partial x}\frac{\partial m_{6,x}}{\partial t}-\frac{\partial W_{n-n_{0}}}{\partial t}\right]$ (5.1.3)
+$=$ $\frac{DS_{C}}{\left(\lambda^{\theta}\mu\right)^{n}}\left\{\frac{n}{\mu}\frac{\partial\mu}{\partial t}+Cn\left(\lambda^{\theta}\mu\right)^{n}\left[\frac{\theta}{\lambda}\frac{\partial\lambda}{\partial t}+\frac{1}{\mu}\frac{\partial\mu}{\partial t}\right]+O\left(\lambda^{\theta}\mu\right)^{n}\right\}$
+$+$ $O\left(n\left(\frac{\lambda}{\mu}\right)^{n-n_{0}}\frac{1}{\left(\lambda^{\theta}\mu\right)^{n}}\right)-\frac{1}{\left(\lambda^{\theta}\mu\right)^{n}}\frac{\partial W_{n-n_{0}}}{\partial t}$
+$=$ $\frac{DS_{C}}{\left(\lambda^{\theta}\mu\right)^{n}}\left\{\frac{n}{\mu}\frac{\partial\mu}{\partial t}+Cn\left(\lambda^{\theta}\mu\right)^{n}\left[\frac{\theta}{\lambda}\frac{\partial\lambda}{\partial t}+\frac{1}{\mu}\frac{\partial\mu}{\partial t}\right]+O\left(\lambda^{\theta}\mu\right)^{n}\right\},$
+and by (4.2.31), (2.2.3), (4.2), Lemma 2.2.5 and (4.1.49) (in the estimation of the order term)
+$\Psi_{1,3}$ $=$ $\frac{1}{\left(\lambda^{\theta}\mu^{2}\right)^{n}}\left[\frac{\partial m_{6,y}}{\partial a}-\frac{\partial W_{n-n_{0}}}{\partial x}\frac{\partial m_{6,x}}{\partial a}-\frac{\partial W_{n-n_{0}}}{\partial a}\right]$ (5.1.4)
+$=$ $\frac{DS_{C}}{\left(\lambda^{\theta}\mu^{2}\right)^{n}}\left\{\mu^{n}+\frac{1}{DS_{C}}+\frac{n}{\mu}\frac{\partial\mu}{\partial a}+Cn\left(\lambda^{\theta}\mu\right)^{n}\left[\frac{\theta}{\lambda}\frac{\partial\lambda}{\partial a}+\frac{1}{\mu}\frac{\partial\mu}{\partial a}\right]+O\left(\left(\lambda^{\theta}\mu\right)^{n}\right)\right\}$
+$+$ $O\left(\frac{\mu^{n}}{\left(\lambda^{\theta}\mu^{2}\right)^{n}}\left(\frac{\lambda}{\mu}\right)^{n-n_{0}}\right)+O\left(\frac{1}{\left(\lambda^{\theta}\mu^{2}\right)^{n}}\frac{n-n_{0}}{\mu^{n-n_{0}}}\right)$
+$=$ $\frac{DS_{C}}{\left(\lambda^{\theta}\mu\right)^{n}}\left[1+O\left(\frac{1}{\left(\lambda^{\theta}\mu^{2}\right)^{n}}\right)\right]+O\left(\left(\lambda^{1-\alpha+\theta}\mu\right)^{n}\right)$
+$=$ $\frac{DS_{C}}{\left(\lambda^{\theta}\mu\right)^{n}}\left[1+O\left(\frac{1}{\left(\lambda^{\theta}\mu^{2}\right)^{n}}\right)\right]+O\left(\left(\lambda^{\theta}\mu\right)^{n}\right)$
+$=$ $\frac{DS_{C}}{\left(\lambda^{\theta}\mu\right)^{n}}\left[1+O\left(\left(\lambda^{\theta}\mu\right)^{n}\right)\right],$
+where we used that $\alpha<1$ and $\lambda^{1-\alpha+\theta}\mu<\lambda^{\theta}\mu$.
+Hence, by the expression for $S_{C}$, see Proposition 4.2.2, $\Psi_{1,2},\Psi_{1,3}>0$.
+Moreover, by Lemma 4.2.36, (4.2) and (2.2.4) we have $\Psi_{2,1}\neq 0$.
+This implies that $D\Psi({0,0,0})$ is onto and we get that the set of secondary tangencies $\Psi^{-1}(0)$ is locally the graph of a ${\mathcal{C}}^{2}$} function $b$.
 Moreover the tangent space of $\Psi$ in zero satisfies $T_{(0,0,0)}\Psi^{-1}(0)=\text{Ker}D\Psi$.
-Hence
+Hence,
 $\left(\lambda^{\theta}\mu\right)^{n}\Psi_{1,2}\Delta t+\left(\lambda^{\theta}\mu^{2}\right)^{n}\Psi_{1,3}\Delta b=0,$
 and
-$\frac{db}{dt}=-\frac{\Psi_{1,2}}{\Psi_{1,3}}\frac{1}{\mu^{n}}=-\frac{n}{\mu^{n+1}}\frac{\partial\mu}{\partial t}-{n\lambda^{\theta n}}v+O\left(\lambda^{\theta n}\right).$ (5.4)
+$\frac{db}{dt}=-\frac{\Psi_{1,2}}{\Psi_{1,3}}\frac{1}{\mu^{n}}=-\frac{n}{\mu^{n+1}}\frac{\partial\mu}{\partial t}-{n\lambda^{\theta n}}C\left[\frac{\theta}{\lambda}\frac{\partial\lambda}{\partial t}+\frac{1}{\mu}\frac{\partial\mu}{\partial t}\right]+O\left(\lambda^{\theta n}\right).$ (5.1.5)
 In particular $db/dt$ is a negative number and $b$ is the graph of a smooth function.
-Furthermore, by the previous estimate on the slope of $b$, by Proposition 3 and by (2.11), the curve $b$ have a angle satisfying the following,
-$\frac{db}{dt}=\frac{da_{n}}{dt}-{n\lambda^{\theta n}}v+O\left(\lambda^{\theta n}\right),$ (5.5)
-with $v$ a positive constant.
-Let $(t_{0},a_{0})$ be any parameter in $\mathcal{B}_{n}$ corresponding to a map with a secondary tangency of type $n_{0}$, $q_{1,n,n_{0}}$.
+Furthermore, by the previous estimate on the slope of $b$, and by (2.2.7), the curve $b$ have a angle satisfying the following,
+$\frac{db}{dt}=\frac{da_{n}}{dt}-{n\lambda^{\theta n}}C\left[\frac{\theta}{\lambda}\frac{\partial\lambda}{\partial t}+\frac{1}{\mu}\frac{\partial\mu}{\partial t}\right]+O\left(\lambda^{\theta n}\right),$ (5.1.6)
+where $\left[\frac{\theta}{\lambda}\frac{\partial\lambda}{\partial t}+\frac{1}{\mu}\frac{\partial\mu}{\partial t}\right]$ is away from zero, see (2.3.3).
+Let $(t_{n_{0}},a_{n_{0}})$ be any parameter in $\mathcal{B}_{n}$ corresponding to a map with a secondary tangency of type $n_{0}$, $q_{1,n,n_{0}}$.
 The proof is exactly the same as the previous case.
-In particular, by (4.2), (2.9), (4.2), (4.17) (in the estimation of the order term)
-$\Psi_{1,2}$ $\geq$ $D(C-E))\frac{n}{\mu}\frac{\partial\mu}{\partial t}+O\left(n\left(\lambda^{\theta}\mu\right)^{n}\right),$
-and by (4.2), (2.9), (4.2), Lemma 2 and (4.17) (in the estimation of the order term)
-$\Psi_{1,3}$ $\geq$ $D(C-E)\left[1+O\left({\left({\lambda}^{1+\theta-\alpha}{\mu}\right)^{n}}\right)\right].$
-Moreover, by (4.43), (2.9), (4.2),
-$\Psi_{1,2}$ $\leq$ $DT\frac{n}{\mu}\frac{\partial\mu}{\partial t}+O\left(\frac{n}{\mu^{n_{0}/2}}\right),$
-and by (4.2), (2.9), (4.2), Lemma 2
-$\Psi_{1,3}$ $\leq$ $DT\left[1+O\left(\frac{1}{\mu^{n-n_{0}/2}}\right)\right].$
-Since, by Proposition 2 and (F6), $D,C-E,T,{\partial\mu}/{\partial t}>0$, then $\Psi_{1,2}>0$, $\Psi_{1,3}>0$ and $\Psi_{2,1}>0$, see Lemma 12.
+In particular, by (4.2), (2.2.3), (4.2), (4.1.49) (in the estimation of the order term)
+$\Psi_{1,2}$ $\geq$ $D(C-E)\frac{n}{\mu}\frac{\partial\mu}{\partial t}\left[1+O\left(\frac{1}{n\mu^{\theta n/2}}\right)\right],$
+and by (4.2.25), (2.2.3), (4.2), Lemma 2.2.5 and (4.1.49) (in the estimation of the order term)
+$\Psi_{1,3}$ $\geq$ $D(C-E)\left[1+O\left(\frac{1}{\mu^{\theta n/2}}\right)\right].$
+Moreover, by (4.2.27), (2.2.3), (4.2),
+$\Psi_{1,2}$ $\leq$ $\frac{DT}{\left(\lambda^{\theta}\mu\right)^{n}\mu^{n_{0}^{*}/2}}\frac{n}{\mu}\frac{\partial\mu}{\partial t}\left[1+O\left(\frac{1}{\mu^{n_{0}^{*}/2}}\right)\right].$
+Since, by Proposition 4.2.2 and Remark 2.1.6, $D,C-E,T,{\partial\mu}/{\partial t}>0$, then $\Psi_{1,2}>0$, $\Psi_{1,3}>0$ and $\Psi_{2,1}>0$, see Lemma 4.2.36.
 As before, because $D\Psi({0,0,0})$ is onto we get that the set of secondary tangencies $\Psi^{-1}(0)$ is locally the graph of a ${\mathcal{C}}^{2}$} function $b$.
-Moreover
-$\frac{(C-E)}{T}\frac{n}{\mu^{n+1}}\frac{\partial\mu}{\partial t}+O\left(n\lambda^{\theta n}\right)\leq\frac{db}{dt}\leq-\frac{T}{(C-E)}\frac{n}{\mu^{n+1}}\frac{\partial\mu}{\partial t}+O\left(\frac{n}{\mu^{n_{0}/2}}\right).$ (5.6)
-The proof is completed.
-5.2 Limitations to the existence of secondary tangencies
-We prove now that for any $n_{0}$ by moving in the direction of the $a$ parameter one can find only one tangency of type $n_{0}$ as well as when one move on the $a_{n}$ curve.
-Lemma 14.
+Moreover,
+$\left|\frac{db}{dt}\right|=\left|\frac{\Psi_{1,2}}{\Psi_{1,3}}\frac{1}{\mu^{n}}\right|=O\left(\frac{n}{\left(\lambda^{\theta}\mu^{2}\right)^{n}\mu^{n_{0}^{*}/2}}\right),$ (5.1.7)
+which goes to zero exponentially fast, see (2.3.1). The proof is complete. ∎
+5.2 Properties of the loci of secondary tangencies
+We prove now that for any $n_{0}$ by moving in the direction of the $a$ parameter one can find at the most one tangency of type $n_{0}$.
+Lemma 5.2.1.
 For $n$ large enough,
-$\frac{\partial z_{3,y}}{\partial a}-\frac{\partial W_{n-n_{0}}}{\partial a}>0$
-and
-$\frac{\partial z_{3,y}}{\partial t}+\frac{\partial z_{3,y}}{\partial a}\frac{da_{n}}{dt}>0.$
+$\frac{\partial z_{3,y}}{\partial a}-\frac{\partial W_{n-n_{0}}}{\partial a}>0.$
 Proof.
-Using the estimates in Proposition 2, one can notice that the first coefficient of $\partial z_{3,y}/{\partial a}$ is strictly larger than one, while the first coefficient of $\partial W_{n-n_{0}}/{\partial a}$ is strictly smaller than one, see Lemma 2.
-The first inequality follows.
-For the second one, if $(t,a)\in\mathcal{B}_{n}$, by Proposition 2 and (2.11),
-$\frac{\partial z_{3,y}}{\partial t}+\frac{\partial z_{3,y}}{\partial a}\frac{da_{n}}{dt}=DC^{2}n\left(\lambda^{\theta}\mu\right)^{2n}\left[\frac{\theta}{\lambda}\frac{\partial\lambda}{\partial t}+\frac{1}{\mu}\frac{\partial\mu}{\partial t}\right]+O\left(\left(\lambda^{\theta}\mu\right)^{2n}\right).$
-The statement is proved using Proposition 2 and (2.18). ∎
+Using the estimates in Proposition 4.2.2, one notices that the leading term of $\partial z_{3,y}/{\partial a}$ is proportional to $\left(\lambda^{\theta}\mu^{2}\right)^{n}$ which is strictly larger than one, while $\partial W_{n-n_{0}}/{\partial a}$ is strictly smaller than one, see Lemma 2.2.5, for $n$ large enough.
+The inequality follows.
+∎
 The following lemma shows that one cannot find a secondary tangency of type $n_{0}$ on the line defining the lower bound of the $\mathcal{B}_{n}$ strip.
-Lemma 15.
-There are no curve secondary tangencies of type $n_{0}$ along the line $a_{n}-E{\lambda^{\theta n}}$ where $E=C-\frac{\epsilon}{\mu^{n_{0}/2}}$, with $\epsilon>0$ small.
+Lemma 5.2.2.
+For all $\epsilon>0$ and for $n$ large enough, there are no secondary tangencies of type $n_{0}$ along the line $a_{n}-E{\lambda^{\theta n}}$ where $E=C-\frac{\epsilon}{\mu^{n_{0}^{*}/2}}$.
 Proof.
-By contradiction, there exists a $n_{0}$ tangency on the line $a_{n}-E{\lambda^{\theta n}}$ with $E=C-\frac{\epsilon}{\mu^{n_{0}/2}}$, $\epsilon>0$ small.
-This means that $z_{3,y}\mu^{n_{0}}=2$.
-By (4.6) and Lemma 7,
-$z_{2,y}=C\left(\lambda^{\theta}\mu\right)^{n}+\mu^{n}(a-a_{n}(t))+O\left(\mu^{n}\lambda^{2\theta n}\right)$
-and by (4.19)
-$z_{3,y}=Q\left(z_{2,y}\right)^{2}+O\left(\frac{1}{\mu^{n}}\right),$
-Hence,
-$z_{3,y}\mu^{n_{0}}$ $=$ $Q(C-E)^{2}\left(\lambda^{\theta}\mu\right)^{2n}\mu^{n_{0}}+O\left(\frac{1}{\mu^{n-n_{0}}}\right)$ (5.7)
-$=$ $Q\frac{\epsilon^{2}}{\mu_{0}^{n_{0}}}\left(\lambda^{\theta}\mu\right)^{2n}\mu_{0}^{n_{0}}+O\left(\frac{1}{\mu^{n-n_{0}}}\right)$
-$=$ $Q{\epsilon^{2}}\left(\lambda^{\theta}\mu\right)^{2n}+O\left(\frac{1}{\mu^{n-n_{0}}}\right).$
-Since $\lambda^{\theta}\mu<1$, for $\epsilon$ small enough, this last quantity cannot be equal to $2$. The proof is completed. ∎
-The following lemma shows that one cannot find a $n_{0}$ tangency on the line $a_{n}+\frac{T}{\mu^{n+n_{0}/2}}$ with $T$ a large constant.
-Lemma 16.
-For $T$ large enough, there are no secondary tangencies of type $n_{0}$ along the line $a_{n}+\frac{T}{\mu^{n+n_{0}/2}}$, with $T$ large.
+By contradiction, there exists a $n_{0}$ tangency on the line $a_{n}-E{\lambda^{\theta n}}$ with $E=C-\frac{\epsilon}{\mu^{n_{0}^{*}/2}}$.
+This means that $z_{y}^{(3)}\mu^{n_{0}}=2+O(\lambda^{n_{0}})$.
+By Lemma 4.1.23,
+$z_{y}^{(3)}(t,a)=QH^{2}\left[1+O\left(H\right)\right]+z_{y}(t,a)=QH^{2}\left[1+O\left(H\right)\right]+a.$
+where $H=C\left(\lambda^{\theta}\mu\right)^{n}+\left(\mu^{n}+O(n)\right)\left(a-a_{n}(t)\right)=(C-E)\left(\lambda^{\theta}\mu\right)^{n}\left(1+O\left({n}/{\mu^{n}}\right)\right)$.
+Hence, by (2.3.1), (4.0.1) and Lemma 4.1.48
+$z_{y}^{(3)}\mu^{n_{0}}$ $=$ $Q(C-E)^{2}\left(\lambda^{\theta}\mu\right)^{2n}\mu^{n_{0}}+O\left(\frac{1}{\mu^{n-n_{0}}}\right)$ (5.2.3)
+$=$ $Q\frac{\epsilon^{2}}{\mu^{n_{0}^{*}}}\left(\lambda^{\theta}\mu\right)^{2n}\mu^{n_{0}}+O\left(\frac{1}{\mu^{n-n_{0}}}\right)$
+$=$ $O\left(\left(\lambda^{\theta}\mu\right)^{2n}\mu^{\alpha n-\frac{1}{2}\alpha n}\right)+O\left(\frac{1}{\mu^{n-n_{0}}}\right)$
+$=$ $O\left(\left(\lambda^{3\theta}\mu^{7/4}\right)^{2n}\right).$
+Since $\lambda^{3\theta}\mu^{7/4}<1$, see (2.3.1), this last quantity cannot be close to $2$. Contraddiction.The proof is complete. ∎
+The following lemma shows that one cannot find a $n_{0}$ tangency on the line $a_{n}+\frac{T}{\mu^{n+n_{0}^{*}/2}}$ with $T$ a large constant.
+Lemma 5.2.4.
+For $T$ large enough, there are no secondary tangencies of type $n_{0}$ along the line $a_{n}+\frac{T}{\mu^{n+n_{0}^{*}/2}}$.
 Proof.
-As in the proof of the previous lemma, by (4.6) and Lemma 7,
-$z_{2,y}=C\left(\lambda^{\theta}\mu\right)^{n}+\mu^{n}(a-a_{n}(t))+O\left(\mu^{n}\lambda^{2\theta n}\right)$
-and by (4.19)
-$z_{3,y}=Q\left(z_{2,y}\right)^{2}+O\left(\frac{1}{\mu^{n}}\right)$ $=$ $Q\left(C\left(\lambda^{\theta}\mu\right)^{n}+\frac{T}{\mu^{n_{0}/2}}\right)^{2}+O\left(\frac{1}{\mu^{n}}\right)$
-$>$ $Q\frac{T^{2}}{\mu^{n_{0}}}+O\left(\frac{1}{\mu^{n}}\right).$
-Hence,
-$z_{3,y}\mu^{n_{0}}$ $>$ $QT^{2}+O\left(\frac{1}{\mu^{n-n_{0}}}\right).$ (5.8)
-For $T$ large enough, this last quantity cannot be equal to $2$, so in particular one cannot have a $n_{0}$ tangency. The proof is completed. ∎
-5.3 Global curves of secondary tangencies
-In the next proposition we prove the existence of global curves of secondary tangencies contained in $\left([-t_{0},t_{0}]\times\mathbb{R}\right)\cap\mathcal{B}_{n}$, see Figure 6.
-Proposition 4.
+As in the proof of the previous lemma, By Lemma 4.1.23,
+$z_{y}^{(3)}(t,a)=QH^{2}\left[1+O\left(H\right)\right]+z_{y}(t,a)=QH^{2}\left[1+O\left(H\right)\right]+a.$
+where $H=C\left(\lambda^{\theta}\mu\right)^{n}+\left(\mu^{n}+O(n)\right)\left(a-a_{n}(t)\right)=\frac{T}{\mu^{n_{0}^{*}/2}}\left(1+O\left(\left(\lambda^{\theta}\mu\right)^{n}\mu^{n_{0}^{*}/2}\right)\right)$.
+Hence, by (2.3.1), and Lemma 4.1.48
+$z_{y}^{(3)}\mu^{n_{0}}>\frac{QT^{2}}{2\mu^{n_{0}}}\mu^{n_{0}}+O\left(\frac{1}{\mu^{n}}\mu^{n_{0}}\right)=\frac{QT^{2}}{2}+O\left(\frac{1}{\mu^{n-n_{0}}}\right).$ (5.2.5)
+For $T$ large enough, this last quantity cannot be close to $2$, so in particular one cannot have a $n_{0}$ tangency. The proof is complete. ∎
+5.3 Curves of secondary tangencies with uniform size
+In the next proposition we prove the existence of curves of secondary tangencies which are graphs of functions defined over the full domain $t_{0},t_{0}]$, see Figure 6.
+Proposition 5.3.1.
 There exists a smooth function $b_{n,n_{0}}:[-t_{0},t_{0}]\to\mathbb{R}$ such that for all $t\in[-t_{0},t_{0}]$ the parameter point $(t,b_{n,n_{0}}(t))$ corresponds to a map with a secondary tangency of type $n_{0}$.
@@
 Proof.
-Let $(t^{*},a^{*})\in\mathcal{B}_{n}$ be a parameter corresponding to a map with a secondary tangency of type $n_{0}$. By Lemma 13 there exist $\epsilon>0$ and a smooth curve
-$b_{n,n_{0}}:[t^{*}-\epsilon,t^{*}+\epsilon]\mapsto\mathbb{R}$
-such that, for all $t\in[t^{*}-\epsilon,t^{*}+\epsilon]$ the parameter $(t,b_{n,n_{0}}(t))$ corresponds to a map with a secondary tangency of type $n_{0}$.
+Let $(t_{n_{0}},a_{n_{0}})\in\mathcal{B}_{n}$ be a parameter corresponding to a map with a secondary tangency of type $n_{0}$. By Lemma 5.1.1 there exist $\epsilon>0$ and a smooth curve
+$b_{n,n_{0}}:[t_{n_{0}}-\epsilon,t_{n_{0}}+\epsilon]\mapsto\mathbb{R}$
+such that, for all $t\in[t_{n_{0}}-\epsilon,t_{n_{0}}+\epsilon]$ the parameter $(t,b_{n,n_{0}}(t))$ corresponds to a map with a secondary tangency of type $n_{0}$.
 We want now to prove that this curve does not stop.
 Suppose it does.
-Consider a sequence $t_{n}\in[t^{*}-\epsilon,t^{*}+\epsilon]$ of secondary tangencies of type $n_{0}$ that converges to a boundary point of $t^{*}-\epsilon,t^{*}+\epsilon]$.
-Since the set of tangencies is a close set, the limit point is again a tangency.
-Because of Lemma 14 one cannot have more than one limit point, indeed one can have only one tangency point while moving in the $a$ direction.
-By lemmas 15 and 16 this unique limit point cannot be on the boundary lines that define $\mathcal{B}_{n}$.
+Let $(t_{-},t_{+})$ be the maximal interval in $t_{0},t_{0}]$ to which this function extends smoothly such that its graph is a locus of secondary tangency of type $n_{0}$.
+Suppose that $t_{-},t_{+}]\subset(-t_{0},t_{0})$.
+Consider a sequence $t_{n}\in(t_{-},t_{+})$ of secondary tangencies of type $n_{0}$ that converges to a boundary point in $t_{-},t_{+}]$.
+Since the set of tangencies is a close set, all limit points are again a tangency.
+Because of Lemma 5.2.1 one cannot have more than one limit point, indeed one can have only one tangency point while moving in the $a$ direction.
+By lemmas 5.2.2 and 5.2.4 this unique limit point cannot be on the boundary lines that define $\mathcal{B}_{n}$.
 Such a limit point is then contained in $\mathcal{B}_{n}$.
-One can then apply again Proposition 13 to this point and build a smooth continuous curve of maps with secondary tangencies of type $n_{0}$.
-This new courve $b^{\prime}_{n,n_{0}}$ coincides with the initial one $b_{n,n_{0}}$, because of the continuity of the implicit function theorem used to extend the curve in Proposition 13.
-In this way one creates a curve which is a graph of a smooth function.
-This is obtained by using the angle estimate in Proposition 4 and Lemma 14.
-Because of lemmas 15 and 16 the curve built cannot exit the interior of $\mathcal{B}_{n}$.
+One can then apply again Proposition 5.1.1 to this point and build a smooth curve of maps with secondary tangencies of type $n_{0}$.
+This new courve $b^{\prime}_{n,n_{0}}$ coincides with the initial one $b_{n,n_{0}}$, because of the uniqueness assured by the implicit function theorem used to create the local curve in Proposition 5.1.1.
+This contradicts the maximality of the interval $(t_{-},t_{+})$.
+In particular the function extends on the all domain $t_{0},t_{0}]$ without exiting $\mathcal{B}_{n}$.
+∎
+In the following lemma we prove that the estimate for the angle between the curve $b_{n,n_{0}}$ and $a_{n}$, see Lemma 5.1.1, persists in a neighborhood of $a_{n}$ of width $O\left(1/\mu^{2n}\right)$.
+Lemma 5.3.2.
+For all $t\in[-t_{0},t_{0}]$, with $(b_{n,n_{0}}(t)-a_{n}(t))=O\left(1/\mu^{2n}\right)$,
+$\frac{db_{n,n_{0}}}{dt}=\frac{da_{n}}{dt}+{n\lambda^{\theta n}}C\left[\frac{\theta}{\lambda}\frac{\partial\lambda}{\partial t}+\frac{1}{\mu}\frac{\partial\mu}{\partial t}\right]+O\left(\lambda^{\theta n}\right).$
+Proof.
+Observe that, by (2.3.1), (4.2.6) holds for all parameters $(t,a)$ with $a=a_{n}+O\left(1/\mu^{2n}\right)$.
+As a consequence (4.2) and (4.2.31) hold as well and this implies that (5.1.3) and (5.1.4) also hold.
+In particular (5.1.6) holds for all parameters $(t,a)$ with $a=a_{n}+O\left(1/\mu^{2n}\right)$.
+The proof is complete.
 ∎
 In the next lemma we prove that there exists a neighborhood of the curve of secondary tangencies built above such that the initial family restricted to this neighborhood is an unfolding of the secondary tangency.
-Lemma 17.
+Lemma 5.3.3.
 Let $b_{n,n_{0}}:[-t_{0},t_{0}]\to\mathbb{R}$ be a curve of secondary tangencies. There exists a neighborhood $U$ of $b_{n,n_{0}}$ such that $F_{|U}$ can be reparametrized to become an unfolding.
@@
 The transversal intersection $q_{2}$ it is still present in the new family.
-As a consequence, conditions (f1),(f2),(f3),(f4),(f5),(f6),(F1),(F2),(F3),(F5)(f1),(f2),(f3),(f4),(f5),(f6),(F1),(F2),(F3),(F5) are automatically satisfied.
-Observe that $f_{n,n_{0}}^{n-n_{0}}\left(q_{1,n,n_{0}}\right)$ converges to $q_{2}$. Property $(f6)$ implies $(f7)$ for the secondary tangency $q_{1,n,n_{0}}$.
-Use the notation $m_{1},m_{2},m_{3},m_{4},m_{5}$ from Proposition 2 and observe that $m_{1}\in W^{u}_{\text{loc}}(0)$ close to $q_{3}$. This proves $(f8)$ for the secondary tangency $q_{1,n,n_{0}}$.
+As a consequence, conditions (f1),(f2),(f3),(f5),(f6),(F1),(F2),(F3),(F5)(f1),(f2),(f3),(f5),(f6),(F1),(F2),(F3),(F5) are automatically satisfied.
+Condition $(f4)$ is a consequence of Lemma 4.2.36.
+By construction, condition $(f6)$ implies $(f7)$ for the secondary tangency $q_{1,n,n_{0}}$.
+Use the notation from Proposition 4.2.2 and observe that $m_{1}\in W^{u}_{\text{loc}}(0)$ close to $q_{3}$ such that $F^{3N+\theta n+n+n_{0}}(m_{1})=q_{1,n,n_{0}}$.
+This proves $(f8)$ for the secondary tangency $q_{1,n,n_{0}}$.
 For proving $(F4)$ use the ${\mathcal{C}}^{2}$} function $b_{n,n_{0}}$ and observe that the maps on the graph of this curve, $b_{n,n_{0}}$, have a non degenerate homoclinic tangency.
-Observe that $(P1)$ follows by the fact that $b_{n,n_{0}}$ is the curve of tangencies and $(P2)$ from the fact that $\Psi_{1,3}\neq 0$, see Lemma 13, Equation (5.1). ∎
+Observe that $(P1)$ follows by the fact that $b_{n,n_{0}}$ is the curve of non-degenerate tangencies, see Lemma 4.2.36 and $(P2)$ from Lemma 5.2.1. ∎
 6 Newhouse phenomenon
-In this section, we identify parameters corresponding to maps that exhibit infinitely many sinks, a phenomenon referred to as the Newhouse phenomenon.
-Although this result itself is not new, our approach is constructive in nature and departs fundamentally from earlier methods, which in turn enables us to establish a novel and unexpected property of unfoldings.
+In this section, we identify parameters corresponding to maps that have infinitely many sinks of arbitrarily high period, a phenomenon referred to as the Newhouse phenomenon.
+Although this result itself is not new, our approach is constructive in nature and departs fundamentally from earlier methods, which in turn enables us to describe the dynamics on the closure of these sinks establishing a novel and unexpected property of unfoldings.
 We start by proving the existence of sinks in two-dimensional unfoldings.
 6.1 Cascades of sinks
-In this section we are going to prove that, for a two dimensional set of parameters, an unfolding contains maps which have a sink of high period.
-Choose $\epsilon_{0}>0$ and define, for $n$ large enough,
+In this section we are going to prove that, for a specific open set of parameters, the corresponding maps have a sink of high period. Choose $\epsilon_{0}>0$ and define, for $n$ large enough,
 $\mathcal{A}_{n}=\left\{(t,a)\in[-t_{0},t_{0}]\times[-a_{0},a_{0}]\left|\right.|a-a_{n}(t)|\leq\frac{\epsilon_{0}}{\mu(t,a_{n}(t))^{2n}}\right\}.$
@@
 The following remark describes the error obtained in the change of the eigenvalues while changing the parameters within $\mathcal{A}_{n}$.
-Remark 11.
-Since $\partial\mu/\partial a$, $\partial\lambda/\partial t$ and $\partial\lambda/\partial a$ are all bounded we have that, for all $(t,\tilde{a})\in\mathcal{A}_{n}$,
-$\left(\frac{\mu(t,a_{n}(t))}{\mu(t,\tilde{a})}\right)^{n}=1+O\left(\frac{n}{\mu(t,a_{n}(t))^{2n}}\right),\left(\frac{\lambda(t,a_{n}(t))}{\lambda(\tilde{t},\tilde{a})|}\right)^{n}=1+O\left(\frac{n}{\mu(t,a_{n}(t))^{2n}}\right).$
+Remark 6.1.1.
+As in Remark 4.0.3, for all $(t,\tilde{a})\in\mathcal{A}_{n}$,
+$\left(\frac{\mu(t,a_{n}(t))}{\mu(t,\tilde{a})}\right)^{n}=1+O\left(\frac{n}{\mu(t,a_{n}(t))^{2n}}\right),\left(\frac{\lambda(t,a_{n}(t))}{\lambda(\tilde{t},\tilde{a})}\right)^{n}=1+O\left(\frac{n}{\mu(t,a_{n}(t))^{2n}}\right).$
 In the following we prove, for a properly chosen $\epsilon_{0}$, that for all parameters $(t,a)\in\mathcal{A}_{n}$, $F_{t,a}$ has a sink of period $n+N$ which we call primary sink.
@@
 When the choice of $(t,a)$ is clear we just use the notation $B^{n}_{\delta}$.
-In the next lemma we prove that $B^{n}_{\delta}$ returns into itself, see Figure 7. Let $\tilde{Q}=\max Q_{x,y}$ where $Q_{x,y}$ as in Lemma 5.
+In the next lemma we prove that $B^{n}_{\delta}$ returns into itself, see Figure 7.
+Let $\tilde{Q}=\max Q_{x,y}$ where $Q_{x,y}$ is as in Lemma 2.2.12.
 Figure 7: Invariant Region
-Lemma 18.
+Lemma 6.1.2.
 Choose $\delta=\frac{1}{4\tilde{Q}},\epsilon_{0}=\frac{1}{32\tilde{Q}}$. Then, for $n$ large enough and $(t,a)\in\mathcal{A}_{n}$,
@@
 Proof.
-[…]
-$|\Delta x|\leq\frac{1}{3},|\Delta y|\leq\frac{\delta}{\mu(t,a_{n}(t))^{2n}}\text{ and }|\Delta a|\leq\frac{\epsilon_{0}}{\mu(t,a_{n}(t))^{2n}}.$ (6.1)
-Denote by $\left(\Delta x^{\prime},\Delta y^{\prime}\right)=F^{n}_{t,a}\left(\Delta x,\Delta y\right)-c_{t,a}$. Using that $F^{n}_{t,a}$ is linear on $B^{n}_{\delta}$, see (2.4), we get
-$|\Delta x^{\prime}|\leq 2\lambda(t,a)^{n},$
+Fix $(t,a)\in\mathcal{A}_{n}$.
+For $n$ large enough, we write $F^{n+N}_{t,a}$ in coordinates centered at $z_{n}$, namely $F^{n+N}_{t,a}\left(\Delta x_{z_{n}},\Delta y_{z_{n}}\right)=\left(\Delta\tilde{x}_{z_{n}},\Delta\tilde{y}_{z_{n}}\right)$.
+Let $\left(\Delta x_{z_{n}},\Delta y_{z_{n}}\right)$ be such that if $z_{n}+\left(\Delta x_{z_{n}},\Delta y_{z_{n}}\right)\in B^{n}_{\delta}$ and $\Delta a=(a-a_{n}(t))$, then
+$|\Delta x_{z_{n}}|\leq\frac{1}{3},|\Delta y_{z_{n}}|\leq\frac{\delta}{\mu(t,a_{n}(t))^{2n}}\text{ and }|\Delta a|\leq\frac{\epsilon_{0}}{\mu(t,a_{n}(t))^{2n}}.$ (6.1.3)
+Denote by $\left(\Delta x_{c_{t,a}},\Delta y_{c_{t,a}}\right)=F^{n}_{t,a}\left(\Delta x_{z_{n}},\Delta y_{z_{n}}\right)-c_{t,a}$.
+Using that $F^{n}_{t,a}$ is linear on $B^{n}_{\delta}$, see (2.1.8), we get
+$|\Delta x_{c_{t,a}}|\leq 2\lambda(t,a)^{n},$
 and
-$|\Delta y^{\prime}|$ $\leq$ $\mu(t,a)^{n}\left|\Delta y\right|+\left|\mu(t,a)^{n}z_{n,y}-c(0,t,a)\right|$
-$\leq$ $\mu(t,a)^{n}\left|\Delta y\right|+\left|\left(\left[1+O\left(\frac{n}{\mu(t,a_{n}(t))^{2n}}\right)\right]\mu(t,a_{n}(t))^{n}z_{n,y}\right)-c_{n,y}\right|$
-$+$ $\left|c_{n,y}-c(0,t,a)\right|,$
-where we used Remark 11. Since $F^{n}_{t,a_{n}(t)}\left(z_{n}\right)\in\Gamma_{t,a_{n}(t)}$, then
-$\mu(t,a_{n}(t))^{n}z_{n,y}=c_{n,y}+O\left(\lambda(t,a_{n}(t))^{n}\right).$
-Moreover $\left|c_{n,y}-c(0,t,a)\right|=O\left(\Delta a\right)$ (see Lemma 1). Hence, by (6.1) and by Remark 11 we get
-$|\Delta y^{\prime}|$ $\leq$ $\mu(t,a)^{n}\left|\Delta y\right|+O\left(\lambda(t,a_{n}(t))^{n}\right)+O\left(\frac{n}{\mu(t,a_{n}(t))^{2n}}\right)+O\left(\frac{1}{\mu(t,a_{n}(t))^{2n}}\right)$
+$|\Delta y_{c_{t,a}}|$ $\leq$ $\mu(t,a)^{n}\left|\Delta y_{z_{n}}\right|+\left|\mu(t,a)^{n}z_{n,y}-c(0,t,a)\right|$
+$\leq$ $\mu(t,a)^{n}\left|\Delta y_{z_{n}}\right|+\left|\left(\left[1+O\left(\frac{n}{\mu(t,a_{n}(t))^{2n}}\right)\right]\mu(t,a_{n}(t))^{n}z_{n,y}\right)-1\right|,$
+where we used Remark 6.1.1. Since $F^{n}_{t,a_{n}(t)}\left(z_{n}\right)\in\Gamma_{t,a_{n}(t)}$, then
+$\mu(t,a_{n}(t))^{n}z_{n,y}=1+O\left(\lambda(t,a_{n}(t))^{n}\right).$
+Hence, by (6.1.3) and by Remark 6.1.1 we get,
+$|\Delta y_{c_{t,a}}|$ $\leq$ $\mu(t,a)^{n}\left|\Delta y_{z_{n}}\right|+O\left(\lambda(t,a_{n}(t))^{n}\right)+O\left(\frac{n}{\mu(t,a_{n}(t))^{2n}}\right)$
 $\leq$ $\frac{\delta}{\mu(t,a_{n}(t))^{n}}\left[1+O\left(\frac{n}{\mu(t,a_{n}(t))^{2n}}\right)\right]+O\left(\frac{n}{\mu(t,a_{n}(t))^{2n}}\right).$
-By Lemma 5 (center $F^{N}_{t,a}$ in coordinates around $c_{t,a}$) extended to include also the Taylor expansion in $\Delta a$ we get, for $n$ large enough
-[…]
+By Lemma 2.2.12 (center $F^{N}_{t,a}$ in coordinates around $c_{t,a}$) extended to include also the Taylor expansion in $\Delta a$ we get, for $n$ large enough
+$|\Delta\tilde{x}_{z_{n}}|$ $=$ $O\left(\Delta x_{c_{t,a}}\right)+O\left(\Delta y_{c_{t,a}}\right)+O\left(\Delta a\right)\leq\frac{1}{3},$
 and
-[…]
-$\leq$ $\frac{\tilde{Q}\delta^{2}}{\mu(t,a_{n}(t))^{2n}}+O\left(\frac{1}{\mu(t,a_{n}(t))^{3n}}\right)+\frac{\epsilon_{0}}{\mu(t,a_{n}(t))^{2n}},$
-where we also used $(F3)$ and Remark 11. By our choice of $\epsilon_{0}$ and $\delta$, for $n$ large enough, the lemma follows. ∎
-We fix $\epsilon_{0},$ and $\delta$ such that Lemma 18 holds.
+$|\Delta\tilde{y}_{z_{n}}|$ $\leq$ $O\left(\lambda(t,a)^{n}\right)+Q_{x,y}\left|\Delta y_{c_{t,a}}\right|^{2}+O\left(\frac{1}{\mu(t,a_{n}(t))^{3n}}\right)+|\Delta a|$
+$\leq$ $\frac{\tilde{Q}\delta^{2}}{\mu(t,a_{n}(t))^{2n}}+O\left(\frac{n}{\mu(t,a_{n}(t))^{3n}}\right)+\frac{\epsilon_{0}}{\mu(t,a_{n}(t))^{2n}},$
+where we also used $(F3)$ and Remark 6.1.1. By our choice of $\epsilon_{0}$ and $\delta$, for $n$ large enough, the lemma follows. ∎
+We fix $\epsilon_{0},$ and $\delta$ such that Lemma 6.1.2 holds.
 We are now ready to prove the existence of a sink.
 This is achieved in the next proposition by proving that $F^{n+N}$ contracts the box $B_{\delta}^{n}$.
-Proposition 5.
-For $n$ large enough and for all $(t,a)\in\mathcal{A}_{n}$, $B^{n}_{\delta}(t)$ has a unique periodic point which is a sink.
+Proposition 6.1.4.
+For $n$ large enough and for all $(t,a)\in\mathcal{A}_{n}$, $B^{n}_{\delta}(t)$ has a unique periodic point which is a sink of period $n+N$.
 Proof.
 Because $F_{t,a}^{n}$ is linear on $B_{\delta}^{n}$, the image $F_{t,a}^{n}\left(B_{\delta}^{n}\right)$ is contained in a neighborhood of $c_{t,a}$ of diameter smaller than $\delta\mu^{-n}\left(1+O\left({n}\mu^{-n}\right)\right)$.
-From this and by differentiating (2.15) with respect to $\Delta x$ and $\Delta y$ we get
+From this and by differentiating (2.2.13) with respect to $\Delta x$ and $\Delta y$ we get
 $DF_{t,a}^{N}$ $=$ $\left(\begin{matrix}O(1)&O(1)\\ O(1)&2{\delta\mu^{-n}Q_{x,y}}\left(1+O\left({n}{\mu^{-n}}\right)\right)\\ \end{matrix}\right).$
 Note that $N$ is fixed. Using again that $F_{t,a}^{n}$ is linear on $B_{\delta}^{n}$ we obtain that
-$DF_{t,a}^{n+N}$ $=$ $\left(\begin{matrix}O(1)&O(1)\\ O(1)&{2\delta}{\mu^{-n}}Q_{x,y}\left(1+O\left({n}{\mu^{-n}}\right)\right)\end{matrix}\right)\left(\begin{matrix}O\left(\lambda^{n}\right)&0\\ 0&\mu^{n}\end{matrix}\right)$
+$DF_{t,a}^{n+N}$ $=$ $\left(\begin{matrix}O(1)&O(1)\\ O(1)&{2\delta}{\mu^{-n}}Q_{x,y}\left(1+O\left({n}{\mu^{-n}}\right)\right)\end{matrix}\right)\left(\begin{matrix}\lambda^{n}&0\\ 0&\mu^{n}\end{matrix}\right)$
 $=$ $\left(\begin{matrix}O\left(\lambda^{n}\right)&O\left(\mu^{n}\right)\\ O\left(\lambda^{n}\right)&2\delta Q_{x,y}\left(1+O\left({n}{\mu^{-n}}\right)\right)\end{matrix}\right).$
@@
 We are now ready to set up an induction procedure to prove the existence of maps having infinitely many sinks.
-This procedure uses the so-called “Newhouse boxes”.
+This procedure uses “Newhouse boxes”.
 In the first generation, the Newhouse boxes are essentially rectangles in $\mathcal{A}_{n}$ whose boundary are defined by the curves of secondary tangencies, see Figure 6.
@@
 We proceed by induction.
-Let $(t_{n,n_{0}},a_{n,n_{0}})$ be the parameters at the intersection point $a_{n}\cap b_{n,n_{0}}$, see Proposition 4, and let
+Let $(t_{n,n_{0}},a_{n,n_{0}})$ be the parameters at the intersection point $a_{n}\cap b_{n,n_{0}}$, see Proposition 5.3.1, and let
 $f_{n,n_{0}}=F_{t_{n,n_{0}},a_{n,n_{0}}}.$
@@
 Consider its restriction to the interval $t^{-}_{n,n_{0}},t^{+}_{n,n_{0}}$] where $t^{-}_{n,n_{0}}$ and $t^{+}_{n,n_{0}}$ are the two intersection points of the curve $b$ with the boundary of the strip $\mathcal{A}_{n}$, see Figure 6.
-The domains
-$\mathcal{P}_{n,n_{0}}=\left\{(t,a)\left|\right.t\in[t^{-}_{n,n_{0}},t^{+}_{n,n_{0}}],|a-a_{n}(t)|\leq\frac{\epsilon_{0}}{\mu(t,a_{n}(t))^{2n}}\right\}$
+The existence of these intersection points is guaranteed by Lemma 5.3.2.
+The domains,
+$\mathcal{P}_{n,n_{0}}=\left\{(t,a)\left|\right.t\in[t^{-}_{n,n_{0}},t^{+}_{n,n_{0}}],|a-a_{n}(t)|\leq\frac{\epsilon_{0}}{\mu(t,a_{n}(t))^{2n}}\right\},$
 are called the Newhouse boxes of first generation.
@@
 The following inductive construction will repeat the discussion of the previous sections starting with the map $f_{n,n_{0}}:\mathcal{M}\to\mathcal{M}$ and an unfolding $F:\mathcal{P}_{n,n_{0}}\times\mathcal{M}\to\mathcal{M}$ which is the restriction of the original family.
-Lemma 19.
+Lemma 6.2.1.
 The domains $\mathcal{P}_{n,n_{0}}$ are pairwise disjoint for $n$ large enough and the diameter goes to zero.
 Proof.
-By Proposition 4 the curve $b_{n,n_{0}}$ is the graph of the function $b$ and by the angle estimate in Lemma 13, there exists a uniform constant $K>0$ such that
-$\frac{1}{K}\frac{1}{n\left(\lambda^{\theta}\mu^{2}\right)^{n}}\leq\left|\left\{t\left|\right.(t,b(t))\in\mathcal{A}_{n}\right\}\right|\leq K\frac{1}{n\left(\lambda^{\theta}\mu^{2}\right)^{n}}.$ (6.2)
-Moreover the proof of Proposition 3 gives $\text{dist}\left(f_{n,n_{0}},f_{n,n_{0}+1}\right)$ is proportional to 1/n{1}/{n}.
-The disjointness follows from this estimates, the fact that $\lambda^{\theta}\mu^{2}\geq\left(\lambda^{2\theta}\mu^{3}\right)^{1/2}>1$, see (2.16) and the fact that $\mathcal{A}_{n}$ are pairwise disjoint.
+By Proposition 5.3.1 the curve $b_{n,n_{0}}$ is the graph of the function $b$ and by the angle estimate in Lemma 5.3.2, there exists a uniform constant $K>0$ such that
+$\frac{1}{K}\frac{1}{n\left(\lambda^{\theta}\mu^{2}\right)^{n}}\leq\left|\left\{t\left|\right.(t,b(t))\in\mathcal{A}_{n}\right\}\right|\leq K\frac{1}{n\left(\lambda^{\theta}\mu^{2}\right)^{n}}.$ (6.2.2)
+Moreover the proof of Proposition 4.2.32 gives $\text{dist}\left(f_{n,n_{0}},f_{n,n_{0}+1}\right)$ is proportional to 1/n{1}/{n}.
+The disjointness follows from this estimates, the fact that $\lambda^{\theta}\mu^{2}\geq\left(\lambda^{2\theta}\mu^{3}\right)^{1/2}>1$, see (2.3.1) and the fact that $\mathcal{A}_{n}$ are pairwise disjoint.
 ∎
-The next proposition ensures that the family restricted to $\mathcal{P}_{n,n_{0}}$ is again an unfolding. This allows an inductive procedure.
-Proposition 6.
-For $n$ large enough, the map $f_{n,n_{0}}$ has a strong homoclinic tangency and the restriction $F:\mathcal{P}_{n,n_{0}}\times\mathcal{M}\to\mathcal{M}$ can be reparametrized to become an unfolding.
+The next proposition ensures that the family restricted to $\mathcal{P}_{n,n_{0}}$ is again an unfoldings. This allows an inductive procedure.
+Proposition 6.2.3.
+For $n$ large enough, the map $f_{n,n_{0}}$ has a strong homoclinic tangency and the restriction $F:\mathcal{P}_{n,n_{0}}\times\mathcal{M}\to\mathcal{M}$ can be reparametrized to become an unfoldings.
 Moreover each map in $\mathcal{P}_{n,n_{0}}$ has a sink of period $n+N$.
 Proof.
-The proof is given in Lemma 17. ∎
+The proof is given in Lemma 5.3.3 and Proposition 6.1.4. ∎
 Inductively we are going to construct parameters with multiple sinks of higher and higher periods and a strong homoclinic tangency.
-Let $\mathfrak{N}^{1}=\left\{(n,n_{0})\left|\right.n_{0}\in\mathfrak{N}_{n}\right\}\subset{\mathbb{N}}^{2}$ be the set of labels of the Newhouse boxes of first generation $\mathcal{P}_{n,n_{0}}$.
-As inductive hypothesis assume that there exist sets $\mathfrak{N}^{k}\in\left({\mathbb{N}}^{2}\right)^{k}$ such that, each $\mathfrak{N}^{k}$ is the set of labels of the Newhouse boxes of generation $k$.
-Moreover the natural projections $\mathfrak{N}^{1}\leftarrow\mathfrak{N}^{2}\leftarrow\dots\leftarrow\mathfrak{N}^{g}$ with $\mathfrak{N}^{k}\in\left({\mathbb{N}}^{2}\right)^{k}$ are countable to 11 and correspond to inclusion of Newhouse boxes of successive generations.
+Let $\mathfrak{N}^{1}$ be the set of labels of the Newhouse boxes of first generation, i.e.
+all possible pairs $(n,n_{0})$ for which there exists a Newhouse box $\mathcal{P}_{n,n_{0}}$.
+As inductive hypothesis assume that there exist sets $\mathfrak{N}^{k}\subset{\mathbb{N}}^{2}$ such that, each $\mathfrak{N}^{k}$ is the set of labels of the Newhouse boxes of generation $k$.
+Moreover the natural projections $\mathfrak{N}^{1}\leftarrow\mathfrak{N}^{2}\leftarrow\dots\leftarrow\mathfrak{N}^{g}$ with $\mathfrak{N}^{k}\in{\mathbb{N}}^{2}$ are countable to $1$ and correspond to inclusion of Newhouse boxes of successive generations.
 They satisfy the following inductive hypothesis.
-For $\underline{n}=\left\{(n^{(k)},n_{0}^{(k)})\right\}_{k=1}^{g}\in\mathfrak{N}^{g}$, there exist sets $\mathcal{P}^{k}_{n^{(k)},n_{0}^{(k)}}\subset\mathcal{P}$, where $n^{(k)}$ labels the sink and $n_{0}^{(k)}$ labels the secondary tangency, such that
+For $\underline{n}=\left\{(n^{(k)},n_{0}^{(k)})\right\}_{k=1}^{g}$, there exist sets $\mathcal{P}^{k}_{n^{(k)},n_{0}^{(k)}}\subset\mathcal{P}$, where $n^{(k)}$ labels the sink and $n_{0}^{(k)}$ labels the secondary tangency, such that
   • $\text{diam}\left(\mathcal{P}^{k}_{n^{(k)},n_{0}^{(k)}}\right)\leq\frac{1}{k}$,
@@
   • there exists a map $f^{k}_{n^{(k)},n_{0}^{(k)}}\in\mathcal{P}^{k}_{n^{(k)},n_{0}^{(k)}}$ which has a strong homoclinic tangency,
-  • the restriction $F:{\mathcal{P}^{k}_{n^{(k)},n_{0}^{(k)}}}\times\mathcal{M}\to\mathcal{M}$ can be reparametrized to become and unfolding of $f^{k}_{n^{(k)},n_{0}^{(k)}}$,
+  • the restriction $F:{\mathcal{P}^{k}_{n^{(k)},n_{0}^{(k)}}}\times\mathcal{M}\to\mathcal{M}$ can be reparametrized to become and unfoldings of $f^{k}_{n^{(k)},n_{0}^{(k)}}$,
   • every map in $\mathcal{P}^{k}_{n^{(k)},n_{0}^{(k)}}$ has at least $k$ sinks of different periods.
-By induction, using Proposition 6, we get an infinite sequence of sets $\mathfrak{N}^{k}$.
+By induction, using Proposition 6.2.3, we get an infinite sequence of sets $\mathfrak{N}^{k}$.
 Figure 8: Newhouse boxes
-Definition 6.
+Definition 6.2.4.
 The sets $\mathcal{P}^{g}_{n^{(g)},n_{0}^{(g)}}$ are called Newhouse boxes of generation $g$, see Figure 8, and
-$NH=\bigcap_{g}\bigcup_{\underline{n}\in\mathfrak{N}^{g}}\mathcal{P}^{g}_{n^{(g)},n_{0}^{(g)}}.$
+$NH=\bigcap_{g}\bigcup_{\mathfrak{N}^{g}}\mathcal{P}^{g}_{n^{(g)},n_{0}^{(g)}}.$
 The set $NH$, consisting of parameters for which the corresponding maps have infinitely many sinks, accumulates on the curve of the original tangency.
@@
 We have the following lemma.
-Lemma 20.
+Lemma 6.2.5.
 $\overline{NH}\supset[-t_{0},t_{0}]\times\left\{0\right\}.$
 In particular,
-$\overline{NH}\cap\mathcal{P}_{n,n_{0}}\supset\text{\rm graph}(b_{n,n_{0}}).$ (6.3)
+$\overline{NH}\cap\mathcal{P}_{n,n_{0}}\supset\text{\rm graph}(b_{n,n_{0}})\cap\mathcal{P}_{n,n_{0}}.$ (6.2.6)
 Proof.
-Given $(t,0)$, by Proposition 3, for every $n$ large enough, $a_{n}$ has a secondary tangency at $(t_{n},a_{n}(t_{n}))$ with $|t-t_{n}|=O\left(\frac{1}{n}\right)$.
+Given $(t,0)$, by Proposition 4.2.32, for every $n$ large enough, $a_{n}$ has a secondary tangency at $(t_{n},a_{n}(t_{n}))$ with $|t-t_{n}|=O\left(\frac{1}{n}\right)$.
 Hence, there exists a sequence of Newhouse boxes in $\mathfrak{N_{1}}$ accumulating at $(t,0)$.
@@
 Given any family $F$ of diffeomorphisms, we define the Newhouse set $NH_{F}$ as the set of parameters having infinitely many sinks. The upper Minkowski dimension is denoted by $MD$.
-Theorem 3.
+Theorem 6.2.7.
 Let $F:\mathcal{P}\times\mathcal{M}\to\mathcal{M}$ be an unfolding then
@@
 Proof.
-The inductive construction, using Proposition 6, implies that all maps in $NH$ have infinitely many sinks. From Lemma 19 we know that
-$\bigcup_{\underline{n}\in\mathfrak{N}^{g}}\mathcal{P}^{g}_{n^{(g)},n_{0}^{(g)}},$
+The inductive construction, using Proposition 6.2.3, implies that all maps in $NH$ have infinitely many sinks. From Lemma 6.2.1 we know that
+$\bigcup_{\mathfrak{N}^{g}}\mathcal{P}^{g}_{n^{(g)},n_{0}^{(g)}},$
 consists of countably many disjoint boxes.
@@
 Consider a sequence of first generation Newhouse boxes $\mathcal{P}_{n,n_{0}}\in\mathfrak{N_{1}}$ accumulating at $(t^{*},0)$.
-This is possible because of Lemma 20.
+This is possible because of Lemma 6.2.5.
 Choose $\epsilon>0$ and let $n$ be maximal such that $\epsilon\leq{\epsilon_{0}}/{\mu(t^{*},a^{*})^{2n}}$.
-Because (6.2), (6.3) and the fact that the vertical size of $\mathcal{P}_{n,n_{0}}$ is ${\epsilon_{0}}/{\mu(t^{*},a^{*})^{2n}}$, we need at least ${K}/{\lambda(t^{*},a^{*})^{\theta n}}$ balls of radius $\epsilon$ to cover $\overline{NH}\cap\mathcal{P}_{n,n_{0}}$.
+Because (6.2.2), (6.2.6) and the fact that the vertical size of $\mathcal{P}_{n,n_{0}}$ is ${\epsilon_{0}}/{\mu(t^{*},a^{*})^{2n}}$, we need at least ${K}/{\lambda(t^{*},a^{*})^{\theta n}}$ balls of radius $\epsilon$ to cover $\overline{NH}\cap\mathcal{P}_{n,n_{0}}$.
 As consequence
 $MD(NH)\geq\frac{\theta}{2}\max_{NH}\left(\frac{\log\frac{1}{\lambda}}{\log\mu}\right),$
-and $MD(NH)\geq{2}/{3}$, where we used (2.17). ∎
-Remark 12.
+and $MD(NH)\geq{2}/{3}$, where we used (2.3.2). ∎
+Remark 6.2.8.
 Observe that the estimate for the upper Minkowski dimension is not sharp. Other dimension estimates for maps with infinitely many sinks were obtained in [11, 55, 60].
 7 Dynamical Partitions, Directed Graphs and Strips
-In this section, we introduce the objects and the induction step (see Proposition 7) that will allow to build later the main player of this work: the lamination and the invariant Cantor set.
+In this section, we introduce the objects and the induction step (see Proposition 7.4.1) that will allow to build later the main player of this work: the lamination and the invariant Cantor set.
 7.1 Dynamical Partitions
-We give now the definition of a dynamical partition associated to a map. The reader can refer to the Figure 9. The definition models the combinatorics as described in Subsection 4.1.
+We give now the definition of a dynamical partition associated to a map. The reader can refer to Figure 9. The definition models the combinatorics as described in Subsection 4.1.
 Figure 9: $(1,N,M)$ dynamical partition of $f$
-Recall the rescaling done in Subsection 2.1.1 while introducing our families of unfoldings and let $Q$ be the domain of linearization.
-We denote by $W^{u}_{\text{loc}}(p)$ a connected component of $W^{u}(p)\cap Q$ and by $W^{s}_{\text{loc}}(p)$ a connected component of $W^{s}(p)\cap Q$.
-Definition 7.
+Recall the rescaling done in Subsection 2.1.1 while introducing our unfoldings and let $Q$ be the domain of linearization.
+We denote by $W^{u}_{\text{loc}}(p)$ the connected component of $W^{u}(p)\cap Q$ and by $W^{s}_{\text{loc}}(p)$ the connected component of $W^{s}(p)\cap Q$ both containing $p$.
+Definition 7.1.1.
 Let $\mathcal{M}$ be a $m$-dimensional manifold. A map $f:\mathcal{M}\to\mathcal{M}$ has a $(1,N,M)$ dynamical partition
@@
     • $f:C^{N-1}\to f\left(C^{N-1}\right)=C^{N}$ is a diffeomorphism,
-    • $f\left(C^{N-1}\right)\subset Q\setminus f(Q)$,
-    • $C^{0}\cap W^{u}_{\text{loc}}(p)\neq\emptyset$, where $W^{u}_{\text{loc}}(p)$ is a connected component of $W^{u}(p)\cap Q$,
-    • $f\left(C^{N-1}\right)\cap W^{s}_{\text{loc}}(p)\neq\emptyset$.
+    • $C^{N}\subset Q\setminus f(Q)$,
+    • $C^{0}\cap f\left(W^{u}_{\text{loc}}(p)\right)\neq\emptyset$,
+    • $C^{N}\cap W^{s}_{\text{loc}}(p)\neq\emptyset$.
   4. 4.
@@
     • $f:H^{M-1}\to f\left(H^{M-1}\right)=H^{M}$ is a diffeomorphism,
-    • $f\left(H^{M-1}\right)\subset Q\setminus f(Q)$,
-    • $H^{0}\cap W^{u}_{\text{loc}}(p)\neq\emptyset$, where $W^{u}_{\text{loc}}(p)$ is a connected component of $W^{u}(p)\cap Q$,
-    • $f\left(H^{M-1}\right)\cap W^{s}_{\text{loc}}(p)\neq\emptyset$,
-    • $f^{M}\left(H_{0}\cap W^{u}_{\text{loc}}(p)\right)$ intersects transversally $W^{u}_{\text{loc}}(p)$ in the point $q_{2}$.
-Definition 8.
+    • $H^{M}\subset Q\setminus f(Q)$,
+    • $H^{0}\cap f\left(W^{u}_{\text{loc}}(p)\right)\neq\emptyset$,
+    • $H^{M}\cap W^{s}_{\text{loc}}(p)\neq\emptyset$,
+    • $f^{M}\left(H_{0}\cap f\left(W^{u}_{\text{loc}}(p)\right)\right)$ intersects transversally $W^{u}_{\text{loc}}(p)$ in the point $q_{2}$.
+Definition 7.1.2.
 Let $\mathcal{M}$ be a $m$-dimensional manifold, $f:\mathcal{M}\to\mathcal{M}$ a map and let $\mathfrak{D}$ be the $(1,N,M)$ dynamical partition of $f$. We define the mesh of $\mathfrak{D}$ as
 $\text{mesh}\left(\mathfrak{D}\right)=\max_{U\in\mathfrak{D}}\left(\text{diam(U)}\right),$
-where diam(U) is the diameter of $U$.
-Moreover we denote by $\mathcal{D}$ the union of elements in $\mathfrak{D}$. Namely,
+where diam(U) is the diameter of $U$. Moreover we denote by $\mathcal{D}$ the union of elements in $\mathfrak{D}$. Namely,
 $\mathcal{D}=\bigcup_{U\in\mathfrak{D}}U.$
@@
 Fixed two positive natural numbers, $N$ and $M$, we define the $(1,N,M)$ directed graph. The reader can refer to Figure 10. As before this models the combinatorics as described in Subsection 4.1.
-Definition 9.
+Definition 7.2.1.
 Given two positive natural numbers, $N$ and $M$, a directed graph $X$ is called the $(1,N,M)$ directed graph if it has $1+N+M$ vertices, $q,c^{0},\dots,c^{N-1},h^{0},\dots,h{M-1}$ and unique edges:
@@
 The sub graph formed by $q,c0,\dots,c^{M-1}$ is called the $c$-loop and is denote by $\lambda^{c}$.
-Remark 13.
-Observe that $q$ has three ingoing and outgoing edges. All other vertices have exactly one ingoing and one outgoing edge.
+The shortest path from the vertex $v_{1}$ to the vertex $v_{2}$ is denoted by $\overrightarrow{v_{1}v_{2}}$.
+Remark 7.2.2.
+Observe that $q$ has three ingoing and outgoing edges.
+All other vertices have exactly one ingoing and one outgoing edge.
+In particular each vertex other than $q$ has a unique predecessor and a unique successor.
 Let $X$ be the $(1,N,M)$ directed graph.
@@
 The loops $\lambda^{q}$, $\lambda^{h}$ and $\lambda^{c}$ are called the prime loops of $X$, respectively $q$-loop, $h$-loop and $c$-loop.
-Observe that a loop $\lambda$ is a concatenation of prime loops $\gamma_{i}$, namely
+Observe that a loop $\lambda$ is a concatenation of prime loops $\gamma_{i}$, namely,
 $\lambda=\gamma_{u}\ast\cdots\ast\gamma_{1}.$
@@
 Graph Morphisms.
-Given $N,M,\tilde{N},\tilde{M}$ positive natural numbers, consider $the (1,N~,M~)(1,\tilde{N},\tilde{M}) directed graph X~\tilde{X}$ and the graph morphism
+Given $N,M,\tilde{N},\tilde{M}$ positive natural numbers, consider $the (1,N~,M~)(1,\tilde{N},\tilde{M}) directed graph X~\tilde{X}$ and assume there is a graph morphism
 $w:\tilde{X}\mapsto X.$
@@
 Projections.
-In the following lemma we introduce the projection of $D$ on $X$.
-Lemma 21.
+In the following lemma we introduce the projection of $D$ onto $X$.
+Lemma 7.2.3.
 Let $f$ be a map.
@@
 7.3 Strips
-We introduce strips in the plane that will enable us to later construct, inductively, the object of primary interest to us: the lamination.
+We introduce strips in the plane that will enable us later to construct, inductively, the object of primary interest to us: the lamination.
 Observe that the strip are an abstract generalization of the $\mathcal{B}_{n}$ strips built earlier to contain the curves of secondary tangencies.
-Definition 10.
+We start by defining $r$-tangency strips, with $r$ a positive number; see Figure 11.
+Definition 7.3.1.
 Let $r>0$, $r\in\mathbb{N}$ and let $F$ be an unfolding. An $r$-tangency strip $S$ is an open region in $t_{0},t_{0}]\times[-a_{0},a_{0}]$ satisfying the following:
   • $S$ is bounded by the graph of two functions $s_{l},s_{u}:[-t_{0},t_{0}]\to\mathbb{R}$ such that, for all $t\in[-t_{0},t_{0}]$,
-    $\left|s_{l}(t)-s_{u}(t)\right|\leq C\rho^{r(1+\alpha)-1},$
-    where $0<\rho<1$, $\rho<\max_{t}\left\{\frac{1}{\mu(t)}\right\}$, $0<\alpha<1$ and $C$ is a positive constant.
-  • $S$ contains $r$ global curves $\left\{b_{k}\right\}_{k=1}^{r}$, whcih are graph of functions $b_{k}:[-t_{0},t_{0}]\mapsto\mathbb{R}$ for all $k$, such that the following holds:
+    $\left|s_{l}(t)-s_{u}(t)\right|\leq C\rho^{r},$
+    where $0<\rho<1$, $\rho<\max_{t}\left\{\frac{1}{\mu(t)}\right\}$, and $C$ is a positive constant.
+  • $S$ contains $r$ curves $\left\{b_{k}\right\}_{k=1}^{r}$, which are graph of functions $b_{k}:[-t_{0},t_{0}]\mapsto\mathbb{R}$ for all $k$, such that the following holds:
     • (i)
@@
 For all $k\in\left\{1,\dots,r\right\}$, the pair $\left\{b_{k},U_{k}\right\}$ is called a $(1,N_{k},M_{k})$ descendant of $S$.
-We start by defining $r$-tangency strips, with $r$ a positive real number; see Figure 11.
+Remark 7.3.2.
+Let $f\in S$, then $f$ has a $(1,N,M)$ dynamical partition.
+We identify $f$ with its restriction to this dynamical partition.
+If in particular $f\in U_{k}$ for some $k$, then it has a finer dynamical partition $(1,N_{k},M_{k})$ which describes the dynamics on a smaller scale.
+For this reason we call the restriction of $f$ to the finer dynamical partition the renormalization of $f$.
 Figure 11: A $r$-tangency strip
@@
 In this subsection we present the induction step that will allow later to build the lamination and the invariant Cantor set.
-Proposition 7.
+Proposition 7.4.1.
 Let $S$ be a $r$-tangency strip and let $\left\{b,U\right\}$ be a $(1,N,M)$ descendant of $S$.
-Then there exist $n\in\mathbb{N}$, $n>>r$ large enough and a countable sequence of regions $\left\{S_{j}\right\}_{j=1}^{\infty}\subset U\subset S$ satisfying the following:
+Then there exist $C>0$ and $n\in\mathbb{N}$, $n>>r$ large enough and a countable sequence of regions $\left\{S_{j}\right\}_{j=1}^{\infty}\subset U\subset S$ satisfying the following:
   • (i)
-    $\forall j$, $S_{j}$ is a $(n+j)$-tangency strip,
+    $\forall j$, $S_{j}$ is a $(\sqrt{n+j})$-tangency strip,
   • (ii)
     $\forall j,j^{\prime}$ with $j\neq j^{\prime}$, $S_{j}\cap S_{j}^{\prime}=\emptyset$.
-Moreover, for all $j$ and for all $k\in\left\{1,\dots,n+j\right\}$ let $\left\{b_{k}^{(j)},U_{k}^{(j)}\right\}$ be a $\left(1,N_{k}^{(j)},M_{k}^{(j)}\right)$ descendant of $S_{j}$.
+Moreover, for all $j$ and for all $k\in\left\{1,\dots,\sqrt{n+j}\right\}$ let $\left\{b_{k}^{(j)},U_{k}^{(j)}\right\}$ be a $\left(1,N_{k}^{(j)},M_{k}^{(j)}\right)$ descendant of $S_{j}$.
 Then the following holds:
   • (iii)
-    $\exists C>0$ such that $\forall t\in\left[-t_{0},t_{0}\right]$,
-    $\frac{1}{C}\frac{1}{\mu(t)^{n+j}}\leq\left|b_{k}^{(j)}(t)-b(t)\right|\leq\frac{C}{\mu(t)^{n+j}},$
+    $\forall t\in\left[-t_{0},t_{0}\right]$,
+    $\left|b_{k}^{(j)}(t)-b(t)\right|\leq\frac{C}{\mu(t)^{n+j}},$
   • (iv)
-    $\exists\tilde{C}>0$ such that
-    $\frac{1}{\tilde{C}}\frac{n+j}{\mu(t)^{n+1+j}}\frac{\partial\mu}{\partial t}\leq\left|\frac{db_{k}^{(j)}}{dt}-\frac{db}{dt}\right|\leq\tilde{C}\frac{n+j}{\mu(t)^{n+1+j}}\frac{\partial\mu}{\partial t},$ (7.1)
+    $\forall t\in\left[-t_{0},t_{0}\right]$,
+    $\left|\frac{db_{k}^{(j)}}{dt}(t)-\frac{db}{dt}(t)\right|\leq C\frac{n+j}{\left(\lambda(t)^{\theta}\mu(t)^{2}\right)^{n+j}},$ (7.4.2)
   • (v)
-    $\forall f\in U_{k}^{(j)}\subset U$, let $\mathfrak{D}^{(j)}_{k}(f)$ be the $\left(1,N_{k}^{j},M_{k}^{j}\right)$ dynamical partition of $f$ and let $\mathfrak{D}(f)$ be the $\left(1,N,M\right)$ dynamical partition of $f$.
+    $\forall f\in U_{k}^{(j)}\subset U$, let $\mathfrak{D}^{(j)}_{k}(f)$ be the $\left(1,N_{k}^{(j)},M_{k}^{(j)}\right)$ dynamical partition of $f$ and let $\mathfrak{D}(f)$ be the $\left(1,N,M\right)$ dynamical partition of $f$.
 Then $\mathfrak{D}^{(j)}_{k}(f)$ is a refinement of $\mathfrak{D}(f)$ and
-    $\text{mesh}\left(\mathfrak{D}^{(j)}_{k}(f)\right)<\left(\frac{1}{k}\right)\text{mesh}\left(\mathfrak{D}(f)\right).$
-Furthermore, if $X$ is the $\left(1,N,M\right)$ directed graph, $X_{k}^{j}$ is the $\left(1,N_{k}^{j},M_{k}^{j}\right)$ directed graph, $i:\mathcal{D}_{k}^{(j)}\mapsto\mathcal{D}$ is the inclusion, $\pi:\mathcal{D}\mapsto X$ is the projection of $\mathcal{D}$ on $X$ and $\pi_{k}^{j}:\mathcal{D}_{k}^{(j)}\mapsto X_{k}^{(j)}$ is the projection of $\mathcal{D}_{k}^{(j)}$ on $X_{k}^{(j)}$, there exists a unique graph morphism $w:X_{k}^{(j)}\mapsto X$ such that
-$\pi\circ i_{k}^{(j)}=w\circ\pi_{k}^{(j)}.$
+    $\text{mesh}\left(\mathfrak{D}^{(j)}_{k}(f)\right)<\frac{1}{2}\cdot\text{mesh}\left(\mathfrak{D}(f)\right).$
+Furthermore, if $X$ is the $\left(1,N,M\right)$ directed graph, $X_{k}^{(j)}$ is the $\left(1,N_{k}^{(j)},M_{k}^{(j)}\right)$ directed graph, $i:\mathcal{D}_{k}^{(j)}\mapsto\mathcal{D}$ is the inclusion, $\pi:\mathcal{D}\mapsto X$ is the projection of $\mathcal{D}$ on $X$ and $\pi_{k}^{j}:\mathcal{D}_{k}^{(j)}\mapsto X_{k}^{(j)}$ is the projection of $\mathcal{D}_{k}^{(j)}$ on $X_{k}^{(j)}$, there exists a unique graph morphism $w_{k}^{(j)}:X_{k}^{(j)}\mapsto X$ such that
+$\pi\circ i_{k}^{(j)}=w_{k}^{(j)}\circ\pi_{k}^{(j)}.$
 Proof.
-Without loss of generality, we can assume that the $r$-strip $S$ is delimitated by straight lines.
-In particular we can make a reparametrization to ensure that the curve $b$ corresponds to a straight line in the $a$-direction..
-For all $j$, consider the strips $\mathcal{B}_{n+j}$ as defined in (4.1) and take $S_{j}=\mathcal{B}_{n+j}$.
+Without loss of generality, we can assume that the $r$-strip $S$ is bounded by straight lines.
+In particular, we can make a reparametrization to ensure that the curve $b$ corresponds to a straight line in the $t$-direction.
+For all $j$, consider the strips $\mathcal{B}_{n+j}$ as defined in (4.0.2) and take $S_{j}=\mathcal{B}_{n+j}$.
 Observe that for $n$ large enough $S_{j}\subset U\subset S$ and by construction $S_{j}\cap S_{j^{\prime}}=\emptyset$ for all $j\neq j^{\prime}$.
-Apply Section $4$ and Section $5$ to ensure that $\mathcal{B}_{n+j}$ is indeed a $n+j$- strip.
-The descendants are the pairs $\left\{b_{k}^{(j)},U_{k}^{(j)}\right\}$ with $k\in\left\{1,\dots,n\right\}$ where $b_{k}^{(j)}$ is the curve of secondary tangencies as constructed in Proposition 4 and $U_{k}^{(j)}$ is a neighborhood of $b_{k}^{(j)}$ as built in Lemma 17.
-In particular the family $F_{|U_{k}^{(j)}}$ is an unfolding of the homoclinic tangency $b_{k}^{(j)}$.
-Given a descendent of $S_{j}$, $\left\{b_{k}^{(j)},U_{k}^{(j)}\right\}$.
-Observe that $a_{n+j},b_{k}^{(j)}\in S_{j}=\mathcal{B}_{n+j}$ and the diameter of $\mathcal{B}_{n+j}$ is smaller than the distance between $a_{n}$ and the previous tangency $b$ that is indeed of then order of $1/\mu^{n+j}$, see 2.14.
+Moreover Section $4$ and Section $5$ ensure also that $\mathcal{B}_{n+j}$ contains $\sqrt{n+j}$ secondary tangency curves.
+The descendants are the pairs $\left\{b_{k}^{(j)},U_{k}^{(j)}\right\}$ with $k\in\left\{1,\dots,\sqrt{n+j}\right\}$ where $b_{k}^{(j)}$ is the curve of secondary tangencies as constructed in Proposition 5.3.1 and $U_{k}^{(j)}$ is a neighborhood of $b_{k}^{(j)}$ as built in Lemma 5.3.3.
+In particular, the family $F_{|U_{k}^{(j)}}$ is an unfolding of the homoclinic tangency $b_{k}^{(j)}$.
+Given a descendent of $S_{j}$, $\left\{b_{k}^{(j)},U_{k}^{(j)}\right\}$, observe that $a_{n+j},b_{k}^{(j)}\subset S_{j}=\mathcal{B}_{n+j}$ and the width of $\mathcal{B}_{n+j}$ is smaller than the distance between $a_{n}$ and the previous tangency $b$ that is indeed of then order of $1/\mu^{n+j}$, see (2.2.10).
 This proves $(iii)$.
-Condition $(iv)$ on the derivative comes from 5.4 and 5.6.
-Condition $(v)$ and the following statement are consequence of the construction.
+Condition $(iv)$ on the derivative comes from (5.1.7).
+The final step is to construct the corresponding dynamical partitions.
+Let $f\in b_{k}^{(j)}$.
+In particular there is a point $m_{1}\in W^{u}_{\text{loc}}(p)$ and a time $s$ such that at $f^{s+M}(m_{1})$ there is a tangency of $W^{s}(p)\cap W^{u}_{\text{loc}}(p)$, see Section $4$.
+Let $Q_{k}^{(j)}\subset Q$ such that
+$\overline{Q_{k}^{(j)}}\cap\left\{m_{1},f(m_{1}),\dots,f^{s+M}(m_{1})\right\}=\emptyset.$
+Let $\overline{\ell}_{k}^{(j)}$ be minimal such that
+$f^{-\overline{\ell}_{k}^{(j)}+1}(q_{2}),f^{-\overline{\ell}_{k}^{(j)}+1}(m_{1})\in\overline{Q_{k}^{(j)}},$
+and let $\underline{\ell}_{k}^{(j)}$ be minimal such that
+$f^{\underline{\ell}_{k}^{(j)}}(f^{M}(q_{2})),f^{\underline{\ell}_{k}^{(j)}}(f^{s+M}(m_{1})).$
+Let $(C^{0})^{(j)}_{k}$ be a small enough neighborhood of $f^{-\overline{\ell}_{k}^{(j)}+1}(m_{1})$ and let $(H^{0})^{(j)}_{k}$ be a small enough neighborhood of $f^{-\overline{\ell}_{k}^{(j)}+1}(q_{2})$ such that
+$\mathfrak{D}_{k}^{(j)}(f)$ $=$ $\left\{\overline{Q}_{k}^{(j)}\right\}\cup$
+$\left\{\overline{f^{i}\left((C^{0})^{(j)}_{k}\right)}\left|\right.i\in\left\{0,\dots,\overline{\ell}_{k}^{(j)}+s+M+\underline{\ell}_{k}^{(j)}-1\right\}\right\}\cup$
+$\left\{\overline{f^{i}\left((H^{0})^{(j)}_{k}\right)}\left|\right.i\in\left\{0,\dots,\overline{\ell}_{k}^{(j)}+M+\underline{\ell}_{k}^{(j)}-1\right\}\right\}$
+forms a pairwise disjoint collection,
+$\text{mesh}\left(\mathfrak{D}^{(j)}_{k}(f)\right)<\frac{1}{3}\cdot\text{mesh}\left(\mathfrak{D}(f)\right).$
+and
+$\text{dist}\left({f^{\overline{\ell}_{k}^{(j)}+s}\left((C^{0})^{(j)}_{k}\right)},q_{2}\right)<\lambda^{n_{0,k}^{(j)}/2},$ (7.4.3)
+where $n_{0,k}^{(j)}$ is defined in Subsection 4.1.
+Observe that the construction of $\mathfrak{D}^{(j)}_{k}(f)$ can be made continuous along the curve $b_{k}^{(j)}$.
+Moreover by restricting the neighborhood $U_{k}^{(j)}$ we can also make the construction of $\mathfrak{D}^{(j)}_{k}(f)$ varying continuously over the neighborhood still satisfying condition (7.4.3).
+In particular, $\mathfrak{D}^{(j)}_{k}(f)$ is a dynamical partition with $N^{(j)}_{k}=\overline{\ell}_{k}^{(j)}+s+\underline{\ell}_{k}^{(j)}-1$ and $M^{(j)}_{k}=\overline{\ell}_{k}^{(j)}+M+\underline{\ell}_{k}^{(j)}-1$.
+This proves condition $(v)$ and the last statement is consequence of this construction.
 ∎
-Remark 14.
-Observe that, using the notation as in the previous proposition and by using the construction in Subsection 4.1, there exist $n_{k}^{(j)},\underline{\ell}_{k}^{(j)},\overline{\ell}_{k}^{(j)},\theta_{k}^{(j)},n_{0,k}^{(j)}$ such that
-$\underline{w}\left(\omega_{k}^{(j)}\right)=\left(q,{q}^{\underline{\ell}_{k}^{(j)}}{c}{q}^{\theta_{k}^{(j)}n_{k}^{(j)}}{c}{q}^{n_{k}^{(j)}}{c}{q}^{n_{0,k}^{(j)}}{h}{q}^{\overline{\ell}_{k}^{(j)}},{q}^{\underline{\ell}_{k}^{(j)}}{h}{q}^{\overline{\ell}_{k}^{(j)}}\right).$
+Remark 7.4.4.
+Observe that, using the notation as in the previous proposition and by using the construction in Subsection 4.1, there exist $n_{k}^{(j)},\underline{\ell}_{k}^{(j)},\overline{\ell}_{k}^{(j)},n_{0,k}^{(j)}$ such that
+$\underline{\omega}\left(w_{k}^{(j)}\right)=\left(q,{q}^{\underline{\ell}_{k}^{(j)}}{h}{q}^{n_{0,k}^{(j)}}{c}{q}^{n_{k}^{(j)}}{c}{q}^{\theta n_{k}^{(j)}}{c}{q}^{\overline{\ell}_{k}^{(j)}},{q}^{\underline{\ell}_{k}^{(j)}}{h}{q}^{\overline{\ell}_{k}^{(j)}}\right).$
 8 Finite Collet-Eckmann Condition
-In this section we set up the induction step needed to prove later the existence of maps with a Collet-Eckmann point.
-More in details we prove the existence of maps having a point satisfying the finite Collet-Eckmann condition.
+In this section we describe the induction step needed to prove later the existence of maps with a Collet-Eckmann point.
+More in detail, we prove the existence of maps having a point satisfying the finite Collet-Eckmann condition.
 Roughly speaking a point $x$ is a $T$-Collet-Eckmann (T-CE) point if there are vectors $v$ in the tangent space at $x$ such that the derivative of the map along the orbit of $x$ in the direction of $v$ grows at the least exponentially fast up to time $T$.
 Below is the formal definition.
-Definition 11.
-Let $f:\mathcal{M}\to\mathcal{M}$ be a map and let $x\in\mathcal{M}$.
-Given constants $\varphi>0$, $0<C<1$, $\rho>1$, we say that $x$ is a $T$-Collet-Eckmann (T-CE) point if the following holds.
-Denote by $K_{\varphi}(x)$ the cone at $x$ of angle $\varphi$}, and by $\text{T}_{x}$ the tangent space at $x$, then for all vectors $v\in K_{\varphi}\subset\text{T}_{x}$,
+As before, without lose of generality, we consider a diffeomorphism $f:\mathcal{M}\to\mathcal{M}$ where the manifold $\mathcal{M}$ is an open set in $\mathbb{R}^{2}$.
+Let $\varphi>0$.
+Denote by $\text{T}_{x}\mathcal{M}$ the tangent space at $x$, and by
+$K_{\varphi}(x)=\left\{(v_{1},v_{2})\in\text{T}_{x}\mathcal{M}\left|\right.|v_{2}|\tan\varphi>|v_{1}|\right\},$
+the cone at $x$ of angle $\varphi$}.
+Definition 8.0.1.
+Given constants $0<C<1$, $\rho>1$, we say that $x\in\mathcal{M}$ is a $T$-Collet-Eckmann (T-CE) point if the following holds. For all vectors $v\in K_{\varphi}(x)$,
 $\left|Df^{s}_{x}v\right|\geq C\rho^{s}\left|v\right|\text{, \hskip 13.942pt }\forall\text{ }0\leq s\leq T.$
 From now on we are going to assume an extra condition on $\theta$, namely we ask that,
-$\lambda^{\theta}\mu^{1+\alpha}>\lambda^{\alpha},$ (8.1)
-where $\alpha$ is defined in (4.17). By (2.17) and (4.17), there exists $\theta\in(\theta_{0},\theta_{1})$ satisfying (8.1), see Appendix.
+$\lambda^{\theta}\mu^{1+\alpha}>\lambda^{\alpha},$ (8.0.2)
+where $\alpha$ is defined in (4.1.49).
+By (2.3.2) and (4.1.49), there exists $\theta\in(\theta_{0},\theta_{1})$ satisfying (8.0.2), see Appendix.
 Take constants $\varphi>0$, $0<C<1$ and $\rho>1$ such that,
@@
 We introduce now some notation.
-Let $x$ be a point.
-If $\Gamma$ is a curve, we denote by $\text{d}_{\text{v}}\left(x,\Gamma\right)$ the vertical distance between a point $x$ and the curve $\Gamma$.
+Let $(x,y)$ be a point and let $\Gamma$ be a curve which is a graph over the $x$-axes.
+If there is a point $(x,y_{0})\in\Gamma$ then denote by $\text{d}_{\text{v}}\left(x,\Gamma\right)=|y-y_{0}|$ the vertical distance between a point $x$ and the curve $\Gamma$.
 The angle between two vectors is denoted by $\angle$.
@@
 For all $f\in U$, let $\mathfrak{D}(f)$ be the $\left(1,N,M\right)$ dynamical partition of $f$.
-Given $T\in\left\{1,\dots,N-1\right\}$, one can define a vector field $\mathfrak{X}\left(f^{T}\right)$ on $C^{N-T}$ as
+Given $T\in\left\{1,\dots,N-1\right\}$, one can define a vector field $\mathfrak{X}^{N-T}$ on $C^{N-T}$ as
 $Df^{T}(x)v(x)=e_{1}\text{ for all }x\in C^{N-T}.$
@@
 We are now ready to present the induction step needed to prove the existence of laminations of maps with a Collet–Eckmann point. We use the notation introduced above.
-Proposition 8.
-Let $S$ be a tangency strip and let $\left\{b,U\right\}$ be a $(1,N,M)$ descendant of $S$.
-Suppose that for all $f\in U$, there exist a positive constant $K$ and $T\in\left\{1,\dots,N-1\right\}$ such that the following holds,
+Proposition 8.0.4.
+Let $S$ be a $r$-tangency strip and let $\left\{b,U\right\}$ be a $(1,N,M)$ descendant of $S$.
+Suppose that for all $f\in U$, there exist a positive constant $K$ and $T\in\left\{1,\dots,N-1\right\}$ such that the $T$-Collet-Eckmann Condition holds,
   • $\text{IH}_{Q}:$
@@
       There exists a curve $\Gamma$ in $C^{0}$ which is a graph over the $x$-axis and
-      $Df_{x}^{N-T}e_{2}\in\text{span}\left(\mathfrak{X}\left(f^{N-T}\right)\right)\iff x\in\Gamma.$
+      $Df_{x}^{N-T}e_{2}\in\text{span}\left(\mathfrak{X}^{N-T}\right)\iff x\in\Gamma.$
     • (ii)
       For all $x\in C^{0}$,
-      $\frac{1}{K}{{d}_{\text{v}}}\left(x,\Gamma\right)\leq\angle\left(Df_{x}^{N-T}e_{2},\mathfrak{X}\left(f^{N-T}\right)\right)\leq Kd_{\text{v}}\left(x,\Gamma\right).$
+      $\frac{1}{K}{{d}_{\text{v}}}\left(x,\Gamma\right)\leq\angle\left(Df_{x}^{N-T}e_{2},\mathfrak{X}^{N-T}\right)\leq Kd_{\text{v}}\left(x,\Gamma\right).$
     • (iii)
@@
 […]
+$\tilde{C}^{\tilde{N}-\tilde{T}}\subset C^{N-T}$
+such that the $\tilde{T}-$Collet-Eckmann Condition holds. Namely,
   • $\text{IH}_{\tilde{Q}}$:
@@
       There exists a curve $\tilde{\Gamma}$ in $\tilde{C}^{0}$ which is a graph over the $x$-axis and
-      $Df^{\tilde{N}-\tilde{T}}(x)e_{2}\in\text{span}\left({\mathfrak{X}}\left(f^{\tilde{N}-\tilde{T}}\right)\right)\iff x\in\tilde{\Gamma}$.
+      $Df^{\tilde{N}-\tilde{T}}(x)e_{2}\in\text{span}\left({\mathfrak{X}}^{\tilde{N}-\tilde{T}}\right)\iff x\in\tilde{\Gamma}$.
     • (ii)
@@
 Proof.
-[…]
-$\tilde{T}=T+n+N-T+T+n_{0}+M+\overline{\ell},$
-and for all $\tilde{f}\in\tilde{U}$ let $\tilde{K}$ be a positive constant assuring that $\text{IH}_{\tilde{H}^{0}}$, $\text{IH}_{\tilde{C}^{0}}$-(iii) and $\text{IH}_{\tilde{C}^{\tilde{N}-\tilde{T}}}-$(iii) hold.
-The existence of such a constant is guaranteed by the initial conditions on the family.
+Let $\tilde{S}=S_{j}\subset U\subset S$ be a tangency strip and let $\left\{\tilde{b},\tilde{U}\right\}=\left\{b_{k}^{(j)},U_{k}^{(j)}\right\}$ be a $\left(1,\tilde{N},\tilde{M}\right)=\left(1,N_{k}^{(j)},M_{k}^{(j)}\right)$ descendant of $\tilde{S}$ as in Proposition 7.4.1.
+In each step of the construction we might have to increase the $j$.
+Consider the orbit from $\tilde{C}^{0}$ to $\tilde{C}^{\tilde{N}}$.
+By construction there exist exactly three moments, $s_{1}$, $s_{2}$, $s_{3}$ smaller than $\tilde{N}$ such that $\tilde{C}^{s_{i}}\subset C^{N-T}$.
+Using the notation from Remark 7.4.4 we get
+$s_{1}$ $=$ $\overline{\ell}_{k}^{(j)}+N-T,$
+$s_{2}$ $=$ $s_{1}+T+\theta n_{k}^{(j)}+N-T,$
+$s_{3}$ $=$ $s_{2}+T+n_{k}^{(j)}+N-T.$
+Define $\tilde{T}$ such that $\tilde{N}-\tilde{T}=s_{2}$ corresponds to the second passage [^1] The second passage is an essential choice.
+Using the third passage will not lead to a Cantor set, see Lemma 9.3.5 and its proof.
+Using the first passage will not give enough expansion to ensure the CE condition.
+trough $C^{N-T}$, see Figure 12.
+Figure 12: Passages in $C^{N-T}$
+In particular,
+$\tilde{T}=T+n_{k}^{(j)}+N-T+T+n_{0,k}^{(j)}+M+\underline{\ell}_{k}^{(j)}.$
+For all $\tilde{f}\in\tilde{U}$, let $\tilde{K}$ be a positive constant describing the bounds as required in $\text{IH}_{\tilde{H}^{0}}$, $\text{IH}_{\tilde{C}^{0}}$-(iii) and $\text{IH}_{\tilde{C}^{\tilde{N}-\tilde{T}}}-$(iii).
+The existence of such a constant is guaranteed by the fact that this conditions consider only a finite number of iterates of the original diffeomorphism.
 Observe that $\text{IH}_{\tilde{Q}}$ is trivially satisfies.
-Define $\tilde{\Gamma}=f^{-\overline{\ell}}\left(\Gamma\right)$, then $\text{IH}_{\tilde{C}^{0}}$-(i) and $\text{IH}_{\tilde{C}^{0}}$-(ii) are satisfied by construction.
+Let $\tilde{\Gamma}$, be the curve associated to the critical point $\tilde{c}$, then $\text{IH}_{\tilde{C}^{0}}$-(i) and $\text{IH}_{\tilde{C}^{0}}$-(ii) are satisfied by construction.
 The rest of the proof is devoted to prove $\text{IH}_{\tilde{C}^{\tilde{N}-\tilde{T}}}-$(i) and $\text{IH}_{\tilde{C}^{\tilde{N}-\tilde{T}}}-$(ii).
 This will be achieved by steps.
-Each step corresponds to a time defining $\tilde{T}$.
+Each step corresponds to a term in the definition of $\tilde{T}$.
 In order to lighten the notation, we denote $\tilde{C}=\tilde{C}^{\tilde{N}-\tilde{T}}$.
@@
   • Step 1:
-    We begin by proving $\text{IH}_{\tilde{C}^{\tilde{N}-\tilde{T}}}-$(i) from time 0 up to time $T$. By $\text{IH}_{C^{N-T}}$-(i),
-    • $\text{IH}_{\tilde{C}}-$(i)P1:
-      $\left|Df_{x}^{s}v\right|\geq C\rho^{s}\left|v\right|\text{ }\forall\text{ }0\leq s\leq T.$
+    We begin by proving $\text{IH}_{\tilde{C}^{\tilde{N}-\tilde{T}}}-$(i) from time 0 up to time $T$. By $\text{IH}_{C^{N-T}}$-(i), we get,
+    $\text{IH}_{\tilde{C}}\text{-(i)P1: }\left|Df_{x}^{s}v\right|\geq C\rho^{s}\left|v\right|\text{ }\forall\text{ }0\leq s\leq T.$
     Moreover, by $\text{IH}_{C^{N-T}}$-(ii)
@@
     it follows that,
-    $v^{(1)}_{2}>\cos\varphi\left|v^{(1)}\right|\text{ and }\frac{v^{(1)}_{1}}{v^{(1)}_{2}}<\tan\varphi.$ (8.3)
+    $v^{(1)}_{2}>\cos\varphi\left|v^{(1)}\right|\text{ and }\frac{v^{(1)}_{1}}{v^{(1)}_{2}}<\tan\varphi.$ (8.0.5)
   • Step 2:
-    We begin by proving $\text{IH}_{\tilde{C}^{\tilde{N}-\tilde{T}}}-$(i) from time $T+1$ up to time $T+n$.
-Indeed, by $\text{IH}_{Q}$ and $\text{IH}_{\tilde{C}}-$(i)P1, for all $1\leq k\leq n$, we get,
-    • $\text{IH}_{\tilde{C}}-$(i)P2:
-      $\left|Df^{T+k}_{x}v\right|$ $=$ $\left|\left(\begin{matrix}\lambda^{k}&0\\ 0&\mu^{k}\end{matrix}\right)\left(\begin{matrix}v^{(1)}_{1}\\ v^{(1)}_{2}\end{matrix}\right)\right|>\mu^{k}v^{(1)}_{2}\geq\mu^{k}\cos\varphi\left|v^{(1)}\right|$
-      $=$ $\mu^{k}\cos\varphi\left|Df^{T}_{x}v\right|\geq\mu^{k}\cos\varphi C\rho^{T}\left|v\right|\geq C\rho^{T+k}\left|v\right|.$
+    Now we prove $\text{IH}_{\tilde{C}^{\tilde{N}-\tilde{T}}}-$(i) from time $T+1$ up to time $T+n_{k}^{(j)}$.
+Indeed, by $\text{IH}_{Q}$, $\text{IH}_{\tilde{C}}-$(i)P1 and (8.0.5), for all $1\leq s\leq n$, we get,
+    $\text{IH}_{\tilde{C}}\text{-(i)P2: }\left|Df^{T+s}_{x}v\right|$ $=$ $\left|\left(\begin{matrix}\lambda^{s}&0\\ 0&\mu^{s}\end{matrix}\right)\left(\begin{matrix}v^{(1)}_{1}\\ v^{(1)}_{2}\end{matrix}\right)\right|>\mu^{s}v^{(1)}_{2}\geq\mu^{s}\cos\varphi\left|v^{(1)}\right|$
+    $=$ $\mu^{s}\cos\varphi\left|Df^{T}_{x}v\right|\geq\mu^{s}\cos\varphi C\rho^{T}\left|v\right|\geq C\rho^{T+s}\left|v\right|.$
     Observe that in the last inequality of $\text{IH}_{\tilde{C}}-$(i)P2 we used that
-    $\mu^{k}\cos\varphi\geq\rho^{k},\text{ }\forall 1\leq k\leq n.$
+    $\mu^{s}\cos\varphi\geq\rho^{s},\text{ }\forall 1\leq s\leq n.$
     This follow from the condition imposed in (8)-(o). In particular, $\text{IH}_{\tilde{C}}-$(i)P2 proves
-    $\left|Df^{T+n}_{x}v\right|\geq\mu^{n}\cos\varphi C\rho^{T}\left|v\right|.$ (8.4)
+    $\left|Df^{T+n_{k}^{(j)}}_{x}v\right|\geq\mu^{n_{k}^{(j)}}\cos\varphi C\rho^{T}\left|v\right|.$ (8.0.6)
     Denote by
-    $v^{(2)}=Df^{T+n}_{x}v=\left(\begin{matrix}v^{(2)}_{1}\\ v^{(2)}_{2}\end{matrix}\right)$.
-    By $\text{IH}_{Q}$ and by (8.3),
-    $\frac{v^{(2)}_{1}}{v^{(2)}_{2}}=\left(\frac{\lambda}{\mu}\right)^{n}\frac{v^{(1)}_{1}}{v^{(1)}_{2}}<\left(\frac{\lambda}{\mu}\right)^{n}\tan\varphi=\left(\frac{\lambda}{\mu}\right)^{n}\varphi.$ (8.5)
+    $v^{(2)}=Df^{T+n_{k}^{(j)}}_{x}v=\left(\begin{matrix}v^{(2)}_{1}\\ v^{(2)}_{2}\end{matrix}\right)$.
+    By $\text{IH}_{Q}$ and by (8.0.5),
+    $\frac{v^{(2)}_{1}}{v^{(2)}_{2}}=\left(\frac{\lambda}{\mu}\right)^{n}\frac{v^{(1)}_{1}}{v^{(1)}_{2}}<\left(\frac{\lambda}{\mu}\right)^{n}\tan\varphi.$ (8.0.7)
   • Step 3:
-    We start proving $\text{IH}_{\tilde{C}^{\tilde{N}-\tilde{T}}}-$(i) from time $T+n+1$ up to time T+n+NTT+n+N-T. Indeed, by $\text{IH}_{C^{0}}$-(iii) and (8.4), for all $1\leq k\leq N-T$, we get,
-    • $\text{IH}_{\tilde{C}}-$(i)P3:
-      $\left|Df^{T+n+k}_{x}v\right|\geq\frac{1}{K}\left|Df^{T+n}_{x}v\right|>\frac{1}{K}\mu^{n}\cos\varphi C\rho^{T}\left|v\right|\geq C\rho^{T+n+k}\left|v\right|.$
+    Now we prove $\text{IH}_{\tilde{C}^{\tilde{N}-\tilde{T}}}-$(i) from time $T+n_{k}^{(j)}+1$ up to time $T+n_{k}^{(j)}+N-T$.
+Indeed, by $\text{IH}_{C^{0}}$-(iii) and (8.0.6), for all $1\leq s\leq N-T$, we get,
+    $\text{IH}_{\tilde{C}}\text{-(i)P3: }\left|Df^{T+n_{k}^{(j)}+s}_{x}v\right|\geq\frac{1}{K}\left|Df^{T+n_{k}^{(j)}}_{x}v\right|>\frac{1}{K}\mu^{n_{k}^{(j)}}\cos\varphi\cdot C\rho^{T}\left|v\right|\geq C\rho^{T+n_{k}^{(j)}+s}\left|v\right|.$
     Observe that in the last inequality of $\text{IH}_{\tilde{C}}-$(i)P3 we used that
-    $\frac{1}{K}\left(\frac{\mu}{\rho}\right)^{n}\cos\varphi\geq\rho^{k},\text{ }\forall 1\leq k\leq N-T$
+    $\frac{1}{K}\left(\frac{\mu}{\rho}\right)^{n_{k}^{(j)}}\cos\varphi\geq\rho^{s},\text{ }\forall 1\leq s\leq N-T$
     which is equivalent to ask
-    $\frac{1}{K}\left(\frac{\mu}{\rho}\right)^{n}\cos\varphi\geq\rho^{N-T}.$
-    This follow from the condition imposed in (8)-(oo) and taking $n$ large enough. In particular, $\text{IH}_{\tilde{C}}-$(i)P3 proves
-    $\left|Df^{T+n+N-T}_{x}v\right|\geq\frac{C}{K}cos\varphi\mu^{n}\ \rho^{T}\left|v\right|.$ (8.6)
-    Moreover by $\text{IH}_{C^{0}}$-(ii) and Lemma 7,
-    $\angle\left(Df_{f^{T+n}(x)}^{N-T}e_{2},\mathfrak{X}\left(f^{T+n+N-T}\right)\right)\geq\frac{1}{K}{\text{d}_{\text{v}}}\left(f^{T+n}(x),\Gamma\right)\geq\frac{1}{K^{2}}\left(\lambda^{\theta}\mu\right)^{n},$
-    and by (8.5),
+    $\frac{1}{K}\left(\frac{\mu}{\rho}\right)^{n_{k}^{(j)}}\cos\varphi\geq\rho^{N-T}.$
+    This follow from the condition imposed in (8)-(oo) and taking $n_{k}^{(j)}$ large enough (which is equivalent to take $j$ large enough). In particular, $\text{IH}_{\tilde{C}}-$(i)P3 proves
+    $\left|Df^{T+n_{k}^{(j)}+N-T}_{x}v\right|\geq\frac{C}{K}\cos\varphi\cdot\mu^{n_{k}^{(j)}}\ \rho^{T}\left|v\right|.$ (8.0.8)
+    Moreover by $\text{IH}_{C^{0}}$-(ii) and Lemma 4.1.21,
+    $\angle\left(Df_{f^{T+n_{k}^{(j)}}(x)}^{N-T}e_{2},\mathfrak{X}^{N-T}\right)\geq\frac{1}{K}{\text{d}_{\text{v}}}\left(f^{T+n_{k}^{(j)}}(x),\Gamma\right)\geq\frac{1}{K^{2}}\left(\lambda^{\theta}\mu\right)^{n_{k}^{(j)}},$
+    and by (8.0.7),
     $\tan\left(v^{(2)},e_{2}\right)\leq\left(\frac{\lambda}{\mu}\right)^{n}\tan\varphi.$
-    The last two inequalities together with $\text{IH}_{C^{0}}$-(iii), gives that, for $n$ large enough,
-    $\angle\left(Df_{x}^{T+n+N-T}v,\mathfrak{X}\left(f^{T+n+N-T}\right)\right)\geq\frac{1}{2K^{2}}\left(\lambda^{\theta}\mu\right)^{n}.$ (8.7)
+    The last two inequalities together with $\text{IH}_{C^{0}}$-(iii), gives that, for $n_{k}^{(j)}$ large enough,
+    $\angle\left(Df_{x}^{T+n_{k}^{(j)}+N-T}v,\mathfrak{X}^{N-T}\right)\geq\frac{1}{2K^{2}}\left(\lambda^{\theta}\mu\right)^{n_{k}^{(j)}}.$ (8.0.9)
   • Step 4:
-    We start proving $\text{IH}_{\tilde{C}^{\tilde{N}-\tilde{T}}}-$(i) from time T+nNT+1T+nN-T+1 up to time T+n+NT+TT+n+N-T+T.
-Indeed, by $\text{IH}_{C^{N-T}}$-(iii) and (8.6), for all $1\leq k\leq T$, we get,
-    • $\text{IH}_{\tilde{C}}-$(i)P4:
-      $\left|Df^{T+n+N-T+k}_{x}v\right|\geq\frac{1}{K}\left|Df^{T+n+N-T}_{x}v\right|\geq\frac{C}{K^{2}}\cos\varphi\mu^{n}\rho^{T}\left|v\right|\geq C\rho^{T+n+N-T+k}\left|v\right|.$
+    Now we prove $\text{IH}_{\tilde{C}^{\tilde{N}-\tilde{T}}}-$(i) from time $T+n_{k}^{(j)}+N-T+1$ up to time $T+n_{k}^{(j)}+N-T+T$.
+Indeed, by $\text{IH}_{C^{N-T}}$-(iii) and (8.0.8), for all $1\leq s\leq T$, we get,
+    $\text{IH}_{\tilde{C}}\text{-(i)P4: }\left|Df^{T+n_{k}^{(j)}+N-T+s}_{x}v\right|$ $\geq$ $\frac{1}{K}\left|Df^{T+n_{k}^{(j)}+N-T}_{x}v\right|$
+    $\geq$ $\frac{C}{K^{2}}\cos\varphi\cdot\mu^{n_{k}^{(j)}}\rho^{T}\left|v\right|\geq C\rho^{T+n_{k}^{(j)}+N-T+s}\left|v\right|.$
     Observe that in the last inequality of $\text{IH}_{\tilde{C}}-$(i)P4 we used that
-    $\frac{1}{K^{2}}\cos\varphi\left(\frac{\mu}{\rho}\right)^{n}\rho^{T}\geq\rho^{N+k},\text{ }\forall 1\leq k\leq T$
+    $\frac{1}{K^{2}}\cos\varphi\left(\frac{\mu}{\rho}\right)^{n_{k}^{(j)}}\rho^{T}\geq\rho^{N+s},\text{ }\forall 1\leq s\leq T$
     which is equivalent to ask
-    $\frac{1}{K^{2}}\cos\varphi\left(\frac{\mu}{\rho}\right)^{n}\geq\rho^{N}.$
-    This follow from the condition imposed in (8)-(oo) and taking $n$ large enough. In particular, $\text{IH}_{\tilde{C}}-$(i)P4 proves
-    $\left|Df^{T+n+N}_{x}v\right|\geq\frac{C}{K^{2}}cos\varphi\mu^{n}\ \rho^{T}\left|v\right|.$ (8.8)
-    Equation (8.7), $\text{IH}_{C}$-(iii) and the definition of the vector field $\mathfrak{X}$ give that, for $n$ large enough,
-    $\angle\left(Df_{x}^{T+n+N}v,e_{1}\right)\geq\frac{1}{2K^{3}}\left(\lambda^{\theta}\mu\right)^{n}.$ (8.9)
+    $\frac{1}{K^{2}}\cos\varphi\left(\frac{\mu}{\rho}\right)^{n_{k}^{(j)}}\geq\rho^{N}.$
+    This holds for $n_{k}^{(j)}$ large enough from the condition imposed in (8)-(oo). In particular, $\text{IH}_{\tilde{C}}-$(i)P4 proves
+    $\left|Df^{T+n_{k}^{(j)}+N}_{x}v\right|\geq\frac{C}{K^{2}}\cos\varphi\cdot\mu^{n_{k}^{(j)}}\ \rho^{T}\left|v\right|.$ (8.0.10)
+    Equation (8.0.9), $\text{IH}_{C}$-(iii) and the definition of the vector field $\mathfrak{X}$ give that, for $n_{k}^{(j)}$ large enough,
+    $\angle\left(Df_{x}^{T+n_{k}^{(j)}+N}v,e_{1}\right)\geq\frac{1}{2K^{3}}\left(\lambda^{\theta}\mu\right)^{n_{k}^{(j)}}.$ (8.0.11)
     In particular, if we denote by
-    $v^{(4)}=Df^{T+n+N}_{x}v=\left(\begin{matrix}v^{(4)}_{1}\\ v^{(4)}_{2}\end{matrix}\right)$,
-    then, by (8.9),
-    $\frac{v^{(4)}_{2}}{\left|v^{(4)}\right|}\geq\sin\frac{1}{2K^{3}}\left(\lambda^{\theta}\mu\right)^{n}\geq\frac{1}{4K^{3}}\left(\lambda^{\theta}\mu\right)^{n}$ (8.10)
+    $v^{(4)}=Df^{T+n_{k}^{(j)}+N}_{x}v=\left(\begin{matrix}v^{(4)}_{1}\\ v^{(4)}_{2}\end{matrix}\right)$,
+    then, by (8.0.11),
+    $\frac{v^{(4)}_{2}}{\left|v^{(4)}\right|}\geq\sin\frac{1}{2K^{3}}\left(\lambda^{\theta}\mu\right)^{n_{k}^{(j)}}\geq\frac{1}{4K^{3}}\left(\lambda^{\theta}\mu\right)^{n_{k}^{(j)}}$ (8.0.12)
   • Step 5:
-    We start proving $\text{IH}_{\tilde{C}^{\tilde{N}-\tilde{T}}}-$(i) from time $T+nN+1$ up to time $T+n+N+n_{0}$. Indeed, by $\text{IH}_{Q}$, (8.10) and (8.8), for all $1\leq k\leq n_{0}$, we get,
-    • $\text{IH}_{\tilde{C}}-$(i)P5:
-      $\left|Df^{T+n+N+k}_{x}v\right|$ $\geq$ $\mu^{k}v^{(4)}_{2}\geq\mu^{k}\frac{1}{4K^{3}}\left(\lambda^{\theta}\mu\right)^{n}\left|Df^{T+n+N}_{x}v\right|$
-      $\geq$ $\frac{C}{4K^{5}}\cos\varphi\mu^{k+n}\rho^{T}\left(\lambda^{\theta}\mu\right)^{n}\left|v\right|\geq C\rho^{T+n+N+k}\left|v\right|.$
+    Now we prove $\text{IH}_{\tilde{C}^{\tilde{N}-\tilde{T}}}-$(i) from time $T+n_{k}^{(j)}+N+1$ up to time $T+n_{k}^{(j)}+N+n_{0,k}^{(j)}$.
+Indeed, by $\text{IH}_{Q}$, (8.0.12) and (8.0.10), for all $1\leq s\leq n_{0,k}^{(j)}$, we get,
+    $\text{IH}_{\tilde{C}}\text{-(i)P5: }\left|Df^{T+n_{k}^{(j)}+N+s}_{x}v\right|$ $\geq$ $\mu^{s}v^{(4)}_{2}\geq\mu^{s}\frac{1}{4K^{3}}\left(\lambda^{\theta}\mu\right)^{n_{k}^{(j)}}\left|Df^{T+n_{k}^{(j)}+N}_{x}v\right|$
+    $\geq$ $\frac{C}{4K^{5}}\cos\varphi\cdot\mu^{s+n_{k}^{(j)}}\rho^{T}\left(\lambda^{\theta}\mu\right)^{n_{k}^{(j)}}\left|v\right|$
+    $\geq$ $C\rho^{T+n_{k}^{(j)}+N+s}\left|v\right|.$
     Observe that in the last inequality of $\text{IH}_{\tilde{C}}-$(i)P5 we used that
-    $\frac{1}{4K^{5}}\cos\varphi\left(\frac{\mu}{\rho}\right)^{n+k}\left(\lambda^{\theta}\mu\right)^{n}\geq\rho^{N},\text{ }\forall 1\leq k\leq n_{0}$
+    $\frac{1}{4K^{5}}\cos\varphi\left(\frac{\mu}{\rho}\right)^{n_{k}^{(j)}+s}\left(\lambda^{\theta}\mu\right)^{n_{k}^{(j)}}\geq\rho^{N},\text{ }\forall 1\leq s\leq n_{0,k}^{(j)}$
     which is equivalent to ask
-    $\frac{1}{4K^{5}}\cos\varphi\left(\frac{\mu}{\rho}\right)\left(\lambda^{\theta}\mu^{2}\right)^{n}\geq\rho^{n+N}.$
-    This follow from the condition imposed in (8)-(oo) and taking $n$ large enough. In particular, $\text{IH}_{\tilde{C}}-$(i)P5 proves
-    $\left|Df^{T+n+N+n_{0}}_{x}v\right|\geq\frac{C}{4K^{5}}\cos\varphi\mu^{n_{0}}\rho^{T}\left(\lambda^{\theta}\mu^{2}\right)^{n}\left|v\right|.$ (8.11)
-    Moreover, by (8.9),
-    $\tan\angle\left(Df_{x}^{T+n+N}v,e_{2}\right)\leq\frac{2K^{3}}{\left(\lambda^{\theta}\mu\right)^{n}}$
-    which together with (8.1) and by taking $n$ large enough imply that,
-    $\tan\angle\left(Df_{x}^{T+n+N+n_{0}}v,e_{2}\right)\leq\left(\frac{\lambda}{\mu}\right)^{n_{0}}\frac{2K^{3}}{\left(\lambda^{\theta}\mu\right)^{n}}<<1.$ (8.12)
+    $\frac{1}{4K^{5}}\cos\varphi\left(\frac{\mu}{\rho}\right)\left(\lambda^{\theta}\mu^{2}\right)^{n_{k}^{(j)}}\geq\rho^{n_{k}^{(j)}+N}.$
+    This follow from the condition imposed in (8)-(oo) and taking $n_{k}^{(j)}$ large enough. In particular, $\text{IH}_{\tilde{C}}-$(i)P5 proves
+    $\left|Df^{T+n_{k}^{(j)}+N+n_{0,k}^{(j)}}_{x}v\right|\geq\frac{C}{4K^{5}}\cos\varphi\cdot\mu^{n_{0,k}^{(j)}}\rho^{T}\left(\lambda^{\theta}\mu^{2}\right)^{n_{k}^{(j)}}\left|v\right|.$ (8.0.13)
+    Moreover, by (8.0.11),
+    $\tan\angle\left(Df_{x}^{T+n_{k}^{(j)}+N}v,e_{2}\right)\leq\frac{2K^{3}}{\left(\lambda^{\theta}\mu\right)^{n_{k}^{(j)}}},$
+    which together with (8.0.2) and by taking $n_{k}^{(j)}$ large enough imply that,
+    $\tan\angle\left(Df_{x}^{T+n_{k}^{(j)}+N+n_{0,k}^{(j)}}v,e_{2}\right)\leq\left(\frac{\lambda}{\mu}\right)^{n_{0,k}^{(j)}}\frac{2K^{3}}{\left(\lambda^{\theta}\mu\right)^{n_{k}^{(j)}}}<<1.$ (8.0.14)
   • Step 6:
-    We start proving $\text{IH}_{\tilde{C}^{\tilde{N}-\tilde{T}}}-$(i) from time $T+nN+n_{0}+1$ up to time $T+n+N+n_{0}+M$. Indeed, by $\text{IH}_{H}$ and (8.11), for all $1\leq k\leq M$, we get,
-    • $\text{IH}_{\tilde{C}}-$(i)P6:
-      $\left|Df^{T+n+N+n_{0}+k}_{x}v\right|$ $\geq$ $\frac{1}{K}\left|Df^{T+n+N+n_{0}}_{x}v\right|$
-      $\geq$ $\frac{C}{4K^{6}}\cos\varphi\mu^{n_{0}}\rho^{T}\left(\lambda^{\theta}\mu^{2}\right)^{n}\left|v\right|$
-      $\geq$ $C\rho^{T+n+N+n_{0}+k}\left|v\right|.$
+    We start proving $\text{IH}_{\tilde{C}^{\tilde{N}-\tilde{T}}}-$(i) from time $T+n_{k}^{(j)}+N+n_{0,k}^{(j)}+1$ up to time $T+n_{k}^{(j)}+N+n_{0,k}^{(j)}+M$.
+Indeed, by $\text{IH}_{H}$ and (8.0.13), for all $1\leq s\leq M$, we get,
+    $\text{IH}_{\tilde{C}}\text{-(i)P6: }\left|Df^{T+n_{k}^{(j)}+N+n_{0,k}^{(j)}+s}_{x}v\right|$ $\geq$ $\frac{1}{K}\left|Df^{T+n_{k}^{(j)}+N+n_{0,k}^{(j)}}_{x}v\right|$
+    $\geq$ $\frac{C}{4K^{6}}\cos\varphi\cdot\mu^{n_{0,k}^{(j)}}\rho^{T}\left(\lambda^{\theta}\mu^{2}\right)^{n_{k}^{(j)}}\left|v\right|$
+    $\geq$ $C\rho^{T+n_{k}^{(j)}+N+n_{0,k}^{(j)}+s}\left|v\right|.$
     Observe that in the last inequality of $\text{IH}_{\tilde{C}}-$(i)P6 we used that
-    $\frac{1}{4K^{6}}\cos\varphi\left(\frac{\mu}{\rho}\right)^{n+n_{0}}\left(\lambda^{\theta}\mu\right)^{n}\geq\rho^{N+k},\text{ }\forall 1\leq k\leq M$
+    $\frac{1}{4K^{6}}\cos\varphi\left(\frac{\mu}{\rho}\right)^{n_{k}^{(j)}+n_{0,k}^{(j)}}\left(\lambda^{\theta}\mu\right)^{n_{k}^{(j)}}\geq\rho^{N+s},\text{ }\forall 1\leq s\leq M$
     which is equivalent to ask
-    $\frac{1}{4K^{6}}\cos\varphi\left(\frac{\mu}{\rho}\right)^{n_{0}}\left(\frac{\lambda^{\theta}\mu^{2}}{\rho}\right)^{n}\geq\rho^{N+M}.$
-    This follow from the condition imposed in (8)-(oo) and taking $n$ large enough. In particular, $\text{IH}_{\tilde{C}}-$(i)P6 proves
-    $\left|Df^{T+n+N+n_{0}+M}_{x}v\right|\geq\frac{C}{4K^{6}}\cos\varphi\mu^{n_{0}}\rho^{T}\left(\lambda^{\theta}\mu^{2}\right)^{n}\left|v\right|.$ (8.13)
+    $\frac{1}{4K^{6}}\cos\varphi\left(\frac{\mu}{\rho}\right)^{n_{0,k}^{(j)}}\left(\frac{\lambda^{\theta}\mu^{2}}{\rho}\right)^{n_{k}^{(j)}}\geq\rho^{N+M}.$
+    This follow from the condition imposed in (8)-(oo) and taking $n_{k}^{(j)}$ large enough. In particular, $\text{IH}_{\tilde{C}}-$(i)P6 proves
+    $\left|Df^{T+n_{k}^{(j)}+N+n_{0,k}^{(j)}+M}_{x}v\right|\geq\frac{C}{4K^{6}}\cos\varphi\cdot\mu^{n_{0,k}^{(j)}}\rho^{T}\left(\lambda^{\theta}\mu^{2}\right)^{n_{k}^{(j)}}\left|v\right|.$ (8.0.15)
     Observe that, there exists $\beta>0$ such that
     $\angle\left(Df^{M}_{q_{2}}e_{2},e_{1}\right)>\beta$
-    and by (8.12),
-    $\text{d}_{\text{v}}\left(f^{T+n+N+n_{0}}(x),q_{2}\right)<<1.$
-    The previous two inequalities together with $\text{IH}_{H}$ gives
-    $\angle\left(Df^{T+n+N+n_{0}+M}_{x}v,e_{1}\right)>\frac{\beta}{2}.$
+    and by (7.4.3),
+    $\text{dist}\left(f^{T+n_{k}^{(j)}+N+n_{0,k}^{(j)}}(x),q_{2}\right)<<1.$
+    The previous two inequalities, (8.0.14) together with $\text{IH}_{H}$ gives
+    $\angle\left(Df^{T+n_{k}^{(j)}+N+n_{0,k}^{(j)}+M}_{x}v,e_{1}\right)>\frac{\beta}{2}.$
     In particular, if we denote by
-    $v^{(6)}=Df^{T+n+N+n_{0}+M}_{x}v=\left(\begin{matrix}v^{(6)}_{1}\\ v^{(6)}_{2}\end{matrix}\right)$,
+    $v^{(6)}=Df^{T+n_{k}^{(j)}+N+n_{0,k}^{(j)}+M}_{x}v=\left(\begin{matrix}v^{(6)}_{1}\\ v^{(6)}_{2}\end{matrix}\right)$,
     it follows that
-    $\frac{v^{(6)}_{2}}{\left|v^{(6)}\right|}\geq\sin\frac{\beta}{2}.$ (8.14)
+    $\frac{v^{(6)}_{2}}{\left|v^{(6)}\right|}\geq\sin\frac{\beta}{2}.$ (8.0.16)
   • Step 7:
-    We start proving $\text{IH}_{\tilde{C}^{\tilde{N}-\tilde{T}}}-$(i) from time $T+nN+n_{0}+M+1$ up to time $T+n+N+n_{0}+M+\overline{\ell}$.
-Indeed, by $\text{IH}_{Q}$, (8.14) and (8.13), for all $1\leq k\leq\overline{\ell}$, we get,
-    • $\text{IH}_{\tilde{C}}-$(i)P7:
-      $\left|Df^{T+n+N+n_{0}+M+k}_{x}v\right|$ $\geq$ $\mu^{k}\sin\frac{\beta}{2}\left|Df^{T+n+N+n_{0}+M}_{x}v\right|$
-      $\geq$ $\frac{C}{4K^{6}}\cos\varphi\sin\frac{\beta}{2}\mu^{n_{0}+k}\rho^{T}\left(\lambda^{\theta}\mu^{2}\right)^{n}\left|v\right|$
-      $\geq$ $C\rho^{T+n+N+n_{0}+M+k}\left|v\right|.$
+    Now we prove $\text{IH}_{\tilde{C}^{\tilde{N}-\tilde{T}}}-$(i) from time $T+n_{k}^{(j)}+N+n_{0,k}^{(j)}+M+1$ up to time $\tilde{T}=T+n_{k}^{(j)}+N+n_{0,k}^{(j)}+M+\underline{\ell}_{k}^{(j)}$.
+Indeed, by $\text{IH}_{Q}$, (8.0.16) and (8.0.15), for all $1\leq s\leq\underline{\ell}_{k}^{(j)}$, we get,
+    $\text{IH}_{\tilde{C}}\text{-(i)P7: }\left|Df^{T+n_{k}^{(j)}+N+n_{0,k}^{(j)}+M+s}_{x}v\right|$ $\geq$ $\mu^{s}\sin\frac{\beta}{2}\left|Df^{T+n_{k}^{(j)}+N+n_{0,k}^{(j)}+M}_{x}v\right|$
+    $\geq$ $\frac{C}{4K^{6}}\cos\varphi\cdot\sin\frac{\beta}{2}\mu^{n_{0,k}^{(j)}+s}\rho^{T}\left(\lambda^{\theta}\mu^{2}\right)^{n_{k}^{(j)}}\left|v\right|$
+    $\geq$ $C\rho^{T+n_{k}^{(j)}+N+n_{0,k}^{(j)}+M+s}\left|v\right|.$
     Observe that in the last inequality of $\text{IH}_{\tilde{C}}-$(i)P7 we used that
-    $\frac{1}{4K^{6}}\cos\varphi\sin\frac{\beta}{2}\left(\frac{\mu}{\rho}\right)^{n_{0}+k}\left(\frac{\lambda^{\theta}\mu^{2}}{\rho}\right)^{n}\geq\rho^{N+M}.$
-    This follow from the condition imposed in (8)-(oo) and taking $n$ large enough. Denote by
-    $v^{(7)}=Df^{T+n+N+n_{0}+M+\overline{\ell}}_{x}v=\left(\begin{matrix}v^{(7)}_{1}\\ v^{(7)}_{2}\end{matrix}\right)$,
-    then, by $\text{IH}_{Q}$ and by (8.14), we get
-    $\frac{v^{(7)}_{1}}{v^{(7)}_{2}}=\left(\frac{\lambda}{\mu}\right)^{\overline{\ell}}\frac{v^{(6)}_{1}}{v^{(6)}_{2}}\leq\left(\frac{\lambda}{\mu}\right)^{\overline{\ell}}\cot\frac{\beta}{2}<\tan\varphi,$
-    with the last inequality achieved by taking $\overline{\ell}$ large enough.
+    $\frac{1}{4K^{6}}\cos\varphi\cdot\sin\frac{\beta}{2}\left(\frac{\mu}{\rho}\right)^{n_{0,k}^{(j)}+s}\left(\frac{\lambda^{\theta}\mu^{2}}{\rho}\right)^{n_{k}^{(j)}}\geq\rho^{N+M}.$
+    This follow from the condition imposed in (8)-(oo) and taking $n_{k}^{(j)}$ large enough. Denote by
+    $v^{(7)}=Df^{T+n_{k}^{(j)}+N+n_{0,k}^{(j)}+M+\underline{\ell}_{k}^{(j)}}_{x}v=\left(\begin{matrix}v^{(7)}_{1}\\ v^{(7)}_{2}\end{matrix}\right)$,
+    then, by $\text{IH}_{Q}$ and by (8.0.16), we get
+    $\frac{v^{(7)}_{1}}{v^{(7)}_{2}}=\left(\frac{\lambda}{\mu}\right)^{\underline{\ell}_{k}^{(j)}}\frac{v^{(6)}_{1}}{v^{(6)}_{2}}\leq\left(\frac{\lambda}{\mu}\right)^{\underline{\ell}_{k}^{(j)}}\frac{1}{\sin\beta/2}<\tan\varphi,$
+    with the last inequality achieved by taking $j$ large enough which implies that $\underline{\ell}_{k}^{(j)}$} is large enough.
 In conclusion, the inequality above proves $\text{IH}_{\tilde{C}}-$(ii) and $\text{IH}_{\tilde{C}}-$(i) is achieved by putting together $\text{IH}_{\tilde{C}}-$(i)P1, $\text{IH}_{\tilde{C}}-$(i)P2, $\text{IH}_{\tilde{C}}-$(i)P3, $\text{IH}_{\tilde{C}}-$(i)P4, $\text{IH}_{\tilde{C}}-$(i)P5, $\text{IH}_{\tilde{C}}-$(i)P6 and $\text{IH}_{\tilde{C}}-$(i)P7.
-9 The main objects: the lamination and the invariant Cantor set
+∎
+9 The lamination and the invariant Cantor set
 In this section we build a lamination in parameter space whose maps have an invariant Cantor set.
@@
 Take $U=S$.
-In Subsection 2.1, we defined a family $F$ which is an unfolding of the tangency $b$ and each map $f\in S$ has a $(1,N_{1},M_{1})$ dynamical partition with $N_{1}=N$ and $M_{1}=M$, where $N$ and $M$ are defined in Definition 2, property $(f10)$.
+In Subsection 2.1, we defined a family $F$ which is an unfolding of the tangency $b$ and each map $f\in S$ has a $(1,N_{1},M_{1})$ dynamical partition with $N_{1}=N$ and $M_{1}=M$, where $N$ and $M$ are defined in Definition 2.1.1, property $(f8)$.
 Denote by
-$\mathfrak{S}_{1}=S$
+$\mathfrak{S}_{1}=\left\{S\right\},$
 and by
-$\mathfrak{S}_{2}=\left\{S_{j,2}\right\}_{j=1}^{\infty}$
-where each $S_{j,2}$ is a $n_{2}+j$-tangency strip contained in $S$ whose existence is guaranteed by Proposition 7.
-By applying Proposition 7 inductively, for all $g\geq 3$ one can define a sequence of tangency strips
-$\mathfrak{S}_{g}=\left\{S_{j,g}\right\}_{j=1}^{\infty}$
+$\mathfrak{S}_{2}=\left\{S_{j,2}\right\}_{j=1}^{\infty},$
+where each $S_{j,2}$ is a $n_{2}+j$-tangency strip contained in $S$ whose existence is guaranteed by Proposition 7.4.1.
+By applying Proposition 7.4.1 inductively, for all generations $g\geq 3$ one can define a sequence of collections of tangency strips
+$\mathfrak{S}_{g}=\left\{S_{j,g}\right\}_{j=1}^{\infty},$
 satisfying the following properties:
   • $\forall j,j^{\prime}$ with $j\neq j^{\prime}$, $S_{j,g}\cap S_{j^{\prime},g}=\emptyset$,
-  • $\forall S_{j,g}\in\mathfrak{S}_{g}$ there exists a unique $S_{i,g-1}\in\mathfrak{S}_{g-1}$ with $S_{j,g}\subset S_{i,g-1}$.
-We are now ready to define the lamination
-$\mathfrak{L}=\bigcap_{g=1}^{\infty}\bigcup_{j=1}^{\infty}S_{j,g}.$ (9.1)
-A leave of the lamination is defined as
-$\ell=\bigcap_{g=1}^{\infty}S_{g}$ (9.2)
+  • $\forall S_{j,g}\in\mathfrak{S}_{g}$ there exists a unique $S_{i,g-1}\in\mathfrak{S}_{g-1}$ with $S_{j,g}\subset S_{i,g-1}$,
+  • $\forall j,$ $S_{j,g}$ is bounded by the graph of two functions $s_{j,g}^{+},s_{j,g}^{-}:[-t_{0},t_{0}]\to\mathbb{R}$ such that,
+    $\left\lVert s_{j,g}^{+}-s_{j,g}^{-}\right\rVert_{0}\leq\frac{1}{2^{g}},$
+    and if $S_{j,g}\subset S_{i,g-1}$, then
+    $\left\lVert\frac{ds_{j,g}^{\pm}}{dt}-\frac{ds_{i,g-1}^{\pm}}{dt}\right\rVert_{0}\leq\frac{1}{2^{g}}.$
+This last property relies on (7.4.2). Observe that $\cup\mathfrak{S}_{g}\subset\cup\mathfrak{S}_{g-1}$. We are now ready to define the lamination
+$\mathfrak{L}=\bigcap_{g=1}^{\infty}\bigcup_{j=1}^{\infty}S_{j,g}\subset[-t_{0},t_{0}]\times[0,a_{0}].$ (9.1.1)
+The connected components of $\mathfrak{L}$ are called leafs of the lamination. Namely, a leaf of the lamination has the form
+$\ell=\bigcap_{g=1}^{\infty}S_{g},$ (9.1.2)
 where, for all $g$, $S_{g}\in\mathfrak{S}_{g}$ and $S_{g}\subset S_{g-1}$.
-Observe that, by Equation (7.1), $\mathfrak{L}$ is a ${\mathcal{C}}^{1}$} lamination, i.e.
-each leave is the graph of a ${\mathcal{C}}^{1}$} function.
-The following proposition proves Theorem A-(v).
-Proposition 9.
-Every leaves of $\mathfrak{L}$ is the graph of a smooth function $\ell:[-t_{0},t_{0}]\mapsto[-a_{0},a_{0}]$.
+Observe that $\mathfrak{L}$ is a ${\mathcal{C}}^{1}$} lamination, i.e.
+each leaf is the graph of a ${\mathcal{C}}^{1}$} function, $\lim_{q\to\infty}s^{+}_{g}$ where the graph of $s^{+}_{g}$ is the upper boundary of $S_{g}$.
+The following proposition proves Theorem A-(v). The proof is a consequence of the construction of the lamination and of the fact that the strips of the same generation are pairwise disjoint.
+Proposition 9.1.3.
+Every leaf of $\mathfrak{L}$ is the graph of a smooth function $\ell:[-t_{0},t_{0}]\mapsto[-a_{0},a_{0}]$.
 Morever if $t\in[-t_{0},t_{0}]$, the closure of the set $L_{t}=\left\{(t,a)\in\mathfrak{L}\right\}$ is a Cantor set.
-Proof.
-The proof follows from the construction of the lamination, and Proposition 7 and Proposition 4.
-The main point is that each strip contains leaves corresponding to maps with a new homoclinic tangency and these leaves are graphs of global curves.
-∎
+Remark 9.1.4.
+As in the proof of Proposition 8.0.4 by taking all $n_{g}^{(j)}$ large enough we may also assume that, for all $g$, $N_{g+1},M_{g+1}\geq 32\max\{N_{g},M_{g}\}$ and $N_{g},M_{g}\geq 3$.
+The following Remark proves Theorem A-(viii)
+Remark 9.1.5.
+Observe that, by construction, given an open neighborhood of $f\in\mathfrak{L}$, it contains in parameter space a curve $b$ of secondary tangencies.
+Hence, by Remark 6.2.6, such open neighborhood contains maps with infinitely many sinks.
 9.2 The invariant Cantor set
-The following discussion is consequence of Proposition 7.
-Given a leave of the lamination $\ell$ and a map $f\in\ell$, one can associate a sequence $\left\{\left(b_{g},U_{g},\mathfrak{D}_{g}\right)\right\}_{g=1}^{\infty}$ and a sequence of numbers $\left\{\left(1,N_{g},M_{g}\right)\right\}_{g=1}^{\infty}$ such that, for all $g$, $(b_{g},U_{g})$ is the $(1,N_{g},M_{g})$ descendant of $S_{g-1}$ and $\mathfrak{D}_{g}$ is the $(1,N_{g},M_{g})$ dynamical partition of $f$.
+Given a leaf $\ell=\cap S_{g}$ of the lamination $\mathfrak{L}$ and a map $f\in\ell$, one can associate a sequence $\left\{\left(b_{g},U_{g},\mathfrak{D}_{g}\right)\right\}_{g=1}^{\infty}$ and a sequence of numbers $\left\{\left(1,N_{g},M_{g}\right)\right\}_{g=1}^{\infty}$ such that, for all $g$, $(b_{g},U_{g})$ is the $(1,N_{g},M_{g})$ descendant of $S_{g-1}$ and $\mathfrak{D}_{g}$ is the $(1,N_{g},M_{g})$ dynamical partition of $f$.
 The sequence $\left\{\left(1,N_{g},M_{g}\right)\right\}_{g=1}^{\infty}$ is called the type of $\ell$.
 Moreover for each $g$, $\mathfrak{D}_{g}$ is a refinement of $\mathfrak{D}_{g-1}$, with mesh going to zero when $g$ goes to infinity. We can now define the set associated to $f$
-$A(f)=\bigcap_{g=1}^{\infty}\mathcal{D}_{g}=\bigcap_{g=1}^{\infty}\bigcup_{D_{g}\in\mathfrak{D}_{g}}{D}_{g}$
+$A(f)=\bigcap_{g=1}^{\infty}\mathcal{D}_{g}=\bigcap_{g=1}^{\infty}\bigcup_{D\in\mathfrak{D}_{g}}{D}$
 where $\mathcal{D}_{g}$ is the union of the elements of $\mathfrak{D}_{g}$.
-This set is going to play a fundamental role in our research.
+This set is going to play a fundamental role.
 Let us begin by studying its properties.
-Following Proposition 7, for all $g$, let $X_{g}$ be the $(1,N_{g},M_{g})$ directed graph and let $w_{g}:X_{g+1}\to X_{g}$ be the projection and $w_{g^{\prime}g}:X_{g}\to X_{g^{\prime}}$ be
+Following Proposition 7.4.1, for all $g$, let $X_{g}$ be the $(1,N_{g},M_{g})$ directed graph, let $w_{g}:X_{g+1}\to X_{g}$ be the projection and $w_{g^{\prime}g}:X_{g}\to X_{g^{\prime}}$ be
 $w_{g^{\prime}g}=w_{g^{\prime}}\circ\cdots w_{g-1}\circ w_{g}.$
 9.3 Properties of the set AA
-In this subsection we collect properties of the set $A$. In particular, the following lemma proves Theorem A-(i).
-Lemma 22.
-Let $\ell$ be a leave of the lamination $\mathfrak{L}$ and let $f\in\ell$. Then the set $A(f)$ is a Cantor set.
+In this subsection we collect properties of the set $A$.
+Lemma 9.3.1.
+Let $\ell$ be a leaf of the lamination $\mathfrak{L}$ and let $f\in\ell$. Then the set $A(f)$ is a Cantor set.
 Proof.
@@
 It is left to prove that $A(f)$ is perfect.
-Let us start by observing that, by Remark 14, each set $D_{g}\in\mathfrak{D}_{g}$ contains at least two sets $D_{g+1},D^{\prime}_{g+1}\in\mathfrak{D}_{g+1}$.
+First observe that, by Remark 7.4.4, each set $D_{g}\in\mathfrak{D}_{g}$ contains at least two sets $D_{g+1},D^{\prime}_{g+1}\in\mathfrak{D}_{g+1}$, see Figure 12.
 Now take $x\in A(f)$, then for every $g$ there exists $D_{g}\in\mathfrak{D}_{g}$ such that $x\in D_{g}$.
@@
 ∎
-In the following lemma we describe the possible limit behavior of points in $A$.
-We denote by $\omega(x)$, the $\omega$-limit set of the point $x$.
-In particular, it implies that the set $A$ is non-hyperbolic.
-Recall that $q_{2}$ is the point with a transversal homoclinic intersection. Let
+In the following lemma we describe the possible limit behavior of points in $A$. In particular, it implies that the set $A$ is non-hyperbolic.
+We denote by $\omega(x)$, the $\omega$-limit set of the point $x$. Recall that $q_{2}$ is the point with a transversal homoclinic intersection. Let
 $O_{2}=\left\{f^{t}(q_{2})\left|\right.t\in\mathbb{Z}\right\}\cup\left\{p\right\}.$
@@
 Denote the corresponding prime loops in $X_{g}$ respectively by $\lambda^{q}_{g}$, $\lambda^{h}_{g}$, and $\lambda^{c}_{g}$.
-We will use frequently the following observation.
-If a loop in $X_{g}$ consists purely of $h$- and $q$-loops of $X_{g}$ then $w_{g^{\prime}g}(\lambda)$ consists purely of $h$- and $q$-loops of $X_{g^{\prime}}$.
-Lemma 23.
-Let $\ell$ be a leave of the lamination $\mathfrak{L}$ and let $f\in\ell$. Each $x\in A(f)$ satisfies one of the following:
+We will use frequently the following observation: if a loop in $X_{g}$ consists purely of $h$- and $q$-loops of $X_{g}$ then $w_{g^{\prime}g}(\lambda)$ consists purely of $h$- and $q$-loops of $X_{g^{\prime}}$.
+Lemma 9.3.2.
+Let $\ell$ be a leaf of the lamination $\mathfrak{L}$ and let $f\in\ell$. Each $x\in A(f)$ satisfies one of the following:
   • $\omega(x)=\left\{p\right\}$,
@@
 We claim that, if for infinitely many $g$,
-$w_{g-1}\left(\overrightarrow{x_{g}q_{g}}\right)\supset\lambda^{c}_{g-1}$
+$w_{g-1}\left(\overrightarrow{x_{g}q_{g}}\right)\supset\lambda^{c}_{g-1}$ (9.3.3)
 then $\omega(x)=A(f)$. Indeed, take an open set $U\subset A(f)$ and let $g_{0}$ be such that there exists $\tilde{D}_{g_{0}}\in\mathfrak{D}_{g_{0}}$ with $\tilde{D}_{g_{0}}\subset U$ and
@@
 Hence, there exists $t$ such that $f^{t}\left(D_{g_{0}+2}\right)\subset\tilde{D}_{g_{0}}$ and in particular $f^{t}(x)\in\tilde{D}_{g_{0}}\subset U$.
+In particular we showed that under if (9.3.3) holds for infinitely many $g$ then $\omega(x)=A(f)$.
 Now consider the opposite situation when there exists $g_{0}\in\mathbb{N}$ such that $w_{g-1}\left(\overrightarrow{x_{g}q_{g}}\right)\subset X_{g-1}$ does not contain $\lambda^{c}_{g-1}$ for every $g\geq g_{0}$.
@@
 $\omega(x)=O_{2}.$
-The proof relies on the following claim. Let $g_{0}\leq g^{\prime}<g$. There exists $t_{g^{\prime}}\geq 0$ such that for every $g>g^{\prime}$
-$w_{g^{\prime}g}(\overrightarrow{y_{g}q_{g}})=\lambda_{g^{\prime}}^{q}\cup\lambda_{g^{\prime}}^{h},$
-where $y=f^{t_{g^{\prime}}}(x)$. In other words, the loop $w_{g^{\prime}g}(\overrightarrow{y_{g}q_{g}})$ is a concatenation of multiple times the $q$-loop and $h$-loop of $X_{g^{\prime}}$.
-The proof of this Claim is by induction. First observe
-$w_{g^{\prime}}(\overrightarrow{x_{g^{\prime}+1}q_{g^{\prime}+1}})=\gamma\lambda,$
+The proof relies on the following claim. Let $g_{0}\leq g^{\prime}$. There exists $t_{g^{\prime}}\geq 0$ such that the point $y=f^{t_{g^{\prime}}}(x)$ has the property that for every $g>g^{\prime}$,
+$w_{g^{\prime}g}(\overrightarrow{y_{g}q_{g}})=\lambda_{g^{\prime}}^{q}\cup\lambda_{g^{\prime}}^{h}.$
+In other words, the path $w_{g^{\prime}g}(\overrightarrow{y_{g}q_{g}})$ is a concatenation of multiple times the $q$-loop and $h$-loop of $X_{g^{\prime}}$.
+The proof of this Claim is by induction.
+First observe
+$w_{g^{\prime}}(\overrightarrow{x_{g^{\prime}+1}q_{g^{\prime}+1}})=\lambda\gamma,$
 where $\gamma=w_{g^{\prime}}(\overrightarrow{x_{g^{\prime}+1}y_{g^{\prime}+1}})\subset\lambda^{c}_{g^{\prime}}$ is a path of length $t_{g^{\prime}}$ and $y=f^{t_{g^{\prime}}}(x)$.
@@
 where $\lambda$ is a loop which is a concatenation of multiple $\lambda_{g}^{q}$ and one $\lambda_{g}^{h}$.
-This is true because $w_{g}(\overrightarrow{y_{g+1}q_{g+1}})$ cannot contain the $c$-loop of $X_{g}$, we are in the second case.
+This is true because $w_{g}(\overrightarrow{y_{g+1}q_{g+1}})\subset w_{g}(\overrightarrow{x_{g+1}q_{g+1}})$ which cannot contain the $c$-loop of $X_{g}$, we are in the second case.
 The induction assumption implies that
@@
 The Claim implies that for all $g^{\prime}\geq g_{0}$,
-$\text{Orbit}(f^{t_{g}^{\prime}}(x))\subset Q_{g^{\prime}}\cup\bigcup_{j=1}^{M_{g^{\prime}}-1}H_{g^{\prime}}(j).$
+$\text{Orbit}(f^{t_{g^{\prime}}}(x))\subset Q_{g^{\prime}}\cup\bigcup_{j=1}^{M_{g^{\prime}}-1}H_{g^{\prime}}^{j}.$
 Hence,
-$\omega(x)\subset\bigcap_{g\geq g_{0}}\left(Q_{g}\cup\bigcup_{j=1}^{M_{g}-1}H_{g}(j)\right)$
-where for all $g$ and for all $j$, $Q_{g}$ and $H_{g}(j)$ are the sets in $\mathfrak{D}_{g}$ as described in Definition 7. In order to complete the proof one should establish that
-$\bigcap_{g}\left(Q_{g}\cup\bigcup_{j=1}^{M_{g}-1}H_{g}(j)\right)=O_{2}.$
-One inclusion is trivial. Let us concentrate on the other one and for this purpose let us take a point
-$x\in\bigcap_{g}\left(Q_{g}\cup\bigcup_{j=1}^{M_{g}-1}H_{g}(j)\right).$
+$\omega(x)\subset\bigcap_{g\geq g_{0}}\left(Q_{g}\cup\bigcup_{j=1}^{M_{g}-1}H_{g}^{j}\right)$
+where for all $g$ and for all $j$, $Q_{g}$ and $H_{g}^{j}$ are the sets in $\mathfrak{D}_{g}$ as described in Definition 7.1.1. In order to complete the proof one should establish that
+$\bigcap_{g}\left(Q_{g}\cup\bigcup_{j=1}^{M_{g}-1}H_{g}^{j}\right)=O_{2}.$
+One inclusion is trivial, since each $h$-loop contains $q_{2}$. Let us concentrate on the other one and for this purpose let us take a point
+$x\in\bigcap_{g}\left(Q_{g}\cup\bigcup_{j=1}^{M_{g}-1}H_{g}^{j}\right).$
 If $x=p$ then $x\in O_{2}$. Assume $x\neq p$, then from a certain moment on
-$x\in\bigcap_{g\geq\tilde{g}}\bigcup_{j=1}^{M_{g}-1}H_{g}(j).$
+$x\in\bigcap_{g\geq\tilde{g}}\bigcup_{j=1}^{M_{g}-1}H_{g}^{j}.$
 Here we are using the fact that if $x_{g}\neq q_{g}$, then the same holds for all next points in the sequence defining the point $x$. Say, for $g\geq\tilde{g}$
-$x\in H_{g}(j_{x}).$
-In particular, there exists a number $t_{g}$ such that
-$f^{t_{g}}(q_{2})\in H_{g}(j_{x}).$
+$x\in H_{g}^{j_{x}}.$
+In particular, there exists a number $t_{g}\in\mathbb{Z}$ such that
+$f^{t_{g}}(q_{2})\in H_{g}^{j_{x}}.$
 Because, $f^{t}(q_{2})\to p$ when $|t|\to\infty$ we have that $t_{g}$ is bounded.
-Observe also that $f^{t_{g}}(q_{2})$ is close to $x$ because the diameter of $H_{g}(j_{x})$ goes to zero when $g$ goes to infinity.
-It follows that $f^{t}(q_{2})=x$ form some $t\in\mathbb{Z}$.
+Observe also that $f^{t_{g}}(q_{2})$ is close to $x$ because the diameter of $H_{g}^{j_{x}}$ goes to zero when $g$ goes to infinity.
+It follows that $f^{t}(q_{2})=x$ for some $t\in\mathbb{Z}$.
 Hence, $x\in O_{2}$ and the proof is complete.
 ∎
-A natural question at this point is: does a point with dense orbit exist? The answer is yes and the proof is the following lemma which also proves Theorem A-(ii).
-Lemma 24.
-Let $\ell$ be a leave of the lamination $\mathfrak{L}$ and let $f\in\ell$. There exists a point $c\in A(f)$ such that $\omega(c)=A(f).$
+A natural question at this point is: does a point with dense orbit exist? The answer is yes. The following lemma also confirms Theorem A-(ii).
+Lemma 9.3.4.
+Let $\ell$ be a leaf of the lamination $\mathfrak{L}$ and let $f\in\ell$. There exists a point $c\in A(f)$ such that $\omega(c)=A(f).$
 Proof.
-Fix $C_{1}\in\mathfrak{D}_{1}$ and consider $c_{1}=\pi_{1}(C_{1})\in X_{1}$ where $\pi_{1}$ is the projection of $\mathcal{D}_{1}$ on $X_{1}$.
-We want to prove that for infinitely many $g$ there is an element $c_{g}$ in the $c$-loop of $X_{g}$ such that
+Chose $C_{1}\in\mathfrak{D}_{1}$ such that $c_{1}=\pi_{1}(C_{1})\in\lambda^{c}_{1}$ where $\pi_{1}$ is the projection of $\mathcal{D}_{1}$ on $X_{1}$.
+We want to prove that for all $g$ there is an element $c_{g}$ in the $c$-loop of $X_{g}$ such that
 $w_{g-1}\left(\overrightarrow{c_{g}q_{g}}\right)\supset\lambda^{c}_{g-1}$
 and $w_{g-1}(c_{g})=c_{g-1}$.
-Suppose we have constructed $c_{1},\dots,c_{g-1}$.
-Consider the unique path $\gamma=\overrightarrow{c_{g-1}q_{g-1}}$ contained in the $c$-loop of $X_{g-1}$.
-Take the first vertex $c_{g}$ in the $c$-loop of $X_{g}$ with the property that $w_{g-1}(c_{g})=c_{g-1}$.
-Observe that
-$w_{g-1}\left(\overrightarrow{c_{g}q_{g}}\right)=\gamma\circ\lambda$
-where $\lambda$ is a loop and, by Remark 14 $\lambda$ contains all primary loops of $X_{g-1}$. In particular $w_{g-1}\left(\overrightarrow{c_{g}q_{g}}\right)\supset\lambda^{c}_{g-1}$.
-For all $g$, consider the set of $\mathfrak{D}_{g}$ defined by $\pi_{g}(C_{g})=c_{g}$, where $\pi_{g}$ is the projection of $D_{g}$ on $X_{g}$. Our selected point is
+Suppose we have constructed $c_{1},\dots,c_{g-1}$. Take the first vertex $c_{g}$ in the $c$-loop of $X_{g}$ with the property that $w_{g-1}(c_{g})=c_{g-1}$. Observe that
+$w_{g-1}\left(\overrightarrow{c_{g}q_{g}}\right)=\lambda\overrightarrow{c_{g-1}q_{g-1}}$
+where $\lambda$ is a loop and, by Remark 7.4.4 $\lambda$ contains all primary loops of $X_{g-1}$. In particular $w_{g-1}\left(\overrightarrow{c_{g}q_{g}}\right)\supset\lambda^{c}_{g-1}$.
+For all $g$, consider the set of $C_{g}$ defined by $\pi_{g}(C_{g})=c_{g}$, where $\pi_{g}$ is the projection of $\mathcal{D}_{g}$ on $X_{g}$. Our selected point is
 $c=\bigcap_{g}C_{g}\in A(f).$
-Following the proof of Lemma 23 one gets that $\omega(c)=A(f)$. ∎
-The following lemma proves Theorem A-(iii).
-Lemma 25.
-Let $\ell$ be a leave of the lamination $\mathfrak{L}$ and let $f\in\ell$. There exists a point $b\in A(f)$ such that $\omega(b)=O_{2}.$
+Observe that $c$ satisfies (9.3.3) for all $g$. Hence, $\omega(c)=A(f)$. ∎
+The following lemma confirms Theorem A-(iii).
+Lemma 9.3.5.
+Let $\ell$ be a leaf of the lamination $\mathfrak{L}$ and let $f\in\ell$. There exists a point $b\in A(f)$ such that $\omega(b)=O_{2}.$
 Proof.
 Observe, for every $x_{g-1}\in\lambda^{c}_{g-1}\subset X_{g-1}$ there exists a unique $x_{g}\in\lambda^{c}_{g}\subset X_{g}$ such that
-$w_{g-1}(\overrightarrow{x_{g}q_{g}})=\overrightarrow{x_{g-1}q_{g-1}}\lambda,$
+$w_{g-1}(\overrightarrow{x_{g}q_{g}})=\lambda\overrightarrow{x_{g-1}q_{g-1}},$
 where $\lambda$ is a concatenation of multiple $\lambda^{q}_{g-1}$ and only one $\lambda^{h}_{g-1}$.
@@
 The proof of the Claim is by induction. It holds for $g=g_{0}+1$. By construction we have
-$w_{g}(\overrightarrow{y_{g+1}q_{g+1}})=\overrightarrow{y_{g}q_{g}}\lambda,$
+$w_{g}(\overrightarrow{y_{g+1}q_{g+1}})=\lambda\overrightarrow{y_{g}q_{g}},$
 where $\lambda$ is a concatenation of multiple $\lambda^{q}_{g}$ and one $\lambda^{h}_{g}$. Then
-$w_{g_{0}g}(\overrightarrow{y_{g+1}q_{g+1}})=w_{g_{0}g}(\overrightarrow{y_{g}q_{g}})w_{g_{0}g}(\lambda).$
+$w_{g_{0}g}(\overrightarrow{y_{g+1}q_{g+1}})=w_{g_{0}g}(\lambda)w_{g_{0}g}(\overrightarrow{y_{g}q_{g}}).$
 The induction hypothesis implies that $w_{g_{0}g}(\overrightarrow{y_{g}q_{g}})$ is a concatenation of multiple $\lambda^{q}_{g_{0}}$ and $\lambda^{h}_{g_{0}}$.
@@
 The Claim implies for all $g\geq 1$
-$\omega(x)\subset Q_{g}\cup\bigcup_{j=1}^{M_{g}-1}H_{g}(j)$
+$\omega(x)\subset Q_{g}\cup\bigcup_{j=1}^{M_{g}-1}H_{g}^{j}$
 and
@@
 and for all $j\leq M_{g}-1$
-$\omega(x)\cap H_{g}(j)\neq\emptyset.$
+$\omega(x)\cap H_{g}^{j}\neq\emptyset.$
 This means that $\omega(x)=O_{2}$ where we use that
-$\bigcap_{g}\left(Q_{g}\cup\bigcup_{j=1}^{M_{g}-1}H_{g}(j)\right)=O_{2}.$
+$\bigcap_{g}\left(Q_{g}\cup\bigcup_{j=1}^{M_{g}-1}H_{g}^{j}\right)=O_{2}.$
 In the next proposition we prove that $A$ is an invariant set.
-Proposition 10.
-Let $\ell$ be a leave of the lamination $\mathfrak{L}$ and let $f\in\ell$. Then the map $f:A(f)\mapsto A(f)$ is an homeomorphism.
+Proposition 9.3.6.
+Let $\ell$ be a leaf of the lamination $\mathfrak{L}$ and let $f\in\ell$. Then the map $f:A(f)\mapsto A(f)$ is an homeomorphism.
 Proof.
@@
 Observe that for all $x$, $f(\omega(x))=\omega(x)$.
-In particular, when $x=c$, $f(\omega(c))=\omega(c)$ and by Lemma 24 we obtain $f(A(f))=A(f)$.
+In particular, when $x=c$, $f(\omega(c))=\omega(c)$ and by Lemma 9.3.4 we obtain $f(A(f))=A(f)$.
 The proof is complete.
@@
 10 Topological Equivalence along the Leaves
-The aim of this section is to prove that two maps which are in the same leave of the lamination are conjugate on their respective Cantor sets, see Theorem 4.
+The aim of this section is to prove that two maps which are in the same leaf of the lamination are conjugate on their respective Cantor sets, see Theorem 10.2.3.
 In order to do that, we need to build a topological model for the invariant Cantor sets.
 10.1 Inverse limits
-Take a leave $\ell$ of the lamination $\mathfrak{L}$ as constructed in (9.1) and (9.2).
-By the construction following Proposition 7 we know that for all $g$ there is a $(1,N_{g},M_{g})$-directed graph and a graph morphism $w_{g}:X_{g+1}\mapsto X_{g}$.
+Take a leaf $\ell$ of the lamination $\mathfrak{L}$ as constructed in (9.1.1) and (9.1.2).
+By the construction following Proposition 7.4.1 we know that for all $g$ there is a $(1,N_{g},M_{g})$-directed graph and a graph morphism $w_{g}:X_{g+1}\mapsto X_{g}$.
 We can then define the inverse limit as
@@
 $\left\{\underline{x}_{n}\right\}_{n\in\mathbb{N}}\mapsto\underline{x}\iff\forall g\text{ }\exists n_{0}\in\mathbb{N}\text{ s.t }\forall n\geq n_{0}\text{, }x_{g,n}=x_{g}.$
-Lemma 26.
-Let $\underline{x}\in X$, $\underline{x}=(x_{g})$. If there exist $g_{0}$ such that $x_{g_{0}}=q_{g_{0}}$ then for all $g\leq g_{0}$, $x_{g}=q_{g}$.
+In particular $X$ is a Cantor set.
+Lemma 10.1.1.
+Let $\underline{x}\in X$, $\underline{x}=(x_{g})$. If there exists $g_{0}$ such that $x_{g_{0}}=q_{g_{0}}$ then for all $g\leq g_{0}$, $x_{g}=q_{g}$.
 Proof.
 Since for all $g$, $w_{g-1}(q_{g})=q_{g-1}$, then $x_{g_{0}-1}=w_{g_{0}-1}(x_{g_{0}})=w_{g_{0}-1}(q_{g_{0}})=q_{g_{0}-1}$. The same can be repeated for all $g\leq g_{0}$. ∎
-In the following proposition we prove the existence of an homeomorphism on the inverse limit defined by the inverse graphs.
-Proposition 11.
+In the following proposition we prove the existence of an homeomorphism on the inverse limit defined by the directed graphs.
+Proposition 10.1.2.
 There exists a unique homeomorphism $\phi:X\to X$ with the following property.
@@
 Let $\underline{x}=(x_{g})\in X$.
-By Lemma 26, either $\underline{x}=\underline{q}=(q_{g})$ or there exists $g_{0}$ such that, for all $g\geq g_{0}$, $x_{g}\neq q_{g}$.
+By Lemma 10.1.1, either $\underline{x}=\underline{q}=(q_{g})$ or there exists $g_{0}$ such that, for all $g\geq g_{0}$, $x_{g}\neq q_{g}$.
 For the case when $\underline{x}=\underline{q}$, define $\phi(\underline{x})=\underline{q}$.
@@
 It is left to prove that $\phi$ is an homeomorphism.
-For the injectivity, take $\underline{x}=(x_{g}),\underline{\tilde{x}}=(\tilde{x}_{g})\in X$ with $\underline{x}\neq\underline{\tilde{x}}$.
 We start proving that if $\underline{x}\neq\underline{q}$ then $\phi(\underline{x})\neq\underline{q}$.
@@
 Since $\underline{x}\neq\underline{q}$, there exists $g_{0}$ such that, for all $g\geq g_{0}$ $x_{g}\neq q_{g}$.
-In particular, for all $g\geq g_{0}$, $x_{g}$ either is the precursor of $q_{g}$ in the $c$-loop or $x_{g}$ is the precursor of $q_{g}$ in the $h$-loop.
+In particular, for all $g\geq g_{0}$, $x_{g}$ either is the predecessor of $q_{g}$ in the $c$-loop or $x_{g}$ is the predecessor of $q_{g}$ in the $h$-loop.
 Denote this loop by $\lambda_{g}$.
 Take $g\geq g_{0}+1$, then there exists $z_{g}\in\lambda_{g}\setminus q_{g}$ such that $\overrightarrow{z_{g}q_{g}}\subset w_{g-1}^{-1}(q_{g-1})\cap\lambda_{g}$.
-Observe that $x_{g}$ cannot be contained in $\overrightarrow{z_{g}q_{g}}$, otherwise $w_{g-1}(x_{g})=q_{g-1}=x_{g-1}$.
-This contradicts the fact that $x_{g-1}\neq q_{g-1}$.
-Hence, if $\underline{x}\neq\underline{q}$, then $\phi(\underline{x})\neq\underline{q}$.
+In particular we can choose $z_{g}$ with the property that the length of $\overrightarrow{z_{g}q_{g}}$ is $\underline{\ell}_{g}$ which is larger than $2$.
+Hence, the predecessor of $q_{g}$ in $\lambda_{g}$ is contained in this path, i.e.
+$x_{g}$ is in $\overrightarrow{z_{g}q_{g}}$.
+This implies that $w_{g-1}(x_{g})=q_{g-1}=x_{g-1}$.
+Contradiction.
+We proved that if $\underline{x}\neq\underline{q}$, then $\phi(\underline{x})\neq\underline{q}$.
 […]
@@
 Suppose $\underline{y}\neq\underline{q}$.
-By Lemma 26, there exists $g_{0}$ such that, for all $g\geq g_{0}$, $y_{g}\neq q_{g}$ and $y_{g}$ has a unique precursor $x_{g}$.
-If $w_{g}(\overrightarrow{x_{g+1}y_{g+1}})=\overrightarrow{z_{g}y_{g}}$, by the uniqueness of the predecessor $z_{g}=x_{g}$ and $w_{g}(x_{g+1})=x_{g}$.
+By Lemma 10.1.1, there exists $g_{0}$ such that, for all $g\geq g_{0}$, $y_{g}\neq q_{g}$ and $y_{g}$ has a unique predecessor $x_{g}$.
+Let $w_{g}(\overrightarrow{x_{g+1}y_{g+1}})=\overrightarrow{z_{g}y_{g}}$.
+Then, by the uniqueness of the predecessor, $z_{g}=x_{g}$ and $w_{g}(x_{g+1})=x_{g}$.
 Hence, the sequence $\underline{x}=(x_{g})\in X$ and $\phi(\underline{x})=\underline{y}$.
@@
 Case 1.
-If $\underline{x}\neq\underline{q}$ then, by Lemma 26 there exists $g_{0}$ such that for all $g\geq g_{0}$ $x_{g}\neq q_{g}$ and $\phi(x_{g})=y_{g}$ is the unique successor of $x_{g}$.
+If $\underline{x}\neq\underline{q}$ then, by Lemma 10.1.1 there exists $g_{0}$ such that for all $g\geq g_{0}$ $x_{g}\neq q_{g}$ and $\phi(x_{g})=y_{g}$ is the unique successor of $x_{g}$.
 Moreover, since $\underline{x}_{n}\to\underline{x}$, there exists $n_{0}$ such that, for all $n\geq n_{0}$, $x_{g,n}=x_{g}$.
@@
 Let $y_{g+1,n}$ be the successor of $x_{g+1,n}=q_{g+1}$.
-Then $y_{g+1,n}\in Q_{g+1}\bigcup C_{g+1}(1)\bigcup H_{g+1}(1)\subset w_{g}^{-1}(q_{g})$.
+Then $y_{g+1,n}\in Q_{g+1}\bigcup C_{g+1}^{1}\bigcup H_{g+1}^{1}\subset w_{g}^{-1}(q_{g})$.
 This implies that $y_{g,n}=w_{g+1}(y_{g+1,n})=q_{g}$.
-Hence, for all $n\geq n_{0}$, $y_{g,n}=q_{g}$ and $\phi(x_{g,n})=q_{g}$.
-In conlcusion $\phi(\underline{x})\to\underline{q}$.
-The construction shows that the homeomorphism is unique.
-∎
+Hence, for all $n\geq n_{0}$, $y_{g,n}=q_{g}$.
+In conlcusion $\phi(\underline{x}_{n})\to\underline{q}$.
+The construction shows that the homeomorphism is unique. ∎
 As for the directed graphs, if for all $g$, $\mathfrak{D}_{g}$ is the $(1,N_{g},M_{g})$ dynamical partition of a map $f\in\ell$, $\mathcal{D}_{g}$ is the union of the elements of $\mathfrak{D}_{g}$ and $i_{g}:\mathcal{D}_{g+1}\to\mathcal{D}_{g}$ is the inclusion, we can define the inverse limit as
@@
 10.2 Diagram
-In this section we build a diagram structure as in Figure 12. This is the the core of the proof of Theorem 4.
+In this section we build a diagram as in Figure 13. This is the the core of the proof of Theorem 10.2.3.
 Figure 13: Diagram Structure
-Orserve that, for all $g$, $\pi_{g+1}\circ w_{g}=i_{g}\circ\pi_{g}$.
-Moreover, for all $g$, let $j_{g}:A\to\mathcal{D}_{g}$ be the inclusion and let $j:A\to\mathcal{D}$.
+Observe that, for all $g$, $\pi_{g+1}\circ w_{g}=i_{g}\circ\pi_{g}$.
+Moreover, for all $g$, let $j_{g}:A\to\mathcal{D}_{g}$ be the inclusion and denote the corresponding limit by $j:A\to\mathcal{D}$.
 By construction the map $j$ is an homeomorphism.
 Similarly define the projection $\pi:\mathcal{D}\to X$.
-Lemma 27.
+Lemma 10.2.1.
 The map $\pi:\mathcal{D}\to X$ is an homeomorphism.
@@
 Let ${a}\in\bigcap U_{g}$.
-Then $(a)\in\mathcal{D}$ and $\pi((a))=\underline{x}$.
+Then $a\in\mathcal{D}$ and $\pi(a)=\underline{x}$.
 This proves that $\pi$ is a surjective map.
-Moreover, since $\text{mesh}(U_{g})$ tends to zero, then $\left\{a\right\}=\bigcap U_{g}$. Hence $\pi^{-1}(\underline{x})=(a)$.
-The proof is complete. ∎
+Moreover, since $\text{mesh}(U_{g})$ tends to zero, then $\left\{a\right\}=\bigcap U_{g}$.
+Hence $\pi^{-1}(\underline{x})=(a)$.
+The proof is complete.
+∎
 We are now ready to define an homeomorphism $h:A\to X$ as
 $h=\pi\circ j.$
-Lemma 28.
+Lemma 10.2.2.
 The following diagram commutes,
@@
 Proof.
-Let $a\in A$. If $a=p$, then $h(f((p))=h(p)=q$ and $\phi(h(p))=\phi(q)=q$.
+Let $a\in A$.
+If $a=p$, then $h(f(p))=h(p)=q$ and $\phi(h(p))=\phi(q)=q$.
 Assume $a\neq p$, $a=\bigcap_{g}U_{g}$ where for all $g$, $U_{g}$ is a set in $\mathcal{D}_{g}$.
@@
 The following theorem proves Theorem A-(vi).
-Theorem 4.
-If $f$ and $\tilde{f}$ are in the same leave $\ell$ of the lamination $\mathfrak{L}$ then they are conjugate on their respective Cantor sets.
+Theorem 10.2.3.
+If $f$ and $\tilde{f}$ are in the same leaf $\ell$ of the lamination $\mathfrak{L}$ then they are conjugate on their respective Cantor sets.
 Proof.
 The dynamical partitions of both maps $f$ and $\tilde{f}$ give rise to the same $(1,N_{g},M_{g})$-directed graph and graph morphism $w_{g}:X_{g+1}\mapsto X_{g}$, for all $g\geq 1$.
-According to Lemma 28 both $A(f)$ and A(f~)A($\tilde${f}) are conjugate to the corresponding $X$.
+According to Lemma 10.2.2 both $A(f)$ and A(f~)A($\tilde${f}) are conjugate to the corresponding $X$.
 ∎
-Remark 15.
-Probably, the Cantor sets of maps on different leaves are not conjugated.
+Remark 10.2.4.
+Probably, the Cantor set of maps on different leaves are not conjugated.
 10.3 Codimension-1 Stability
-In this section we prove that the leaves of the lamination constructed above are the loci of infinitely renormalizable maps in the following sense.
-Recall, from the proof of Lemma 13, there exists a smooth function
-$m_{1}:S_{g}\mapsto\mathbb{R}^{2}$
+In this section, following the Remark 7.3.2, we prove that the leaves of the lamination constructed above are the loci of infinitely renormalizable maps in the following sense.
+Let $\ell=\cap S_{g}$ be a leaf of $\mathfrak{L}$.
+Recall, from the proof of Lemma 5.1.1, there exists a smooth function
+$m_{1}:S_{g}\to\mathbb{R}^{2},$
 such that
-  • $m_{1}\in W^{u}_{\text{loc}}(p)\cap C_{g}(0)$,
+  • $m_{1}\in W^{u}_{\text{loc}}(p)\cap C_{g}^{0}$,
   • $m_{7}=F^{3N_{g}+(1+\theta_{g})n_{g}+n_{0,g}}(m_{1})\in W^{u}(p)$
 and $W^{u}(p)$ has horizontal tangency at $m_{7}$.
-Definition 12.
-Let $\ell$ be a leave of type $\left\{\left(1,N_{g},M_{g}\right)\right\}_{g=1}^{\infty}$.
-A map $f\in S_{g}$ is renormalizable of type $(1,N_{g+1},M_{g+1})$ if $m_{7}(f)\in H_{g}(0)(f)$.
+Definition 10.3.1.
+Let $\ell=\cap S_{g}$ be a leaf of type $\left\{\left(1,N_{g},M_{g}\right)\right\}_{g=1}^{\infty}$.
+A map $f\in S_{g}$ is renormalizable of type $(1,N_{g+1},M_{g+1})$ if $m_{7}(f)\in H_{g}^{0}(f)$.
 Moreover, $f$ is infinitely renormalizable of type $\left\{\left(1,N_{g},M_{g}\right)\right\}_{g=1}^{\infty}$ if it is renormalizable of type $(1,N_{g+1},M_{g+1})$ for all $g\geq 1$.
-Definition 13.
+Definition 10.3.2.
 Let $f$ be an infinitely renormalizable map of type $\left\{\left(1,N_{g},M_{g}\right)\right\}_{g=1}^{\infty}$.
-The Cantor set $A(f)$ associated to $f$ is codimension-$k$ stable if there exists a codimension-$k$ manifold $W$ in parameter space containing $f$ such that every map in $W$ is infinitely renormalizable of type $\left\{\left(1,N_{g},M_{g}\right)\right\}_{g=1}^{\infty}$ and there does not exists a manifold of codimension k1k-1 with the same property.
+The Cantor set $A(f)$ associated to $f$ is codimension-$k$ stable if there exists a codimension-$k$ manifold $W$ in parameter space containing $f$ such that every map in $W$ is infinitely renormalizable of type $\left\{\left(1,N_{g},M_{g}\right)\right\}_{g=1}^{\infty}$ and there does not exists a manifold of codimension-k1k-1 with the same property.
 The manifold $W$ is called the manifold of stability.
-The following theorem proves Theorem A-(vii).
-Theorem 5.
-Let $\ell$ be a leave of type $\left\{\left(1,N_{g},M_{g}\right)\right\}_{g=1}^{\infty}$.
+The following theorem confirms Theorem A-(vii).
+Theorem 10.3.3.
+Let $\ell$ be a leaf of type $\left\{\left(1,N_{g},M_{g}\right)\right\}_{g=1}^{\infty}$.
 Every map $f\in\ell$ is infinitely renormalizable of type $\left\{\left(1,N_{g},M_{g}\right)\right\}_{g=1}^{\infty}$ and $A(f)$ is codimension-$1$ stable.
@@
 11 Invariant Measures
-In this section we prove that each set $A$ has a unique invariant probability measure. In particular, the following theorem proves Theorem A-(iv).
-Theorem 6.
+In this section we prove that each set $A$ has a unique invariant probability measure. In particular, the following theorem confirms Theorem A-(iv).
+Theorem 11.0.1.
 For each function $f\in\ell$, the set $A(f)$ has a unique invariant probability measure. This invariant measure is the point mass on the fixed point in $A(f)$ .
-The following corollary proves Theorem A-(i).
-Corollary 1.
+Corollary 11.0.2.
 The set $A$ is not hyperbolic.
@@
 it has a continuum of ergodic invariant measures.
-This contradicts Theorem 6.
+This contradicts Theorem 11.0.1.
 (Another reason is that $A$ has only one periodic orbit.) ∎
-The proof of Theorem 6 requires some preparation.
+The proof of Theorem 11.0.1 requires some preparation.
 Let $X$ be the $(1,N,M)$ directed graph.
@@
 $w_{g*}:\mathfrak{M}_{g+1}\to\mathfrak{M}_{g}.$
-As in Remark 14, the winding of $w_{g}$ is given by
-$\underline{w}\left(w_{g}\right)=\left(q,{q}^{\underline{\ell}_{g}}{c}{q}^{\theta_{g}n_{g}}{c}{q}^{n_{g}}{c}{q}^{n_{0,g}}{h}{q}^{\overline{\ell}_{g}},{q}^{\underline{\ell}_{g}}{h}{q}^{\overline{\ell}_{g}}\right).$
-Then the matrix for $w_{g*}$, in the standard basis of $\mathfrak{M}_{g}$ and $\mathfrak{M}_{g+1}$ is
-$W_{g}=\left(\begin{matrix}1&\underline{\ell}_{g}+\theta_{g}n_{g}+n_{g}+n_{0,g}+\overline{\ell}_{g}&\underline{\ell}_{g}+\overline{\ell}_{g}\\ 0&3&0\\ 0&1&1\end{matrix}\right)$
+As in Remark 7.4.4, the winding of $w_{g}$ is given by
+$\underline{\omega}\left(w_{g}\right)=\left(q,{q}^{\underline{\ell}_{g}}{c}{q}^{\theta n_{g}}{c}{q}^{n_{g}}{c}{q}^{n_{0,g}}{h}{q}^{\overline{\ell}_{g}},{q}^{\underline{\ell}_{g}}{h}{q}^{\overline{\ell}_{g}}\right).$
+Then the matrix for $w_{g*}$, in the standard basis of $\mathfrak{M}_{g}$ and $\mathfrak{M}_{g+1}$ can be easily read off from $\underline{\omega}\left(w_{g}\right)$ and it is
+$W_{g}=\left(\begin{matrix}1&\underline{\ell}_{g}+\theta n_{g}+n_{g}+n_{0,g}+\overline{\ell}_{g}&\underline{\ell}_{g}+\overline{\ell}_{g}\\ 0&3&0\\ 0&1&1\end{matrix}\right).$
 The simplex $\mathcal{P}_{g}\subset\mathfrak{M}_{g}$ consisting of probability measures is
@@
 $\mathcal{P}=\underleftarrow{\lim}\mathcal{P}_{g}.$
-Let $\mathcal{I}$ be the set of invariant probability measures of $f:X\to X$. If $\mu\in\mathcal{I}$ denote the restriction of the measure to the $\varsigma$-algebra of $X_{g}$ by
+Let $\mathcal{I}$ be the set of invariant probability measures of $f:X\to X$. If $\mu\in\mathcal{I}$ then the restriction of the measure to the $\sigma$-algebra of $X_{g}$ is denoted by
 $\mu_{g}=w_{g,\infty*}(\mu)\in\mathcal{P}_{g}.$
 Observe, $w_{g*}(\mu_{g+1})=\mu_{g}$. Hence, $\mu=(\mu_{g})\in\mathcal{P}$. By construction, the map
-$\mathcal{I}\ni\mu\mapsto(\mu_{g})\in\mathcal{P}$
-is a bijection. For more details on this construction see LABEL:GM.
+$\mathcal{I}\ni\mu\mapsto(\mu_{g})\in\mathcal{P}$ (11.0.3)
+is a bijection. For more details on this construction see [32].
 Denote the vertices of $\mathcal{P}_{g}$ by
-$\underline{\delta}^{c}_{g}=\frac{1}{N_{g}},\underline{\delta}^{h}_{g}=\frac{1}{M_{g}},\underline{\delta}^{q}_{g}=\delta^{q}_{g}.$
+$\underline{\delta}^{c}_{g}=\frac{1}{N_{g}}{\delta}^{c}_{g},\text{ }\underline{\delta}^{h}_{g}=\frac{1}{M_{g}}{\delta}^{h}_{g},\text{ }\underline{\delta}^{q}_{g}=\delta^{q}_{g}.$
 Observe,
 $\mathcal{P}_{g}=\text{hull}\{\underline{\delta}^{c}_{g},\underline{\delta}^{h}_{g},\underline{\delta}^{q}_{g}\}\subset\mathfrak{M}_{g}.$
-Define the norm $|.|$ on $\mathfrak{M}_{g}$ by
-$\mu=\max\{|x|,|y|,|z|\},$
+Define the norm $|\mu|$ of $\mu\in\mathfrak{M}_{g}$ by
+$|\mu|=\max\{|x|,|y|,|z|\},$
 where
@@
 $\mu(\lambda^{h}_{g}\setminus\lambda^{q}_{g})=y\cdot\frac{M_{g}-1}{M_{g}}.$
-The map $W_{g}:\mathcal{P}_{g+1}\to\mathcal{P}_{g}.$ contracts towards $\underline{\delta}^{q}_{g}$. Namely,
-Lemma 29.
+The map $W_{g}:\mathcal{P}_{g+1}\to\mathcal{P}_{g}$ contracts towards $\underline{\delta}^{q}_{g}$. Namely,
+Lemma 11.0.4.
 If $N_{g+1},M_{g+1}\geq 32\max\{N_{g},M_{g}\}$ and $N_{g},M_{g}\geq 3$ then for all $\mu\in\mathcal{P}_{g+1}$
@@
 Observe,
-$W_{g}\underline{\delta}^{c}_{g+1}(\lambda^{q}_{g})=\frac{\underline{\ell}_{g}+\theta_{g}n_{g}+n_{g}+n_{0,g}+\overline{\ell}_{g}}{N_{g+1}}$
+$W_{g}\underline{\delta}^{c}_{g+1}(\lambda^{q}_{g})=\frac{\underline{\ell}_{g}+\theta n_{g}+n_{g}+n_{0,g}+\overline{\ell}_{g}}{N_{g+1}}.$
 Because,
-$N_{g+1}=\underline{\ell}_{g}+N_{g}+\theta_{g}n_{g}+N_{g}+n_{g}+N_{g}+n_{0,g}+M_{g}+\overline{\ell}_{g}$
+$N_{g+1}=\underline{\ell}_{g}+N_{g}+\theta n_{g}+N_{g}+n_{g}+N_{g}+n_{0,g}+M_{g}+\overline{\ell}_{g}$
 we get
@@
 $|W_{g}\underline{\delta}^{c}_{g+1}-\underline{\delta}^{q}_{g}|\leq\frac{1}{2}.$
-Proof of Theorem 6: Let $\mu\in\mathcal{P}$, say $\mu=(\mu_{g})$.
-Recall, this measure corresponds to an invariant probability of $f$.
+Proof of Theorem 11.0.1: Observe that the conditions in Lemma 11.0.4 are guaranteed by Remark 9.1.4.
+Let $\mu\in\mathcal{I}$ be an invariant probability measure of $f$.
+In particular, $\mu$ corresponds to $(\mu_{g})\in\mathcal{P}$, see (11.0.3).
 We will show that $\mu_{g}=\underline{\delta}^{q}_{g}$, for all $g\geq 1$.
@@
 $|\mu_{g}-\underline{\delta}^{q}_{g}|\leq\frac{1}{2^{s}}|\mu_{g+s}-\underline{\delta}^{q}_{g+s}|\leq\frac{1}{2^{s}}.$
-Hence, $\mu_{g}=\underline{\delta}^{q}_{g}$, for all $g\geq 1$. This means that the only invariant measure of $f:A\to A$ is the unit mass on its fixed point.
-12 Properties of the Lamination: Collet-Eckmann and Newhouse points
-In this section, we prove the existence of a sub-lamination whose leaves have an interesting property. In particular we prove Theorem B and Theorem C.
-The invariant non-hyperbolic Cantor set, associated to each map in the leaves of this sublamination have a Collet-Eckmann point, which is dense in the Cantor set.
+Hence, $\mu_{g}=\underline{\delta}^{q}_{g}$, for all $g\geq 1$. This means that the only invariant measure of $f:A\to A$ is the unit mass on its fixed point. ∎
+12 Collet-Eckmann Lamination and Newhouse points
+In the previous sections we observed many different phenomena.
+In this section we show that all these phenomena occur in one and the same lamination.
+In particular there exists a lamination with a stable non-hyperbolic Cantor set which contains a Collet-Eckmann point with a dense orbit.
+Moreover each leave has also a Newhouse point.
+At this Newhouse points there is an intriguing coexistence of stable and unstable dynamics.
+This section confirms Theorem B and Theorem C.
 Let us start by giving the formal definition of a Collet-Eckmann point.
-Definition 14.
+Definition 12.0.1.
 Let $f:\mathcal{M}\to\mathcal{M}$ be a map and let $x\in\mathcal{M}$. We say that $x$ is a Collet-Eckmann (CE) point if there exists a vector $v$, in the tangent space at $x$, such that
@@
 where $C$ and $\rho$ are two constants, $C>0$ and $\rho>1$.
-The following Theorem proves Theorem B.
-Theorem 7.
+The following Theorem confirms Theorem B.
+Theorem 12.0.2.
 There exists a sub-lamination $\mathfrak{L}_{CE}\subset\mathfrak{L}$ satisfying the following property.
 For each function $f\in\ell\subset\mathfrak{L}_{CE}$, the invariant non-hyperbolic Cantor set $A(f)$ contains a Collet-Eckmann point $c$ such that $\omega(c)=A(f)$.
-Moreover the closure $\overline{\mathfrak{L}_{CE}}=\overline{\mathfrak{L}}$.
+Moreover, every leaf of $\mathfrak{L}_{CE}$ is the graph of a smooth function $\ell:[-t_{0},t_{0}]\mapsto[-a_{0},a_{0}]$ and if $t\in[-t_{0},t_{0}]$, the closure of the set $L^{CE}_{t}=\left\{(t,a)\in\mathfrak{L}_{CE}\right\}$ is a Cantor set.
 Proof.
 The proof goes by induction on the generations $g$ defining the lamination.
-The induction step is contained in Proposition 8.
+The induction step is contained in Proposition 8.0.4.
 Using the notation in that proposition we prove the base of induction, i.e., we prove that the conditions are verified at generation $g=1$.
@@
 Take $U=S$.
-In Subsection 2.1, we defined a family $F$ which is an unfolding of the tangency $b$ and each map $f\in S$ has a $(1,N_{1},M_{1})$ dynamical partition with $N_{1}=N$ and $M_{1}=M$, where $N$ and $M$ are defined in Definition 2, property $(f10)$.
+In Subsection 2.1, we defined a family $F$ which is an unfolding of the tangency $b$ and each map $f\in S$ has a $(1,N_{1},M_{1})$ dynamical partition with $N_{1}=N$ and $M_{1}=M$, where $N$ and $M$ are defined in Definition 2.1.1, property $(f8)$.
 Define $T=0$ and for all $f\in U$ let $K$ be a positive constant assuring that $\text{IH}_{H^{0}}$, $\text{IH}_{C^{0}}$-(iii) and $\text{IH}_{C^{N}}$-(iii) hold.
@@
 This concludes the base of induction.
-As said above, the induction step is Proposition 8.
+As said above, the induction step is Proposition 8.0.4.
 We can now draw our conclusions.
@@
 $c=\bigcap_{g}C^{N_{g}-T_{g}}\in A(f).$
-Since $T_{g}$ goes to infinity, $c$ is a Collet-Eckmann point. The proof of the density is the same as the proof of Lemma 24. ∎
+Since $T_{g}$ goes to infinity, $c$ is a Collet-Eckmann point. The proof of the density is the same as the proof of Lemma 9.3.4. ∎
 Even more surprisingly, there exists a sub-lamination of $\mathfrak{L}_{CE}$ whose leaves contains a map with infinitely many sinks of higher and higher order accumulating at the orbit of the Collet-Eckmann point.
-The following Theorem proves Theorem C.
-Theorem 8.
+The following Theorem confirms Theorem C.
+Theorem 12.0.3.
 There exists a sub-lamination $\mathfrak{L}_{N}\subset\mathfrak{L}_{CE}\subset\mathfrak{L}$ satisfying the following property.
-Each leave $\ell\subset\mathfrak{L}_{N}$ contains a map $f_{\ell}$ with infinitely many attracting periodic points, $p_{k}$ of arbitrarily high period, such that,
-$\bigcap_{n}\overline{\bigcup_{k\geq n}\text{Orb}(p_{k})}=A(f_{\ell}).$
-Moreover if $NC=\left\{f_{\ell}\right\}_{f_{\in}\ell\subset\mathfrak{L}_{N}}$, then the closure $\overline{{NC}}=\overline{\mathfrak{L}_{CE}}=\overline{\mathfrak{L}}$.
+Each leaf $\ell\subset\mathfrak{L}_{N}$ contains a map $f_{\ell}$ with infinitely many attracting periodic points, $p_{g}$ of arbitrarily high period, such that,
+$\bigcap_{n}\overline{\bigcup_{g\geq n}\text{Orb}(p_{g})}=A(f_{\ell}).$
+Moreover, every leaf of $\mathfrak{L}_{N}$ is the graph of a smooth function $\ell:[-t_{0},t_{0}]\mapsto[-a_{0},a_{0}]$ and if $t\in[-t_{0},t_{0}]$, the closure of the set $L^{N}_{t}=\left\{(t,a)\in\mathfrak{L}_{N}\right\}$ is a Cantor set.
 Proof.
-The theorem is consequence of Theorem 7 and Theorem 3. ∎
+The theorem is a consequence of Theorem 12.0.2 and Theorem 6.2.7.
+More in details, using the terminology from Subsection 6.2, given a point $(t,a)\in NH$ there exists a sequence of Newhouse boxes such that
+$\left\{(t,a)\right\}=\bigcap_{g\geq 1}\mathcal{P}^{g}_{n^{(g)},n_{0}^{(g)}}.$
+Associated to each Newhouse box $\mathcal{P}^{g}_{n^{(g)},n_{0}^{(g)}}$ there is a secondary tangency curve which is the graph of $b^{g}_{n^{(g)},n_{0}^{(g)}}$.
+By taking $n^{(g)}$ large enough we can assure that these curves converge in ${\mathcal{C}}^{1}$} to a leaf $\ell_{(t,a)}$ which is also the graph of a ${\mathcal{C}}^{1}$} function over $t_{0},t_{0}]$.
+In particular $(t,a)\in\ell_{(t,a)}$.
+Let
+$\mathfrak{L}_{N}=\bigcup_{(t,a)\in NH}\ell_{(t,a)}.$
+By taking $n^{(g)}$ large enough we can assure that $\mathfrak{L}_{N}\subset\mathfrak{L}_{CE}$.
+Let $p_{g}$ be the periodic sink of generation $g$.
+By construction the orbit of $p_{g}$ is contained in the $c$- and $q$-loop of the dynamical partition $\mathfrak{D}_{g}$.
+Hence, the sinks accumulate at the Cantor set $A$.
+∎
 13 Appendix
-We prove that there exists a $\theta$ satisfying (8.1).
-Lemma 30.
+We prove that there exists a $\theta$ satisfying (8.0.2).
+Lemma 13.0.1.
 There exists $\theta\in\left(\theta_{0},\theta_{1}\right)$ such that
@@
 Proof.
-By (4.17), there exists $K>1$ such that
-$\alpha=\frac{\log K}{\log\mu}+2\theta\frac{\log\lambda}{\log\mu}+3.$
+By (4.1.47),
+$\alpha=2\theta\frac{\log\lambda}{\log\mu}+3.$
 We need to prove that
@@
 We substitute the value of $\alpha$ and we get
-$\log K\left(1-\frac{\log\lambda}{\log\mu}\right)-3\log\lambda+4\log\mu>\theta\left(2\frac{(\log\lambda)^{2}}{\log\mu}-3\log\lambda\right).$
+$3\log\lambda+4\log\mu>\theta\left(2\frac{(\log\lambda)^{2}}{\log\mu}-3\log\lambda\right).$
 The inequality is satisfied if one substitute the value of $\theta_{0}=-4\log\mu/3\log\lambda$.